### v2

# Potential weights and implicit causal designs in linear regression

Abstract.
When we interpret linear regression estimates as causal effects justified by quasi-experiments, what do we mean? This paper characterizes the necessary implications when researchers ascribe a design-based interpretation to a given regression. To do so, we define a notion of potential weights, which encode counterfactual decisions a given regression makes to unobserved potential outcomes. A plausible design-based interpretation for a regression estimand implies linear restrictions on the true distribution of treatment; the coefficients in these linear equations are exactly potential weights. Solving these linear restrictions leads to a set of implicit designs that necessarily include the true design if the regression were to admit a causal interpretation. These necessary implications lead to practical diagnostics that add transparency and robustness when design-based interpretation is invoked for a regression. They also lead to new theoretical insights: They serve as a framework that unifies and extends existing results, and they lead to new results for widely used but less understood specifications.
1. Introduction
Much of economic research, particularly in applied microeconomics, emphasizes research design and causal identification. The distinguishing feature of this school of thought—which we will refer to as “design-based”—is an experimentalist focus on reasoning with plausible randomization of treatment, so as to mimic some randomized control trial. [^1] By May 2024, over 50% of papers in the NBER working paper series mentions experimental or quasi-experimental methods (Goldsmith-Pinkham, 2024; Currie et al., 2020). We refer to “design-based” as a focus on modeling treatment assignment, in contrast to modeling potential outcomes (which is often called “model-based”). This is consistent with the terminology adopted in Borusyak et al. (2024a), Borusyak and Hull (2024), and Angrist and Pischke (2010). In particular, Angrist and Pischke (2010) observe, “In applied micro fields such as development, education, environmental economics, health, labor, and public finance, researchers seek real experiments where feasible, and useful natural experiments if real experiments seem (at least for a time) infeasible. In either case, a hallmark of contemporary applied microeconometrics is a framework that highlights specific sources of variation. These studies can be said to be design based in that they give the research design underlying any sort of study the attention it would command in a real experiment.” In observational studies, researchers often argue that the treatment assignment is plausibly as-if random, so that causal effects are identified as they are in randomized experiments. This design-based mode of identification has certain virtues: It is valid without functional form assumptions on counterfactual outcomes, which may be harder to motivate than restrictions on treatment assignment; it also allows causal identification for different outcome variables subjected to a common intervention, without modeling each individually.
Despite the popularity of this design-based reasoning in terms of identification, the same is not true for estimation. Economic research continues to overwhelmingly prefer linear regression methods to other methods that are arguably more “natively” design-based (Imbens, 2015). [^2] These other methods—including inverse propensity-score weighting, matching (Zubizarreta et al., 2014; Abadie and Imbens, 2011), doubly robust (Bang and Robins, 2005), or balancing methods (Ben-Michael et al., 2021; Li et al., 2018)—explicitly use the probability of treatment assignment to target some explicit treatment effect. If one defends identification on design-based grounds, it seems natural that estimation should model the process of treatment assignment and target an estimand defined in terms of unrestricted potential outcomes. In contrast, linear regression does neither—at least not explicitly. Indeed, Freedman (2008) criticizes regressions in experimental settings precisely because treatment randomization does not imply the linear model.
To address this tension, a modern perspective views linear regressions as simply estimators to be analyzed under certain designs (i.e., process of treatment assignment). In this view, linear regression is essentially a machine that compares different groups of units. Under the right design, these comparisons recover causal effects—often weighted average treatment effects with weights implicitly chosen by the regression specification.
Much has been written on the properties of various linear specifications under known restrictions on treatment assignment (among others, Imbens and Angrist, 1994; Angrist, 1998; Lin, 2013; Słoczyński, 2022, 2020; Blandhol et al., 2022; Aronow and Samii, 2016; Goldsmith-Pinkham et al., 2024; Borusyak and Hull, 2024; Athey and Imbens, 2022; Kline, 2011; Bugni et al., 2023; Mogstad and Torgovitsky, 2024; Arkhangelsky and Imbens, 2023; Arkhangelsky et al., 2021; Chetverikov et al., 2023; Kolesár and Plagborg-Møller, 2024). This line of research analyzes particular regression specifications under a known set of designs and examines whether a specification makes sensible comparisons under the design. It often provides sufficient conditions for design-based interpretation of linear estimands—or cautionary tales for the lack thereof. A limitation is that this literature produces a hodge-podge of results: Certain specifications have design-based interpretations under certain restrictions, but other configurations of specification and design may produce nonsensical estimands (e.g., negatively weighted causal effects).
Complementary to this literature, this paper instead asks what are the necessary implications when researchers ascribe a design-based interpretation to a particular regression specification. Suppose a researcher presents a regression and argues that certain coefficients estimate causal effects under some design—some process of random treatment assignment—without restrictions on potential outcomes. However, as is frequently the case, our colleague does not explicitly specify the treatment randomization process nor the target estimand. For their design-based interpretation to be tenable given the data, what must our colleague be prepared to defend?
Specifically, this paper provides generic tools that answer the following questions:
  1. (Q1)
    Is our colleague obviously mistaken and the design-based interpretation indefensible?
  2. (Q2)
    If a design-based interpretation is at least plausible, for it to be appropriate, what must be true of the design when it is not explicitly specified?
  3. (Q3)
    Given a plausible designs, what is the estimand that the specification targets?
Our approach is constructive, computational, and general: It computes concrete designs and estimands that answer (Q2) and (Q3) without assuming much structure on the specification (allowing for panels, cross-sections, binary, or multi-valued treatments). In leading cases, it reveals at most one design and at most one estimand; design-based interpretation is indefensible (Q1) if no design and estimand is revealed. As a result, our diagnostics are highly informative on what design and economic target that regression-users must justify.
Our results arise from two simple observations. First, linear regression estimands $\bm{\tau}$ under a given design are linear aggregations of individual potential outcomes $\mathbf{y}_{i}(\cdot),i=1,\ldots,n$. In this aggregation, the weight placed on $\mathbf{y}_{i}(\mathbf{w})$ can be written in the form $\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})$, which does not involve the outcomes: That is, $\bm{\tau}$ can be written as
$\underbrace{\bm{\tau}}_{\text{Regression estimand, }\mathbb{R}^{k}}=\frac{1}{n}\sum_{i=1}^{n}\underbrace{\sum_{\mathbf{w}}}_{\begin{subarray}{c}\text{Summing over}\\ \text{treatment values $\mathbf{w}$}\end{subarray}}\underbrace{\pi_{i}^{*}(\mathbf{w})}_{\text{true design, }\mathbb{R}}\cdot\underbrace{\bm{\rho}_{i}(\mathbf{w})}_{\text{potential weights, }\mathbb{R}^{k\times T}}\cdot\underbrace{\mathbf{y}_{i}(\mathbf{w})}_{\text{potential outcomes, }\mathbb{R}^{T}}.$
These weights depend both on the design itself—the unknown probability $\pi_{i}^{*}(\mathbf{w})$ that individual $i$ receives treatment $\mathbf{w}$—and on certain numbers $\bm{\rho}_{i}(\mathbf{w})$, which we call potential weights. These potential weights encode how a regression specification makes comparisons over counterfactual treatment assignments. Importantly, the potential weights $\bm{\rho}_{i}(\mathbf{w})$ are known up to the population Gram matrix of the linear regression (i.e. the “$\mathbb{E}[x_{i}x_{i}^{\prime}]$” matrix). They are thus known in the population and estimable from the sample.
Second, since not all aggregations of potential outcomes are bona fide causal effects, interpreting this aggregation (1) as causal necessarily restricts the weights $\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})$ placed on the potential outcomes. This in turn implies restrictions on $\pi_{i}^{*}$. Suppose by a causal contrast, we mean comparisons among potential outcomes that satisfy level irrelevance under the true treatment assignment process (Blandhol et al., 2022). That is, an estimand is a causal contrast if shifting every potential outcome of any unit by the same number never mutates the estimand.
Crucially, it turns out that if $\bm{\tau}$ were to be a causal contrast under $\bm{\pi}^{*}$, then for every individual, over all treatment values $\mathbf{w}$, the weights on $\mathbf{y}_{i}(\mathbf{w})$ necessarily sum to zero
For every $\text{For every $i=1,\ldots,n$, }\sum_{\mathbf{w}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=0.$
Since the potential weights $\bm{\rho}_{i}(\mathbf{w})$ are known, these restrictions are linear restrictions for the true design $\pi_{i}^{*}(\mathbf{w})$. They represent precisely what we ought to defend about $\pi^{*}_{i}(\mathbf{w})$ if we would like to interpret the linear estimand as a causal contrast. Treating (1) as linear equations in the unknown $\bm{\pi}_{i}^{*}(\mathbf{w})$ with known coefficients $\bm{\rho}_{i}(\mathbf{w})$, we call the solutions to these equations the implicit design of the specification, and distinguish them from the true assignment process $\bm{\pi}_{i}^{*}(\mathbf{w})$ that Nature chooses. In leading settings, there are more equations in (1) than there are unknowns, meaning that the set of implicit designs is often either empty or a singleton.
Implied by these observations, our core result is simply: The estimand $\bm{\tau}$ is a vector of causal contrasts under the true design $\left\{{\pi^{*}_{i}(\cdot)}\right\}$ if and only if the true design is equal to some implicit design $\left\{{\pi_{i}(\cdot)}\right\}$. Because the potential weights in (1) are known up to an estimable Gram matrix, we can simply compute the set of implicit designs $\pi_{i}(\mathbf{w})$. This in turn is highly informative of the true design $\pi_{i}^{*}(\mathbf{w})$ when the estimand $\bm{\tau}$ is thought to be causal contrasts. Thus, returning to (Q1)(Q3):
  1. (A1)
    The causal interpretation is indefensible if (1) cannot be solved with probability distributions.
  2. (A2)
    Implicit designs are exactly those candidate designs that one must defend while defending a design-based interpretation.
  3. (A3)
    Plugging plausible implicit designs as $\pi_{i}^{*}(\cdot)$ into (1) reveals the causal estimands targeted by the regression, and the researcher should likewise defend their economic relevance.
Our results inform the debate between critics of regression and those who are more sanguine on regression’s utility. [^3] At risk of oversimplification, the critics argue that empiricists should target clearly defined economic parameters; thus, the practice of fixing an estimation method (typically regression) and defining the estimand through the estimation method is dubious (This sentiment is present in, e.g., Blandhol et al., 2022; Mogstad and Torgovitsky, 2024; Yitzhaki, 1996; Freedman, 2008; Imbens, 2015; Heckman and Urzua, 2010; Manski and Tamer, 2019; Haile, 2018). In contrast, others argue that regressions provide practical approximations to structural quantities—whether as outcome models or as implicit models of design—and in many cases target reasonable and nonparametrically defined economic quantities; moreover, understanding the properties of regression without imposing a linear model on the outcomes is valuable (This sentiment is present in, e.g., Goldsmith-Pinkham et al., 2024; Angrist and Pischke, 2008; Angrist, 2001; Imbens and Angrist, 1994; Kline, 2011; Imbens, 2010; Lin, 2013; Borusyak and Hull, 2024). The heart of this debate is whether it is harmless or reckless for researchers to let regression make certain choices on autopilot. This paper advances this debate with the simple observation that these implicit choices—how a regression models the assignment process and chooses the economic parameter—are in part knowable and computable. Depending on the context, these choices may be difficult to defend or they may be reasonable and robust to minor changes. At a minimum—and especially since regressions are overwhelmingly popular—these choices should be made transparent to scrutiny.
To that end, we propose a set of practical tools to assess and improve potential weights, implicit designs, and implicit estimands, which we illustrate with an application to Blakeslee et al. (2020). If a plausible implicit design is found and the corresponding estimand sensible, then the regression specification plausibly makes reasonable choices—thereby dispelling concerns like “negative weights” altogether. If the regression reveals a defensible design but not an economically motivated estimand, the researcher can conveniently construct estimates for alternative estimands using the implicit design; these alternative estimates are comparable to the regression estimates because they have the same implicit design. Finally, the researcher can also assess whether the outcomes and treatment effects correlate with the implicit design in a given dataset. If there is substantial treatment effect heterogeneity correlated with implicit designs, then different treatment effect estimands are likely to be substantially different. In such cases, treatment effect heterogeneity would be a pertinent concern for the interpretation of the regression estimand.
If no defensible implicit designs are found, empirical researchers are better served either by “design-native” methods that explicitly model the process of treatment assignment (see Wager, 2024, for a review of the modern literature on estimating causal effects), or by defending the regression specification as restrictive models of potential outcomes. Interpreting the regression estimates as design-based is no longer tenable. Nevertheless, if the original regression encodes substantive knowledge of the assignment process that practitioners are unwilling to jettison, it is possible to at least nudge things in a design-based direction by recalibrating its implicit design. Such an approach, which we term design patching, aims to produce an estimator that is at least more defensible than the original regression, while remaining close to the original regression—and thus doing no harm. As the name we choose connotes, patching’s inelegance should not impugn its occasional practicality.
When applied to some common specifications, our framework uncovers new theoretical insights. First, we unify and extend many results in the literature in starkly distinct settings, by simply computing potential weights and implicit designs (Angrist, 1998; Blandhol et al., 2022; Goldsmith-Pinkham et al., 2024; Kline, 2011; Athey and Imbens, 2022). For these specifications, our framework can supply a converse of sorts: We show that the only designs justifying causal interpretations are exactly those analyzed in the literature. Second, we show that specifications that interact a binary treatment with covariates have delicate design-based interpretations (Lin, 2013; Miratrix et al., 2013; Imbens and Wooldridge, 2009; Kline, 2011). Different designs are necessary for causal interpretations of different contrasts of the coefficients (Chattopadhyay and Zubizarreta, 2023). These designs are often mutually incompatible: Requiring all contrasts to be causal necessitates generally logically inconsistent restrictions on the design. Lastly, we show that two-way fixed effects (TWFE) specifications have fragile and stringent design-based interpretations under staggered adoption. Adding time-varying covariates or estimating on imbalanced panels can render the equations (1) unsolvable; even without these features, the only solution to (1) involves totally randomized treatment paths.
This paper is particularly related to Chattopadhyay and Zubizarreta (2023), who characterize and interpret the weights on observed outcomes for several leading linear regression specifications. Our perspective departs from theirs in at least two ways: First, we treat their weights as the observed counterpart to the potential weights, much like observed outcomes are potential outcomes evaluated at the observed treatment. Second, we consider necessary conditions for interpreting the regression as estimating some causal effect, whereas Chattopadhyay and Zubizarreta (2023) focus on sufficient conditions for interpreting regression as estimating average treatment effect.
This paper proceeds as follows. Section 2 contains our main results. To build intuition, Section 2.1 starts with the simple setting with cross-sectional data and binary treatments. Section 2.2 contains our main results, which analyze regressions in the population. Section 2.3 discusses a few practical takeaways, and Section 2.4 connects the population to the sample with consistency and inference results. Section 3 then applies our framework to a litany of regression specifications, yielding new theoretical results. Section 4 illustrates our diagnostics with Blakeslee et al. (2020), who study the economic impact of water loss in rural India, using plausibly as-if random variation. Section 5 concludes.
2. Potential weights and implicit designs
Consider a finite population of units $i\in[n]\equiv\left\{{1,\ldots,n}\right\}$. Each individual receives one of $J+1$ treatments $\mathbf{w}\in\mathcal{W}$. $\mathcal{W}$ is some generic finite set with $J+1<\infty$ elements. We denote by $\mathbf{W}_{i}$ the realized treatment of unit $i$. Each unit is associated with a vector of covariates $\mathbf{x}_{i}$ and $J+1$ vector-valued potential outcomes of length $T$, $\left\{{\mathbf{y}_{i}(\mathbf{w})\in\mathbb{R}^{T}:\mathbf{w}\in\mathcal{W}}\right\}$, without restrictions. [^4] For expositional clarity, we assume that the dimension of the outcome vector is the same across individuals (i.e., balanced panels). Our results in Section 3.3 do discuss imbalanced panels. Upon treatment realization, we observe a corresponding realized outcome $\mathbf{Y}_{i}=\mathbf{y}_{i}(\mathbf{W}_{i})$.
We isolate the variation coming from the process of treatment assignment by thinking of $(\mathbf{x}_{i},\mathbf{y}_{i}(\cdot))$ as fixed numbers and only considering the randomness in $\mathbf{W}_{i}$. [^5] One could imagine drawing $(\mathbf{x}_{i},\mathbf{y}_{i}(\cdot))\overset{\mathrm{i.i.d.}}{\sim}P_{0}$ from some population and then subsequently conditioning on the drawn values. To translate between the finite-population and superpopulation perspectives, it is usually sufficient to replace averaging over the finite population ($\frac{1}{n}\sum_{i=1}^{n}$) with the population expectation over $\mathbb{E}_{\mathbf{x}_{i},\mathbf{y}_{i}(\cdot)\sim P_{0}}$. Unfortunately, from a statistical inference perspective, this finite population setup that isolates variation from treatment assignment is often termed “design-based” as well (Abadie et al., 2020; Li and Ding, 2017; Rambachan and Roth, 2020). This is distinct from the sense that we mean. Our focus is on justifying identification through variation in treatment rather than on characterizing statistical uncertainty. To that end, let $\Pi^{*}$ denote the joint distribution of $(\mathbf{W}_{1},\ldots,\mathbf{W}_{n})$ and let $\bm{\pi}^{*}$ denote the associated marginal treatment assignment probabilities (i.e., propensity scores):
$\bm{\pi}^{*}=(\pi^{*}_{1},\ldots,\pi^{*}_{n})\text{ where }\pi_{i}^{*}(\mathbf{w})=\operatorname{P}_{\Pi^{*}}(\mathbf{W}_{i}=\mathbf{w}).$
We refer to such a collection $\bm{\pi}$ of marginal treatment probabilities as a design. We call $\bm{\pi}^{*}$ the true design and $\Pi$ a joint design. Because we focus on linear estimators, only $\bm{\pi}^{*}$ matters for population properties; $\Pi^{*}$ in turn matters for the statistical properties of the regression estimator.
We think of $(\Pi^{*},\bm{\pi}^{*})$ as unknown and chosen by Nature; they are also unknowable since only a single draw $\mathbf{W}_{i}$ from $\pi_{i}^{*}(\cdot)$ is ever observed. $\bm{\pi}^{*}$ could in principle be in arbitrary configuration with the covariates and the potential outcomes—meaning that units select into treatment in an arbitrary fashion. However, if a researcher would like to defend a regression estimand on design-based grounds, they should then be prepared to defend restrictions on $\Pi^{*}$ consistent with data. We precisely uncover these restrictions.
Linear regressions are generically written as $\mathbf{Y}_{it}$ regressed on some known transformations of covariates and treatment $z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})\in\mathbb{R}^{K}$ with coefficients $\beta\in\mathbb{R}^{K}$:
$\mathbf{Y}_{it}=z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}\beta+\epsilon_{it}.$
We are interested in interpreting a subvector of regression contrasts $\bm{\tau}=\Lambda\beta\in\mathbb{R}^{k}$ as a vector of causal effects, for a known matrix $\Lambda\in\mathbb{R}^{k\times K}$, $k\leq K$. Importantly, the regression is not treated as a model for $\mathbf{y}_{i}(\cdot)$, but simply as an algorithm that—in the population—produces some $\bm{\tau}$ given $(\Pi^{*},(\mathbf{y}_{i}(\cdot),\mathbf{x}_{i})_{i=1}^{n})$.
This setup is sufficiently general to encompass cross-sectional ($T=1$), panel ($T>1$), binary treatment ($J+1=2$), multivalued treatment ($J+1>2$), scalar contrast ($k=1$), and multiple contrasts ($k>1$) settings. [^6] Our results do extend to settings with infinitely-valued $\mathbf{w}$ (e.g. continuous treatments), but they become much less powerful, essentially since there are only finitely many restrictions for infinitely many objects. To flesh out the intuition in (1) and (1), we start informally with the binary-treatment, scalar-outcome, and scalar-contrast case $(T=J=k=1)$. Our main results push this essential intuition to the more general case and formalize the exercise.
2.1. Leading example and core intuition
We do not bold variables in this familiar setting. To further fix ideas, a popular regression in this case is:
$Y_{i}=\alpha+\tau W_{i}+x_{i}^{\prime}\gamma+\epsilon_{i},$
for which $z(x_{i},W_{i})=[1,W_{i},x_{i}^{\prime}]^{\prime},\beta=[\alpha,\tau,\gamma^{\prime}]^{\prime},\Lambda=[0,1,0_{\dim(x)}^{\prime}].$ Since treatment is binary, for a design $\bm{\pi}$, we write $\pi_{i}=\pi_{i}(1)$ and $1-\pi_{i}=\pi_{i}(0)$.
As an empirical motivation for our exercise, consider Blakeslee et al. (2020), who study the impact of water loss in rural India on employment and income. Water loss is proxied by a binary $W_{i}$, indicating whether the first borewell household $i$ drilled has failed at the time of the authors’ survey. Blakeslee et al. (2020) appear to argue for identification via design-based reasoning, writing, e.g., “a household’s access to groundwater depends on highly irregular and quasi-random properties of the subsurface beneath its land” (p.206, Blakeslee et al., 2020).
The true design $\bm{\pi}^{*}$—the natural process of borewell failure—is unknown, but the authors argue that it nevertheless does not preclude valid causal inference. Since borewell failure—a function of complex hydrogeological features—is difficult to predict, households with failed wells are otherwise similar to their more fortunate counterparts. To buttress this argument, the authors painstakingly and carefully describe hydrogeological features that determine borewell failure, even having inserted specialized camera equipment to failed wells.
As is common practice, however, the authors estimate treatment effects via regressions of the form (2.1), for many different outcome variables. This appears somewhat in tension with the rich descriptions of treatment assignment: These regressions are not connected—at least not explicitly—to those rich descriptions of $\bm{\pi}^{*}$ in the paper. What should a reader believe about borewell failure in order to interpret the authors’ regression estimates as causal effects?
Consider a regression of $Y_{i}$ on $z(x_{i},W_{i})$. The population regression coefficient is defined as an analogue of the familiar regression formula (cf. $\mathbb{E}[x_{i}x_{i}^{\prime}]^{-1}\mathbb{E}[x_{i}y_{i}]$)
$\beta=\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{z(x_{i},W_{i})z(x_{i},W_{i})^{\prime}}\right]}\right)^{-1}\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}[z(x_{i},W_{i})y_{i}(W_{i})]}\right).$
The two terms in $\beta$ average over the finite population ($\frac{1}{n}\sum_{i=1}^{n}$) and integrate over the only random objects $W_{1},\ldots,W_{n}$ in our setup.
Critically, let
$G_{n}\equiv G_{n}(\bm{\pi}^{*})\equiv\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{z(x_{i},W_{i})z(x_{i},W_{i})^{\prime}}\right]$
denote the Gram matrix of this regression. $G_{n}$ is a population object that depends on $\bm{\pi}^{*}$. However, since $G_{n}$ is an aggregation of the individually unknowable $\pi_{i}^{*}$, the Gram matrix can usually be estimated well with its sample counterpart $\hat{G}_{n}=\frac{1}{n}\sum_{i=1}^{n}z(x_{i},W_{i})z(x_{i},W_{i})^{\prime}$. [^7] In particular, since the regression estimator replaces $G_{n}$ with $\hat{G}_{n}$, it is implausible that the regression estimator is satisfactory (e.g., consistent) but $\hat{G}_{n}$ is far from $G_{n}$. Sections 2.4 and B.2 provide formal guarantees for $\hat{G}_{n}$. Therefore, much like in analyses of identification, we abstract away from the estimation noise and treat $G_{n}$ as known. This known-$G_{n}$ perspective is crucial for our results: We shall provide additional arguments for it at various points later.
With these preparations, observe that the estimand $\tau$ can be written in the form of (1):
$\tau=\Lambda\beta=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}\underbrace{\Lambda G_{n}^{-1}z(x_{i},1)}_{\rho_{i}(1)}y_{i}(1)+(1-\pi_{i}^{*})\underbrace{\Lambda G_{n}^{-1}z(x_{i},1)}_{\rho_{i}(0)}y_{i}(1).$
Here, the potential weights $\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)$ are known up to $G_{n}$. Just like potential outcomes $y_{i}(w)$, the potential weights are indexed by the unit and the treatment assignment.
Potential weights are readily connected to the weighting in the regression estimator. The least-squares estimator—itself linear in $Y_{i}$—can be written as
$\hat{\tau}=\Lambda\hat{\beta}=\Lambda\hat{G}_{n}^{-1}\frac{1}{n}\sum_{i=1}^{n}z(x_{i},W_{i})Y_{i}=\frac{1}{n}\sum_{i=1}^{n}\underbrace{\Lambda\hat{G}_{n}^{-1}z(x_{i},W_{i})}_{\hat{\rho}_{i}(W_{i})}y_{i}(W_{i}).$
The weight on $Y_{i}$ in $\hat{\tau}$ corresponds exactly to the empirical counterpart of $\rho_{i}(W_{i})$—the potential weights for $i$ evaluated at $i$’s observed outcome. Comparing with (2.1) and ignoring the difference $\hat{G}_{n}\neq G_{n}$, we can think of the potential weight $\rho_{i}(1-W_{i})$ of the counterfactual treatment $1-W_{i}$ as what the regression estimator would have done to the potential outcome $y_{i}(1-W_{i})$, had unit $i$ received the opposite assignment instead. Thus, in this sense, the potential weights encode how the regression compares the two potential outcomes.
Specializing (2.1) to (2.1), the potential weights are
$\rho_{i}(w)=\frac{w-\tilde{x}_{i}^{\prime}\delta}{\frac{1}{n}\sum_{i=1}^{n}\tilde{x}_{i}^{\prime}\delta(1-\tilde{x}_{i}^{\prime}\delta)}\text{ for }\tilde{x}_{i}\equiv[1,x_{i}^{\prime}]^{\prime},\,\delta\equiv\left({\frac{1}{n}\sum_{i=1}^{n}\tilde{x}_{i}\tilde{x}_{i}^{\prime}}\right)^{-1}\left({\frac{1}{n}\sum_{i=1}^{n}\tilde{x}_{i}\pi_{i}^{*}}\right).$
The $\rho_{i}(w)$ is proportional to $w$ minus the population projection of $\pi_{i}^{*}$ onto the covariates, whose projection coefficient is $\delta$. Analogously, we may compute that $\hat{\rho}_{i}(W_{i})$ replaces $w$ with $W_{i}$ and $\delta$ with its sample counterpart $\hat{\delta}=\left({\frac{1}{n}\sum_{i=1}^{n}\tilde{x}_{i}\tilde{x}_{i}^{\prime}}\right)^{-1}\left({\frac{1}{n}\sum_{i=1}^{n}\tilde{x}_{i}W_{i}}\right)$.
The representation (2.1) simply writes $\tau$ as an aggregation of potential outcomes. Not all such aggregations deserve to be called causal effects. Equivalently, to interpret (2.1) as “causal,” we must defend certain restrictions on $\pi_{i}^{*}\rho_{i}(1),\pi_{i}^{*}\rho_{i}(0)$. For instance, were one to assert that (2.1) is the average treatment effect (ATE), then one must defend $\pi_{i}^{*}\rho_{i}(1)=1$ and $(1-\pi_{i}^{*})\rho_{i}(0)=-1$. Since $\rho_{i}(1),\rho_{i}(0)$ are known given the estimable $G_{n}$, these assertions have empirical content. Now, rather than demanding that $\tau$ estimates the ATE, we consider the following strictly weaker property that all causal comparisons should satisfy. Namely, they should be invariant to transformations of potential outcomes that only shift underlying levels but not the individual treatment effects.
Definition 2.1 (Level irrelevance, (Blandhol et al., 2022)).
For $\tau$ some function of $\left({\bm{\pi}^{*},(y_{i}(\cdot),x_{i})_{i=1}^{n}}\right)$, we say that $\tau$ is a causal contrast under $\bm{\pi}^{*}$ if $\tau$ is always unchanged when we replace all potential outcomes $y_{i}(w)$ with $y_{i}(w)+c_{i}$ for arbitrary $c_{i}\in\mathbb{R}$.
For a $\tau$ that is linear in the binary-treatment potential outcomes, 2.1 is equivalent to $\tau$ being a weighted average treatment effect, allowing for negative unit-level weights. This is a natural and weak requirement for causal objects. Allowing for negative weights is perhaps too lenient for some treatment effects, but it accommodates for, e.g., differences in treatment effects between subgroups.
On the other hand, imposing 2.1 is also strong in the sense that it rules out any modeling restrictions on $y_{i}(0)$, as the constants $c_{i}$’s can be arbitrarily and adversarially chosen. [^8] It does not in fact rule out restrictions on individual treatment effects, since adding $c_{i}$’s leave initial treatment effects unchanged. In this sense, 2.4 reflects the fact that identification of $\tau$ is purely on design-based grounds. Had we imposed, say, that $y_{i}(0)$ is a linear function of $x_{i}$, we could reasonably decide that certain aggregations $\tau$ that do not satisfy 2.1 are causal as well. That said, even when we are willing to impose outcome models, we might retain an interest in whether these outcome restrictions can be relaxed under certain designs.
Importantly, observe that $\tau$ satisfies 2.1 under $\bm{\pi}^{*}$ if and only if
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0\text{ for all $i=1,\ldots,n$.}$
By (2.1), the left-hand side of (2.1) represents the effect on $\tau$ of shifting unit $i$’s potential outcomes by $c_{i}=n$. By 2.1, this effect must be zero. Thus, the equation (2.1) must hold for $\bm{\pi}^{*}$ if we are to interpret $\tau$ as a causal contrast. Of course, there is no intrinsic reason why Nature obeys (2.1) when casting the proverbial die $\bm{\pi}^{*}$. However, were we to argue that Nature randomizes treatment in such a manner that our regression estimates a causal contrast regardless of potential outcomes, then we must defend (2.1).
To make the restrictions in (2.1) explicit, we can solve for $\pi_{i}^{*}$, at least when $\rho_{i}(0)\neq\rho_{i}(1)$. Let $\pi_{i}$ denote the solution to (2.1) viewed as equations in $\pi_{i}^{*}$, and we call $\bm{\pi}=(\pi_{1},\ldots,\pi_{n})$ an implicit design:
$\pi_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}.$
The restriction (2.1) can then be equivalently stated as $\pi_{i}^{*}=\pi_{i}$—that the true design is equal to the implicit design. Since the potential weights $(\rho_{i}(1),\rho_{i}(0))$ are known, solving (2.1) exactly pinpoints a putative design $\bm{\pi}$ that must equal $\bm{\pi}^{*}$ if $\tau$ were a causal contrast. This makes explicit the necessary conditions of interpreting $\tau$ causally.
Furthermore, pinpointing $\bm{\pi}$ also pinpoints the corresponding estimand, which additionally adds transparency. To see this, if (2.1) holds, then $\tau$ is equal to a weighted average treatment effect with weights $\omega_{i}\equiv\pi_{i}\rho_{i}(1)=-(1-\pi_{i})\rho_{i}(0)$
$\tau=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}\rho_{i}(1)(y_{i}(1)-y_{i}(0))=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0)).$
These weights are non-negative ($\omega_{i}\geq 0$) when $\rho_{i}(1)\geq 0$ and $\rho_{i}(0)\leq 0$. Therefore, if we are to interpret $\tau$ as a causal contrast under $\bm{\pi}^{*}$, then, simply by computing the solution to linear equations (2.1), we would recover both what the design $\bm{\pi}^{*}$ and what our target $\tau$ must necessarily be—and thus what we must defend economically and statistically. This answers (Q2) and (Q3) posed in the introduction.
For (Q1), there are at least two ways in which $\pi_{i}$ cannot possibly equal $\pi^{*}_{i}$ and are thus indefensible. The more obvious one is if $\pi_{i}\not\in[0,1]$ for any $i$ or if $\rho_{i}(1)=\rho_{i}(0)\neq 0$. This occurs when both $\rho_{i}(1)$ and $\rho_{i}(0)$ are on the same side of zero. When this happens, the implicit design is not a probability distribution, and cannot equal the true design. More subtlely, $\pi_{i}$ is also indefensible if it implies a different Gram matrix:
$G_{n}(\bm{\pi})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}z(x_{i},1)z(x_{i},1)^{\prime}+(1-\pi_{i})z(x_{i},0)z(x_{i},0)^{\prime}\neq G_{n}(\bm{\pi}^{*}).$
This restriction is useful when we analyze specifications theoretically under this framework. It is harder to implement when we do not know and have to estimate $G_{n}$—though one could use it as a basis for inference (Section 2.4).
Returning to the specification (2.1), solving (2.1) shows that the implicit design is precisely the population projection of $\pi_{i}^{*}$ onto the covariates:
$\pi_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}=\frac{\tilde{x}_{i}^{\prime}\delta}{1-\tilde{x}_{i}^{\prime}\delta+\tilde{x}_{i}^{\prime}\delta}=\tilde{x}_{i}^{\prime}\delta,$
where the corresponding estimand is a weighted average treatment effect
$\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))\quad\omega_{i}=\frac{\pi_{i}(1-\pi_{i})}{\frac{1}{n}\sum_{j=1}^{n}\pi_{j}(1-\pi_{j})}.$
Thus, to interpret (2.1) as estimating an economically interesting causal effect, one must defend (a) that $\pi_{i}^{*}=\pi_{i}=\tilde{x}_{i}^{\prime}\delta$ and (b) that one is interested in the $\pi_{i}(1-\pi_{i})$-weighted average treatment effect. On the other hand, a design-based interpretation is obviously untenable if we observe $\pi_{i}\not\in[0,1]$, which happens when the linear probability model for $\pi_{i}^{*}$ produces nonsensical fitted values. [^9] For this specification, one could verify that $G_{n}(\bm{\pi})=G_{n}(\bm{\pi}^{*})$ always holds regardless of whether $\bm{\pi}=\bm{\pi}^{*}$, and thus there is no additional power in the Gram matrix criterion (2.1).
The specification (2.1) is analyzed in Angrist (1998) and Blandhol et al. (2022). Angrist (1998) essentially shows that if $\pi_{i}^{*}=\delta^{\prime}\tilde{x}_{i}$ for some $\delta$, then $\tau$ in (2.1) estimates the weighted treatment effect (2.1). In turn, Blandhol et al. (2022) show that the linearity of the propensity score is also necessary. [^10] Both Angrist (1998) and Blandhol et al. (2022) analyze slightly different setups from ours. Both consider a superpopulation sampling setup in contrast to our finite population setup. Angrist (1998) considers a binary $x_{i}$ in his equation (9), but the argument can be easily generalized, e.g., in Borusyak and Hull (2024); Goldsmith-Pinkham et al. (2024). Blandhol et al. (2022)’s Proposition 7, when applied to (2.1), shows that assuming unconfoundedness and constant treatment effects, $\tau$ is a positively weighted average treatment effect if and only if the propensity score is linear. Our result drops the (finite-population analogues of) unconfoundedness and constant effects assumptions. That said, the underlying mechanics and takeaways are very similar, and so we view our result (formally in Theorem 3.1(1)) as a reinterpretation of theirs. Our framework recovers these results by simply mechanically computing $\rho_{i}(1)$ and $\rho_{i}(0)$. We echo Blandhol et al. (2022) in that if one would like to interpret $\tau$ as some causal contrast (2.1), then one must be willing to believe the linearity $\pi_{i}^{*}=\delta^{\prime}\tilde{x}_{i}$ and that one is targeting (2.1). There is no other configuration that justifies a design-based interpretation of this regression.
Before moving on to the general case with any $(k,T,J)$, we distill the key elements from this leading example and discuss a few subtleties.
  1. i.
    We take the perspective that, by observing a regression specification in the population, the triplet ($(\left\{{z(x_{i},\cdot)}\right\}_{i=1}^{n},\Lambda,G_{n})$ is known. Here, $(z(x_{i},\cdot),\Lambda)$ describes the regression specification, and $G_{n}$ encodes a set of observable implications of the data.
  2. ii.
    We write the population regression estimand $\tau$ in the form (1) and (2.1). Because we treat $G_{n}$ as known, the potential weights $\rho_{i}(w)$ are known for all units.
  3. iii.
    We observe that imposing level irrelevance (2.1) amounts to imposing that $\pi_{i}^{*}$ satisfies linear restrictions, in which the coefficients are the potential weights.
    Here, there is one restriction and one unknown for every unit. In general, there are $kT$ restrictions for $J$ unknowns.
  4. iv.
    We call the solutions to these linear equations implicit designs. Computing them also yields the corresponding estimands by (2.1). If (and only if) $\tau$ is a causal contrast under $\bm{\pi}^{*}$, then $\bm{\pi}^{*}$ must equal one of the implicit designs.
The downstream ramification of treating $G_{n}$ as known in i. could be somewhat subtle. Alternatively, we can think of (2.1) as a set of nonlinear equations in $\pi_{i}^{*}$:
$\pi_{i}^{*}\rho_{i}(1;\bm{\pi}^{*})+(1-\pi_{i}^{*})\rho_{i}(0;\bm{\pi}^{*})=0$
where $\rho_{i}(\cdot;\bm{\pi}^{*})$ depends on $\bm{\pi}^{*}$ through $G_{n}$. These equations then define a nonlinear surface of designs $\left\{{\bm{\pi}:\pi_{i}\rho_{i}(1;\bm{\pi})+(1-\pi_{i})\rho_{i}(0;\bm{\pi})=0\,\forall\,i}\right\}.$ This surface is defined solely as a function of the regression $(\Lambda,z(x_{i},\cdot))$ and thus does not reflect any information in the data.
The implicit designs can then be thought of as choosing a point near [^11] When a implicit design generates the same Gram matrix as the true design $G_{n}(\bm{\pi})=G_{n}(\bm{\pi}^{*})$, then the implicit design is on the surface. this surface that is “most consistent” with the data by fixing $(G_{n},\rho_{i}(\cdot))$ and solving the corresponding equations. For the regression (2.1), this surface is the span of the column space of the covariates $\left\{{\bm{\pi}:\pi_{i}=\tilde{x}_{i}^{\prime}\delta,\delta\in\mathbb{R}^{\dim(x_{i})+1}}\right\}$; the implicit design chooses the point on the surface by projecting $\bm{\pi}^{*}$ to the covariates, since all other points are ruled out by large datasets. When $\bm{\pi}^{*}$ belongs to the surface, doing this exactly recovers $\bm{\pi}^{*}$.
Finally, in the context of (2.1), we can also concretely understand what it means to ignore the estimation noise in $\hat{G}_{n}$. We can numerically write the regression estimator as an inverse propensity-weighting estimator: For $\hat{\pi}_{i}=\frac{-\hat{\rho}_{i}(0)}{\hat{\rho}_{i}(1)-\hat{\rho}_{i}(0)}$ and $\hat{\omega}_{i}=\hat{\pi}_{i}\hat{\rho}_{i}(1)=(1-\hat{\pi}_{i})\hat{\rho}_{i}(0)$ the sample analogues of $(\pi_{i},\omega_{i})$,
$\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\hat{\rho}_{i}(W_{i})y_{i}(W_{i})$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{\omega}_{i}\left({\frac{W_{i}}{\hat{\pi}_{i}}y_{i}(1)+\frac{1-W_{i}}{1-\hat{\pi}_{i}}y_{i}(0)}\right)$
[…]
Representation (2.6) makes apparent that, from a design-based perspective, we can think of a regression estimator as doing two things simultaneously. It first decides on a model for the propensity score and on a target estimand—in (2.1), the model is a linear probability model and the target is the $\pi^{*}(1-\pi^{*})$-weighted average treatment effect. Second, it then estimates the propensity score model and plugs in the estimated coefficients from that model—in (2.1), replacing $\delta$ with $\hat{\delta}$. For design-based interpretation, we are primarily interested in the first. To that end, the Gram matrix $G_{n}$ exactly encodes how the regression models propensity scores and chooses the estimand. Abstracting away from the estimation noise in $\hat{G}_{n}$ exactly ignores the estimation noise in the propensity model. Indeed, in (2.1), knowing $G_{n}$ is equivalent to knowing the population regression coefficient $\delta$.
2.2. Main results
Returning to the general setup, consider a regression of $\mathbf{Y}_{it}$ on a covariate transform $z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})\in\mathbb{R}^{K}$. The population Gram matrix is now
$G_{n}(\bm{\pi}^{*})\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{W}_{i}\sim\pi_{i}^{*}}\left[{z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right],$
whose sample counterpart $\hat{G}_{n}$ omits the expectation over $\mathbf{W}_{i}$. We proceed by formalizing i. through iv. and add additional commentary along the way.
Definition 2.2 (Population regression specification).
Let $G_{n}=G_{n}(\bm{\pi}^{*})$, assumed to be invertible. Let $\mathbf{z}(\mathbf{x}_{i},\cdot)$ be a $T\times K$ matrix whose rows collect $z_{t}(\mathbf{x}_{i},\cdot)^{\prime}$. A population regression specification collects the objects
$(\Lambda,G_{n},\mathbf{z}(\mathbf{x}_{1},\cdot),\ldots,\mathbf{z}(\mathbf{x}_{n},\cdot)).$
The corresponding population regression estimand $\bm{\tau}$ is
$\bm{\tau}\equiv\Lambda G_{n}^{-1}\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{W}_{i}\sim\pi_{i}^{*}}[z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})\mathbf{y}_{it}(\mathbf{W}_{i})].$
This definition formalizes the analogue to item i.: If an empirical researcher shows us a regression specification, as a modeling device, what we take to mean is that the researcher reveals how the regression is implemented ($\mathbf{z}(\mathbf{x}_{i},\cdot)$), which coefficients to interpret causally ($\Lambda$), and how the regression models and estimates the process of treatment assignment ($G_{n}$). By analogy with analyses of identification (Manski, 2022; Koopmans and Reiersol, 1950), we abstract away from the estimation noise in $G_{n}$, returning to it in Section 2.4.
We pause for three remarks before moving on to ii.
Remark 2.1 (Fixed effects).
To justify $\hat{G}_{n}\approx G_{n}$, it is important to include unit-level fixed effects via the within-transformation. That is, unit fixed effects should be parametrized by ensuring that the covariate transform $z_{t}(\mathbf{x}_{i},\mathbf{w})$ sums to zero: $\sum_{t=1}^{T}z_{t}(\mathbf{x}_{i},\mathbf{w})=0$ for all $i$ and $\mathbf{w}\in\mathcal{W}$. It may be tempting to instead parametrize fixed effects as unit-level dummy variables. Doing so makes the dimension of $z_{t}$ grow with the size of the population. However, the Gram matrix $G_{n}$ associated with the many-dummy parametrization is in general no longer consistently estimable. [^12] For unit $i$, let $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{W}_{i})$ denote the covariate transforms that exclude the unit dummy. Then $\sum_{t=1}^{T}\mathbb{E}[\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{W}_{i})]$ is in the Gram matrix (it is the interaction between $\tilde{z}_{t}$ and the unit-$i$ dummy variable). However, this quantity is not consistently estimable as unit $i$ is only observed once. One can view this as a finite-population manifestation of the incidental parameter problem. $\blacksquare$
Remark 2.2 (Fixed Gram designs).
Under certain joint designs $\Pi^{*}$, it is possible that $\hat{G}_{n}=G_{n}(\bm{\pi}^{*})$ with probability one. Under such a design, there is simply no estimation noise to abstract away. As an example, suppose that we consider $Y_{i}=\alpha+\tau W_{i}+\epsilon_{i}$ with binary $W_{i}$, but $W_{i}$ is assigned in a completely randomized experiment. [^13] That is, there are $n_{1}$ treated units, selected uniformly at random from the set of $\binom{n}{n_{1}}$ subsets. Then, the sample and the population Gram matrices are equal and fixed in every realization of the treatment,
$\hat{G}_{n}=G_{n}=\begin{bmatrix}1&n_{1}/n\\ n_{1}/n&n_{1}/n\end{bmatrix}.$
We term these designs fixed Gram designs for a particular specification. Our results in the population can alternatively be read as exact results for fixed Gram designs. [^14] Regression estimators in design-based statistical inference are frequently analyzed under fixed Gram designs, as they are unbiased for the estimand (see, e.g., Rambachan and Roth, 2020; Athey and Imbens, 2022; Zhao and Ding, 2022; Neyman, 1923/1990; Lin, 2013). $\blacksquare$
Remark 2.3 (Fixed covariates).
By assuming that $\mathbf{z}(\mathbf{x}_{i},\cdot)$ is known, we restrict the right-hand side of the regression to only include objects unchanged by treatment. If we include mediators or lagged outcomes, then we no longer know counterfactual values of the covariate transform. We subsequently cannot “complete” the representation of $\bm{\tau}$ in terms of potential outcomes in (1). On the other hand, known functions of the treatment, e.g. lagged treatments, do not introduce any difficulty.
Because of this difficulty, our framework does not apply immediately to instrumental variable (IV) settings for local average treatment effect-type interpretations. Most IV regression estimands can be written as the ratio of two regression estimands. Our analysis can apply to the numerator and the denominator separately to yield an interpretation as ratios of causal contrasts, in terms of the randomly assigned instrument. Whether this ratio is the causal effect of the endogenous treatment of any subpopulation relies on assumptions about how endogenous treatments respond to the instrument. Incorporating these monotonicity-type restrictions in this framework is interesting future work. $\blacksquare$
Returning to ii., exactly analogous to (2.1), we can write
$\bm{\tau}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\cdot\underbrace{\left[{\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right]}_{\bm{\rho}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}}\mathbf{y}_{i}(\mathbf{w})$ (2.7)
which verifies the representation (1). Relative to the simple case (2.1), here we sum over $J+1$ values, and the potential weights are now matrices of dimension $k\times T$.
Definition 2.3 (Potential weights).
Given a population regression specification, the potential weights for individual $i$ are matrices defined by $\bm{\rho}_{i}(\mathbf{w})\equiv\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}.$
Remark 2.4 (Invariances to parametrization).
Section B.1 shows that the potential weights for a given contrast do not depend on how the regression is parametrized. For instance, it does not matter if we write $Y_{i}=\alpha+\tau W_{i}+\epsilon_{i}$ instead as $Y_{i}=\mu_{1}\mathbbm{1}(W_{i}=1)+\mu_{0}\mathbbm{1}(W_{i}=0)+\epsilon_{i}$ and consider $\tau=\mu_{1}-\mu_{0}$. Section B.1 also shows that the potential weights are suitably invariant under the Frisch–Waugh–Lovell transform. $\blacksquare$
Next, we turn to level irrelevance (iii.). A natural generalization is the following, where we now add arbitrary numbers $c_{it}$ to each of the $J+1$ potential outcomes.
Definition 2.4 (Causal contrast).
We say that $\bm{\tau}$ is a vector of causal contrasts under $\bm{\pi}^{*}$ if it satisfies level irrelevance: That is, $\bm{\tau}$ is unchanged if we replaced all potential outcomes $\mathbf{y}_{it}(\mathbf{w})$ with $\mathbf{y}_{it}(\mathbf{w})+c_{it}$, for any $i\in[n],t\in[T]$, and $c_{it}\in\mathbb{R}$.
Again, we think of level irrelevance as a minimal requirement for causal quantities whose identification is purely design-based and outcome model-agnostic (Blandhol et al., 2022). Imposing 2.4 amounts to encoding that, in each entry of $\bm{\tau}$, each individual’s time-$t$ potential outcomes are aggregated by some weights that sum to zero:
$\bm{\tau}=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}{\bm{\omega}_{i}(\mathbf{w})}\mathbf{y}_{i}(\mathbf{w})\text{ for some $\bm{\omega}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}$ where }0=\sum_{\mathbf{w}\in\mathcal{W}}\bm{\omega}_{i}(\mathbf{w}).$
These aggregations are thus considerably more complex than weighted average treatment effects for 2.1.
Moving on, the analogue to iv. is the following set of equations
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=0,\quad\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})=1.$
Absorbing the sum-to-one condition, there are $kT$ restrictions for $J$ unknowns. We formally define them as implicit designs.
Definition 2.5 (Implicit design).
Given the potential weights, we call $\bm{\pi}=(\pi_{1},\ldots,\pi_{n})$ that satisfies (2.2) for every $i$ an (improper) implicit design. If each $\pi_{i}(\mathbf{w})\geq 0$, then we call $\bm{\pi}$ a proper implicit design. If $\bm{\pi}$ generates the same Gram matrix $G_{n}$,
$G_{n}=G_{n}(\bm{\pi})\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})z_{t}(\mathbf{x}_{i},\mathbf{w}_{i})z_{t}(\mathbf{x}_{i},\mathbf{w}_{i})^{\prime},$
then we say that $\bm{\pi}$ is Gram-consistent.
We can often expect that implicit designs are unique, if they exist. This is because the number of equations is often at least the number of unknowns ($kT\geq J$). For instance, with $J+1$ treatments, we often have $k=J$ contrasts; panels under staggered adoption admit fewer unique treatment times $(J+1)$ than time horizon $T$. Lemma B.8 formally establishes the uniqueness of implicit designs in cross-sectional settings ($T=1$).
Relative to an implicit design $\bm{\pi}$, the regression specification estimates the $\bm{\pi}$-weighted average of the potential outcomes, which we formally define as the implicit estimand, indexed by weights $\bm{\omega}_{i}(\bm{\pi},\mathbf{w})=\pi_{i}\rho_{i}(\mathbf{w})$.
Definition 2.6 (Implicit estimand).
Let $\bm{\pi}$ denote an implicit design of a given regression specification. The corresponding implicit estimand is
$\bm{\tau}(\bm{\pi})=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\omega}_{i}(\bm{\pi},\mathbf{w})\mathbf{y}_{i}(\mathbf{w}),$
where $\bm{\omega}_{i}(\bm{\pi},\mathbf{w})=\pi_{i}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}$. Without introducing ambiguity, we also use “implicit estimand” to refer to the weights $\left\{{\bm{\omega}_{i}(\bm{\pi},\mathbf{w}):i\in[n],\mathbf{w}\in\mathcal{W}}\right\}$.
Having formalized i. through iv., finally, we collect our observations in the following theorem, which is the main result of this paper.
Theorem 2.1.
$\bm{\tau}$ is a vector of causal contrasts under $\bm{\pi}^{*}$ in the sense of 2.4 if and only if $\bm{\pi}^{*}$ is equal to some implicit design. When this happens, if, furthermore, there is a unique implicit design $\bm{\pi}$, then $\bm{\pi}$ must be proper and Gram-consistent.
Theorem 2.1 implies our result in Section 2.1 as a corollary.
Corollary 2.2 ($k=T=J=1$).
When $k=T=J=1$, $\tau$ is a causal contrast under the true design $\bm{\pi}^{*}$ if and only if $\rho_{i}(1)\rho_{i}(0)\leq 0$ for all $i$ and $\pi_{i}^{*}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}$ for all units $i$ with one of $\rho_{i}(1)$ and $\rho_{i}(0)$ nonzero. When this happens, the implicit estimand is uniquely defined by $\omega_{i}(\bm{\pi}^{*},1)=\pi_{i}^{*}\rho_{i}(1)=-\omega_{i}(\bm{\pi}^{*},0)$, and thus
$\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}^{*}(y_{i}(1)-y_{i}(0))\text{ for }\omega_{i}^{*}\equiv\omega_{i}(\bm{\pi}^{*},1).$
The unit-level weights $\omega_{i}^{*}$ are negative if and only if $\rho_{i}(1)<0<\rho_{i}(0)$.
Since the set of implicit designs is known in the population and estimable in sample, Theorem 2.1 shows that they are informative of $\bm{\pi}^{*}$ under the presumption that the regression admits a causal interpretation. Thus, the implicit designs reveal the necessary implications on $\bm{\pi}^{*}$ of interpreting $\bm{\tau}$ as a design-based estimand. If a researcher so interprets, then they must be prepared to defend why the true assignment process $\bm{\pi}^{*}$ is equal to one of the implicit designs $\bm{\pi}$.
In the simple case ($k=T=J=1$), 2.2 formalizes our discussion in Section 2.1. If it exists, the implicit design is uniquely and explicitly defined, up to units with zero potential weights and do not factor into the estimand. The implicit estimand is a weighted average treatment effect. All unit-level weights $\omega_{i}$ in the implicit estimand are nonnegative if no unit has the “wrongly ordered” potential weights $\rho_{i}(1)<0<\rho_{i}(0)$.
2.3. Discussion
Theorem 2.1 constructively yields plausible designs and estimands that answer questions (Q1)(Q3) posed in the introduction. In our view, computing the implicit design and estimand adds transparency and connects estimation to identification. When researchers point to design-based identification arguments but do not explicitly model the design or describe the estimand, our results make clear what choices a regression maintains on the model for treatment and on the target estimand. When researchers treat regression as an outcome model, a reasonable implicit design and implicit estimand make the regression interpretable to those who challenge the outcome model but are willing to entertain restrictions on treatment assignment.
Since we aim to understand regressions from the necessary implications of their causal interpretation, our arguments do resemble trying to “reverse engineer” an interpretation for the regression estimator—in fact our arguments show that reverse engineering is ubiquitous and mechanical. Like Mogstad and Torgovitsky (2024), who critique careless reverse engineering in instrumental variable settings, we caution that the mere fact that one could obtain an implicit design does not by itself justify the statistical and economic relevance of the design or the estimand. Whether this is the case always involves judgment from the researcher and the audience. To that end, we provide a checklist of diagnostics that assess and improve the implicit design and implicit estimand. Section 4 illustrates these in action.
To implement these diagnostics, practitioners can explicitly compute potential weights and implicit designs—most straightforwardly by replacing $G_{n}$ with $\hat{G}_{n}$. If they are additionally willing to assume that the true design $\Pi^{*}$ is fixed Gram (2.2), then Theorem 2.1 is applicable as-is. In general, $\hat{G}_{n}$ involves noise from realized treatment assignments, and we defer a discussion to Section 2.4. To give some intuition that may assuage concerns, given a confidence set for $G_{n}$, it is in principle possible to project this confidence set to the set of potential weights and the set of proper and Gram-consistent implicit designs. Moreover, if one has a candidate design $\bm{\pi}$ in mind such that $\bm{\pi}$ and $G_{n}(\bm{\pi})$ satisfy (1), it is also straightforward to check whether the implied $G_{n}(\bm{\pi})$ is statistically close enough to $\hat{G}_{n}$.
A checklist for the implicit design. Abstracting away from sampling noise, suppose (1) reveals a proper implicit design $\bm{\pi}$, we may consider the following checks:
  • Is the implicit design well-calibrated? That is, among those with $\pi_{i}\approx\pi$, do approximately $\pi$-proportion of units receive treatment? [^15] When we need $\pi^{*}_{i}$ to be linear, Blandhol et al. (2022) suggest using the Ramsey RESET test against the hypothesis that $\mathbb{E}[W_{i}\mid x_{i}]$ is linear. For fitted values $\hat{\pi}_{i}$, the RESET test effectively tests $\beta_{2}=\cdots=\beta_{p}=0$ in the regression $W_{i}=\alpha_{0}+\alpha_{1}^{\prime}x_{i}+\beta_{1}(\hat{\pi}_{i})^{2}+\cdots+\beta_{p}(\hat{\pi}_{i})^{p}+\epsilon_{i}.$ If we replace $\alpha_{0}+\alpha_{1}^{\prime}x_{i}$ with $\hat{\pi}_{i}$, the RESET test is thus similar to a test of the linearity—and thus calibration—of $\mathbb{E}[W_{i}\mid\hat{\pi}_{i}]$.
  • The variance of $\pi_{i}$ describes how predictable $W_{i}$ is from observable information. Is the variance of $\pi_{i}$ sufficiently small for the treatment to be credibly quasi-random?
  • Qualitatively, relative to a plausible model of treatment selection, do $\pi_{i}$ covary positively with features that predict selection into treatment?
  • Are covariates (and transformations thereof) balanced by $\pi_{i}$? In the binary-treatment case, if $\pi_{i}=\pi_{i}^{*}$, then we should expect
    $\frac{1}{n}\sum_{i=1}^{n}\left({\frac{\mathbf{W}_{i}}{\pi_{i}}-\frac{1-\mathbf{W}_{i}}{1-\pi_{i}}}\right)f(\mathbf{x}_{i})\approx 0$
    for any function $f(\mathbf{x}_{i})$.
If these checks all pass, then it continues to be plausible that $\bm{\pi}=\bm{\pi}^{*}$. While one can never be sure that $\bm{\pi}^{*}$ is truly equal to $\bm{\pi}$, at least there lacks evidence to the contrary.
A checklist for the implicit estimand. We may also compute the implicit estimand via 2.6. Usually, the estimand would involve weights that do not correspond to treatment effect contrasts from first principles. Similarly, we may check the following:
  • Are any of the weights unintentionally negative (i.e., $\tau$ is not intended as a difference of treatment effects)? Is it possible to justify non-uniform weights as describing economically or statistically interesting objects? [^16] Section 4.2 in Goldsmith-Pinkham et al. (2024) provides such an argument on variance weighting.
  • How large is the implicit subpopulation whose treatment effect the estimand is measuring (Poirier and Słoczyński, 2024)?
  • What types of units are getting more weight?
  • For particular realizations of the outcomes, does the weighting matter? One could use the implicit design to construct alternative estimates via inverse propensity weighting
    $\hat{\bm{\tau}}(\bm{\pi};\bm{\omega})=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\frac{\mathbbm{1}(\mathbf{W}_{i}=\mathbf{w})}{\pi_{i}(\mathbf{w})}\bm{\omega}_{i}(\mathbf{w})\mathbf{y}_{i}(\mathbf{w}),$
    for user-chosen weights $\bm{\omega}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}$. [^17] In the simple case ($k=T=J=1$), a popular choice may be the trimmed ATE (Crump et al., 2009): $\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left({\frac{W_{i}}{\pi_{i}}-\frac{1-W_{i}}{1-\pi_{i}}}\right)Y_{i}\quad\omega_{i}\propto\mathbbm{1}(\epsilon<\pi_{i}<1-\epsilon)\quad\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=1.$ This estimator targets the user-chosen estimand $\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\omega}_{i}(\mathbf{w})\mathbf{y}_{i}(\mathbf{w})$; it solely alters the estimand but not the propensity model for the regression, and is thus very comparable to the regression estimate.
  • Are the outcomes studied correlated with the implicit design? One can, for instance, plot estimates of $\mathbb{E}[Y_{i}\mid W_{i}=w,\pi_{i}=\pi]$. If these functions are flat in terms of $\pi$, then it is likely that the regression-chosen weights make little difference and alternative user-chosen weights would result in similar estimates.
If these checks all pass—we find a plausible implicit design under which the regression estimates a reasonable causal contrast, then we are reassured that there could be a plausible treatment assignment process under which our regression aggregates a sensible set of comparisons. This dispels concerns about “negative weights” when we are willing to defend the implicit design as reflecting reality. Since it is usually straightforward to interpret regressions as outcome models, having a valid and outcome-agnostic design-based interpretation adds an additional robustness.
Alternatives and design patching. If no proper and Gram-consistent implicit design is found or if any of the above checks fails, then this exercise calls design-based interpretation into question. Defending the regression estimand on design-based grounds is no longer logically tenable. Practitioners essentially face two options. We may maintain that causal identification comes from design-based reasoning, but we need to better model the design and target a better estimand—adopting what Mogstad and Torgovitsky (2024) call “forward engineering.” There is a large and growing literature in causal inference that provides such methods; rarely are such methods implementable as a simple linear regression (see monographs and reviews by Wager, 2024; Imbens and Rubin, 2015; Ben-Michael et al., 2021; Chernozhukov et al., 2024). Alternatively, we may retreat by defending functional form restrictions on the potential outcomes and treatment effects, sacrificing some agnosticism. [^18] Even with outcome models, identification typically still requires assumptions on the process of treatment assignment. These restrictions are typically of the form that treatment assignment does not predict values of some “structural residual” defined by an outcome model.
A frequent case where design-based interpretation is rejected (for $k=J=T=1$) is when the implicit design has a small proportion of units with $\pi_{i}\not\in[0,1]$. This occurs, for instance, in our empirical application to Blakeslee et al. (2020) as well as in Blandhol et al. (2022)’s illustration of Dube and Harish (2020). It is possible that the implicit propensity model suffers from some misspecification, but is otherwise close to something defensible. In that case, practitioners may find the initial regression a flawed but reasonable starting point. Carefully modeling treatment assignment is a principled alternative, but may not be worth the hassle. A simple patch—one that at least does no harm—could be desired.
If we cast the problem as simply trying to improve a given predictor $\pi_{i}$, one could consider recalibrating the implicit design. Specifically, one could consider discretizing $\pi_{i}$ into bins, and setting $\hat{\pi}_{i}^{\text{patched}}$ as the empirical frequency of $W_{i}=1$ within the bin that $i$ is in. For user-chosen unit-level weights $\omega_{i}$ such that $\omega_{i}=0$ whenever $\hat{\pi}_{i}^{\text{patched}}\in\left\{{0,1}\right\}$, we can then compute the design-patched estimate
$\hat{\tau}^{\text{patched}}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left({\frac{W_{i}}{\hat{\pi}_{i}^{\text{patched}}}-\frac{1-W_{i}}{1-\hat{\pi}_{i}^{\text{patched}}}}\right)Y_{i}.$
Here, it is sensible to make observations with $\hat{\pi}_{i}<0$ and $\hat{\pi}_{i}>1$ their own bins.
Design patching can be viewed as applying histogram binning, a popular method in the machine learning literature, to calibrate the predictor $\pi_{i}$ (Zadrozny and Elkan, 2001). Loosely speaking, one could view recalibration as marginally improving the quality of predictions. [^19] When $\pi_{i}$ is linear, this approach can be heuristically understood as perturbing the predicted values in the direction that makes rejection of Ramsey’s RESET test less severe. This method can also be viewed as treating $\pi_{i}$ as a propensity score estimate and subclassifying on the propensity score (see Chapter 17 in Imbens and Rubin, 2015). Imbens and Rubin (2015) also provide a data-driven method for choosing the number of bins. The patched estimate $\hat{\tau}^{\text{patched}}$ has an explicit design-based interpretation when the true design is approximately constant within the bins. It behaves similarly to the regression estimator when the regression specification is close to correct.
2.4. Estimating potential weights
We conclude our main results by tying up the loose ends in the sampling uncertainty for $G_{n}$. Naturally, we estimate potential weights and implicit designs by plugging in the sample Gram matrix $\hat{G}_{n}$. When we fail to find valid estimated implicit designs, we may wonder whether that is due to sampling noise in $\hat{G}_{n}$ rather than the fault of the regression specification. When we do find a valid estimated implicit design, we may also ask whether $\hat{\bm{\pi}}$ is close to $\bm{\pi}$. To answer these questions, we provide a set of simple estimation and inference guarantees. [^20] Since we view the estimated implicit designs and potential weights as diagnostics and sanity checks, we do not provide particularly strong or optimal guarantees and leave those to future work. In our view, if a practitioner fails to find valid estimated implicit designs, even if that is due to noise, it should nevertheless raise alarms for the interpretation of the regression estimator.
In terms of inference, conceptually, for $\hat{\mathcal{G}}_{n}$ a $(1-\alpha)$-confidence set for $G_{n}$, we may form corresponding confidence sets ^$\hat{\mathcal{R}}_{i}$ for $\bm{\rho}_{i}$ and $\hat{\mathcal{P}}$ for $\bm{\pi}^{*}$ by projection:
^$\hat{\mathcal{R}}_{i}$ =$=\left\{{\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}:G_{n}\in\hat{\mathcal{G}}_{n}}\right\}$
$\hat{\mathcal{P}}=\bigg{\{}\bm{\pi}:\,\,$ $G\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\in\hat{\mathcal{G}}_{n}$ (Gram-consistent)
$\forall i,\,\pi_{i}\in\Delta(\mathcal{W})$ (Proper)
$\forall i,\,\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\Lambda G^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}=0\bigg{\}}.$ (Level irrelevance)
It is easy to see that, when $\bm{\pi}^{*}$ satisfies (1), $\operatorname{P}[\bm{\pi}^{*}\in\hat{\mathcal{P}}]=\operatorname{P}[G_{n}(\bm{\pi}^{*})\in\hat{\mathcal{G}}_{n}]\geq 1-\alpha$. $\hat{\mathcal{G}}_{n}$ is readily available from, e.g., a simultaneous confidence set for all entries of $G_{n}$. These confidence sets are usually available upon verification of high-probability bounds for $\lVert\hat{G}_{n}-G_{n}\rVert$ or central limit theorems for $\sqrt{n}(\hat{G}_{n,jk}-G_{n,jk})$. Unfortunately, fully computing $\hat{\mathcal{P}}$ is a difficult nonlinear optimization problem. Nonetheless, checking whether a candidate $\bm{\pi}\in\hat{\mathcal{P}}$ is straightforward. Thus, if the researcher has a particular design in mind—perhaps in response to the red flags raised by estimated implicit designs—they can check whether such a design is a plausible justification for the specification in question.
[…]
  1. (1)
    The estimated potential weights $\hat{\bm{\rho}}_{i}(\mathbf{w})$ are consistent for $\bm{\rho}_{i}(\mathbf{w})$.
  2. (2)
    In many cases, the consistency of potential weights implies consistency of estimated implied designs and estimands, though the precise statement depends on further regularity conditions. A key obstruction to a general result is that small perturbations to coefficients in linear equations may result in large perturbations to their solutions.
To discuss asymptotics, we embed our finite population in a sequence of populations. Formally, let $\mathbf{y}_{1}(\cdot),\mathbf{x}_{1},\mathbf{y}_{2}(\cdot),\mathbf{x}_{2},\ldots$ denote a sequence of potential outcomes and covariates. Let $\Pi_{n}^{*}$ denote the joint distribution of treatment assignments for the first $n$ units. We shall consider the behavior as $n\to\infty$, relative to these sequences. We impose a high-level condition on the sequence of populations such that the estimated Gram matrix is consistent for the population Gram matrix. [^21] Assumption 2.1 holds whenever the underlying populations are such that a law of large numbers hold. If the treatments are independently assigned, for instance, then Assumption 2.1 is true under standard laws of large number (e.g., Theorem 2.2.6 in Durrett (2019)), which we state in Lemma B.4. Section B.2.1 verifies Assumption 2.1 for designs that involve sampling without replacement (Hájek, 1964; Rambachan and Roth, 2020).
Assumption 2.1.
[…]
Assumption 2.2.
The sequence of population Gram matrices $G_{n}$ is such that their minimum eigenvalues are bounded below: For some $\epsilon>0$, for all $n$, $\lambda_{\min}(G_{n})\geq\epsilon>0$.
Assumption 2.2 makes sure that the population regression specification is strongly identified, in the sense that the Gram matrix is bounded away from singularity. Under Assumption 2.1 and Assumption 2.2, we can show that the estimated potential weights are consistent for their population counterparts.
Proposition 2.3.
[…]
where $|\cdot|_{\infty}$ takes the entrywise maximum absolute value.
Naturally, we define the set of estimated implicit designs as the set of solutions to the sample counterpart of (2.2): $\hat{\bm{\pi}}$ is an estimated implicit design if for every $i$,
$\sum_{\mathbf{w}\in\mathcal{W}}\hat{\pi}_{i}(\mathbf{w})\hat{\bm{\rho}}_{i}(\mathbf{w})=0,\quad\sum_{\mathbf{w}\in\mathcal{W}}\hat{\pi}_{i}(\mathbf{w})=1.$
Unfortunately, in general, the estimated implicit designs are not always well-behaved. This is because they are the solutions to a system of overdetermined linear equations with noisy coefficients, and small perturbations to the coefficients can result in large changes in the solutions—or even to their existence in the first place. [^23] For instance, it is possible for noise in the coefficients of a underdetermined system to turn the system into an overdetermined one. Thus, it is possible that a population implicit design exists—meaning that (2.2) is under or exactly determined—but no estimated implicit designs do. Failing a general result, we present weaker consistency results that apply in at least in the cross-section ($T=1$). [^24] Proposition B.5 shows that if an estimated implicit design $\hat{\bm{\pi}}$ exists with bounded entries, then it is consistent in the sense that it approximately solves the population level irrelevance equations: $\sum_{\mathbf{w}}\hat{\pi}_{i}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})$ tends to zero.
In particular, if the population implicit design is unique and equal to $\bm{\pi}^{*}$, the estimated implicit designs are consistent for $\bm{\pi}^{*}$. We illustrate with the following result in the cross-sectional case ($T=1$) where the system of equations (2.2) is exactly determined ($k=J$). To state this result, let $R_{i}=[\bm{\rho}_{i}(0),\ldots,\bm{\rho}_{i}(J)]\in\mathbb{R}^{k\times(J+1)}$ be the matrix whose columns are potential weights at each treatment level. Let $R_{i}=\sum_{r=1}^{J}\sigma_{ir}u_{ir}v_{ir}^{\prime}$ be its singular value decomposition, for $\sigma_{1i}\geq\ldots\geq\sigma_{iJ}\geq 0$, $u_{i1},\ldots,u_{iJ}$ orthonormal vectors in $\mathbb{R}^{J}$ and $v_{i1},\ldots,v_{iJ}$ orthonormal vectors in $\mathbb{R}^{J+1}$.
Proposition 2.4.
Assume that $k=J$. Suppose that:
  1. (i)
    The estimand $\bm{\tau}$ is a vector of causal contrasts in the sense of 2.4 under $\bm{\pi}^{*}$.
  2. (ii)
    Given $\rho_{i}(\cdot)$, the smallest singular value of $R_{i}$ is uniformly bounded away from zero: For some $\epsilon>0$, $\liminf_{n\to\infty}\min_{i\in[n]}\sigma_{iJ}>\epsilon>0.$
  3. (iii)
    The estimated potential weights are consistent in the sense of (2.3).
Then, for some $C>0$, a unique estimated implicit design $\hat{\bm{\pi}}$ that is bounded by $C$ (that is, $\max_{i}\max_{\mathbf{w}\in\mathcal{W}}|\hat{\pi}_{i}(\mathbf{w})|\leq C$) exists with probability tending to one. Let $\hat{\bm{\pi}}$ be an estimated implicit design if it exists, and otherwise let $\hat{\bm{\pi}}$ be an arbitrary probability vector. Then $\hat{\bm{\pi}}$ is consistent for $\bm{\pi}^{*}$:
[…]
The key condition in Proposition 2.4 is (ii), which ensures that the solution to the population level-irrelevance restrictions (2.2) is unique and robust to small perturbations of the potential weights. In the binary treatment, cross-sectional setting, (ii) is equivalent to that the potential weights are bounded away from zero: $\rho_{i}(1)^{2}+\rho_{i}(0)^{2}>\epsilon^{2}$. Lemma B.8 further provides sufficient conditions for (ii) to hold in the $T=1$ case.
3. Theoretical applications and examples
Having established the machinery of potential weights and implicit designs, we now reap its fruits. Our results are useful both for theoretical investigation and for practical diagnostics with a given dataset. This section illustrates the former by computing the potential weights and implicit designs for a number of population regression specifications; Section 4 illustrates how our results apply to an empirical setting. To emphasize, for both purposes, our results essentially reduce the problem to computing the potential weights and the set of implicit designs. This makes them broadly applicable to generate a litany of results in starkly distinct settings, with relatively little specialized analysis. Reasoning through $\bm{\pi}^{*}=\bm{\pi}$ theoretically is sometimes bespoke, but can often be aided by numerical methods.
Throughout this section, we assume that the population Gram matrix is invertible. Some of these specifications are well-studied in the literature and have known causal interpretations under certain design assumptions (Angrist, 1998; Goldsmith-Pinkham et al., 2024; Kline, 2011; Imbens and Wooldridge, 2009; Lin, 2013; Athey and Imbens, 2022). We show that the corresponding implicit designs essentially rediscover these properties and—moreover—supply a converse of sorts. Specifically, we show that the design-based interpretations analyzed in these settings are tenable only under those designs assumed in the literature; they target exactly those estimands found in the literature. There are no other designs or estimands that are consistent with a design-based interpretation of these specifications.
When applied to a number of examples that are less known, our results uncover some new insights. We show that two types of specifications have design-based interpretations only under stringent conditions. The first class contains cross-sectional regressions that involve the interaction between $W$ and $x$. For these regressions, if all interactions between $W$ and $x$ are causal contrasts, then essentially, $x$ must be saturated discrete covariates or $W$ must be randomly assigned independently of $x$. The second class contains two-way fixed effect (TWFE) regressions that either have time-varying covariates or imbalanced panels. For these regressions, if the treatment pattern covaries with the covariates or the imbalance pattern, then implicit designs do not exist.
3.1. A unified analysis of design-based interpretation in regression
Theorem 3.1.
We compute the implicit designs and estimands of the following regression specifications. In all of the following scenarios, the implicit design exists and is unique. The implicit design is Gram-consistent regardless of whether $\bm{\pi}=\bm{\pi}^{*}$ for all scenarios except for (4).
  1. (1)
    (Angrist, 1998; Blandhol et al., 2022, Section 2.1, $k=T=J=1$) Suppose $x_{i}$ includes a constant. For $Y_{i}=\tau W_{i}+x_{i}^{\prime}\gamma+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      $\pi_{i}=x_{i}^{\prime}\delta$ for $\delta=\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}=x_{i}^{\prime}\delta$
    3. (c)
      $\omega_{i}\equiv\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=\frac{\pi_{i}(1-\pi_{i})}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(1-\pi_{i})}$. When $\pi_{i}^{*}=\pi_{i}$, $\omega_{i}\geq 0$.
  2. (2)
    (Goldsmith-Pinkham et al., 2024, $k=J>1,T=1$) Suppose $\mathcal{W}=\left\{{0,1,\ldots,J}\right\}$ for $J>1$ and $x_{i}$ includes a constant. Let $W_{ij}=\mathbbm{1}(W_{i}=j)$. For $Y_{i}=\sum_{j=1}^{J}\tau_{j}W_{ij}+x_{i}^{\prime}\gamma+\epsilon_{i}$ and $\bm{\tau}=[\tau_{1},\ldots,\tau_{J}]^{\prime}$ the coefficient of interest:
    1. (a)
      $\pi_{i}(j)=x_{i}^{\prime}\delta_{j}$ for $\delta_{j}=\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(j)x_{i}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}(j)=x_{i}^{\prime}\delta_{j}$ for all $j\in[J]$
    3. (c)
      The implicit estimand is shown in (2c). This estimand is generally contaminated (that is, $\omega_{ij}(\bm{\pi},\ell)\neq 0$ for some $j\in[J]$ and $\ell\not\in\left\{{0,j}\right\}$).
  3. (3)
    (Miratrix et al., 2013; Imbens and Wooldridge, 2009; Lin, 2013, $k=T=J=1$) Suppose $x_{i}$ represents saturated discrete covariates taking values in $\left\{{0,\ldots,L}\right\}$, in the sense that $x_{i}=[x_{i1},\ldots,x_{iL}]^{\prime}$ for $x_{i\ell}=\mathbbm{1}(x_{i}=\ell)$. Let $\overline{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$. For $Y_{i}=\alpha_{0}+\gamma_{1}^{\prime}x_{i}+\tau W_{i}+W_{i}(x_{i}-\overline{x})^{\prime}\gamma_{2}+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      For all units $i$ whose covariate value is $\ell$ (i.e., $x_{i\ell}=1$), the implicit design $\pi_{i}=\pi_{\ell}$, where $\pi_{\ell}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(x_{i}=\ell)$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}$ is the same for all units with the same $x_{i}$-value
    3. (c)
      The implicit estimand is the ATE. That is, $\omega_{i}=\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=1$.
  4. (4)
    (Kline, 2011, $k=T=J=1$) Let $\overline{x}_{1}=\frac{\sum_{i=1}^{n}\pi_{i}^{*}x_{i}}{\sum_{i=1}^{n}\pi_{i}^{*}}$. For $Y_{i}=\alpha_{0}+\gamma_{1}^{\prime}x_{i}+\tau W_{i}+W_{i}(x_{i}-\overline{x}_{1})^{\prime}\gamma_{2}+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      $\pi_{i}=\frac{\delta_{0}+(x_{i}-\overline{x})^{\prime}\delta_{1}}{1+\delta_{0}+(x_{i}-\overline{x})^{\prime}\delta_{1}}$, where $\delta_{0},\delta_{1}$ are equal to the population projection coefficients of $\pi_{i}^{*}/(1-\pi_{i}^{*})$ on $x_{i}-\overline{x}$ and a constant, weighted by $1-\pi_{i}^{*}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x})$
    3. (c)
      When $\bm{\pi}^{*}=\bm{\pi}$, the implicit estimand is the ATT. That is, $\omega_{i}=\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=\frac{\pi_{i}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}}$.
  5. (5)
    (Athey and Imbens, 2022, $T>1$) Let $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$ be the set of treatment paths such that its nonzero elements are linearly independent vectors whose span excludes the constant vector $1_{T}=[1,\ldots,1]^{\prime}$. [^25] Staggered adoption corresponds to the case $\mathbf{w}_{t}\leq\mathbf{w}_{t+1}$ for all $t$. If $\mathcal{W}$ satisfies staggered adoption and excludes always-treated units, then it contains linearly independent vectors whose span excludes $1_{T}$. For the two-way fixed effects (TWFE) specificaiton $\mathbf{Y}_{it}=\alpha_{i}+\mu_{t}+\tau\mathbf{W}_{it}+\epsilon_{it}$ and the coefficient of interest $\tau$:
    1. (a)
      The implicit design is constant in $i$ and is unique, $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}$ is the same for all $i$
    3. (c)
      The implicit estimand is shown in (5c), which matches Theorem 1(ii) in Athey and Imbens (2022) under staggered adoption.
Theorem 3.1 computes implicit designs and estimands for specifications considered in the literature. The literature analyzes each specification under some restrictions on the design and provides the corresponding estimand. In each case, the implicit design exists, is unique, and takes exactly the form studied; the implicit estimand is exactly the corresponding estimand. In this sense, Theorem 3.1 is a set of converses to these results—the regression estimand is a causal contrast under $\bm{\pi}^{*}$ only if $\bm{\pi}^{*}=\bm{\pi}$ and the desired causal contrast is the implicit estimand. Theorem 3.1 computes these quantities in the population; the corresponding estimated implicit designs are all straightforward sample analogues of the population quantities.
We now discuss Theorem 3.1(2) through (5) in detail. Theorem 3.1(2) analyzes the analogue of (2.1) with multivalued treatments, which are saturated in the specification. If this regression were a model for outcomes, the coefficient $\tau_{j}$ is the contrast between treatment $w=j$ and the control treatment $w=0$. Without an outcome model, the only design that makes $\tau_{1},\ldots,\tau_{J}$ all causal contrasts is the design that is linear in $x_{i}$. Since there are multiple treatments, this means that each entry of the generalized propensity score $\pi_{i}(j)$ is linear in the covariates. However, under this design, $\tau_{j}$ involves potential outcomes $y_{i}(\ell)$ for some $\ell\neq j$. This means that $\tau_{j}$ does not capture only treatment effects of treatment $j$ versus the control condition. Goldsmith-Pinkham et al. (2024) call this feature “contamination bias” and show it exists under linear generalized propensity scores (their Assumption 2). We recover this result and additionally show that no other treatment assignment model would de-contaminate the estimand while retaining level irrelevance. [^26] A leading case in which the estimand is not contaminated is unconditional random assignment (i.e., constant $\bm{\pi}^{*}$ in $i$). Interestingly, there does exist setups with non-constant $\bm{\pi}^{*}$ such that the estimand is uncontaminated, which are corner cases that Goldsmith-Pinkham et al. (2024) (p.4028) informally rule out. See Example A.2.
The specification Theorem 3.1(3) enriches (2.1) by saturating in both $x_{i}$ and $W_{i}$, and de-means the interaction term appropriately so that $\tau$ targets the average treatment effect when treated as an outcome model. [^27] The citations to Lin (2013) and Imbens and Wooldridge (2009) are somewhat improper. Lin (2013) studies this regression in a design-based context, but assumes complete random assignment of $W_{i}$ and does not restrict to saturated covariates. Imbens and Wooldridge (2009) (section 5.3) study this regression while assuming unconfoundedness and that conditional means of potential outcomes are linear in the covariates, also not restricting to saturated covariates; under saturation, the outcome-modeling assumptions are without loss. Since the specification is saturated, as long as treatment assignment satisfies unconfoundedness, we expect $\tau$ to estimate a causal effect. Indeed, Theorem 3.1(3) shows that unconfoundedness is necessary and the causal effect is the average treatment effect.
Theorem 3.1(4) studies a similar regression to Theorem 3.1(3). Compared to Theorem 3.1(3), Theorem 3.1(4) does not saturate the covariates and de-means the interaction term at $\overline{x}_{1}$ instead of $\overline{x}$. If the regression were an outcome model, the contrast in Theorem 3.1(4) would be the average treatment effect on the treated. [^28] There are two subtleties in our presentation of this regression. First, the average treatment effect on the treated for this finite population is defined to be $\frac{1}{n}\sum_{i=1}^{n}\frac{\pi_{i}^{*}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}}(y_{i}(1)-y_{i}(0)).$ One could view this as the finite-population analogue of the superpopulation ATT under unconfoundedness: $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid W_{i}=1]=\mathbb{E}\left[{\frac{\pi(X_{i})}{\mathbb{E}[\pi(X_{i})]}(Y_{i}(1)-Y_{i}(0))}\right].$ Second, this regression specification has a covariate transform that depends on $\overline{x}_{1}$, which involves $\pi_{i}^{*}$. Thus assuming that the covariate transform is known in population also amounts to ignoring the uncertainty in estimating $\overline{x}_{1}$ when we estimate it using $\sum_{i}W_{i}x_{i}/\sum_{i}W_{i}$. Kline (2011) shows that $\tau$ is also the ATT model-agnostically: If the odds for the propensity score ($\pi_{i}^{*}/(1-\pi_{i}^{*})$) is linear in the covariates, then this regression estimates the ATT. Theorem 3.1(4) provides the converse to this result.
Theorem 3.1(4) also shows how this regression implicitly estimates the linear model for the odds, especially since there is no obvious sample analogue to $\pi_{i}^{*}/(1-\pi_{i}^{*})$. The coefficients $\delta_{0},\delta_{1}$ in $\pi_{i}/(1-\pi_{i})$ are population regression coefficients of the odds $\pi_{i}^{*}/(1-\pi_{i}^{*})$ on the covariates, weighted by $1-\pi_{i}^{*}$: For instance, for $\overline{x}_{0}=\sum_{i}(1-\pi_{i}^{*})x_{i}/\sum_{i}(1-\pi_{i}^{*})$, $\delta_{1}$ is given by
$\delta_{1}=\left({\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})\frac{\pi_{i}^{*}}{1-\pi_{i}^{*}}(x_{i}-\overline{x}_{0}).$
Thanks to the weighting by $1-\pi_{i}^{*}$, $(\delta_{0},\delta_{1})$ can nevertheless be consistently estimated without involving $\pi_{i}^{*}/(1-\pi_{i}^{*})$: In sample, the odds for the estimated implicit design has coefficients replaced by their sample counterparts: For $\hat{x}_{0}=\sum_{i}(1-W_{i})x_{i}/\sum_{i}(1-W_{i})$,
$\hat{\delta}_{1}=\left({\frac{1}{n}\sum_{i=1}^{n}(1-W_{i})(x_{i}-\hat{x}_{0})(x_{i}-\hat{x}_{0})^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}W_{i}(x_{i}-\hat{x}_{0}).$
Theorem 3.1(3) and (4) leave a few puzzles, which we resolve in the next subsection. Both can be viewed as taking a contrast in the interacted regression $Y_{i}=\gamma_{0}+\gamma_{1}^{\prime}x+\tau_{0}W_{i}+\tau_{1}^{\prime}x_{i}W_{i}+\epsilon_{i}$heorem 3.1(3) takes $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ while Theorem 3.1(4) takes $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{1}$. Curiously, Theorem 3.1(3) asks that the covariates be saturated, while Theorem 3.1(4) has no such requirement. Moreover, Theorem 3.1(4) can be applied—by symmetry of $W_{i}$ and $1-W_{i}$—to generate a result for the model-based average treatment effect on the untreated (ATU) estimand $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{0}$, where $\overline{x}_{0}=\frac{\sum_{i}(1-\pi_{i}^{*})x_{i}}{\sum_{i}(1-\pi_{i}^{*})}$. Doing so shows that $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{0}$ is a causal contrast only if the reciprocal propensity odds $(1-\pi_{i}^{*})/\pi_{i}^{*}$ is linear in $x_{i}$. Worryingly, despite generated by the same specification, the ATU thus seems to require a different design than the ATT. Our results in the next subsection detail this conflict.
Finally, Theorem 3.1(5) echoes and complements Athey and Imbens (2022), who analyze TWFE under staggered adoption and totally randomized treatment timing. Theorem 3.1(5) shows a converse: $\tau$ fails to satisfy level irrelevance under any other design. TWFE is commonly used under difference-in-differences assumptions. In these settings, the key identifying assumption is typically parallel trends, which certainly involves outcome modeling and is thus not purely design-based. This is partly why a strict design-based interpretation of $\tau$ is only available under pure random assignment. That said, since these estimators are often narratively justified by random timing assumptions [^29] For instance, Martinez-Bravo et al. (2022) write (emphasis ours), “we study the introduction of elections in the 1980s and 1990s […] This reform was mandated by the central government and rolled out in a top-down manner. We document that the timing of the first election is uncorrelated with a large set of village characteristics. This suggests that timing was quasi-random, which is consistent with the descriptive literature. Thus, we exploit the staggered timing of the introduction of elections across villages to estimate a difference-in-difference effect of the introduction of elections.” and applied to many outcomes at once (Roth and Sant’Anna, 2023), studying the designs under which the estimand has a outcome-model-free interpretation is nevertheless valuable. Section 3.3 enriches this result further by showing that the design-based interpretation of TWFE, already only possible under a stringent design, is fragile against two common complications—time-varying covariates and imbalanced panel. Section B.5 further provides extensions to one-way fixed effects and event-study specifications.
Since the implicit design corresponds exactly to the design studied by Athey and Imbens (2022), the estimand corresponds to theirs (their Theorem 1(ii)) when $\mathcal{W}$ represents staggered adoption. Furthermore, we can often inspect the weights put on $y_{it}(\mathbf{w})$, $\bm{\omega}_{t}(\bm{\pi}^{*},\mathbf{w})$ and verify if the weights satisfy some additional requirements. For instance, one might wish to impose that the post-treatment weights are nonnegative (i.e., $\bm{\omega}_{t}(\bm{\pi}^{*},\mathbf{w})\geq 0$ if $\mathbf{w}_{t}=1$). Failure of this condition implies that post-treatment units are severely used as comparisons for newly treated units, echoing the “forbidden comparison” issue in the recent difference-in-differences literature (Roth et al., 2023; Borusyak et al., 2024b; De Chaisemartin and d’Haultfoeuille, 2020; Goodman-Bacon, 2021). Proposition B.14 shows that when $\mathcal{W}$ only has two elements and includes a never treated unit, all weights post treatment are non-negative, but such forbidden comparisons are possible in all other cases.
3.2. Forbidden interactions?
Suppose $T=J=1$. Let $x_{1i},x_{2i}$ be subvectors of $x_{i}$, which may be disjoint, nested, identical, or overlapping. Consider the specification
$Y_{i}=\gamma_{0}+\tau_{0}W_{i}+\tau_{1}^{\prime}W_{i}x_{1i}+\gamma_{1}^{\prime}x_{2i}+\epsilon_{i}.$
Read as a linear model for potential outcomes (i.e., $\mathbb{E}[Y_{i}(w)\mid x_{i}]=\gamma_{0}+\gamma_{1}^{\prime}x_{2i}+w(\tau_{0}+\tau_{1}^{\prime}x_{1i})$), this specification models the baseline potential outcome as linear in $x_{2i}$ and the conditional average treatment effect as linear in $x_{1i}$, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid x_{i}]=\tau_{0}+\tau_{1}^{\prime}x_{1i}$. Under this outcome model, $\tau_{0}+\tau_{1}^{\prime}x_{1i}$ is a causal contrast for all $x_{1i}$. If the covariates $(1,x_{1i}^{\prime})$ are not collinear, then this means that $\bm{\tau}=(\tau_{0},\tau_{1}^{\prime})$ are themselves vectors of causal contrasts.
It is tempting to think that the causal contrast interpretation of $\bm{\tau}$ exists independently of the outcome model. Perhaps $x_{1}\mapsto\tau_{0}+\tau_{1}^{\prime}x_{1}$ is some linear approximation of the conditional average treatment effects. Thus, one might conjecture that without the outcome model but with the right design, $\tau_{0}+\tau_{1}^{\prime}x_{1}$ is nevertheless some weighted average of conditional average treatment effects evaluated at values around $x_{1}$.
Surprisingly, the next proposition shows that this is in fact implausible without a randomized experiment or saturating the covariates. It turns out that the causal interpretation of $\bm{\tau}$ necessitates that both $\pi_{i}^{*}$ and $\pi_{i}^{*}x_{1i}$ are linear in $x_{2i}$. When this fails, some contrast $\tau_{0}+\tau_{1}^{\prime}x_{1}$ does not satisfy level irrelevance.
Proposition 3.2.
Consider the specification (3.2) and let $\bm{\tau}=(\tau_{0},\tau_{1}^{\prime})^{\prime}$ be the coefficients of interest. Then its corresponding implicit design exists if and only if, for some conformable $\Gamma_{0},\Gamma_{1}$, for all $i$,
$(\delta_{0}+\delta_{1}^{\prime}x_{2i})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$
where $\delta_{0},\delta_{1}$ are population projection coefficients of $\pi_{i}^{*}$ on $x_{2i}$: $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{2i}+\eta_{i}$ and $\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}+\upsilon_{i}$. When this happens, the implicit design $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$ uniquely. Therefore, if $\bm{\tau}$ is a vector of causal contrasts under $\bm{\pi}^{*}$, then $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{2i}$ and $\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$.
The necessary condition for interpreting $\bm{\tau}$ as a vector of causal contrasts is that both the propensity score $\pi_{i}^{*}$ and its interaction with the covariates $\pi_{i}^{*}x_{1i}$ are linear functions of $x_{2i}$. When $x_{1i}$ is included in $x_{2i}$, this condition is unlikely to hold in general, as $\pi_{i}^{*}x_{1i}$ would involve nonlinear transformations of $x_{1i}$ and thus cannot be linear. Two common exceptions are available. First, this condition can hold if the propensity score is a constant function of $x_{2i}$ (i.e. $\delta_{1}=0$). Second, if $x_{1i}$ represents a saturated categorical variable and $x_{2i}$ contains all other covariates interacted with $x_{1i}$, [^30] That is, $x_{1i}$ contains mutually exclusive binary random variables, and $x_{2i}$ contains $x_{1i}$, some set of other covariates $x_{3i}$, and all interactions $x_{3ik}x_{1i\ell}$. then entries of $x_{1i}x_{2i}^{\prime}$ are in the linear span of $x_{2i}$ and the condition reduces to $\pi_{i}^{*}$ being linear in $x_{2i}$.
When $x_{1i}$ is a scalar variable, $\tau_{1}$ is often interpreted as some approximation of treatment effect heterogeneity comparing units with larger $x_{1i}$ to units with smaller $x_{1i}$, and $\tau_{0}$ is often interpreted as a baseline treatment effect for units with $x_{1i}=0$. Proposition 3.2 circumscribes the design-based interpretation of these treatment effect contrasts. If $x_{1i}$ is binary, and $x_{2i}$ includes $x_{1i}$ and its interactions with all other covariates, then $\bm{\tau}$ is a causal contrast only under linear propensity scores. On the other hand, if $x_{1i}$ is a continuous covariate and included in $x_{2i}$, then design-based interpretation requires both linear propensity scores and—implausibly—that the nonlinear quantity $\delta_{2}^{\prime}x_{2i}x_{1i}$ is in the linear span of $x_{2i}$.
One way to understand this difficulty is by reasoning through why $\tau_{0}+\tau_{1}^{\prime}x_{1i}$ is not a best linear approximation of the conditional average treatment effect, even when we think of the units as sampled from a superpopulation. One could think of (3.2) as two regressions, one on the treated $W=1$ and one on the untreated $W=0$. Both regressions are indeed best linear approximations to $\mathbb{E}[Y(1)\mid x,W=1]$ and $\mathbb{E}[Y(0)\mid x,W=0]$, which are equal to the mean potential outcomes $\mathbb{E}[Y(1)\mid x],\mathbb{E}[Y(0)\mid x]$ under unconfoundedness. The contrast $\tau_{0}+\tau_{1}^{\prime}x_{1}$ is then the difference of the fitted values of these two regressions. Importantly, however, the two regressions are only best linear approximations with respect to different distributions of the covariates ($x_{i}\mid W=1$ vs. $x_{i}\mid W=0$). Thus, their difference is not a best linear approximation to the conditional average treatment effect under any particular distribution of $x$. Shifting $Y(1)$ and $Y(0)$ by the same arbitrary amount therefore causes asymmetric behavior in the two regressions, leading to a failure of level irrelevance.
When $x_{1i}=x_{2i}=x_{i}$, this result supplements Theorem 3.1(3)–(4) by cementing the intuition that different contrasts of the interacted regression have design-based interpretations under different designs. These designs are often logically incompatible, unless the covariates are appropriately saturated or if the propensity score is constant. Indeed, insisting on all such contrasts being causal imposes a knife-edge condition for the design. As a manifestation of this phenomenon, Theorem 3.1(4) shows that the model-based ATT $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{1}$ maintains its interpretation under the design whose odds is linear, but the ATU requires that the reciprocal propensity odds is instead linear. Theorem 3.1(3) evades this difficulty by requiring that the covariates are saturated.
One might wonder what can be said of the model-based ATE $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ without saturating the covariates (cf. Theorem 3.1(3)). Theorem 1 in Chattopadhyay and Zubizarreta (2023) shows that, for this contrast to actually be the ATE under a design $\bm{\pi}^{*}$, $\bm{\pi}^{*}$ needs both $(1-\pi_{i}^{*})/\pi_{i}^{*}$ and $\pi_{i}^{*}/(1-\pi_{i}^{*})$ to be linear in the covariates, which is implausible unless $\pi_{i}^{*}$ is constant or the covariates are saturated. If we simply require $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ to be a causal contrast—but not necessarily the ATE—then the implicit design does exist and is equal to some fractional linear function $\pi_{i}=\frac{\theta_{0}+\theta_{1}^{\prime}(x-\overline{x})}{1-\Gamma_{2}^{\prime}(x-\overline{x})}$. The requirement that $\pi_{i}=\pi_{i}^{*}$ then imposes additional restrictions on $(\theta_{0},\theta_{1},\Gamma_{2})$. These contraints are unpleasant in the sense that they depend on the values $x_{1:n}$: The same design $\pi_{i}^{*}=\frac{\theta_{0}+\theta_{1}^{\prime}(x_{i}-\overline{x})}{1-\Gamma_{2}^{\prime}(x_{i}-\overline{x})}$ may be consistent with design-based interpretation of $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ for one set of covariates $x_{1:n}$ but not for another $\tilde{x}_{1:n}$, even though $\pi_{i}^{*}\in(0,1)$ under both sets of covariates. [^31] Note that the linear-odds parametrization of Kline (2011) is not unpleasant in this way. These issues are formalized in B.1 in Section B.4.
One way to read these results is that, without a randomized experiment, the interacted regression is less robust than the uninteracted regression (2.1), in terms of its design-based interpretation. [^32] This is in contrast to Lin (2013)’s arguments against Freedman (2008). Lin (2013) shows that the interacted regression is in many ways superior in a randomized experiment. The uninteracted regression introduces variance weighting for the estimand, but maintains validity under a simple design. The interacted regression removes the weighting when the covariates are saturated, but estimates a weighted average treatment effect under a bizarre design in general. Is there a simple regression that both removes the variance weighting and is design-based under linear propensity scores? Unfortunately, Proposition B.11 shows that the answer is no, at least not with right-hand-side variables that are linear in $[1,x_{i},W_{i},W_{i}x_{i}]$. [^33] One could estimate the uninteracted regression and weigh by $1/(\pi_{i}(1-\pi_{i}))$ to remove the variance weighting, but this approach requires estimating $\pi_{i}$ separately. As a result, targeting the average treatment effect under the same implicit design as (2.1) necessitates moving beyond simple linear regression methods.
3.3. The fragility of design-based TWFE
Theorem 3.1(5) shows that the standard TWFE regression has design-based interpretation only under total randomization of the treatment paths. This design is restrictive, and we may wonder whether it is possible to modify the regression so that designs are allowed to be functions of observable information. The most straightforward idea—indeed common empirical practice—is to include time-varying covariates in a TWFE regression.
Surprisingly, doing so can often destroy the design-based interpretation of the corresponding estimand altogether. In panel settings, usually the equations defining the implicit design are overdetermined, and so small modifications to the potential weights can make the system of equations unsolvable. With time-varying covariates, this happens precisely when the covariates matter for identification and change the estimand. On the other hand, when adding the covariates do not change the estimand, constant propensity scores remain the unique implicit design, and thus the covariates do not enrich the design either.
Proposition 3.3.
Assume $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$. Consider the regression $\mathbf{Y}_{it}=\alpha_{i}+\gamma_{t}+\tau\mathbf{W}_{it}+\delta^{\prime}\mathbf{x}_{it}$ where $\tau$ is the coefficient of interest. Let $\beta_{w\to x}$ be the population projection coefficient of $W_{it}$ on $x_{it}$ under $\bm{\pi}^{*}$, including individual and time fixed effects. For the set of implicit designs to be non-empty, a necessary condition is
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\})\,\text{ for all $i\in[n]$}.$
When $\beta_{w\to x}=0$, if $\mathcal{W}$ contains linearly independent vectors whose span excludes $1_{T}$, the implicit design is uniquely equal to $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$ for all $i$ as in Theorem 3.1(5).
Proposition 3.3 states that a necessary condition for the set of implicit designs under $\bm{\pi}^{*}$ to be nonempty is that a particular linear combination of the demeaned covariates sits in the linear span of $\mathcal{W}$ and $1_{T}$ for every unit. This condition arises because we essentially need that the mean treatment $\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\mathbf{w}$ is exactly described by two-way fixed effects with time-varying covariates, analogous to the intuition for (2.1). This then restricts the space of the covariate vectors, since they need to generate vectors that lie in the linear span of $\mathcal{W}$.
Under staggered adoption, $\operatorname{span}\left({\mathcal{W}\cup\left\{{1_{T}}\right\}}\right)$ is the set of vectors that are piecewise constant between adjacent adoption dates. These vectors are peculiar if there are relatively few adoption dates. If $\beta_{w\to x}$ is nonzero under $\bm{\pi}$, it is thus knife-edge that $\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}$ happens to be located in that subspace, unless columns of $\mathbf{x}_{i}$ happens to be piecewise constant over $t$ as well. On the other hand, if $\beta_{w\to x}$ is zero under $\bm{\pi}^{*}$, including the covariates makes no difference to the coefficient on $W_{it}$ in population. Therefore, Proposition 3.3 implies that TWFE with time-varying covariates does not in general have a design-based interpretation. Researchers using such a specification either believe that the covariates do not affect treatment assignment and are thus immaterial to identification from a design-based perspective, or they are implicitly embedding certain outcome modeling assumptions.
Proposition 3.3 shows that the design-based interpretation for TWFE is delicate, and seemingly perfunctory modifications to the regression can destory it altogether. In a similar vein, TWFE on an imbalanced panel also has this feature. Such a regression is only design-based when the missingness pattern is uncorrelated with the treatment assignment pattern, in which case the design must again be total randomization of treatment paths. To introduce this result, suppose units are observed in some subperiod $\mathcal{T}_{i}\subset\left\{{1,\ldots,T}\right\}$. We consider the TWFE specification
$\mathbf{Y}_{it}=\alpha_{i}+\gamma_{t}+\tau\mathbf{W}_{it}+\epsilon_{it}$
over $i=1,\ldots,n$ and $t\in\mathcal{T}_{i}$. The objects $\mathcal{T}_{1},\ldots,\mathcal{T}_{n}$ are considered fixed and observed.
[…]
Proposition 3.4.
Let the population regression specification be characterized by (3.3) under $\bm{\pi}^{*}$, where $\tau$ is the coefficient of interest. Suppose $\mathcal{W}$ satisfies staggered adoption and excludes always-treated units. Assume further that there is a commonly observed period $\mathcal{T}\subset\bigcap_{i=1}^{n}\mathcal{T}_{i}$ such that the treatment paths have rich variation in $\mathcal{T}$.
Let $\overline{\pi}(\mathbf{w})\equiv\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$. Let $Q_{i}(\mathbf{w})=\frac{\sum_{t\in\mathcal{T}_{i}}\mathbf{w}_{t}}{|\mathcal{T}_{i}|}$ be the proportion of treated periods within unit $i$’s observed period for treatment path $\mathbf{w}$. Then an implicit design exists if and only if $\bm{\pi}^{*}$ is uncorrelated with the missingness pattern in the sense that for all $t\in[T]$:
$\underbrace{\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\sum_{i:t\in\mathcal{T}_{i}}(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))}_{\sum_{i:t\in\mathcal{T}_{i}}\mathbb{E}_{\mathbf{w}\sim\overline{\pi}}[(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))]}=\underbrace{\sum_{\mathbf{w}\in\mathcal{W}}\sum_{i:t\in\mathcal{T}_{i}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))}_{\sum_{i:t\in\mathcal{T}_{i}}\mathbb{E}_{\mathbf{w}\sim\pi_{i}^{*}}[(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))]}.$
When this happens, the implicit design is unique and satisfies $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ for all $i$.
Proposition 3.4 gives a necessary and sufficient condition for an implicit design to exist, and characterizes the implicit design when it does. In short, an implicit design exists if and only if the missingness patterns are uncorrelated with the treatment timing in a particular sense (3.4). When it exists, the implicit design is the same for all units. Therefore, if $\bm{\pi}^{*}$ justifies $\tau$ as a design-based estimand in the sense of 2.4, then $\bm{\pi}^{*}$ must equal to the implicit design and thus must describe random assignment.
The condition (3.4) describes a situation where for any given time period $t$, among units for whom $t$ is observed, the average expected de-meaned treatment path under $\bm{\pi}^{*}$, $\mathbb{E}_{\mathbf{w}\sim\pi_{i}^{*}}[\mathbf{w}_{t}-Q_{i}(\mathbf{w})]$, is equal to its analogue under the random assignment $\overline{\pi}$. This condition is difficult to satisfy unless the design is constant $\pi_{i}^{*}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ in the first place. As a result, Proposition 3.4 shows that TWFE is fragile to imbalanced panels.
4. Empirical application
We illustrate several empirical recommendations (Section 2.3) by reanalyzing Blakeslee et al. (2020). Blakeslee et al. (2020) study the economic impact of water loss in rurual India, using borewell failure ($W_{i}$) as a quasi-experimentally assigned treatment. Blakeslee et al. (2020) consider a range of income and employment outcomes and conclude that (i) well failure causes a decline in agricultural income and employment, but reallocation to off-farm offsets the lost income, and (ii) those living in high economic development areas adapt more easily.
The main empirical analyses of Blakeslee et al. (2020) contain estimates of the following regression (their equation (1) reproduced in our notation): For a household $i$, let $v(i)$ denote its village and $t(i)$ denote the year when it drilled its first borewell. Consider the coefficient $\tau$ in
$Y_{i}=\tau W_{i}+\underbrace{\alpha_{0}+x_{i}^{\prime}\gamma+\eta_{1v(i)}+\eta_{2t(i)}}_{\check{x}_{i}^{\prime}\mu}+\epsilon_{i}.$
where $x_{i}$ collects a vector of household characteristics and includes “the age, caste, and literacy of the household head, as well as the total land inherited by the household” (p.208 Blakeslee et al., 2020). By Theorem 3.1(1) (Blandhol et al., 2022; Angrist, 1998), $\tau$ is a causal contrast only if $\pi_{i}^{*}$ is linear in $\check{x}_{i}$. When that happens, the estimand is a weighted average treatment effect with weights proportional to $\pi_{i}^{*}(1-\pi_{i}^{*})$.
In their heterogeneity analysis, Blakeslee et al. (2020) consider the regression that interacts (4) with an indicator $h_{i}$ of whether the village $v(i)$ has high economic development
$Y_{i}=\tau_{0}W_{i}+\tau_{1}W_{i}h_{i}+\underbrace{(\alpha_{0}+x_{i}^{\prime}\gamma_{0}+\eta_{1v(i)}+\eta_{2t(i)}+\alpha_{1}h_{i}+h_{i}x_{i}^{\prime}\gamma_{1}+\kappa_{2t(i)}h_{i})}_{\tilde{x}_{i}^{\prime}\mu}+\epsilon_{i},$
where $\tau_{0}$ is interpreted as a treatment effect for those with $h_{i}=0$ and $\tau_{1}$ is interpreted as a difference of treatment effects among $h_{i}=1$ versus $h_{i}=0$. Our result in Proposition 3.2 show that $\tau_{0}$ and $\tau_{1}$ in (4) are both causal contrasts only if $\pi_{i}^{*}$ is linear in $\tilde{x}_{i}$ and $\pi_{i}^{*}h_{i}$ is also linear in $\tilde{x}_{i}$. Because $h_{i}$ is in the linear span of the village fixed effects, it is easy to check that $\pi_{i}^{*}h_{i}$ is linear if $\pi_{i}^{*}$ is linear—and thus the implicit design exists. In this case, we may also compute and verify that the corresponding estimand $\tau_{1}$ takes the form of a difference of two variance-weighted estimands:
$\tau_{1}=\frac{1}{n_{1}}\sum_{i:h_{i}=1}\omega_{1i}(y_{i}(1)-y_{i}(0))-\frac{1}{n_{0}}\sum_{i:h_{i}=0}\omega_{0i}(y_{i}(1)-y_{i}(0))$
such that $\frac{1}{n_{1}}\sum_{i:h_{i}=1}\omega_{1i}=\frac{1}{n_{0}}\sum_{i:h_{i}=0}\omega_{0i}=1$, and $\omega_{1i},\omega_{0i}\propto\pi_{i}^{*}(1-\pi_{i}^{*})$.
Interestingly, the implicit designs underlying (4) and (4) are different, as (4) involves additional covariates. Since the specifications are nested, one can test $H_{0}:\pi_{i}^{*}$ is linear in $\check{x}_{i}$ against $H_{1}:\pi_{i}^{*}$ is linear in $\tilde{x}_{i}$. A standard Wald test does reject $H_{0}$ in favor of $H_{1}$, providing evidence against the implicit design for (4). In the remainder of the section, we focus on the specification (4), as well as the version of (4) that includes $\tilde{x}_{i}$ instead of $\check{x}_{i}$ as the covariates:
$Y_{i}=\tau W_{i}+\tilde{x}_{i}^{\prime}\mu+\epsilon_{i}.$
Specifications (4) and (4) in turn have the same implicit design.
[…]
These diagnostics on the implicit design undermine a design-based interpretation of the specification (4), answering (Q1). However, at least as an illustration, it may be interesting to understand the regression specification further. Even ignoring the various deficiencies of $\hat{\bm{\pi}}$, we may still be concerned with various choices in the implicit estimand. For instance, the specification (4) includes variance weighting, and we may ask—holding the propensity model fixed—how much the variance weighting alters the estimate. The estimand $\tau_{1}$ in (4) is a difference of two variance-weighted estimands. A priori, we cannot rule out that the difference is driven by the weighting scheme compared to the difference in average treatment effects. The remainder of this section investigates these questions.
We can simply computing a battery of alternative estimates, using the implicit design. These are shown in LABEL:fig:alt_coefficients for both the treatment effect $\tau$ in (4) and the treatment effect difference $\tau_{1}$ in (4). First, the simplest assessment is whether the units with out-of-bounds implicit design feature prominently in the regression estimate. To that end, the variance-weighted estimates $\bm{\times}$ uses the same implicit design and weighting scheme as the implicit estimand, but removes the out-of-bounds units. For both $\tau_{0}$ and $\tau_{1}$, these $\bm{\times}$ estimates are almost identical to the regression estimates, indicating that the regression estimates put relatively little weight on out-of-bounds units. Removing out-of-bounds units by patching the implicit design $\star$ also does not meaningfully alter the estimate.
Second, we may assess whether the variance-weighting in the implicit estimand matters by considering weighting schemes that treat units more equally. Since the propensity score estimates are often close or equal to zero and one, overlap violations make estimating the average treatment effect infeasible. Thus, we trim the propensity scores to $[0.02,0.98]$ and construct corresponding estimators for the trimmed average treatment effect (Crump et al., 2009). These estimates—especially the patched estimates $\star$ [^36] The patched implicit design has no observations within $(0,0.02)\cup(0.98,1)$; thus the trimming only trims out those with overlap failure.—are more different from the regression estimates, though not substantially so compared to sampling noise. On average, the difference between $\star$ and the regression estimates is about 0.5–0.7 times the standard error of the regression estimate.
Finally, moving entirely away from the implicit designs in the regression, we also compute estimates by augmented inverse propensity weighting ($\bm{A}$IPW) by using a simple logit model for the propensity scores and a linear model for the outcome means. These alternative estimates are again similar to the regression estimates, indicating that the outcomes in this application are not so adversarially configured: The implicit design, while clearly rejected, nevertheless produces estimates that are reasonable.
(a) Treatment effect $\tau$ in (4) Refer to caption
(b) Heterogeneous treatment effect $\tau_{1}$ in (4) Refer to caption
Figure 1. The correlations of outcomes and implicit designs
We can also unpack why the alternative weighting schemes do not appear to make a difference by examining the correlation between the outcomes and the implicit design. Figure 1(a) partitions the implicit designs into 7 bins by quantile, so that each bin within $[0,1]$ contains the same number of units. On each bin, it displays the mean treated and mean untreated outcomes as well as the weight placed on each bin by the implicit estimand. As the bin size become small, computing the difference between the blue and orange curves, weighted by the weights, approximates the regression estimate. As expected, the regression upweights units with implicit design closer to 0.5. In particular, the out-of-bounds units receive a total of about $-3\%$ weighting, explaining their apparent irrelevance to the bottom line. Remarkably, for agricultural employment, the treatment effects hardly vary by the implicit design. Households with low and high propensity to lose water access appear to be affected similarly, with about 10% fewer members engaged in agricultural work, suggesting borewell failure is orthogonal to vulnerability to water loss.
Figure 1(b) considers the analogous decomposition for $\tau_{1}$. [^37] In Figure 1(b), if everyone is treated in a bin, we treat the mean control outcome as zero, and vice versa. Thus the “treatment effects” for $\hat{\pi}_{i}\leq 0$ represent negative mean control outcomes, and the “treatment effects” for $\hat{\pi}_{i}>1$ represent mean treated outcomes. The weighting schemes for $h_{i}=1$ versus $h_{i}=0$ are indeed different in $\tau_{1}$: High employment area ($h_{i}=1$) puts larger weight for households more likely to lose water access—peaking at the bin $(0.605,0.719]$ as opposed to at $(0.435,0.605]$. Thus the apparent treatment effect heterogeneity reflects in part the difference in weighting. But since the differences in conditional average treatment effects are effectively constant and zero, the different weighting again makes little difference to the bottom line estimate.
Our exercise demonstrates a few ways to use implicit designs and estimands. One can examine the distribution and calibration properties of the estimated implicit design $\hat{\bm{\pi}}$ (LABEL:fig:dist_implicit_design and LABEL:fig:calibration). One can use $\hat{\bm{\pi}}$ to construct implement alternative estimands that maintain the implicit model of treatment assignment (LABEL:fig:alt_coefficients). One can also examine how outcomes covary with observed selection (Figure 1). These exercises add transparency and evaluate robustness of the regression estimates, especially if practitioners, like Blakeslee et al. (2020), stake their claim on design-based causal identification.
To summarize our substantive findings, the implicit design for (4) and (4) appears to be untenable given the data. The problems with the implicit design aside, the corresponding weighting in the implicit estimand plays a minor role in the estimates. Because the outcomes are remarkably constant across values of implicit designs, implementations of alternative estimands under minor modifications of the estimated design are quite similar to the regression estimates. Modeling the design via a simple logit model and using a doubly robust estimator also does not substantively alter conclusions.
5. Conclusion
Linear regressions are ubiquitous. Interpreting their results as causal, thanks to random assignment, is similarly commonplace. This paper studies the necessary implications of so interpreting. We do so by studying the comparisons that regression estimands make, in terms of potential outcomes, under random assignment. For these comparisons to admit a causal interpretation—in the absence of assumptions on the outcomes—the true process of treatment assignment must satisfy certain restrictions. Given the Gram matrix of a regression, these restrictions are then linear in the design and often exactly pin down the distribution of treatment. The resulting set of designs that satisfy these restrictions can be thought of as models of treatment assignment that the regression implicitly estimates. Each design also pinpoints a corresponding estimand that the regression implicitly chooses.
Understanding design-based interpretation of regressions in this way essentially reduces to mechanical computations that can be scaled and automated. These computations can aid in examining new theoretical properties of particular specifications, itself the subject of a highly influential recent literature. In several theoretical vignettes, these computations unify and strengthen disparate strands of the literature. Additionally, we find that regressions with interactions and with two-way fixed effects have fragile design-based interpretations that call for caution and nuance when using them and presenting their results.
Directly computing implicit designs and estimands in practice provides a set of simple diagnostics for practitioners who wish to understand the design-based properties of a given regression. Doing so makes transparent the statistical and economic choices masked by a regression specification—especially when the regression is read simply as an algorithm for comparisons. Having opened up the black box, we can examine each of its components: e.g., evaluating whether the implicit design is plausible, improving the implicit design if it has obvious deficiencies, assessing whether the regression targets an economically interesting estimand, and constructing estimates for alternative estimands. Additionally, making these implicit choices transparent may nudge design-minded practitioners to wrest the wheel from familiar methods and choose methods that are more natively design-based.
We conclude by speculating on design-based statistical inference following a regression. If the regression has a unique implicit design and is interpreted under random treatment assignment, it may be attractive to consider design-based inference—especially in settings where superpopulation sampling is difficult to conceptualize (e.g. spatial settings). Design-based inference allows one to quantify uncertainty without appealing to sampling, and sometimes without asymptotic arguments. Design-based inference usually assumes a known assignment process, which is often difficult to articulate and specify in observational studies. For instance, Borusyak and Hull (2023) specify counterfactual assignment of Chinese high-speed rail lines by permuting the ordering of line completion in time. Pollmann (2020) proposes to find counterfactual locations for a spatial treatment using convolutional neural networks. Our approach potentially provides an additional option in observational settings to elicit the assignment process by tethering it to a regression specification that one contemplates using. Since the implicit design is often unique, when a regression-using researcher is internally consistent, then they are implicitly specifying the design.
In practice, we could consider design-based inference taking the estimated implicit design $\hat{\pi}_{i}$ as the true design, ignoring the fact that $\hat{\pi}_{i}$ is estimated from data. We anticipate doing so is justifiable when the $\Pi^{*}$ is a fixed Gram design. When it is not, we nevertheless suspect doing so to be justifiable in large samples when $\hat{\bm{\pi}}$ is consistent, but we leave formal justification to future work.
Appendix A Proofs
See 2.1
Proof.
We first show that 2.4 is equivalent to the design satisfying (2.2) for a population regression estimand. By (2.7), $\bm{\tau}$ satisfies 2.4 if and only if for all $i,t$
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})c_{it}=0\iff\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=0.$
The sum-to-one condition is automatically satisfied by $\pi_{i}^{*}$. This is then equivalent to (2.2).
[…]
When $\bm{\pi}^{*}$ is an implicit design, if there is only one implicit design, then $\bm{\pi}^{*}$ must be that design. As a result, that design is proper and Gram-consistent. ∎
See 2.2
Proof.
For the $\implies$ direction, if $\tau$ is a causal contrast, then $\bm{\pi}^{*}$ is an implicit design. Hence
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0.$
[…]
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0$
for all $i$ where one of $\rho_{i}(1),\rho_{i}(0)$ is nonzero by construction. For the other units (i.e., those with $\rho_{i}(1)=\rho_{i}(0)=0$), the condition holds trivially, irrespective of $\pi_{i}^{*}$, since both potential weights are zero. Thus $\tau$ satisfies level irrelevance under $\bm{\pi}^{*}$.
Note that, when this happens, $\omega_{i}(\bm{\pi}^{*},0)=(1-\pi_{i}^{*})\rho_{i}(0)=-\pi_{i}^{*}\rho_{i}(1)$. This proves the representation of $\tau$ by (2.7). Moreover,
$\pi_{i}^{*}\rho_{i}(1)<0\iff\rho_{i}(1)<0\text{ and }\rho_{i}(0)\neq 0\iff\rho_{i}(1)<0<\rho_{i}(0)$
since $\rho_{i}(1)\rho_{i}(0)\leq 0$. ∎
See 2.3
Proof.
Take $\eta>0$. For the result with fixed $i$, it suffices to show that $\operatorname{P}[\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}>\eta]\to 0$. Note that
$\operatorname{P}\left({\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}>\eta}\right)$ $\leq\operatorname{P}[\lambda_{\min}(\hat{G}_{n})\leq\epsilon/2]$
$+\operatorname{P}\left[{\lVert\Lambda(\hat{G}_{n}^{-1}-G_{n}^{-1})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\rVert_{F}>\eta,\lambda_{\min}(\hat{G}_{n})>\epsilon/2}\right].$
The first term converges to zero by Assumptions 2.1 and 2.2 and the Hoffman–Wielandt inequality. Note that, by the submultiplicativity of the Frobenius norm, when $\lambda_{\min}(\hat{G}_{n})>\epsilon/2$,
[…]
Thus, for some $C$ dependent on $\epsilon$, $\Lambda$, and $\lVert z_{t}(\mathbf{x}_{i},\mathbf{w})\rVert_{\infty}$,
$\operatorname{P}\left[{\lVert\Lambda(\hat{G}_{n}^{-1}-G_{n}^{-1})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\rVert_{F}>\eta,\lambda_{\min}(\hat{G}_{n})>\epsilon/2}\right]\leq\operatorname{P}\left[{C\lVert G_{n}-\hat{G}_{n}\rVert_{F}>\eta}\right]\to 0.$
[…]
To prove convergence that is uniform in $i$, note that, if $\lambda_{\min}(\hat{G}_{n})>\epsilon/2$, up to constants,
$\max_{i}|\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})|_{\infty}$ $\lesssim\max_{i}\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}$
$\lesssim\max_{i}\lVert\mathbf{z}(\mathbf{x}_{i},\mathbf{w})\rVert_{F}\cdot\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}$
$\lesssim\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}\cdot\max_{i}\max_{t}\lVert z_{t}(\mathbf{x}_{i},\mathbf{w})\rVert_{\infty}$
$\lesssim\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}$
$\lesssim\lVert G_{n}-\hat{G}_{n}\rVert_{F}.$
Therefore,
$\operatorname{P}\left({\max_{i}|\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})|_{\infty}>\eta}\right)$ $\leq\operatorname{P}[\lambda_{\min}(\hat{G}_{n})\leq\epsilon/2]+\operatorname{P}\left[{\lVert G_{n}-\hat{G}_{n}\rVert_{F}\gtrsim\eta}\right]\to 0.$
See 2.4
Proof.
Let $\hat{R}_{i}=[\hat{\bm{\rho}}_{i}(0),\ldots,\bm{\rho}_{i}(J)]$ be the sample analogue of $R_{i}$. Both $\hat{R}_{i}$ and $R_{i}$ may depend on $n$. Assume $n$ is sufficiently large such that $\min_{i\in[n]}\sigma_{iJ}>\epsilon/2>0$. Thus, each $R_{i}$ is a $J\times(J+1)$ matrix with rank $J$ such that $R_{i}\pi_{i}^{*}=0$. As a minor abuse of notation, let $\pi_{i},\pi_{i}^{*}$ be represented as vectors in $\mathbb{R}^{J+1}$. Let $u_{i}=\pi_{i}^{*}/\lVert\pi_{i}^{*}\rVert$ be the unit vector in the (one-dimensional) null space of $R_{i}$. Note that $11/(J+1)\leq\lVert\pi_{i}^{*}\rVert_{\infty}\leq\lVert\pi_{i}^{*}\rVert\leq\lVert\pi_{i}^{*}\rVert_{1}=$ 1, and thus $1^{\prime}u_{i}=1/\lVert\pi_{i}^{*}\rVert\in[1,J+1]$ is bounded above and below.
Fix some $C$ to be chosen. Let $E$ be complement to the event that there exists a unique bounded estimated implicit design. Note that $E=\bigcup_{i\in[n]}E_{i}$, where each $E_{i}$ is the complement to the event that $\hat{R}_{i}\hat{\pi}_{i}=0$ has a unique solution $\hat{\pi}_{i}$ with $11^{\prime}\hat{\pi}_{i}=$ 1 where $\lVert\hat{\pi}_{i}\rVert_{\infty}\leq C$. Let $\hat{\sigma}_{i1}\geq\cdots\geq\hat{\sigma}_{iJ}\geq 0$ be the singular values of $\hat{R}_{i}$. Suppose $\hat{\sigma}_{iJ}>0$, then $\hat{R}_{i}$ has full rank and the equation $\hat{R}_{i}\hat{\pi}_{i}=0$ has a unique one-dimensional space of solutions. If that space exists, let $\hat{u}_{i}$ be a unit vector in that space, unique up to sign. If $1^{\prime}\hat{u}_{i}\neq 0$, then we choose the sign of $\hat{u}_{i}$ so that $01^{\prime}\hat{u}_{i}>$ 0. Note that if $|\hat{u}_{i}^{\prime}1|>1/C$, then $\hat{\pi}_{i}=\hat{u}_{i}/\hat{u}_{i}^{\prime}1$ is a bounded estimated implicit design with $\lVert\hat{\pi}_{i}\rVert_{\infty}\leq C$.
Note that
$E_{i}\subset\left\{{\hat{\sigma}_{iJ}\leq\epsilon/4}\right\}\cup\left\{{\hat{\sigma}_{iJ}>\epsilon/4,|\hat{u}_{i}^{\prime}1|<1/C}\right\}.$
We will show that $E_{i}\subset\left\{{\lVert\hat{R}_{i}-R_{i}\rVert_{F}>\kappa}\right\}$ for some $\kappa$. For the event $\left\{{\hat{\sigma}_{iJ}\leq\epsilon/4}\right\}$, the Hoffman–Wielandt inequality implies that
$|\hat{\sigma}_{iJ}-\sigma_{iJ}|^{2}\leq\sum_{k=1}^{J}|\hat{\sigma}_{ik}-\sigma_{ik}|^{2}\leq\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2},$
and thus $\hat{\sigma}_{iJ}\geq\epsilon/2-\lVert R_{i}-\hat{R}_{i}\rVert_{F}$, which further implies that $\left\{{\hat{\sigma}_{iJ}<\epsilon/4}\right\}\subset\left\{{\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\epsilon/4}\right\}.$
Let $\hat{u}_{i}=\hat{c}_{i}u_{i}+\sqrt{1-\hat{c}_{i}^{2}}\hat{u}_{\perp,i}$ where $\hat{c}_{i},\hat{u}_{\perp,i}$ are uniquely chosen so that $\hat{u}_{\perp,i}$ is a unit vector orthogonal to $u_{i}$. Then
$|\hat{u}_{i}^{\prime}1|\geq|\hat{c}_{i}1^{\prime}u_{i}|-\sqrt{1-\hat{c}_{i}^{2}}|1^{\prime}\hat{u}_{\perp,i}|\geq|\hat{c}_{i}|-\sqrt{1-\hat{c}_{i}^{2}}\sqrt{J+1}.$
Now, note that
$\lVert R_{i}\hat{u}_{i}\rVert=\lVert(R_{i}-\hat{R}_{i})\hat{u}_{i}\rVert\leq\lVert R_{i}-\hat{R}_{i}\rVert_{F}.$
On the other hand,
$\lVert R_{i}\hat{u}_{i}\rVert=\sqrt{1-\hat{c}_{i}^{2}}\lVert R_{i}\hat{u}_{\perp,i}\rVert\geq\sqrt{1-\hat{c}_{i}^{2}}\sigma_{iJ}\geq\sqrt{1-\hat{c}_{i}^{2}}\frac{\epsilon}{2},$
and thus
$1-\hat{c}_{i}^{2}\leq\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}\iff|\hat{c}_{i}|\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}}.$
Hence,
$|\hat{u}_{i}^{\prime}1|\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}}-\frac{2\sqrt{J+1}}{\epsilon}\lVert\hat{R}_{i}-R_{i}\rVert_{F}.$
Suppose $\lVert R_{i}-\hat{R}_{i}\rVert_{F}<\epsilon/(4\sqrt{J+1})$, then
$|\hat{u}_{i}^{\prime}1|\geq\sqrt{1-\frac{1}{J+1}}-\frac{1}{2}\geq\sqrt{1/2}-1/2\geq 0.207.$
Thus, if we pick $C>1/0.207$, then
$E_{i}\subset\left\{{\hat{\sigma}_{iJ}<\epsilon/4}\right\}\cup\left\{{\hat{\sigma}_{iJ}>\epsilon/4,|\hat{u}_{i}^{\prime}1|<1/C}\right\}\subset\left\{{\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\frac{\epsilon}{4\sqrt{J+1}}}\right\}.$
As a result,
$\operatorname{P}\left[{\bigcup_{i}E_{i}}\right]\leq\operatorname{P}\left[{\max_{i\in[n]}\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\frac{\epsilon}{4\sqrt{J+1}}}\right]\to 0$
by assumption. This proves the first part regarding the existence of a unique bounded implicit design.
For the second part, note that when $\hat{\sigma}_{iJ}>0$,
$\hat{\pi}_{i}-\pi_{i}=\frac{\hat{u}_{i}}{1^{\prime}\hat{u}_{i}}-\frac{u_{i}}{1^{\prime}u_{i}}=\frac{1^{\prime}}{1^{\prime}\hat{u}_{i}}(u_{i}-\hat{u}_{i})\frac{\hat{u}_{i}}{1^{\prime}u_{i}}+\frac{1}{1^{\prime}u_{i}}(\hat{u}_{i}-u_{i}).$
Thus
$\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}\leq\frac{1}{0.207}\sqrt{J+1}\lVert u_{i}-\hat{u}_{i}\rVert_{2}+\lVert\hat{u}_{i}-u_{i}\rVert_{2}$
if $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\epsilon/(4\sqrt{J+1})$.
To bound $\lVert\hat{u}_{i}-u_{i}\rVert$, we first show that $\hat{c}_{i}>0$ when $R_{i}$ is close to $\hat{R}_{i}$. When $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, we note that by (A),
$|\hat{c}_{i}|\geq\sqrt{1-\frac{1}{4(J+1)}}\geq\sqrt{7/8}.$
When $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, we also have that, because we pick $\hat{u}_{i}$ to have $1^{\prime}\hat{u}_{i}\geq 0$,
$0.207\leq 1^{\prime}\hat{u}_{i}=\hat{c}_{i}+\sqrt{1-\hat{c}_{i}^{2}}1^{\prime}\hat{u}_{\perp,i}\leq\hat{c}_{i}+\sqrt{1-\hat{c}_{i}^{2}}\sqrt{J+1}\leq\hat{c}_{i}+\frac{2}{\epsilon}\frac{\epsilon}{4\sqrt{J+1}}\sqrt{J+1}=\hat{c}_{i}+\frac{1}{2}.$
Thus $\hat{c}_{i}>0.207-0.5$. Since $|\hat{c}_{i}|>\sqrt{7/8}$, we conclude that $\hat{c}_{i}>\sqrt{7/8}>0$. Thus, when $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, (A) implies
$c_{i}\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}}\geq 1-\frac{4}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}.$
Now, when $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$,
$\lVert\hat{u}_{i}-u_{i}\rVert_{2}^{2}$ $\leq(1-\hat{c}_{i})^{2}+(1-\hat{c}_{i}^{2})$
$=2-2\hat{c}_{i}\leq\frac{8}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}$
Therefore, for some constant $C(J,\epsilon)$,
$\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}\leq C(J,\epsilon)\lVert R_{i}-\hat{R}_{i}\rVert_{F}.$
Hence, for $\eta>0$,
$\left\{{\max_{i\in[n]}\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}>\eta}\right\}\subset\left\{{\max_{i\in[n]}\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\max\left({\frac{\epsilon}{4\sqrt{J+1}},\frac{1}{C(J,\epsilon)}}\right)}\right\}$
The probability on the right-hand side converges to zero by assumption. This proves the convergence of $\hat{\pi}_{i}$ to $\pi_{i}$ uniformly in $i$ in $\lVert\cdot\rVert_{2}$. Since $\lVert\cdot\rVert_{\infty}\leq\lVert\cdot\rVert_{2}$ in $\mathbb{R}^{J+1}$, this concludes the proof. ∎
See 3.1
Proof.
We prove the following:
  1. (1)
    By Theorem B.1, the potential weights of this regression are the same as those for the Frisch–Waugh–Lovell-transformed regression of $Y_{i}$ on $W-\delta^{\prime}x_{i}$. For the FWL-transformed regression, its Gram matrix is a positive scalar
    $\tilde{G}_{n}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\delta^{\prime}x_{i})^{2}+(1-\pi_{i}^{*})(\delta^{\prime}x_{i})^{2}>0$
    Thus the potential weights are
    $\rho_{i}(w)=\tilde{G}_{n}^{-1}(w-x_{i}^{\prime}\delta).$
    Since $\tilde{G}_{n}>0$, $\rho_{i}(1)\neq\rho_{i}(0)$ for all $i$.
    1. (a)
      By 2.2, the implicit design is equal to
      $\pi_{i}=-\frac{\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}=x_{i}^{\prime}\delta.$
      To show Gram-consistency of $\pi_{i}$, note entries of the Gram matrix of the original regression take the form
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}z_{k}(x_{i},1)z_{\ell}(x_{i},1)+(1-\pi_{i})z_{k}(x_{i},0)z_{\ell}(x_{i},0)$
      where $z_{k}(x_{i},w)z_{\ell}(x_{i},w)$ take one of the following forms: $1,w,wx_{i\ell},x_{ik},x_{i\ell}x_{ik}$. Thus, it suffices to check that
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}$
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}x_{i}.$
      Both are true since $\pi_{i}$ is the projection of $\pi_{i}^{*}$ onto $x_{i}$, which contains a constant.
    2. (b)
      Immediately, $\pi_{i}^{*}=\pi_{i}$ implies $\pi_{i}^{*}=x_{i}^{\prime}\delta$. On the other hand, if $\pi_{i}^{*}=x_{i}^{\prime}\delta$, then the corresponding implicit design is equal to $x_{i}^{\prime}\delta$ by (a).
    3. (c)
      We may compute
      $\omega_{i}=\pi_{i}\rho_{i}(1)=\tilde{G}_{n}^{-1}x_{i}^{\prime}\delta(1-x_{i}^{\prime}\delta).$
      and that
      $\tilde{G}_{n}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{W_{i}^{2}-2W_{i}\pi_{i}+\pi_{i}^{2}}\right]$
      $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}}_{=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}\text{ since $x_{i}$ includes a constant}}-2\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}\pi_{i}+\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{2}$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}-2\frac{1}{n}\sum_{i=1}^{n}(\underbrace{\pi_{i}^{*}-\pi_{i}}_{\text{orthogonal to $\pi_{i}$}}+\pi_{i})\pi_{i}+\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{2}$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(1-\pi_{i}).$
      Lastly, when $\pi_{i}=\pi_{i}^{*}\in(0,1)$, $\omega_{i}>0$ since $\tilde{G}_{n}>0$.
  2. (2)
    By Theorem B.1, the potential weights of this regression are the same as those for the FWL-transformed regression
    $Y_{i}=\sum_{j}\tau_{j}(W_{ij}-\delta_{j}^{\prime}x_{i})+\epsilon_{i}$
    The Gram matrix for this regression is a positive definite matrix whose $k\ell$th entry is:
    $\tilde{G}_{n,k\ell}=\begin{cases}-\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(k)(1-\delta_{k}^{\prime}x_{i})\delta_{\ell}^{\prime}x_{i}+\pi_{i}^{*}(\ell)\delta_{k}^{\prime}x_{i}(1-\delta_{\ell}x_{i})+(1-\pi_{k}^{*}-\pi_{\ell}^{*})(\delta_{k}^{\prime}x_{i})(\delta_{\ell}^{\prime}x_{i}),&\text{if }k\neq\ell\\ \frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(k)(1-\delta_{k}^{\prime}x_{i})^{2}+(1-\pi_{i}^{*}(k))(\delta_{k}^{\prime}x_{i})^{2}&\text{if }k=\ell.\end{cases}$
    As a result, the potential weights are of the form
    $\rho_{i}(j)=\tilde{G}_{n}^{-1}\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)$
    where $e_{j}$ is the $j$th standard basis vector and $e_{0}=0_{J}$ is the vector of zeros.
    1. (a)
      Set $\pi_{i}(0)=1-\sum_{j=1}^{J}\pi_{i}(j)$. The equations (2.2) are then equivalent to
      $\tilde{G}_{n}^{-1}\sum_{j=1}^{J}\pi_{i}(j)\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)=0.$
      Since $\tilde{G}_{n}$ is positive definite, we must then have
      $\sum_{j=1}^{J}\pi_{i}(j)\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)=0$
      The $j$th coordinate of the above equations is of the form
      $0=\pi_{i}(j)-\pi_{i}(j)\delta_{j}^{\prime}x_{i}-\delta_{j}^{\prime}x_{i}\sum_{\ell\in\left\{{0,\ldots,J}\right\}\setminus\left\{{j}\right\}}\pi_{i}(\ell)=\pi_{i}(j)-\delta_{j}^{\prime}x_{i}.$
      The implicit design is thus $\pi_{i}(j)=\delta_{j}^{\prime}x_{i}$.
      The proof for Gram consistency is analogous to the case in (1)(a).
    2. (b)
      Analogous to (1)(b)
    3. (c)
      Inspecting $\omega_{i}(\bm{\pi},\ell)$, we compute that
      $\omega_{i}(\bm{\pi},\ell)=\tilde{G}_{n}^{-1}\pi_{i}(\ell)\left({e_{\ell}-\begin{bmatrix}\pi_{i}(1)\\ \vdots\\ \pi_{i}(J)\end{bmatrix}}\right)$
      Thus the implicit estimand is
      $\begin{bmatrix}\tau_{1}\\ \vdots\\ \tau_{J}\end{bmatrix}=\frac{1}{n}\sum_{i}\sum_{\ell=0}^{J}\omega_{i}(\bm{\pi},\ell)y_{i}(\ell).$
      For the contamination statement, it suffices to show an example. The setup in Example A.1 features a contaminated estimand.
  3. (3)
    The specification is equivalent to the following specification:
    $Y_{i}=\sum_{\ell=0}^{L}\alpha_{\ell}+\tau_{\ell}W_{i}$
    where $\tau=\left({1-\sum_{\ell=1}^{L}\overline{x}_{\ell}}\right)\tau_{0}+\sum_{\ell=1}^{L}\overline{x}_{\ell}\tau_{\ell}\equiv\sum_{\ell=0}^{L}\overline{x}_{\ell}\tau_{\ell}$.
    For this specification (3), let us order the covariates to be
    $[x_{i0},x_{i0}W_{i},x_{i1},x_{i1}W_{i},\ldots,x_{iL},x_{iL}W_{i}].$
    Note that the population Gram matrix is of the form
    $G_{n}=\begin{bmatrix}G_{0}\\ &G_{1}\\ &&\ddots\\ &&&G_{L}\end{bmatrix}$
    where
    $G_{\ell}=\begin{bmatrix}\overline{x}_{\ell}&\overline{x}_{\ell}\pi(\ell)\\ \overline{x}_{\ell}\pi(\ell)&\overline{x}_{\ell}\pi(\ell)\end{bmatrix}$
    where $\pi(\ell)$ is the mean of $\pi_{i}^{*}$ among those with $c_{i}=\ell$. Since $G_{n}$ is block-diagonal, its inverse is similarly block-diagonal with
    $G_{\ell}^{-1}=\frac{1}{\overline{x}_{\ell}\pi(\ell)(1-\pi(\ell))}\begin{bmatrix}\pi(\ell)&-\pi(\ell)\\ -\pi(\ell)&1\end{bmatrix}$
    on the diagonal. Note that $\Lambda=[0,\overline{x}_{0},0,\overline{x}_{1},\ldots,0,\overline{x}_{L}]$. Thus
    $\Lambda G_{n}^{-1}=\left[{\frac{-\pi(0)}{\pi(0)(1-\pi(0))},\frac{1}{\pi(0)(1-\pi(0))},\ldots,\frac{-\pi(L)}{\pi(L)(1-\pi(L))},\frac{1}{\pi(L)(1-\pi(L))}}\right].$
    For someone with $x_{i}=\ell$, the covariate transform is of the form
    $z(x_{i},w)=[0,\ldots,0,1,w,0,\ldots,0]^{\prime}$
    where it is nonzero at the $\ell$th pair. Thus
    $\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)=\frac{w-\pi(\ell)}{\pi(\ell)(1-\pi(\ell))}.$
    1. (a)
      By 2.2, the only implicit design sets $\pi_{i}=\pi(\ell)$ for $c_{i}=\ell$. This design is Gram-consistent by our explicit calculation of the Gram matrix.
    2. (b)
      Analogous to (1)(b).
    3. (c)
      Note that the implicit estimand is described by $\omega_{i}=\pi(\ell)\rho_{i}(1)=1.$ Therefore the implicit estimand is the average treatment effect.
  4. (4)
    The proof for this result is rather long and is delegated to self-contained claims. In the notation setup by Proposition B.9, the potential weights are
    $\rho_{i}(w)=\tilde{G}_{n}^{-1}(w-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{21}^{\prime}(w(x_{i}-\overline{x}_{1})-\Gamma_{01}-\Gamma_{11}(x_{i}-\overline{x}_{1})))$
    where
    $\tilde{G}_{n}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{\left\{{W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{21}^{\prime}(W_{i}(x_{i}-\overline{x}_{1})-\Gamma_{01}-\Gamma_{11}(x_{i}-\overline{x}_{1}))}\right\}^{2}}\right].$
    1. (a)
      The implicit design is shown in Proposition B.10. Proposition B.10 shows that the odds for the implicit design is equal to
      $\pi_{i}/(1-\pi_{i})=a_{0}+a_{1}^{\prime}(x-\overline{x}_{0})$
      where $a_{0},a_{1}$ are population projection coefficients of $\pi_{i}^{*}/(1-\pi_{i}^{*})$ on $x-\overline{x}_{0}$, weighted by $1-\pi_{i}^{*}$. Since it does not matter where we center $x_{i}$ for this regression, as it includes an intercept, we can reparametrize to show that $\pi_{i}/(1-\pi_{i})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$ as described.
    2. (b)
      Immediately, $\pi_{i}^{*}=\pi_{i}$ implies that $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$. On the other hand, if $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$, then the corresponding implicit design has odds equal to $\pi_{i}/(1-\pi_{i})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})=\pi_{i}^{*}/(1-\pi_{i}^{*})$.
    3. (c)
      We calculate, in the notation of Proposition B.9
      $\tilde{G}_{n}\rho_{i}(1)\pi_{i}=\left\{{(1-\Gamma_{21}^{\prime}(x_{i}-\overline{x}_{1}))-\theta_{01}-\theta_{11}^{\prime}(x_{i}-\overline{x})}\right\}\pi_{i}.$
      By the proof of Proposition B.10,
      $(1-\Gamma_{21}^{\prime}(x_{i}-\overline{x}_{1}))-\theta_{01}-\theta_{11}^{\prime}(x_{i}-\overline{x})$
      does not depend on $i$, since $\Gamma_{21}=-\theta_{11}$. As a result, $\rho_{i}(1)\pi_{i}$ is proportional to $\pi_{i}$.
      It suffices to show that $\frac{1}{n}\sum_{i}\rho_{i}(1)\pi_{i}=1$. Showing this is rather tedious by computing $\tilde{G}_{n}$. Alternatively, we note that $\omega_{i}=\rho_{i}(1)\pi_{i}$ do not depend on individual outcomes. When $\pi_{i}^{*}=\pi_{i}$, the regression estimates $\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))$ by 2.2. If the outcomes are such that the model is well-specified: $y_{i}(w)=vw$, then the regression estimates $\tau=v$ exactly for all $v\in\mathbb{R}$. But for this to happen, we must have
      $\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))=v\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=v.$
      As a result $\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=1$.
  5. (5)
    Let $\pi(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$. By Theorem B.1, the potential weights of this regression are equal to those of the regression
    $\mathbf{Y}_{it}=\ddot{\mathbf{W}}_{it}+\epsilon_{it}$
    where
    $\ddot{\mathbf{w}}=\mathbf{w}-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{w}-\sum_{\mathbf{w}_{1}\in\mathcal{W}}\pi(\mathbf{w}_{1})\mathbf{w}_{1}+1_{T}\frac{1_{T}^{\prime}}{T}\sum_{\mathbf{w}_{1}\in\mathcal{W}}\pi(\mathbf{w}_{1})\mathbf{w}_{1}.$
    For the FWL-transformed regression, the Gram matrix $\tilde{G}_{n}$ is a positive scalar by Theorem B.1.
    1. (a)
      Any implicit design $\tilde{\pi}_{i}$ then satisfies
      $0=\tilde{G}_{n}^{-1}\cdot\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{w}}=\tilde{G}_{n}^{-1}\left\{{\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\mathbf{w}-1_{T}\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\frac{1_{T}^{\prime}\mathbf{w}}{T}-\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\mathbf{w}+1_{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\frac{1_{T}^{\prime}}{T}\mathbf{w}}\right\}$
      Note that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$ solves this equation, and so $\pi(\mathbf{w})$ is an implicit design. We now show its uniqueness.
      The preceding display implies that the vector
      $\sum_{\mathbf{w}\in\mathcal{W}}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}\in\operatorname{span}(1_{T})$
      However, the only vector in $\operatorname{span}(1_{T})$ and $\operatorname{span}(\mathcal{W})$ is the zero vector by assumption. Thus $\sum_{\mathbf{w}\in\mathcal{W}}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}=\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}=0$. Since $\mathcal{W}\setminus\left\{{0}\right\}$ is a linearly independent collection of vectors, we conclude that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$ for all $\mathbf{w}\neq 0$. Since both probability vectors sum to one, we conclude that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$. Therefore, $\pi(\mathbf{w})$ is the unique implicit design.
      For the Gram consistency of $\pi(\cdot)$, observe that the Gram matrix of the original regression is equal to
      $G_{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\mathbf{z}(\mathbf{w})^{\prime}\mathbf{z}(\mathbf{w})=\sum_{\mathbf{w}\in\mathcal{W}}\left\{{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})}\right\}\mathbf{z}(\mathbf{w})^{\prime}\mathbf{z}(\mathbf{w})$
      and thus depends on $\bm{\pi}_{i}^{*}$ only through $\pi(\mathbf{w}).$ Therefore $G_{n}(\bm{\pi})=G_{n}$.
    2. (b)
      Immediately, if $\pi_{i}^{*}$ is equal to the implicit design, then it must be constant for all $i$. Conversely, if $\pi_{i}^{*}$ is constant for all $i$, then the implicit design in (a) is equal to it.
    3. (c)
      For $\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$, define
      $\mathbf{w}_{\pi}=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\mathbf{w}\quad g_{t}(\mathbf{w})=\ddot{\mathbf{w}}_{t}=\mathbf{w}_{t}-\mathbf{w}_{\pi,t}-\frac{1}{T}(1_{T}^{\prime}\mathbf{w}-1_{T}^{\prime}\mathbf{w}_{\pi}).$
      Note that
      $\tilde{G}_{n}=\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g_{t}(\mathbf{w})^{2}.$
      Then by definition,
      $\bm{\omega}_{t}(\bm{\pi},\mathbf{w})=\pi(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=\frac{\pi(\mathbf{w})g_{t}(\mathbf{w})}{\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g_{t}(\mathbf{w})^{2}}.$
      Inspecting Theorem 1(ii) in Athey and Imbens (2022), in their expression, for $\mathbf{w}$ having adoption time $a\in\left\{{2,\ldots,T,\infty}\right\}$, the weight put on the potential outcome $y_{t}(\mathbf{w})$ is equal to
      $\gamma_{t,a}=\frac{\pi(\mathbf{w})g(t,a)}{\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g(t,a(\mathbf{w}))^{2}}$
      for the expression in their (3.1)
      $g(t,a(\mathbf{w}))$ $=\left({\mathbbm{1}(a(\mathbf{w})\leq t)-\sum_{\mathbf{w}\in\mathcal{W}:a(\mathbf{w})\leq t}\pi(\mathbf{w})}\right)$
      $\quad+\frac{1}{T}\left({a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)-\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})}\right)+\frac{T+1}{T}(\mathbbm{1}(a=\infty)-\pi(0_{T})).$
      Here, we let $a(\mathbf{w})$ denote the adoption time of a treatment path $\mathbf{w}$.
      Now, observe that
      $\mathbbm{1}(a(\mathbf{w})\leq t)$ $=\mathbf{w}_{t}$
      $\sum_{\mathbf{w}\in\mathcal{W}:a(\mathbf{w})\leq t}\pi(\mathbf{w})$ $=\sum_{\mathbf{w}\in\mathcal{W}}\mathbf{w}_{t}\pi(\mathbf{w})=\mathbf{w}_{\pi,t}$
      $a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)$ =$=\begin{cases}0&\mathbf{w}=0_{T}\\ \text{number of untreated periods $+1$}=T+1-1_{T}^{\prime}\mathbf{w}&\text{ otherwise}.\end{cases}$
      $\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})$ $=T+1-1_{T}^{\prime}\mathbf{w}_{\pi}-(T+1)\pi(0_{T})$
      Therefore, $\left({a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)-\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})}\right)$
      $=1_{T}^{\prime}\mathbf{w}_{\pi}+\begin{cases}(T+1)(\pi(0_{T})-1)&\mathbf{w}=0_{T}\\ -1_{T}^{\prime}\mathbf{w}+(T+1)\pi(0_{T})&\text{ otherwise}\end{cases}$
      Therefore,
      $g(t,a)=\mathbf{w}_{t}-\mathbf{w}_{\pi,t}-\frac{1}{T}\left({1_{T}^{\prime}\mathbf{w}-1_{T}^{\prime}\mathbf{w}_{\pi}}\right)=g_{t}(\mathbf{w})$
      and these expressions match.∎
Example A.1.
It is simple to analyze Theorem 3.1(2) fully numerically, at least given a concrete data-generating process. To do so, we follow the numerical example in Section 2.2 of Goldsmith-Pinkham et al. (2024). Consider the specification
$Y_{i}=\alpha_{0}+\tau_{1}W_{i1}+\tau_{2}W_{i2}+\gamma^{\prime}x_{i}+\epsilon_{i}$
with binary $x_{i}$ and $J=2$. Suppose exactly half the units have $x_{i}=1$. For the units with $x_{i}=0$, $\pi^{*}_{i}(0)=0.5$, $\pi^{*}_{i}(1)=0.05$, and $\pi^{*}_{i}(2)=0.45$. For the units with $x_{i}=1$, $\pi_{i}^{*}(0)=0.1$, $\pi_{i}^{*}(1)=0.45$, and $\pi_{i}^{*}(2)=0.45$. The coefficients of interest are $\bm{\tau}=[\tau_{1},\tau_{2}]^{\prime}$.
Numerically, we can verify that the assignment probabilities $\bm{\pi}^{*}$ are the only solution to (2.2), and thus $\bm{\pi}^{*}$ is the only implicit design for this regression specification. The corresponding implicit estimand for $\tau_{j}$ is then
$\tau_{j}=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{J}\underbrace{\pi_{i}(k)\rho_{ij}(k)}_{\omega_{ij}(\bm{\pi},k)}y_{i}(k),$
which is the sum of potential outcomes weighted by $\omega_{ij}(\bm{\pi},k)$.
We can calculate the implicit estimand as well, and inspecting the implicit estimand allows us to recover the contamination bias in Goldsmith-Pinkham et al. (2024). For each $x_{i}$ value and each estimand, we tabulate $\omega_{ij}(\cdot)$ for individuals with those $x_{i}$:
$\tau_{1}$ $\tau_{2}$
$\omega_{i1}(0)$ $\omega_{i1}(1)$ $\omega_{i1}(2)$ $\omega_{i2}(0)$ $\omega_{i2}(1)$ $\omega_{i2}(2)$
$x_{i}=0$ $-140/106$ $41/106$ $99/106$ $-160/106$ $9/106$ $151/106$
$x_{i}=1$ $-72/106$ $171/106$ $-99/106$ $-52/106$ $-9/106$ $61/106$
We find that the implicit estimand for $\tau_{j}$ does not solely involve $y_{i}(0)$ and $y_{i}(j)$, contrary to intuition; moreover, $\tau_{j}$ does not measure the same estimand for individuals with $x_{i}=0$ and $x_{i}=1$. This echoes the result by Goldsmith-Pinkham et al. (2024). In fact, under $\bm{\pi}^{*}$, the estimand $\tau_{1}$ is
$\tau_{1}=\frac{1}{2}\left[{\frac{41}{106}\overline{\tau}_{1\mid x=0}+\frac{99}{106}\overline{\tau}_{2\mid x=0}}\right]+\frac{1}{2}\left[{\frac{171}{106}\overline{\tau}_{1\mid x=1}-\frac{99}{106}\overline{\tau}_{2\mid x=1}}\right]$
where $\overline{\tau}_{k\mid x=j}$ is the mean of $y_{i}(k)-y_{i}(0)$ among those with $x_{i}=j$. This decomposition exactly matches the decomposition (7) in Goldsmith-Pinkham et al. (2024). $\blacksquare$
Example A.2.
It is possible for there to be no contamination under a non-constant $\pi_{i}^{*}$. [^38] The discussion in Goldsmith-Pinkham et al. (2024) after their Proposition 1 (p.4028) focuses on the case where $\pi_{i}^{*}$ is constant. This example can thus be thought of as a corner case that is implicitly ruled out by that discussion. Their discussion focuses on when the linear model (in a superpopulation setup) for $\mathbb{E}[W_{ik}\mid W_{i,-k},x_{i}]$ is well specified, which implicitly rules out cases where for some values of $x_{i}$, $W_{i,-k}$ has limited support. Numerically, consider a setup where $x_{i}$ is saturated and takes 3 values, and there are exactly the same number of units with each $x$-value. There are 3 treatments. The generalized propensity scores are as follows:
$w=1$ $w=2$
$x=0$ 00 $1/5$
$x=1$ $1/2$ 00
$x=2$ 00 $0.6$
It is simple to show numerically that, for a specification analogous to Example A.1, the Gram matrix is full rank and that the estimand is not contaminated:
$\tau_{1}$ $\tau_{2}$
$\omega_{i1}(0)$ $\omega_{i1}(1)$ $\omega_{i1}(2)$ $\omega_{i2}(0)$ $\omega_{i2}(1)$ $\omega_{i2}(2)$
$x_{i}=0$ 00 00 00 $-1.2$ 00 $1.2$
$x_{i}=1$ $-3$ $3$ 00 00 00 00
$x_{i}=2$ 00 00 00 $-1.8$ 00 $1.8$
This example is not wholly due to the fact that on each $x$-value, only two treatments have positive probability. One could verify that
$w=1$ $w=2$
$x=0$ $1/6$ $1/6$
$x=1$ $1$ 00
$x=2$ 00 00
similarly leads to uncontaminated estimands. $\blacksquare$
See 3.2
Proof.
By Theorem B.1, we can analyze the regression
$Y_{i}=\tau_{0}(W_{i}-\delta_{0}-\delta_{1}^{\prime}x_{2i})+\tau_{1}^{\prime}\left({W_{i}x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}}\right)+\epsilon_{i}$
instead. Here, $\delta_{0},\delta_{1}$ are population projection coefficients of $\pi_{i}^{*}$ on $1,x_{2i}$ and $\Gamma_{0},\Gamma_{1}$ are population projection coefficients of $\pi_{i}^{*}x_{1i}$ on $1,x_{2i}$. The potential weights for this FWL-transformed regression is
$\rho_{i}(w)=\tilde{G}_{n}^{-1}\begin{bmatrix}w-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ wx_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix},$
where $\tilde{G}_{n}$ is a positive definite matrix by Theorem B.1. Now, the implicit design solves (2.2)
$\tilde{G}_{n}^{-1}\left({\pi_{i}\begin{bmatrix}1-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}+(1-\pi_{i})\begin{bmatrix}-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ -\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}}\right)=0.$
This equation admits a solution if and only if
$\pi_{i}\begin{bmatrix}1-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}+(1-\pi_{i})\begin{bmatrix}-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ -\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}=0,$
which immediately rearranges into
$\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}\quad\pi_{i}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}.$
If the implicit design exists, then $\pi_{i}$ satisfies the above display, meaning that $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$ for all $i$. On the other hand, if $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$, then setting $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$ would solve these equations. Hence an implicit design exists iff $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$.
When this happens, $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$. Applying Theorem 2.1, if $\bm{\tau}$ is a vector of causal contrasts, then $\pi_{i}^{*}=\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$ and $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$, since an implicit design exists. ∎
See 3.3
Proof.
By Theorem B.1, we can focus on the population regression of $\mathbf{y}_{i}$ on $\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x}$. Here,
$\ddot{\mathbf{w}}_{t}=\mathbf{w}_{t}-\frac{1^{\prime}\mathbf{w}}{T}-\frac{1}{n}\sum_{i=1}^{n}\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi_{i}^{*}(\tilde{\mathbf{w}})\tilde{\mathbf{w}}_{t}+\frac{1}{n}\sum_{i=1}^{n}\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi_{i}^{*}(\tilde{\mathbf{w}})\frac{1^{\prime}\tilde{\mathbf{w}}_{t}}{T}$
and
$\ddot{\mathbf{x}}_{it}=\mathbf{x}_{it}-\frac{1^{\prime}\mathbf{x}_{i}}{T}-\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{it}+\frac{1}{n}\sum_{i=1}^{n}\frac{1^{\prime}\mathbf{x}_{i}}{T}.$
The potential weight vector is then
$\bm{\rho}_{i}(\mathbf{w})=\frac{\ddot{\mathbf{w}}-\ddot{\mathbf{x}}_{i}\beta_{w\to x}}{\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{W}_{i}\sim\pi_{i}^{*}(\cdot)}\left[{(\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x})^{\prime}(\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x})}\right]}.$
where the denominator is strictly positive by Theorem B.1.
The condition (2.2) is then
00 0 $=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\rho_{i}(\mathbf{w})\implies\ddot{\mathbf{x}}_{i}\beta_{w\to x}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\ddot{\mathbf{w}}$
The right-hand side is a linear combination of the columns of $\mathcal{W}$ and $1_{T}$. Thus, a necessary condition for the existence of implicit designs is that $\ddot{\mathbf{x}}_{i}\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\}).$ The left-hand side is further more
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}+c1_{T}$
for some scalar $c$. Thus, a necessary condition is furthermore
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\}).\qed$
When we know that $\beta_{w\to x}=0$, then the condition (2.2) is
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\ddot{\mathbf{w}}=0.$
By the same argument as for Theorem 3.1(5)(a), the unique implicit design is $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$.
See 3.4
Proof.
For a given unit $i$ and $t\in\mathcal{T}_{i}$, let $L_{i}=|\mathcal{T}_{i}|$ and let $R_{i}(\mathbf{w})=\sum_{s\in\mathcal{T}_{i}}\mathbf{w}_{s}$ so that $Q_{i}(\mathbf{w})=L_{i}^{-1}R_{i}(\mathbf{w})$. The covariate transform—where we specify the unit fixed effect via within transformation—is
$z_{it}(\mathbf{w})^{\prime}=[\dot{\mathbbm{1}}_{it1},\ldots,\dot{\mathbbm{1}}_{itT},\dot{\mathbf{w}}_{t}]$
where
$\dot{\mathbbm{1}}_{its}=\begin{cases}0&s\not\in\mathcal{T}_{i}\\ 1-\frac{1}{L_{i}}&s=t\\ -\frac{1}{L_{i}}&s\in\mathcal{T}_{i}\setminus\left\{{t}\right\}\end{cases}$
is the demeaned time dummy and
$\dot{\mathbf{w}}_{t}=\mathbf{w}_{t}-Q_{i}(\mathbf{w}).$
By Theorem B.1, it suffices to analyze the potential weights of regressing $Y_{it}$ on $\ddot{W}_{it}$, where
$\ddot{W}_{it}=\dot{W_{it}}-\sum_{s=1}^{T}\delta_{s}\dot{\mathbbm{1}}_{its}=\dot{W_{it}}-\sum_{s\in\mathcal{T}_{i}}\delta_{s}\dot{\mathbbm{1}}_{its}.$
and $\delta_{s}$ are the population projection coefficients of $\dot{W}_{it}$ on $\dot{\mathbbm{1}}_{its}$. Let $V\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{t\in\mathcal{T}_{i}}\mathbb{E}[\ddot{W}_{it}^{2}]$, which is strictly positive by Theorem B.1. The potential weights are then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({\dot{\mathbf{w}}_{t}-\sum_{s\in\mathcal{T}_{i}}\delta_{s}\dot{\mathbbm{1}}_{its}}\right)=V^{-1}\left({\dot{\mathbf{w}}_{t}-\delta_{t}+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}}\right).$
An implicit design $\bm{\pi}$ exists if and only if there is some $\pi_{i}(\cdot)$ and $\sum_{\mathbf{w}}\pi_{i}(\mathbf{w})=1$ such that
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=0\text{ for all $i$ and $t\in\mathcal{T}_{i}$}.$
This is further equivalent to that for all $i$, $t\in\mathcal{T}_{i}$,
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}_{t}-\frac{R_{i}(\mathbf{w})}{L_{i}}}\right)=\delta_{t}-\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}.$
Suppose first that such a $\bm{\pi}$ exists. Fix $t\in\mathcal{T}$ and consider two units $i,j$, where $\mathcal{T}_{i}\cap\mathcal{T}_{j}\supset\mathcal{T}$ by assumption. Then, by (A), we have that
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\mathbf{w}_{t}-\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\frac{R_{i}}{L_{i}}+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{j}(\mathbf{w})\mathbf{w}_{t}-\sum_{\mathbf{w}\in\mathcal{W}}\pi_{j}(\mathbf{w})\frac{R_{j}}{L_{j}}+\frac{1}{L_{j}}\sum_{s\in\mathcal{T}_{j}}\delta_{s}.$
Thus, rearranging, we have for all $t\in\mathcal{T}$,
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi_{i}(\mathbf{w})-\pi_{j}(\mathbf{w}))\mathbf{w}_{t}=M(i,j)$
where $M(i,j)$ does not depend on $t$. Viewed as linear combinations for vectors in $\mathcal{W}_{\mathcal{T}}$, the left-hand side collecting over all $t\in\mathcal{T}$ lies in the span of $\mathcal{W}_{\mathcal{T}}$. The right-hand side lies in the span of the constant vector on $|\mathcal{T}|$ dimensions. By assumption, these two spans contain only the zero vector in common. Thus, both sides are equal to zero:
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi_{i}(\mathbf{w})-\pi_{j}(\mathbf{w}))\mathbf{w}_{t}=0\text{ for all $t\in\mathcal{T}$}.$
By linear independence of vectors in $\mathcal{W}_{\mathcal{T}}$, we conclude that $\pi_{i}(\mathbf{w})=\pi_{j}(\mathbf{w})$. Therefore, the implicit design must be constant across units:
$\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$
for some $\pi(\cdot)$.
Next, we show that if $\bm{\pi}$ exists, then it must equal to $\overline{\pi}(\cdot)$. The fact that $\delta_{t}$ are projection coefficients is equivalent to the following orthogonality conditions holding for all $t\in[T]$: For all $s\in\left\{{1,\ldots,T}\right\}$,
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\sum_{t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\ddot{\mathbf{w}}_{it}\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}-\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{it\ell}}\right)\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}}\right)\dot{\mathbbm{1}}_{its}-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{t\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{it\ell}\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}}\right)\mathbbm{1}(t=s)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{t\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\mathbbm{1}(t=\ell)\dot{\mathbbm{1}}_{its}$ ($\dot{\mathbf{w}}$ and $\dot{\mathbbm{1}}_{i}$ sum to zero over $t\in\mathcal{T}_{i}$)
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-R_{i}(\mathbf{w})/L_{i}}\right)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{i\ell s}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-R_{i}(\mathbf{w})/L_{i}}\right)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}(\mathbbm{1}(\ell=s)-1/L_{i})$
Rearranging, we have that the orthogonality condition is equivalent to
$\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-Q_{i}(\mathbf{w})}\right)=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\left({\delta_{s}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}}\right)$
for all $s\in[T]$.
Pick $t\in\mathcal{T}$ and consider the corresponding (A). Note that $t\in\mathcal{T}_{i}$ for all $i$, and hence
$\delta_{t}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}+\sum_{\mathbf{w}\in\mathcal{W}}\underbrace{\left({\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})}\right)}_{\overline{\pi}(\mathbf{w})}\mathbf{w}_{t}-\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})Q_{i}(\mathbf{w}).$
Given (A), we can plug in
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}.$
Rearranging, we obtain that
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi(\mathbf{w})-\overline{\pi}(\mathbf{w}))\mathbf{w}_{t}=M^{\prime}(i)$
for some $M^{\prime}(i)$ that does not depend $t\in\mathcal{T}$. Again, the left-hand side, over $t\in\mathcal{T}$, is in the span of $\mathcal{W}_{\mathcal{T}}$, whereas the right-hand side is in the span on $1_{\mathcal{T}}$. As a result, both sides are equal to zero by assumption. Since vectors in $\mathcal{W}_{\mathcal{T}}$ are linearly independent, we have that
$\pi(\mathbf{w})=\overline{\pi}(\mathbf{w})\text{ for all $\mathbf{w}\neq 0$}\implies\pi(\mathbf{w})=\overline{\pi}(\mathbf{w})\text{ for all $\mathbf{w}\in\mathcal{W}$}.$
Now, we plug $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ into (A), and then plug the expression into (A). We obtain that for all $t\in[T]$,
$\sum_{i:t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))=\sum_{i:t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w})).$
This is (3.4). Thus, we proved that if an implicit design exists, then the above condition holds and it must be uniquely equal to $\overline{\pi}(\cdot)$.
For the reverse direction, let us assume that (3.4) holds. For a given $t$, set
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\mathbf{w}_{t},$
and thus
$\delta_{t}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}=\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right).$
This means that $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ satisfies (A). Thus $\overline{\pi}(\cdot)$ is an implicit design upon checking that (A) defines the projection coefficients.
Summing over $i$ where $t\in\mathcal{T}_{i}$:
$\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\left\{{\delta_{s}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}}\right\}=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w})),$
where the last equality follows from (3.4). This verifies the orthogonality condition (A) for $t$. Therefore, our choice of (A) does indeed equal the projection coefficients. This concludes the proof. ∎
Appendix B Additional results
B.1. Invariances
Theorem B.1 (Frisch–Waugh–Lovell, in population).
Consider a population regression specification in the sense of 2.2. Partition $z_{t}(\mathbf{x}_{i},\mathbf{w})$ into $z_{t1}(\cdot)$ and $z_{t2}(\cdot)$. Suppose $\Lambda=[\Lambda_{1},0]$ loads solely on entries in $z_{t1}$: $\Lambda z_{t}=\Lambda_{1}z_{t1}$. Let
$\Gamma=\left({\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t1}(\mathbf{x}_{i},\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right)\left({\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right)^{-1}$
be the population projection matrix of $z_{t1}$ onto $z_{t2}$. Define $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=z_{t1}(\mathbf{x}_{i},\mathbf{w})-\Gamma z_{t2}(\mathbf{x}_{i},\mathbf{w})$. Consider the regression specification defined by $\Lambda_{1},\tilde{z}_{t}(\cdot),$ and
$\tilde{G}_{n}=\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})^{\prime}.$
Then:
  1. (i)
    $\Gamma,\tilde{G}_{n}$ are functions of the original Gram matrix $G_{n}$. Since $G_{n}$ is assumed to be invertible, $\Gamma$ is well-defined and $\tilde{G}_{n}$ is positive definite.
  2. (ii)
    The potential weights associated with the two population regression specifications are the same.
Proof.
Note that we can partition
$G_{n}=\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t}(\mathbf{x}_{i},\mathbf{w})z_{t}(\mathbf{x}_{i},\mathbf{w})^{\prime}=\begin{bmatrix}G_{11}&G_{12}\\ G_{21}&G_{22}\end{bmatrix}$
conformably where $G_{11}$ is the Gram matrix associated with $z_{t1}$ and $G_{22}$ is the Gram matrix associated with $z_{t2}$. Since $G_{n}$ is assumed to be invertible and thus positive definite, $G_{11},G_{22}$ are invertible and positive definite as well. This implies that $\Gamma=G_{12}G_{22}^{-1}$ and $\tilde{G}_{n}=G_{11}-\Gamma G_{21}$ are well-defined. This also implies that the Schur complement $\tilde{G}_{n}=G_{11}-\Gamma G_{21}$ is invertible (and positive definite). This proves (i).
By the block matrix inversion formula,
$G_{n}^{-1}=\begin{bmatrix}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}&-(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}G_{12}G_{22}^{-1}\\ -G_{22}^{-1}G_{21}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}&G_{22}^{-1}+G_{22}^{-1}G_{21}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}G_{12}G_{22}^{-1}\end{bmatrix}.$
Hence
$\Lambda G_{n}^{-1}=\left[{\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\quad-\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\Gamma}\right]$
Now, the potential weights for the original regression specification are
$\bm{\rho}_{it}(\mathbf{w})$ $=\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}z_{t1}(\mathbf{x}_{i},\mathbf{w})-\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\Gamma z_{t2}(\mathbf{x}_{i},\mathbf{w})$
$=\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})$
$=\Lambda_{1}\tilde{G}_{n}^{-1}\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w}).$
This proves (ii). ∎
Theorem B.2 (Frisch–Waugh–Lovell, in sample).
Consider a population regression specification in the sense of 2.2. Partition $z_{t}(\mathbf{x}_{i},\mathbf{w})$ into $z_{t1}(\cdot)$ and $z_{t2}(\cdot)$. Suppose $\Lambda=[\Lambda_{1},0]$ loads solely in entries in $z_{t1}$: $\Lambda z_{t}=\Lambda_{1}z_{t1}$. Suppose the sample Gram matrix is invertible. Let
$\hat{\Gamma}=\left({\frac{1}{n}\sum_{i,t}z_{t1}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right)\left({\frac{1}{n}\sum_{i,t}z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right)^{-1}$
be the sample projection matrix of $z_{t1}$ onto $z_{t2}$. Define $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=z_{t1}(\mathbf{x}_{i},\mathbf{w})-\hat{\Gamma}z_{t2}(\mathbf{x}_{i},\mathbf{w})$. Consider the regression specification defined by $\Lambda_{1},\tilde{z}_{t}(\cdot)$. Then the estimated potential weights associated with the two regression specifications are the same.
Proof.
The proof entirely follows from the proof of Theorem B.1 by setting $\pi_{i}^{*}(\mathbf{w})=\mathbbm{1}(\mathbf{w}=\mathbf{W}_{i})$. ∎
Theorem B.3 (Invariance under reparameterization).
Consider a population regression specification in the sense of 2.2. Consider an invertible matrix $M$ and another population specification defined by $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=Mz_{t}(\mathbf{x}_{i},\mathbf{w})$ and $\tilde{\Lambda}=\Lambda M^{\prime}$, so that the two vectors of coefficients represent the same underlying contrasts. Then the potential weights associated with the two regression specifications are the same. The estimated potential weights are also the same.
Proof.
We can easily see that the second regression specification has Gram matrix $\tilde{G}_{n}=MG_{n}M^{\prime}$ (and $\widehat{\tilde{G}_{n}}=M\hat{G}_{n}M^{\prime}$). Thus the potential weights for the second specification are
$\tilde{\bm{\rho}}_{it}(\mathbf{w})=\tilde{\Lambda}(M^{\prime})^{-1}G_{n}^{-1}M^{-1}Mz_{t}(\mathbf{x}_{i},\mathbf{w})=\bm{\rho}_{it}(\mathbf{w}).$
The corresponding equalities similarly hold for the estimated potential weights. ∎
B.2. Additional consistency results
Recall the sequence of populations setup in Section 2.4. Additionally, let $\pi_{1,n}^{*},\ldots,\pi_{n,n}^{*}$ denote a triangular array of marginal treatment assignment probabilities, and let $\Pi_{n}^{*}$ denote the joint distribution of the treatment assignments $\mathbf{W}_{1},\ldots,\mathbf{W}_{n}$ under the $n$th population with marginals equal to $\pi_{1,n}^{*},\ldots,\pi_{n,n}^{*}$.
Lemma B.4.
Suppose that in each $\Pi_{n}^{*}$, treatments are independently assigned according to $\pi_{i}^{*}=\pi_{i,n}^{*}$. Let $G_{i}=\sum_{t=1}^{T}z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}$. Assume that for all $1\leq j\leq K,1\leq\ell\leq K$, the average second moment of $G_{i,j\ell}$ grows slower than $n$: as $n\to\infty$, $\frac{1}{n}\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\pi_{i,n}^{*}}[G_{i,j\ell}^{2}]}\right)\to 0.$ Then Assumption 2.1 holds.
Proof.
Fix a coordinate $(j,k)$. Let $S_{n}=\sum_{i=1}^{n}G_{i,jk}$ and note that $\hat{G}_{n,jk}=\frac{1}{n}S_{n}$. Define $\mu_{n}=\mathbb{E}[S_{n}]$ and $\sigma_{n}^{2}=\operatorname{Var}(S_{n})$. Theorem 2.2.6 in Durrett (2019) states that if $\sigma_{n}^{2}/n^{2}\to 0$ then
[…]
Note that
$\sigma_{n}^{2}\leq\sum_{i=1}^{n}\mathbb{E}[G_{i,jk}^{2}].$
[…]
Proposition B.5.
Suppose that, for some $0<C<\infty$, with probability tending to one, an estimated implicit design $\hat{\bm{\pi}}$ exists and is bounded in the sense that:
$\max_{i}\max_{\mathbf{w}\in\mathcal{W}}|\hat{\pi}_{i}(\mathbf{w})|<C.$
[…]
Proof.
Fix $\eta>0$. Note that
[…]
$+\operatorname{P}\left[{\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\rho}_{i}(\mathbf{w})\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}>\eta,\text{a bounded implicit design exists}}\right]$
Note that if $\hat{\pi}_{i}$ is a bounded implicit design, then
$\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\rho}_{i}(\mathbf{w})\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}=\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}(\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w}))\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}\leq C(J+1)\max_{\mathbf{w}\in\mathcal{W}}|\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w})|_{\infty}.$
Thus,
[…]
$+\operatorname{P}\left[{C(J+1)\max_{\mathbf{w}\in\mathcal{W}}|\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w})|_{\infty}>\eta}\right]\to 0.$
by assumption. The claim of uniform-in-$i$ convergence follows by essentially the same argument, where we take maximum over $i$. ∎
B.2.1. Consistency of $\hat{G}_{n}$ to $G_{n}$ under rejective sampling
This subsection considers a law of large numbers under a version of sampling with replacement (what Hájek, 1964, calls “rejective sampling”), and largely follows Rambachan and Roth (2020). In particular, we assume that the treatment is binary, and there is a sequence of unconditional probabilities $p_{1},p_{2},\ldots$ as well as a sequence of sample sizes $N_{n}$ such that $\Pi^{*}_{n}$ describes the joint distribution of $W_{1},\ldots,W_{n}$ conditional on the event $\sum_{i=1}^{n}W_{i}=N_{n}$, where $W_{i}\sim\operatorname{Bern}(p_{i})$ independently unconditionally. Correspondingly, let $\pi_{i,n}$ be the probability that $W_{i}=1$ under $\Pi^{*}_{n}$.
Lemma B.6.
Write $\pi_{i}(w)=\pi_{i,n}(w)$ and omit the $n$ subscript. Consider $\overline{y}_{w}=\frac{1}{n}\sum_{i}\mathbbm{1}(W_{i}=w)y_{i}$ where $\mathbb{E}[\overline{y}_{w}]=\frac{1}{n}\sum_{i}\pi_{i}(w)y_{i}(w)$. Suppose $C_{n}\equiv\sum_{i=1}^{n}\pi_{i}(w)(1-\pi_{i}(w))\to\infty$ as $n\to\infty$. Then, as $n\to\infty$
$\operatorname{Var}(\overline{y}_{w})=\frac{(1+o(1))}{n}\frac{C_{n}}{n}\sum_{i=1}^{n}\frac{\pi_{i}(w)(1-\pi_{i}(w))}{C_{n}}\left({y_{i}(w)-\frac{\sum_{j=1}^{n}\pi_{j}(w)(1-\pi_{j}(w))y_{j}(w)}{C_{n}}}\right)^{2}$
Proof.
This is a restatement of Theorem 6.1 in Hájek (1964). The notation $y_{i}$ in the theorem corresponds to $\pi_{i}(w)y_{i}(w)$ in our notation. ∎
A sufficient condition for the variance to tend to zero is the following:
Assumption B.1.
For all entries $k$, $z_{k}(x_{i},w)$ is uniformly bounded by $0<M<\infty$ and $C_{n}(w)\equiv\sum_{i=1}^{n}\pi_{i}(w)(1-\pi_{i}(w))\to\infty$ as $n\to\infty$.
The boundedness condition for $z_{k}$ is stronger than needed. In particular, what is needed is that the $\pi_{i}(w)(1-\pi_{i}(w))$-weighted variance of $y_{i}(w)$ is $O(1)$.
Lemma B.7.
[…]
Proof.
We prove this claim by showing all entries converge in probability, since there are finitely many entries. Fix some $j,k$. Note that
$\hat{G}_{n,jk}=\sum_{w\in\left\{{0,1}\right\}}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(W_{i}=w)z_{j}(x_{i},w)z_{k}(x_{i},w)}_{\overline{z}_{n}(w)}.$
By Lemma B.6 and Assumption B.1 applied to $\overline{z}_{n}$ where $y_{i}(w)$ is taken to be $z_{j}(x_{i},w)z_{k}(x_{i},w)$, we have that $\operatorname{Var}(\overline{z}_{n}(w))\to 0$ as $n\to\infty$, since $C_{n}(w)/n\leq 1$. Since $\mathbb{E}[\overline{z}_{n}(w)]=\frac{1}{n}\sum_{i}\pi_{i}(w)z_{j}(x_{i},w)z_{k}(x_{i},w)$, we have that
[…]
by Chebyshev’s inequality. ∎
B.3. Uniqueness of implicit designs in cross-sections
Consider a cross-sectional setup with $J+1$ treatments where the potential weights are
$\rho_{i}(w)=\tilde{\Lambda}\tilde{G}_{n}^{-1}\tilde{z}(x_{i},w)$
and $\tilde{\Lambda}$ is a $J\times K$ matrix with rank $J$. It is possible to reparametrize the regressors (i.e. choose an invertible $M$ such that $z(\cdot)=M\tilde{z}(\cdot)$) such that
$\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)$
and $\Lambda=[I_{J},0]$. Without essential loss of generality, let us assume the sequence of reparametrized specifications satisfy Assumption 2.2.
Partition $z$ into $z_{1}$ and $z_{2}$ where $z_{1}$ is $J$-dimensional. By Theorem B.1, the potential weights are further equivalent to
$\rho_{i}(w)=G_{n,1|2}^{-1}z_{1|2}(x_{i},w)$
where $z_{1|2}(x_{i},w)=z_{1}(x_{i},w)-\Gamma^{\prime}z_{2}(x_{i},w)$ for population projection coefficients $\Gamma$ and
$G_{n,1|2}=\frac{1}{n}\sum_{i,w}\pi_{i}^{*}(w)z_{1|2}(x_{i},w)z_{1|2}(x_{i},w)^{\prime}.$
Assumption B.2.
The residualized covariate transform is non-singular in the sense that the $J\times(J+1)$ matrix
$Z_{i}\equiv[z_{1|2}(x_{i},0),\ldots,z_{1|2}(x_{i},J)]$
whose columns are covariate transforms evaluated at a treatment level has minimum singular value (that is, the $J$th singular value) greater than some $\eta>0$, uniformly for all $i$.
Assumption B.3.
The maximum eigenvalue of $G_{n}$ is bounded above by $M<\infty$ for all $n$.
Lemma B.8.
Under Assumptions B.2, 2.2, and B.3, condition (ii) in Proposition 2.4 is satisfied for some lower bound $\epsilon>0$ on the minimum (i.e. $J$th) singular value.
Note that since the potential weight matrix $G_{n,1|2}^{-1}Z_{i}$ is $J\times(J+1)$ and has $J$ positive singular values bounded below, if there exists an implicit design, then it must be unique.
Proof.
We first show that Assumption 2.2 implies that the minimum eigenvalue of $G_{n,1|2}$ is bounded away from zero and maximum bounded by $M$. Note that $G_{n,1|2}$ is a Schur complement of a submatrix of $G_{n}$, and so $G_{n,1|2}^{-1}$ is a principal submatrix of $G_{n}^{-1}$. The eigenvalues of $G_{n,1|2}^{-1}$ interlace the eigenvalues of $G_{n}^{-1}$ by Cauchy’s interlace theorem (Hwang, 2004). In particular, the spectrum of $G_{n,1|2}^{-1}$ is included in the range of the spectrum of $G_{n}^{-1}$. Hence $\lambda_{\min}(G_{n,1|2})>\epsilon$ by Assumption 2.2 and $\lambda_{\max}(G_{n,1|2})<M$ by Assumption B.3.
Next, it suffices to show that $G_{n,1|2}^{-1}Z_{i}$ has minimum singular value bounded below. Write the compact SVD as
$Z_{i}=U\Sigma V^{\prime}$
for diagonal $J\times J$ matrix $\Sigma$ with $UU^{\prime}=U^{\prime}U=V^{\prime}V=I_{J}$. Similarly, write
$G_{n,1|2}^{-1}=QDQ^{\prime}$
for an orthogonal matrix $Q$ and diagonal $D$. Then
$G_{n,1|2}^{-1}Z_{i}Z_{i}^{\prime}G_{n,1|2}^{-1}=QDQ^{\prime}U\Sigma^{2}U^{\prime}QDQ^{\prime}.$
is a real symmetric matrix. The spectrum of this matrix is the same as the spectrum of
$D\underbrace{Q^{\prime}U\Sigma^{2}U^{\prime}Q}_{\text{positive definite matrixwith spectrum $\Sigma^{2}$}}D.$
Since $\lambda_{\min}(AB)\geq\lambda_{\min}(A)\lambda_{\min}(B)$ for two positive definite matrices $A,B$, [^39] To see this, note that the operator norm (largest eigenvalue of a positive semidefinite matrix) is submultiplicative $\lVert AB\rVert_{op}\leq\lVert A\rVert_{op}\lVert B\rVert_{op}$. Apply this inequality to $A^{-1}B^{-1}$. we have that the minimum eigenvalue of $G_{n,1|2}^{-1}Z_{i}Z_{i}^{\prime}G_{n,1|2}^{-1}$ is bounded below by $\frac{\eta^{2}}{\lambda_{\max}(G_{n,1|2})^{2}}$. Thus, the minimum singular value of $G_{n,1|2}^{-1}Z_{i}$ is bounded below by $\eta/M$. ∎
B.4. Additional results for interacted regression
Proposition B.9 provides a unified analyses of the regression (B.9). It implies our results in Theorem 3.1(4) when $\overline{x}_{t}=\overline{x}_{1}$ or $\overline{x}_{0}$, but allows for evaluating on other points in the line connecting $\overline{x}_{1}$ and $\overline{x}_{0}$. The implicit design for the model-based ATE is analyzed when we set $t=\alpha_{0}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}$.
Inspecting the resulting implicit designs (B.9), we find that they are in general fractional linear, encompassing the linear-odds result of $t=1$ as a special case. Moreover, a necessary implication of $\pi_{i}^{*}=\pi_{i}$ is the condition (B.9), which is very difficult to satisfy unless $t=0$ or $t=1$. In this sense, the model-based ATU and the model-based ATT are the only estimands whose implicit design is reasonable. See B.1 for an example where the implicit model for the ATE depends on the support of $x_{i}$.
Proposition B.9.
Consider the specification indexed by $t\in\mathbb{R}$,
$Y_{i}=\gamma_{0}+W_{i}\tau_{t}+\gamma_{0}+\gamma_{1}^{\prime}x_{i}+W_{i}(x_{i}-\overline{x}_{t})^{\prime}\gamma_{2}+\epsilon_{i}$
where $\overline{x}_{t}=t\overline{x}_{1}+(1-t)\overline{x}_{0}$ for $\overline{x}_{1}=\frac{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}}$ and $\overline{x}_{0}=\frac{\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})x_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})}$.
Let
$A$ $=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x})(x_{i}-\overline{x})^{\prime}$
$V_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}\quad U_{t}=\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}$
$\alpha_{0}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}$
$\Gamma_{0t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})$
$\alpha_{1}$ $=A^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x})=A^{-1}(\Gamma_{0t}+\alpha_{0}(\overline{x}_{t}-\overline{x}))$
$\Gamma_{1t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x})^{\prime}=\left({V_{t}+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}}\right)A^{-1}$
$\Gamma_{2t}$ $=\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[{W_{i}(x-\overline{x}_{t})\left({W_{i}(x-\overline{x}_{t})-\Gamma_{0t}-\Gamma_{1t}(x_{i}-\overline{x})}\right)^{\prime}}\right]}\right)^{-1}$
×$\quad\quad\times\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[{(W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x}))\left({W_{i}(x_{i}-\overline{x}_{t})-\Gamma_{0t}-\Gamma_{1t}(x-\overline{x})}\right)}\right]$
$=\left({V_{t}-\Gamma_{0t}\Gamma_{0t}^{\prime}-V_{t}\Gamma_{1t}^{\prime}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}}\right)^{-1}\left({\Gamma_{0t}-\alpha_{0}\Gamma_{0t}-V_{t}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}}\right)$
Then, the potential weights satisfy
$\rho_{i}(w)=\tilde{G}_{n}^{-1}\left({w(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))-\underbrace{(\alpha_{0}-\Gamma_{2t}^{\prime}\Gamma_{0t})}_{\theta_{0t}}-\underbrace{(\alpha_{1}-\Gamma_{1t}^{\prime}\Gamma_{2t})^{\prime}}_{\theta_{1t}^{\prime}}(x_{i}-\overline{x})}\right),$
for some scalar $\tilde{G}_{n}>0$. The implicit design $\pi_{i}$ satisfies
$\left({1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t})}\right)\pi_{i}=\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x}).$
Moreover, if $\pi_{i}^{*}=\pi_{i}$, then a necessary condition is that
$tV_{1}(\theta_{1t}+\Gamma_{2t})=(1-t)U_{0}\theta_{1t}.$
Proof.
Observe that by the Frisch–Waugh–Lovell transform, the coefficient $\tau_{t}$ is the same as in the regression
$Y_{i}=\tau_{t}\left({W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{2}^{\prime}\left({W_{i}(x_{i}-\overline{x}_{t})-\Gamma_{0}-\Gamma_{1}(x_{i}-\overline{x})}\right)}\right)+\epsilon_{i}.$
The quantities $\alpha_{0},\alpha_{1},\Gamma_{0},\Gamma_{1},\Gamma_{2}$ are defined exactly as these auxiliary regression coefficients. Thus, by Theorem B.1, the potential weights are equivalent to the potential weights of (B.4), which rearrange exactly to the potential weights displayed. The constraint on the implicit design corresponds exactly to
$\rho_{i}(1)\pi_{i}+(1-\pi_{i})\rho_{i}(0)=0.$
For the “moreover” part, observe that we can rewrite $\Gamma_{2t}$ as satisfying
$V_{t}\Gamma_{2t}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}\Gamma_{2t}-\Gamma_{0t}\Gamma_{0t}^{\prime}\Gamma_{2t}-V_{t}\Gamma_{1t}^{\prime}\Gamma_{2t}=\Gamma_{0t}-\alpha_{0}\Gamma_{0t}-V_{t}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}.$
Thus,
00 0 $=\left({\alpha_{0}\Gamma_{0t}-\Gamma_{0t}\Gamma_{0t}^{\prime}\Gamma_{2t}}\right)+\left({V_{t}\alpha_{1}-V_{t}\Gamma_{1t}^{\prime}\Gamma_{2t}}\right)$
$\quad\quad+\left({\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}\Gamma_{2t}}\right)+V_{t}\Gamma_{2t}-\Gamma_{0t}$
$=\Gamma_{0t}\theta_{0t}+\underbrace{V_{t}\theta_{1t}+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\theta_{1t}}_{\Gamma_{1t}A\theta_{1t}}+V_{t}\Gamma_{2t}-\Gamma_{0t}$
Observe that
$A\theta_{1t}=\Gamma_{0t}+\alpha_{0}(\overline{x}_{t}-\overline{x})-V_{t}\Gamma_{2t}-(\overline{x}_{t}-\overline{x})\Gamma_{0t}^{\prime}\Gamma_{2t}$
and hence
$V_{t}\Gamma_{2t}-\Gamma_{0t}=-A\theta_{1t}+\alpha_{0}(\overline{x}_{t}-\overline{x})-(\overline{x}_{t}-\overline{x})\Gamma_{0t}^{\prime}\Gamma_{2t}=-A\theta_{1t}+\theta_{0t}(\overline{x}_{t}-\overline{x})$
Therefore
00 0 $=\left({\Gamma_{0t}+(\overline{x}_{t}-\overline{x})}\right)\theta_{0t}+(V_{t}-A+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime})\theta_{1t}$
$=\frac{1}{n}\sum_{i=1}^{n}\{(\pi_{i}^{*}(x_{i}-\overline{x}_{t})-(x_{i}-\overline{x}_{t}))\theta_{0t}$
$\quad\quad+\left({\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}-(x_{i}-\overline{x})(x_{i}-\overline{x})^{\prime}+\pi_{i}^{*}(x_{i}-\overline{x}_{t})(\overline{x}_{t}-\overline{x})^{\prime}}\right)\theta_{1t}\}$
$=-\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x}))$
Thus, if $\pi_{i}^{*}=\pi_{i}$,
$\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x})=(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))\pi_{i}^{*}.$
We thus have that
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))$
$=t\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})(1-\pi_{i}^{*})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))+(1-t)\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{1})\pi_{i}^{*}(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))$
where we may plug in
$\pi_{i}^{*}(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))=\theta_{0t}+\theta_{1t}^{\prime}(x-\overline{x})\quad(1-\pi_{i}^{*})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))=1-\theta_{0t}-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t})-\theta_{1t}^{\prime}(x_{i}-\overline{x}).$
Simplifying by noting that $\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})=0=\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})$ yields the condition
$tV_{1}(\theta_{1t}+\Gamma_{2t})=(1-t)U_{0}\theta_{1t}.$
Proposition B.10.
In Proposition B.9, for $t=1$, the implicit design is equal to
$\pi_{i}=\frac{\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x})}{1+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x}_{1})}.$
The associated odds is
$\frac{\pi_{i}}{1-\pi_{i}}=\frac{\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x})}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}\equiv\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x}_{0}).$
Here,
$\delta_{1}$ $=\left({\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})\frac{\pi_{i}^{*}}{1-\pi_{i}^{*}}$
$\delta_{0}$ $=\frac{\alpha_{0}}{1-\alpha_{0}}=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-\pi_{i}^{*})}{1-\alpha_{0}}\frac{\pi_{i}^{*}}{1-\pi_{i}^{*}}$
Proof.
Here, we omit the $t$ subscript for objects defined in Proposition B.9. Observe that $\Gamma_{0}=0$. Hence
$\Gamma_{2}=-(I-\Gamma_{1}^{\prime})^{-1}\alpha_{1}\text{ and }\theta_{0}=\alpha_{0}$
and thus
$\theta_{1}=\alpha_{1}-\Gamma_{1}^{\prime}\Gamma_{2}=(I+\Gamma_{1}^{\prime}(I-\Gamma^{\prime})^{-1})\alpha_{1}=(I-\Gamma^{\prime})^{-1}\alpha_{1}.$
It remains to show that $\delta_{0},\delta_{1}$ take the form stated. First, observe that
$\theta_{1}=(I-\Gamma_{1}^{\prime})^{-1}\alpha_{1}=(A-V_{1})^{-1}\alpha_{0}(\overline{x}_{1}-\overline{x})=(A-V_{1})^{-1}\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0}).$
by plugging in $\alpha_{1}=A^{-1}(\alpha_{0}(\overline{x}_{1}-\overline{x}))$. Here,
$A$ $=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x})(x_{i}-\overline{x}_{0})^{\prime}$
$=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}-(\overline{x}-\overline{x}_{0})(\overline{x}-\overline{x}_{0})^{\prime}$
$=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}-\alpha_{0}^{2}(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
$V_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})(x_{i}-\overline{x}_{0})$
$=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}+(\overline{x}_{0}-\overline{x}_{1})\cdot\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{0})^{\prime}$
$=V_{0}+\alpha_{0}(\overline{x}_{0}-\overline{x}_{1})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
Therefore,
$A-V_{1}=U_{0}+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
By Sherman–Morrison,
$(A-V_{1})^{-1}=U_{0}^{-1}-\alpha_{0}(1-\alpha_{0})\frac{U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
Then,
$\frac{\theta_{1}}{\alpha_{0}(1-\alpha_{0})}$ $=(A-V_{1})^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$=\left({1-\frac{\alpha_{0}(1-\alpha_{0})\cdot(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}}\right)U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$=\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$\implies\theta_{1}^{\prime}(\overline{x}_{1}-\overline{x}_{0})$ $=\frac{\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
and
$\theta_{1}^{\prime}(\overline{x}-\overline{x}_{1})$ $=-(1-\alpha_{0})\theta_{1}^{\prime}(\overline{x}_{1}-\overline{x}_{0})$
$=-(1-\alpha_{0})\frac{\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
Hence,
$1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})=(1-\alpha_{0})\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})},$
and therefore
$\delta_{1}=\frac{\theta_{1}}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}=U_{0}^{-1}\alpha_{0}\left({\overline{x}_{1}-\overline{x}_{0}}\right)=U_{0}^{-1}\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i})\frac{\pi_{i}}{1-\pi_{i}}(x_{i}-\overline{x}_{0}),$
as claimed.
Now,
$\delta_{0}$ $=\frac{\alpha_{0}+\theta_{1}^{\prime}(\overline{x}_{0}-\overline{x})}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}$
$=\frac{\alpha_{0}-\theta_{1}^{\prime}\alpha_{0}(\overline{x}_{1}-\overline{x}_{0})}{(1-\alpha_{0})\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}}$
$=\frac{\alpha_{0}}{1-\alpha_{0}},$
as claimed. ∎
Remark B.1.
Here we consider the ramifications of (B.9) for the model-based ATE ($\tau_{\alpha_{0}}$). For some true design $\bm{\pi}^{*}$ to justify causal interpretation, we must have that this design is fractional linear:
$\pi_{i}^{*}=\frac{\theta_{0}+\theta_{1}^{\prime}(x_{i}-\overline{x})}{1-\Gamma_{2}^{\prime}(x_{i}-\overline{x})}$
for some $\theta_{0},\theta_{1},\Gamma_{2}$. Moreover, the choices $\theta_{0},\theta_{1},\Gamma_{2}$ needs to satisfy (B.9), which depends on the covariates $x_{1},\ldots,x_{n}$. We may thus consider the parameter space as
$\Theta(t;x_{1:n})=\left\{{(\theta_{0},\theta_{1},\Gamma_{2}):\text{ $\tau_{t}$ is a causal contrast under the corresponding $\bm{\pi}^{*}$ under $(\theta_{0},\theta_{1},\Gamma_{2})$ }}\right\}.$
Note that for $t=0$ and $t=1$, this parameter space does not depend on $x_{1:n}$ aside from the obvious support restriction. For instance, if $x_{1:n}$ and $\tilde{x}_{1:n}$ are two sets of covariate values, and
$\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x}_{1})\geq 0\quad\tilde{\pi}_{i}^{*}/(1-\tilde{\pi}_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}}_{1})\geq 0,$
then the corresponding parameters are in $\Theta(1;x_{1:n})\cap\Theta(1;\tilde{x}_{1:n})$—similarly for $t=0$. This is true by Proposition B.10.
However, for other $t$, and in particular for $t=\alpha_{0}$, it is possible that some $(\theta_{0},\theta_{1},\Gamma_{2})\in\Theta(\alpha_{0};x_{1:n})$, still does not belong to $\Theta(\alpha_{0};\tilde{x}_{1:n})$, even though no support restriction is violated
$\tilde{\pi}_{i}^{*}\equiv\frac{\theta_{0}+\theta_{1}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}})}{1-\Gamma_{2}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}})}\in[0,1]\text{ for all $i$}.$
It is not very difficult to come up with a numerical example. See LABEL:listing:ex. $\blacksquare$
Listing 1: Example for B.1
1import numpy as np
2
3n = 100
4x = (np.arange(n) / n) ** 3
5
6lin_form = 1 + (x - 1 / 2) * 0.25
7
8# Search for a valid pi* from this value
9pi_init = lin_form / (1 + 2 * lin_form)
10
11# Regression is y ~ 1 + w + x + wx
12z0 = np.c_[np.ones(n), np.zeros(n), x, np.zeros(n)]
13z1 = np.c_[np.ones(n), np.ones(n), x, x]
14
15gram = ((1 - pi_init[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
16 (pi_init[:, None, None]) * z1[:, :, None] * z1[:, None, :]
17).mean(0)
18lmbda = np.array([0, 1, 0, x.mean()]) # for the model-based ATE
19rho0 = lmbda @ np.linalg.inv(gram) @ z0.T
20rho1 = lmbda @ np.linalg.inv(gram) @ z1.T
21pi = -rho0 / (rho1 - rho0)
22
23# Iterate to find a fixed point
24for _ in range(1000):
25 new_gram = ((1 - pi[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
26 (pi[:, None, None]) * z1[:, :, None] * z1[:, None, :]
27 ).mean(0)
28 new_rho0 = lmbda @ np.linalg.inv(new_gram) @ z0.T
29 new_rho1 = lmbda @ np.linalg.inv(new_gram) @ z1.T
30 new_pi = -new_rho0 / (new_rho1 - new_rho0)
31 pi = new_pi
32
33# Decompose to find the fractional linear parameters
34A = ((x - x.mean()) ** 2).mean()
35v = (pi * (x - x.mean()) ** 2).mean()
36gamma1 = v / A
37gamma0 = (pi * (x - x.mean())).mean()
38a0 = pi.mean()
39a1 = gamma0 / A
40gamma2 = 1 / (v - v * gamma1 - gamma0 * gamma0) * (1 - a0 - v * a1 / gamma0) * gamma0
41theta0 = a0 - gamma2 * gamma0
42theta1 = a1 - gamma1 * gamma2
43
44print(theta0, theta1, gamma2) # 0.3254446337090774 0.014146176384094272 0.044263163071495276
45
46# pi $\in$ [0,1]
47assert (pi > 0).all() and (pi < 1).all()
48
49# the fractional linear representation is correct
50assert (
51 np.abs(pi - (theta0 + theta1 * (x - x.mean())) / (1 - gamma2 * (x - x.mean()))) < 1e-11
52).all()
53
54# Now, consider a new set of x’s, will show that the parameters theta0, theta1, gamma2 no
55# longer form a design justifying level irrelevance for the corresponding estimand
56new_x = np.arange(n) / n
57new_z0 = np.c_[np.ones(n), np.zeros(n), new_x, np.zeros(n)]
58new_z1 = np.c_[np.ones(n), np.ones(n), new_x, new_x]
59new_design = (theta0 + theta1 * (new_x - new_x.mean())) / (1 - gamma2 * (new_x - new_x.mean()))
60
[…]
62assert (new_design > 0).all() and (new_design < 1).all()
63
64gram_new_design = ((1 - new_design[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
65 new_design[:, None, None] * z1[:, :, None] * z1[:, None, :]
66).mean(0)
67new_lmbda = np.array([0, 1, 0, new_x.mean()])
68new_rho0 = new_lmbda @ np.linalg.inv(gram_new_design) @ new_z0.T
69new_rho1 = new_lmbda @ np.linalg.inv(gram_new_design) @ new_z1.T
70
71# Failure of level irrelevance
72print(
73 np.abs(new_design * new_rho1 + (1 - new_design) * new_rho0).max()
74) # Equal to 0.08293240507391064, not zero
Proposition B.11.
Let $n\geq 3$. Consider binary $W_{i}$, covariates $x_{i}\in\mathbb{R}^{d}$, and scalar outcomes. Suppose the true design is linear $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{i}$ for some $\delta_{0}\in\mathbb{R},\delta_{1}\in\mathbb{R}^{d}$. There is no regression $(\Lambda,z(x,w))$—where $\Lambda$ may [^40] This is to accommodate for estimands like the model-based ATT, where we may consider contrasts that depend on $\overline{x}_{1}=\sum_{i}\pi_{i}^{*}x_{i}/\sum_{i}\pi_{i}^{*}$ depend on $x_{1:n},\pi^{*}_{1:n}$—such that:
  1. (1)
    (Regression is linear in covariates) For all $m$ and all $w$, the $m$th entry of $z(x_{i},w)$ is of the form $a_{m}(w)+b_{m}(w)^{\prime}x_{i}$ for some fixed conformable $a_{m}(\cdot),b_{m}(\cdot)$.
  2. (2)
    ($\Lambda\beta$ is the ATE) The corresponding estimand $\Lambda\beta\in\mathbb{R}$ is equal to the ATE, regardless of the configuration of $d,x_{1:n},\delta_{0},\delta_{1}$ (such that $\pi_{i}^{*}\in[0,1]$ for all $i$)
Proof.
It suffices to show that for certain given choices of $x_{1:n},\delta_{0},\delta_{1}$, there is no regression that satisfies (1) and estimates the ATE.
Suppose $x_{i}\in[-1,1]$ is scalar $(d=1)$ and suppose $\left\{{x_{1},\ldots,x_{n}}\right\}$ includes $\left\{{-1,0,1}\right\}$. Set $\delta_{0}=1/2,\delta_{1}=1/4$ so that $\pi_{i}^{*}\in(0,1)$ for all $i$.
Fix some covariate transform $z(x_{i},w)$. By assumption, its $m$th entry is of the form $a_{mw}+b_{mw}x_{i}$. Suppose this regression estimates the ATE. In particular, it estimates a causal contrast. Thus
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0.$
Because the estimand is the ATE, we also have that $\omega_{i}=\rho_{i}(1)\pi_{i}^{*}=1$. Hence, by (B.4), we have that for all $i$
$1+\left({1-\frac{1}{\rho_{i}(1)}}\right)\rho_{i}(0)=0\iff\rho_{i}(1)\rho_{i}(0)=\rho_{i}(1)+\rho_{i}(0)$
Now, the potential weights are of the form
$\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)=c_{0w}+c_{1w}x_{i}$
where the coefficients $c_{0w},c_{1w}$ may depend on $x_{1:n},\pi_{1:n}^{*}$, but are the same for all $i$. As a result, for the regression to estimate the ATE, we must have
$(c_{00}+c_{10}x)(c_{01}+c_{11}x)=(c_{00}+c_{10}x)+(c_{01}+c_{11}x)$
for $x=-1,0,1$: That is,
$c_{00}c_{01}=c_{00}+c_{01}$ ($x=0$)
$c_{00}c_{01}+c_{10}c_{11}+c_{10}c_{01}+c_{00}c_{11}=c_{00}+c_{01}+c_{10}+c_{11}$ ($x=1$)
$c_{00}c_{01}+c_{10}c_{11}-c_{10}c_{01}-c_{00}c_{11}=c_{00}+c_{01}-c_{10}-c_{11}$ ($x=-1$)
Now, adding the last two equations yield
$c_{00}c_{01}+c_{10}c_{11}=c_{00}+c_{01}=c_{00}c_{01}\implies c_{10}c_{11}=0\implies\text{ one of $c_{10},c_{11}$ is zero}$
Suppose $c_{10}=0$. Then the bottom two equations simplify to
$c_{00}c_{11}=c_{11}.$
This implies that $c_{00}=1$ or $c_{11}=0$. $c_{00}=1$ contradicts the $x=0$ equation. Thus, we must have $c_{11}=0$. Suppose instead we assumed $c_{11}=0$, by a symmetric argument we would derive $c_{10}=0$. Thus, $c_{10}=c_{11}=0$.
However, in that case, $\rho_{i}(w)=c_{0w}$ does not depend on $x$. For (B.4) to hold (for $x=-1,0,1$) then, we must have $c_{0w}=0$, since $\pi_{i}^{*}$ does depend on $x$. This means $\rho_{1}(w)=0$ regardless of $x$, but this contradicts the fact that we estimate the ATE ($\rho_{1}(1)\pi_{i}^{*}=1$). Thus, this regression does not estimate the ATE. ∎
B.5. Additional results for panel specifications
Proposition B.12 (One-way fixed effects).
Consider the regression $Y_{it}=\alpha_{i}+\tau W_{it}+\epsilon_{it}$. If the span of $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$ excludes $1_{T}$, then the set of implicit designs is empty.
Proof.
Let $\dot{\mathbf{w}}=\mathbf{w}-1_{T}\frac{1_{T}^{\prime}\mathbf{w}}{T}$. Then, by Theorem B.1, the potential weights are equal to
$\rho_{i}(w)=\frac{\dot{\mathbf{w}}}{\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\dot{\mathbf{w}}^{\prime}\dot{\mathbf{w}}}.$
Thus, if $\pi_{i}$ corresponds to some implicit design $\bm{\pi}$,
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}-1_{T}\frac{1_{T}^{\prime}\mathbf{w}}{T}}\right)=0.$
In particular, this implies that $1_{T}$ is in the linear span of $\mathcal{W}$. By assumption, this is not the case. Therefore the set of implicit designs must be empty. ∎
Intuitively, this is because one-way fixed effects fail to account for the case where the treatment path correlates with underlying potential outcomes. If there are no treatment effects whatsoever, but the baseline potential outcome correlates with the treatment path (e.g., later potential outcomes tend to be larger in staggered adoption), then one-way fixed effects would in general estimate a nonzero coefficient, violating level irrelevance.
These results contrast with the results in Arkhangelsky and Imbens (2023), where the one-way fixed effect regression admits a design-based interpretation under within-unit random assignment of $W_{it}$. To reconcile, Arkhangelsky and Imbens (2023) consider a sampling-based setup and the unconfoundedness restriction
$W_{it}\indep(Y_{it}(0),Y_{it}(1))\mid\frac{1}{T}\sum_{s=1}^{T}W_{is},$
which is their (2.2) in our notation. Under staggered adoption, for instance, the associated propensity score $\operatorname{P}\left({W_{it}=1\mid\frac{1}{T}\sum_{s=1}^{T}W_{is}}\right)$ is always degenerate, since $\frac{1}{T}\sum_{s=1}^{T}W_{is}$ perfectly distinguishes which treatment path unit $i$ is assigned. Conversely, if $W_{it}$ is randomly assigned within a unit and that permuting the time index results in valid counterfactual assignments, then $\mathcal{W}$ is large enough to contain $1_{T}$.
Proposition B.13 (Event-study and more complex TWFE).
Consider the regression $Y_{it}=\alpha_{i}+\gamma_{t}+\beta^{\prime}f_{t}(\mathbf{W}_{i})$ with target estimand $\bm{\tau}=\Lambda\beta$ under some design $\bm{\pi}^{*}$. We have:
  1. (i)
    Whether or not $\bm{\tau}$ satisfies level irrelevance under the true design $\bm{\pi}^{*}$, one proper and Gram-consistent implicit design is
    $\pi_{i}(\mathbf{w})\equiv\pi(\mathbf{w})\equiv\frac{1}{n}\sum_{j=1}^{n}\pi_{j}^{*}(\mathbf{w}).$
    Correspondingly, one estimated implicit design sets $\hat{\pi}_{i}(\mathbf{w})$ to be the empirical frequency of treatment path $\mathbf{w}$ for all $i$.
  2. (ii)
    Suppose $\Lambda=I$. If any column $k$ of $f_{t}(\mathbf{W}_{i})$ is such that the multiset
    $\left({\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}:\mathbf{w}\in\mathcal{W}}\right)$
    (a) contains the zero vector at most once, (b) has nonzero vectors that are linearly independent, and (c) does not span the subspace $\left\{{c1_{T}:c\in\mathbb{R}}\right\}$, then $\overline{\pi}$ is also the unique implicit design.
Proof.
  1. (i)
    By Theorem B.1, it suffices to compute potential weights for the two-way residualized specification. That is, let
    $\mathbf{z}_{i}(\mathbf{w})=\begin{bmatrix}f_{1}(\mathbf{w})^{\prime}\\ \vdots\\ f_{T}(\mathbf{w})^{\prime}\end{bmatrix}\equiv\mathbf{z}(\mathbf{w})$
    be the covariate transform. Note that the population residual of projecting $\mathbf{z}_{i}(\mathbf{W}_{i})$ on unit and time fixed effect is
    $\ddot{\mathbf{z}}(\mathbf{w})$ $=\mathbf{z}(\mathbf{w})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w})-\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+1_{T}\frac{1_{T}^{\prime}}{T}\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})$
    $=\mathbf{z}(\mathbf{w})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w})-\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+1_{T}\frac{1_{T}^{\prime}}{T}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})$
    Thus the potential weights are
    $\bm{\rho}_{i}(\mathbf{w})=\Lambda\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\ddot{\mathbf{z}}(\mathbf{w})^{\prime}$
    Note that
    $\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=\Lambda\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\underbrace{\left({\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}}\right)}_{=0}=0.$
    Therefore $\bm{\pi}$ defined by $\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$ is a valid implicit design. It is proper by definition. Now, let $\dot{\mathbf{z}}(\mathbf{w})$ be the within-transformed covariate transform for this regression (which includes the time fixed effects), which does not depend on $i$. Note that the Gram matrix is
    $G_{n}(\bm{\pi}^{*})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\dot{\mathbf{z}}(\mathbf{w})^{\prime}\dot{\mathbf{z}}(\mathbf{w})]=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\dot{\mathbf{z}}(\mathbf{w})^{\prime}\dot{\mathbf{z}}(\mathbf{w})=G_{n}(\bm{\pi}).$
    Thus $\bm{\pi}$ is Gram-consistent.
    The argument for the estimated implicit design is analogous. The sample residuals of $\mathbf{z}(\mathbf{W}_{i})$ regressing on unit and time fixed effects is
    $\mathbf{z}(\mathbf{W}_{i})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{W}_{i})-\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\hat{\pi}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\hat{\pi}(\mathbf{w}^{\prime})1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w}^{\prime})$
    where $\hat{\pi}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(\mathbf{W}_{i}=\mathbf{w})$. The estimated potential weights are proportional to these residuals evaluated at $\mathbf{W}_{i}=\mathbf{w}$. Thus $\hat{\bm{\pi}}$ defined by $\hat{\pi}_{i}(\cdot)=\hat{\pi}(\cdot)$ solves (2.4).
  2. (ii)
    In this case, any implicit design $\tilde{\pi}_{i}$ needs to satisfy
    $0=\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\left({\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}}\right)$
    Since $\Lambda=I$ and $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]$ is positive definite by Theorem B.1, this implies that
    $\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}=0$
    Inspecting this expression, we have
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\mathbf{z}(\mathbf{w})=1_{T}\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\frac{1_{T}^{\prime}\mathbf{z}(\mathbf{w})}{T}$
    Column $k$ of this expression is in turn
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=1_{T}C_{k}$
    for some scalar $C_{k}$. The left hand side is in the span of $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}$, which excludes $1_{T}$ by assumption. Thus,
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=0.$
    Since at most only one $\mathbf{w}$ leads to $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=0$, and nonzero vectors $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}$ are linearly independent, we have that
    $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$
    for all but one $\mathbf{w}$ by linear independence. Since both designs sum to 1, they also equal for remaining $\mathbf{w}$.
Proposition B.14 (Forbidden comparisons).
Consider the TWFE specification $Y_{it}=\alpha_{i}+\gamma_{t}+\tau W_{it}$ under staggered adoption ($\mathbf{w}_{t}\leq\mathbf{w}_{t+1}$ for all $\mathbf{w}\in\mathcal{W}$ and all $t\in[T]$). Suppose the treatment time is randomly assigned so that $\pi_{i}^{*}(\mathbf{w})=\pi^{*}(\mathbf{w})$. If there are two treatment paths with one being never-treated, $\mathbf{W}=\left\{{0,\mathbf{w}}\right\}$, then $\bm{\rho}_{it}(\mathbf{w})\geq 0$ for all post treatment periods $\mathbf{w}_{t}=1$ and $i\in[n]$. Otherwise, in all other configurations where $|\mathcal{W}|\geq 2$, there exists a choice of $\pi^{*}(\mathbf{w})$ such that the corresponding $\bm{\rho}_{it}(\mathbf{w})<0$ for some treatment path $\mathbf{w}$ with positive assignment probability ($\pi^{*}(\mathbf{w})>0$) and some post-treatment period $t$ ($\mathbf{w}_{t}=1$).
Proof.
The potential weight can be computed in closed form. For some $V>0$,
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left[{\mathbf{w}_{t}-\frac{1^{\prime}}{T}\mathbf{w}-\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi^{*}(\tilde{\mathbf{w}})\tilde{\mathbf{w}}_{t}+\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi^{*}(\tilde{\mathbf{w}})\frac{1^{\prime}}{T}\tilde{\mathbf{w}}}\right].$
When $\mathcal{W}=\left\{{0,\mathbf{w}}\right\}$, then for a post-treatment $t$,
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}(1-\pi^{*}(\mathbf{w}))(1-1^{\prime}\mathbf{w}/T)\geq 0.$
Otherwise, let $\mathbf{w}$ be the treatment path with the earliest adoption date, and consider $t=T$. By assumption, $\mathcal{W}$ contains a path that adopts later than $\mathbf{w}$. Then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({(1-\pi^{*}(\mathbf{w}))(1-1^{\prime}\mathbf{w}/T)-\sum_{\tilde{\mathbf{w}}\neq\mathbf{w}}\pi^{*}(\tilde{\mathbf{w}})\left({1-1^{\prime}\tilde{\mathbf{w}}/T}\right)}\right)$
Pick $\pi^{*}(\mathbf{w})=1/2=\pi^{*}(\tilde{\mathbf{w}})$ where $\tilde{\mathbf{w}}$ adopts later than $\mathbf{w}$. Then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({1/2(1-1^{\prime}\mathbf{w}/T)-1/2\left({1-1^{\prime}\tilde{\mathbf{w}}/T}\right)}\right)=\frac{1}{2}V^{-1}1^{\prime}(\tilde{\mathbf{w}}-\mathbf{w})/T<0$
since $\tilde{\mathbf{w}}$ adopts later than $\mathbf{w}$. ∎