### v3

# Potential weights and implicit causal designs in linear regression

Abstract.
When we interpret linear regression as estimating causal effects justified by quasi-experimental treatment variation, what do we mean? This paper characterizes the necessary implications when linear regressions are interpreted causally. A minimal requirement for causal interpretation is that the regression estimates some contrast of individual potential outcomes under the true treatment assignment process. This requirement implies linear restrictions on the true distribution of treatment. Solving these linear restrictions leads to a set of implicit designs. Implicit designs are plausible candidates for the true design if the regression were to be causal. The implicit designs serve as a framework that unifies and extends existing theoretical results across starkly distinct settings (including multiple treatment, panel, and instrumental variables). They lead to new theoretical insights for widely used but less understood specifications.
1. Introduction
Linear regression is overwhelmingly popular in applied microeconomics for estimating causal effects. Users frequently justify it by arguing that a treatment variable is quasi-experimentally assigned, rather than that it correctly specifies a structural model for potential outcomes. [^1] Under this view, when one runs $Y_{i}=W_{i}\tau+X_{i}^{\prime}\gamma+\epsilon_{i}$ for treatment $W$ and covariates $X$, one does not think of the regression as a structural model $Y(w,x,\epsilon)=w\tau+x^{\prime}\gamma+\epsilon$; the regression equation simply specifies an estimator. This view reflects in Angrist and Pischke (2010), “With the growing focus on research design, it’s no longer enough to adopt the language of an orthodox simultaneous-equations framework […] The new emphasis on a credibly exogenous source of variation has also filtered down to garden-variety regression estimates, in which researchers are increasingly likely to focus on sources of omitted-variables bias, rather than a quixotic effort to uncover the ‘true model’ generating the data.” This view is also articulated in the preface to Mostly Harmless Econometrics, “Most econometrics texts appear to take econometric models very seriously […Instead,] a principle that guides our discussion is that estimators in common use almost always have a simple interpretation that is not heavily model-dependent” (Angrist and Pischke, 2008). In this view, regression is essentially a machine that compares units under different treatment conditions, potentially adjusting for other observables. For a causal interpretation to stand, then, the treatment must have been assigned in such a way that these comparisons recover causal effects—in a given application, this leaves questions: assigned in which ways? And which causal effects?
A large applied econometrics literature studies these questions, usually for particular specifications. [^2] Among others, Imbens and Angrist (1994); Angrist (1998); Lin (2013); Słoczyński (2022, 2024); Blandhol et al. (2025); Aronow and Samii (2016); Goldsmith-Pinkham et al. (2024); Borusyak and Hull (2024); Athey and Imbens (2022); Kline (2011); Bugni et al. (2023); Mogstad and Torgovitsky (2024); Arkhangelsky and Imbens (2023); Arkhangelsky et al. (2021); Chetverikov et al. (2023); Kolesár and Plagborg-Møller (2024); Zhao et al. (2025). This literature produces many results under specialized analysis: Certain specifications have quasi-experimental interpretations under certain assignment schemes, but other configurations of specification and treatment assignment may produce uninterpretable estimands (e.g., negatively weighted causal effects). However, without a general principle, researchers may struggle to check whether prior results cover a specification and to examine its quasi-experimental properties if not.
This paper develops generic theoretical and computational tools to answer these questions. We ask instead: what necessary conditions must hold for a regression to admit a quasi-experimental interpretation? We show that these conditions are computable for an arbitrary regression specification and are informative of its causal interpretation. Indeed, many conclusions from the literature emerge from mechanically computing these conditions on particular specifications.
To do so, we first propose a minimal requirement for causal interpretation:
  1. (MC)
    Under the true treatment assignment process—henceforth the true design—the regression estimates a contrast of individual potential outcomes [^3] If there are $J$ treatments $\left\{{0,\ldots,J-1}\right\}$ and $n$ units $\left\{{1,\ldots,n}\right\}$, we mean such a contrast of individual potential outcomes $y_{i}(j)$ to be a parameter of the form $\frac{1}{n}\sum_{i=1}^{n}\sum_{j=0}^{J-1}\omega_{i}(j)y_{i}(j)$, for weights that sum to zero across $j$: For all units $i$, $\sum_{j=0}^{J-1}\omega_{i}(j)=0$. With binary treatments, these are parameters of the form $\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))$, where the $\omega_{i}$’s are permitted to be negative. for any potential-outcome distribution—even a worst-case one.
The condition (MC)—for minimally causal—captures two ideas: First, if a regression yields a quasi-experimental estimate for one outcome, it should do so all hypothetical outcomes exposed to the same treatment. Randomized experiments identify causal effects for any outcome by the same comparisons of units under different treatments: (MC) extends this logic to quasi-experiments. (MC) thus imposes that causal identification comes purely from treatment assignment, not from modeling potential outcomes—taking the credibility revolution’s emphasis on treatment variation very seriously (Angrist and Pischke, 2010). [^4] Thus, (MC) would rule out identification strategies that rely on parallel trends (Roth and Sant’Anna, 2023). Nevertheless, studying (MC) is still informative for users of difference-in-differences: Researchers often appeal to treatment variation as justifying parallel trends. For instance, Martinez-Bravo et al. (2022) write (emphasis ours), “we study the introduction of elections in the 1980s and 1990s […] This reform was mandated by the central government and rolled out in a top-down manner. We document that the timing of the first election is uncorrelated with a large set of village characteristics. This suggests that timing was quasi-random, which is consistent with the descriptive literature. Thus, we exploit the staggered timing of the introduction of elections across villages to estimate a difference-in-difference effect of the introduction of elections.”
Second, (MC) imposes that reasonable estimands should at least take the form of a contrast of individual potential outcomes with unit-level weights that sum to zero—called level independence by Blandhol et al. (2025). If not, shifting all potential outcomes of any unit by a constant (holding treatment effects fixed) would change the estimand. Such sensitivity to baseline outcome heterogeneity undermines causal interpretation. However, estimands that meet level independence can still lack substantive relevance, so (MC) remains only a minimal requirement.
To check (MC), we can separate it into a computable and a substantive question:
  1. (MC1)
    Does any design (i.e., distribution of treatment), not rejected by the data, satisfy the condition in (MC). What is the estimand of the regression under that design?
  2. (MC2)
    Is any design found by (MC1) plausibly how treatment was assigned in the application at hand? Is the regression estimand substantively informative?
For (MC) to stand, both must be answered affirmatively. This paper focuses on (MC1): (MC2) depends on the precise empirical context, and is thus not generically answerable. [^5] In observational settings, one also cannot confirm (MC2), since one only observes one draw of treatment from the true design.
We characterize (MC1) constructively. It turns out that answers to (MC1) are characterized by simple linear equations. The characterization covers panels, cross-sections, and binary or multi-valued—but finite—treatments and extends to two-stage least-squares. In leading cases, these linear equations either have no solution or a unique one, making them highly informative of (MC1). These are informative—though not dispositive—of (MC2).
Two simple observations underpin the analysis of (MC1). First, linear regression estimands $\bm{\tau}$ are linear aggregations of individual potential outcomes $\mathbf{y}_{i}(\cdot),i=1,\ldots,n$. The weight placed on $\mathbf{y}_{i}(\mathbf{w})$, for a treatment value $\mathbf{w}$, can be written as $\pi_{i}^{*}(\mathbf{w})$$\bm{\rho}_{i}(\mathbf{w})$:
[…]
$=\frac{1}{n}\sum_{i=1}^{n}\bigg{\{}{\sum_{\mathbf{w}\neq\mathbf{w}_{0}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})\underbrace{(\mathbf{y}_{i}(\mathbf{w})-\mathbf{y}_{i}(\mathbf{w}_{0}))}_{\text{individual $\mathbf{w}$-$\mathbf{w}_{0}$ treatment effect}}}\bigg{\}}+\left({\sum_{\mathbf{w}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})}\right)\mathbf{y}_{i}(\mathbf{w}_{0}).$ (for a choice of baseline treatment $\mathbf{w}_{0}$)
The weight put on $\mathbf{y}_{i}(\mathbf{w})$ depends on the true design $\pi_{i}^{*}(\mathbf{w})$, which is the unknown probability that individual $i$ receives treatment $\mathbf{w}$. It also involves certain numbers $\bm{\rho}_{i}(\mathbf{w})$ related to the specification. We call $\bm{\rho}_{i}(\mathbf{w})$ potential weights: Like potential outcomes, they vary by unit and treatment. Potential weights encode the specification’s counterfactual comparisons. Importantly, $\bm{\rho}_{i}(\mathbf{w})$ are known up to the population Gram matrix [^6] For a regression of random variables $Y_{1},\ldots,Y_{n}$ on $Z_{1},\ldots,Z_{n}$, the population Gram matrix is $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[Z_{i}Z_{i}^{\prime}]$. of the linear regression, which is consistently estimable.
Second, (MC1) requires that $\bm{\tau}$ be a comparison of individual potential outcomes. With (1.1), this requirement is equivalent to:
For every $\text{For every $i=1,\ldots,n$, }\sum_{\mathbf{w}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=0.$
Because $\bm{\rho}_{i}(\mathbf{w})$ are known, these restrictions form linear restrictions in the unknown design $\pi_{i}^{*}(\mathbf{w})$. Solving these equations then answers (MC1); (MC2) examines whether these solutions are plausibly the true design.
We label any solution an implicit design $\pi_{i}(\cdot)$, distinguishing it from the true design $\pi_{i}^{*}(\cdot)$. Often the system (1) is over-identified, so the implicit design is either unique or nonexistent. Substituting $\pi_{i}(\mathbf{w})$ into (1.1) yields the corresponding implicit estimand—which is what the regression targets if the true design were $\pi_{i}(\cdot)$. The implicit design and the implicit estimand—both easily computable with an estimate of the population Gram matrix—characterize (MC1).
As a simple example, consider applying this approach to a regression of an outcome on a binary treatment without other covariates. There, the unique implicit design has constant propensity scores $\pi_{i}(1)=\pi$, for $\pi$ the population mean of the binary treatment. The corresponding implicit estimand is the average treatment effect. This is intuitive: If we would like to interpret this outcome-on-treatment regression quasi-experimentally, then we must believe that the treatment was uniformly randomized and that we should target the average treatment effect. There is no alternative design-estimand pair that satisfies (MC).
We can think of the implicit design as how a regression implicitly estimates the generalized propensity score and of the implicit estimand as its chosen target. Making these choices transparent clarifies assumptions underlying a regression specification and opens those to further scrutiny. Since these choices are often unique, computing them is helpful for evaluating (MC2) by narrowing the set of plausible designs to a single candidate. While the implicit design can never verify (MC2), it could falsify it.
To that end, we offer practical diagnostics—illustrated with Blakeslee et al. (2020)—to assess and refine implicit designs and estimands. They examine statistical plausibility and economic relevance for the design and the estimand. Passing these diagnostics bolsters (MC2) and—by analogy with failing to reject a null—supports a causal interpretation. For instance, confirming that unit weights are non-negative dispels worries of “negative weights.” If the design is plausible but the estimand is not, researchers can retarget alternative estimands while keeping the same treatment model. If these diagnostics fail, a causal interpretation based on (MC) is rejected. Researchers should then adopt explicitly causal methods (see Wager, 2024, for an introduction) or defend the regression as a restrictive outcome model.
Despite its simplicity, reasoning through (MC) unifies prior work and yields new theory. First, (MC) captures the core logic behind many existing results in starkly distinct settings (Angrist, 1998; Blandhol et al., 2025; Goldsmith-Pinkham et al., 2024; Kline, 2011; Athey and Imbens, 2022). For these specifications, simply by computing the implicit design and reasoning through (MC2), we can supply a converse: We show that the only designs justifying causal interpretations, in the sense of (MC), are exactly those analyzed in the literature.
Second, we show that specifications that interact a binary treatment with covariates have delicate quasi-experimental interpretations (Lin, 2013; Miratrix et al., 2013; Imbens and Wooldridge, 2009; Kline, 2011; Zhao et al., 2025). Different designs are necessary for causal interpretations of different contrasts of the coefficients. These designs are often mutually incompatible: Requiring all contrasts to be causal necessitates generally logically inconsistent restrictions on the design.
Third, we show that two-way fixed effects (TWFE) specifications have fragile and stringent quasi-experimental interpretations under staggered adoption. Adding time-varying covariates or estimating on imbalanced panels can render the equations (1) unsolvable; even absent these features, the only solution to (1) involves totally randomized treatment paths. To our knowledge, these latter two sets of results are not converses of known results.
While the bulk of our results concern linear regression, we extend these results to two-stage least-squares. There, the implicit estimand is additionally helpful for deriving necessary conditions on compliance types (e.g. monotonicity Imbens and Angrist (1994)). These calculations recover converses to results in Blandhol et al. (2025); Imbens and Angrist (1994); Słoczyński (2024); Bhuller and Sigstad (2024)—they even help clarify a small gap in recent work on TSLS with multiple treatments.
This paper is particularly related to Chattopadhyay and Zubizarreta (2023), who characterize and interpret the weights on observed outcomes for several leading linear regression specifications. Their results can be thought of as analyzing estimation based on the observed potential weights $\bm{\rho}_{i}(\mathbf{W}_{i})$. Additionally, we consider necessary conditions for interpreting a generic regression as estimating some causal effect, whereas Chattopadhyay and Zubizarreta (2023) focus on sufficient conditions for interpreting particular regressions as estimating the average treatment effect.
This paper proceeds as follows. Section 2 contains our main results. To build intuition, Section 2.1 starts with the simple setting with cross-sectional data and binary treatments. Section 2.2 then formalizes (MC), (MC1), and (MC2) and their relation to implicit designs. Section 3 applies our framework to a litany of regression specifications, yielding new theoretical results. Section 4 extends the framework to two-stage least-squares. Section 5 illustrates our diagnostics with Blakeslee et al. (2020), who study the economic impact of water loss in rural India. Section 6 concludes.
2. Potential weights and implicit designs
Consider a finite population of units $i\in[n]\equiv\left\{{1,\ldots,n}\right\}$. Each unit receives one treatment $\mathbf{w}$ from a finite set $\mathcal{W}$ of size $J$. We denote by $\mathbf{W}_{i}$ the realized treatment. Each unit has covariates covariates $\mathbf{x}_{i}$ and vector-valued potential outcomes of length $T$, $\left\{{\mathbf{y}_{i}(\mathbf{w})\in\mathbb{R}^{T}:\mathbf{w}\in\mathcal{W}}\right\}$, without restrictions. [^7] For expositional clarity, we assume that the dimension of the outcome vector is the same across individuals (i.e., balanced panels). Section D.6.1 discusses imbalanced panels. After assignment, we observe a corresponding realized outcome $\mathbf{Y}_{i}=\mathbf{y}_{i}(\mathbf{W}_{i})$.
To emphasize that identification comes from variation in treatment assignment, we isolate this variation by thinking of $(\mathbf{x}_{i},\mathbf{y}_{i}(\cdot))$ as fixed numbers and only considering the randomness in $\mathbf{W}_{i}$. This design-based perspective aligns with the identification arguments and with computation; it permits correlated assignment, but is not substantively different from a sampling setup: 2.1 derives the analogues of our results in the sampling setup where potential outcomes, covariates, and treatments are i.i.d.
To that end, let $\bm{\pi}^{*}$ denote the marginal treatment assignment probabilities (i.e., propensity scores):
$\bm{\pi}^{*}=(\pi^{*}_{1},\ldots,\pi^{*}_{n})\text{ where }\pi_{i}^{*}(\mathbf{w})=\operatorname{P}(\mathbf{W}_{i}=\mathbf{w}).$
We shorthand $\bm{\pi}$ as a design. Since we consider estimands for linear estimators, the joint distribution of treatment matters only through the marginal distributions $\bm{\pi}$. In principle, these probabilities may be arbitrarily different across units, though only certain ones are plausible for (MC1).
Write a regression generically as $\mathbf{Y}_{it}=z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}\beta+\epsilon_{it}$ with $z_{t}(\cdot,\cdot)$ known. We would like to interpret certain regression contrasts $\bm{\tau}=\Lambda\beta\in\mathbb{R}^{k}$ as a vector of causal effects, for a known matrix $\Lambda\in\mathbb{R}^{k\times K}$, $k\leq K$. We think of this regression simply as an algorithm that—in the population—produces some estimand $\bm{\tau}$ given $(\Pi^{*},(\mathbf{y}_{i}(\cdot),\mathbf{x}_{i})_{i=1}^{n})$.
This setup is sufficiently general to encompass cross-sectional ($T=1$), panel ($T>1$), binary treatment ($J+1=2$), multivalued treatment ($J+1>2$), scalar contrast ($k=1$), and multiple contrasts ($k>1$) settings. [^8] Our results do extend to continuous treatments, but they become much less powerful, essentially since there are only finitely many restrictions for infinitely many objects. Our results also cover two-stage least-squares, which we discuss in Section 4.
To illustrate (1.1) and (1), we start informally with the binary-treatment, scalar-outcome, and scalar-contrast case $(T=J=k=1)$. Our main results push this essential intuition to the more general case.
2.1. Leading example and core intuition
We do not bold variables in this familiar setting. To fix ideas, consider the control-for-$x_{i}$ specification: For $x_{i}$ that includes a constant
$Y_{i}=\tau W_{i}+x_{i}^{\prime}\gamma+\epsilon_{i}.$
Here, $z(x_{i},W_{i})=[W_{i},x_{i}^{\prime}]^{\prime},\beta=[\tau,\gamma^{\prime}]^{\prime},\Lambda=[1,0_{\dim(x)}^{\prime}].$ Since treatment is binary, for a design $\bm{\pi}$, we write $\pi_{i}=\pi_{i}(1)$ and $1-\pi_{i}=\pi_{i}(0)$. We consider potentially more intricate specifications, returning to (2.1) as illustration.
To motivate, Blakeslee et al. (2020) study the impact of water loss in rural India on employment and income. Water loss is measured by a binary $W_{i}$, indicating whether the first borewell household $i$ drilled has failed. The authors justify identification quasi-experimentally, noting that, e.g., “[well failure] …depends on highly irregular, quasi-random subsurface properties” (p. 206).
The true design $\bm{\pi}^{*}$—the natural process of borewell failure—is unknown, but the authors argue that it nevertheless does not preclude causal inference. Since borewell failure is difficult to predict, treated and untreated households are plausibly comparable. To buttress this argument, the authors painstakingly document hydrogeological features that determine borewell failure, even having inserted specialized camera equipment into failed wells.
Nonetheless, the study simply estimates (2.1) across multiple outcomes. That choice appears in tension with their detailed discussion of $\bm{\pi}^{*}$, which the regression model does not explicitly incorporate. What should a reader believe about borewell failure to interpret the authors’ regression estimates as causal effects?
To answer these questions, let us return to a regression of $Y_{i}$ on $z(x_{i},W_{i})$. The population regression coefficient is defined as follows (cf. $\mathbb{E}[x_{i}x_{i}^{\prime}]^{-1}\mathbb{E}[x_{i}y_{i}]$):
[…]
Under these definitions, $\tau$ admits the representation (1.1):
$\tau=\Lambda\beta=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}\underbrace{\Lambda G_{n}^{-1}z(x_{i},1)}_{\rho_{i}(1)}y_{i}(1)+(1-\pi_{i}^{*})\underbrace{\Lambda G_{n}^{-1}z(x_{i},0)}_{\rho_{i}(0)}y_{i}(0).$
Here, the potential weights $\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)$ are known up to $G_{n}$. Like potential outcomes $y_{i}(w)$, the potential weights are indexed by unit and treatment.
Potential weights are readily connected to the weighting in the regression estimator. The least-squares estimator—itself linear in $Y_{i}$—can be written as
$\hat{\tau}=\Lambda\hat{\beta}=\Lambda\hat{G}_{n}^{-1}\frac{1}{n}\sum_{i=1}^{n}z(x_{i},W_{i})Y_{i}=\frac{1}{n}\sum_{i=1}^{n}\underbrace{\Lambda\hat{G}_{n}^{-1}z(x_{i},W_{i})}_{\hat{\rho}_{i}(W_{i})}y_{i}(W_{i}).$
The weight on $Y_{i}$ in $\hat{\tau}$ corresponds exactly to the empirical counterpart of $\rho_{i}(W_{i})$—the potential weights for $i$ evaluated at $i$’s observed outcome. Comparing with (2.1), we can think of the potential weight $\rho_{i}(1-W_{i})$ of the counterfactual treatment $1-W_{i}$ as what the regression estimator would have done to the potential outcome $y_{i}(1-W_{i})$, had unit $i$ received the opposite assignment instead. In this sense, the potential weights encode how the regression makes comparisons.
Specializing to (2.1), the potential weights are
$\rho_{i}(w)=\frac{w-x_{i}^{\prime}\delta}{\frac{1}{n}\sum_{i=1}^{n}x_{i}^{\prime}\delta(1-x_{i}^{\prime}\delta)}\text{ for }x_{i}\equiv[1,x_{i}^{\prime}]^{\prime},\,\delta\equiv\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}\pi_{i}^{*}}\right).$
$\rho_{i}(w)$ is proportional to the partialled out treatment $w-x_{i}^{\prime}\delta$, where $\delta$ is the projection coefficient of $\pi^{*}$ on $x_{i}$.
For the regression to satisfy (MC), we require that the true design $\bm{\pi}^{*}$ is such that the estimand (2.1) satisfies level independence (Blandhol et al., 2025). This is formalized in the following definition.
Definition 2.1.
We say that $\tau$ is minimally causal if $\tau$ is always unchanged when we replace all potential outcomes $y_{i}(w)$ with $y_{i}(w)+c_{i}$ for arbitrary $c_{i}\in\mathbb{R}$, fixing $\bm{\pi}^{*},x_{1},\ldots,x_{n}$. Since $\tau$ is a linear aggregation, equivalently, $\tau$ is minimally causal if there is some $\omega_{1},\ldots,\omega_{n}\in\mathbb{R}$, not dependent on $y_{i}(\cdot)$, such that $\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))$ for all choices of $y_{i}(1),y_{i}(0)\in\mathbb{R}$.
2.1 provides a natural minimal requirement for quasi-experimental estimands. It simply imposes that a quasi-experimental estimand should be invariant to any changes to the potential outcomes that do not change individual treatment effects. For linear estimands, this condition is equivalent to $\tau$ being a weighted average treatment effect—these weights $\omega_{i}$ may be negative. [^10] Allowing for negative weights is sometimes too lenient for some treatment effects, but it accommodates for, e.g., interpreting differences in treatment effects between subgroups as minimally causal. It is also easier to extend to multiple treatments. What Blandhol et al. (2025) term “weakly causal” additionally requires that $\omega_{i}\geq 0$. Since we could recover the estimand only requiring minimally causality, we could also check whether the estimand satisfies convex weighting if desired.
Importantly, 2.1 excludes identification arguments that rely on modeling $y_{i}(0)$. The thought experiment therein keeps $\pi^{*}$ fixed while varying potential outcomes. Hence $\pi^{*}$ must not depend on $y_{i}(\cdot)$—implicitly assuming unconfoundedness. Thus, 2.1 is stringent in that identification rests purely on variation from the treatment. [^11] For instance, it rules out identification arguments that, say, only impose $Y_{i}(0)\indep W_{i}\mid X_{i}$ (in which case $\pi_{i}(w)$ may change if $y_{i}(1)$ changes) or those that impose $\mathbb{E}[Y_{i}(0)\mid X_{i}]$ is linear (which rules out $c_{i}$ that are not linear in $X_{i}$). These arguments may well justify causal interpretation, but they are not strictly arguments about treatment assignment.
Observe that $\tau$ satisfies 2.1 under $\bm{\pi}^{*}$ if and only if
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0\text{ for all $i=1,\ldots,n$.}$
By (2.1), the left-hand side of (2.1) is $\tau$’s change when unit $i$’s potential outcomes shifts by $c_{i}=n$. 2.1 forces this effect to zero. Hence $\pi^{*}$ must satisfy (2.3) for $\tau$ to be minimally causal.
Although $\bm{\pi}^{*}$ is unknown, any design that leads to a minimally causal estimand must satisfy (2.1). We therefore separately ask which designs satisfy (2.1) (i.e., (MC1)) and whether $\bm{\pi}^{*}$ can plausibly be one of those designs (i.e., (MC2)). For the first question, we can solve for $\pi_{i}^{*}$ in (2.1), at least when $\rho_{i}(0)\neq\rho_{i}(1)$. Let $\pi_{i}$ denote the solution to (2.1) viewed as equations in $\pi_{i}^{*}$, and we call $\bm{\pi}=(\pi_{1},\ldots,\pi_{n})$ an implicit design:
$\pi_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}.$
In turn, (MC2) requires $\pi_{i}^{*}=\pi_{i}$—the true and implicit designs coincide. This makes explicit the necessary conditions of interpreting $\tau$ causally in the sense of (MC).
Identifying $\bm{\pi}$ immediately identifies the estimand. If $\pi_{i}$ were the true design, then the estimand $\tau$ is a weighted average treatment effect with weights $\omega_{i}\equiv\pi_{i}\rho_{i}(1)=-(1-\pi_{i})\rho_{i}(0)$
$\tau=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}\rho_{i}(1)(y_{i}(1)-y_{i}(0))=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0)).$
These weights are non-negative when $\rho_{i}(0)\leq 0\leq\rho_{i}(1)$. Thus, simply solving (2.1) yields both the candidate design $\bm{\pi}$ and its estimand. The pair answers (MC1); defending their plausibility addresses (MC2).
There are at least two ways in which $\pi_{i}$ cannot possibly equal $\pi^{*}_{i}$, leading to a rejection of (MC1). The more obvious one is if $\pi_{i}\not\in[0,1]$ for any $i$ or if $\rho_{i}(1)=\rho_{i}(0)\neq 0$. This occurs when both $\rho_{i}(1)$ and $\rho_{i}(0)$ are on the same side of zero. When this happens, the implicit design is not a probability distribution, and cannot equal the true design. More subtlely, $\pi_{i}$ is also indefensible if it generates a different Gram matrix:
$G_{n}(\bm{\pi})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}z(x_{i},1)z(x_{i},1)^{\prime}+(1-\pi_{i})z(x_{i},0)z(x_{i},0)^{\prime}\neq G_{n}(\bm{\pi}^{*}).$
This restriction is useful when we analyze specifications theoretically under this framework. It is harder to implement when we do not know and have to estimate $G_{n}$—though one could use it as a basis for inference (Appendix B).
Returning to (2.1), solving (2.1) shows that the implicit design is precisely the population projection of $\pi_{i}^{*}$ onto the covariates:
$\pi_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}=\frac{\tilde{x}_{i}^{\prime}\delta}{1-\tilde{x}_{i}^{\prime}\delta+\tilde{x}_{i}^{\prime}\delta}=\tilde{x}_{i}^{\prime}\delta,$
where the corresponding estimand is a weighted average treatment effect
$\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))\quad\omega_{i}=\frac{\pi_{i}(1-\pi_{i})}{\frac{1}{n}\sum_{j=1}^{n}\pi_{j}(1-\pi_{j})}.$
Thus, to interpret (2.1) as estimating an economically interesting causal effect, one must defend (a) that $\pi_{i}^{*}=\pi_{i}=\tilde{x}_{i}^{\prime}\delta$ and (b) that one is interested in (2.1). On the other hand, a quasi-experimental interpretation immediately fails if we observe $\pi_{i}\not\in[0,1]$, which happens when the linear probability model for $\pi_{i}^{*}$ produces nonsensical fitted values. [^12] For this specification, Theorem 3.1 verifies that $G_{n}(\bm{\pi})=G_{n}(\bm{\pi}^{*})$ always holds regardless of whether $\bm{\pi}=\bm{\pi}^{*}$, and thus there is no additional power in the Gram matrix criterion (2.1).
Specification (2.1) is studied by Angrist (1998) and Blandhol et al. (2025). [^13] Both Angrist (1998) and Blandhol et al. (2025) consider a superpopulation sampling setup. Angrist (1998) considers a binary $x_{i}$ in his equation (9), but the argument can be easily generalized, e.g., in Borusyak and Hull (2024); Goldsmith-Pinkham et al. (2024). Corollary 1 in Blandhol et al. (2025) shows that assuming unconfoundedness, $\tau$ is a positively weighted average treatment effect if and only if the propensity score is linear. This is effectively what we find, and thus we view our result (formally in Theorem 3.1(1)) as a reinterpretation of theirs. Additionally, Section D.2 clarifies how our results relate to Theorem 1 in Blandhol et al. (2025). Our framework recovers these results by mechanically computing $\rho_{i}(1)$ and $\rho_{i}(0)$. We echo Blandhol et al. (2025) in that if one would like to interpret $\tau$ as minimally causal, then one must believe the linearity $\pi_{i}^{*}=\delta^{\prime}\tilde{x}_{i}$ and that one is targeting (2.1). No alternative design-estimand pair supports a causal interpretation under (MC).
We can also connect the regression estimator to inverse propensity weighting (IPW): For $\hat{\pi}_{i}=\frac{-\hat{\rho}_{i}(0)}{\hat{\rho}_{i}(1)-\hat{\rho}_{i}(0)}$ and $\hat{\omega}_{i}=\hat{\pi}_{i}\hat{\rho}_{i}(1)=(1-\hat{\pi}_{i})\hat{\rho}_{i}(0)$ the sample analogues of $(\pi_{i},\omega_{i})$, we can write the regression estimator as an IPW estimator
$\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\hat{\rho}_{i}(W_{i})y_{i}(W_{i})$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{\omega}_{i}\left({\frac{W_{i}}{\hat{\pi}_{i}}y_{i}(1)-\frac{1-W_{i}}{1-\hat{\pi}_{i}}y_{i}(0)}\right)$
[…]
With $\hat{\pi}_{i}\in(0,1)$, this becomes a standard IPW estimator with estimated propensity scores. Accordingly, in effect, (MC1) asks whether such an IPW representation exists, and (MC2) whether its propensity model is correctly specified. Through this representation, treating $\hat{G}_{n}$ as known exactly abstracts away from estimating this propensity model.
To summarize, there are four key elements to this analysis, which we generalize in the next subsection:
  1. (i)
    We treat the triplet ($(\left\{{z(x_{i},\cdot)}\right\}_{i=1}^{n},\Lambda,G_{n})$ as known (at least in the population).
  2. (ii)
    We write the population regression estimand $\tau$ in the form (1.1) and (2.1). Because we treat $G_{n}$ as known, the potential weights $\rho_{i}(w)$ are known for all units.
  3. (iii)
    We observe that imposing (MC) amounts to imposing that $\pi_{i}^{*}$ satisfies linear restrictions, in which the coefficients are the potential weights.
  4. (iv)
    Following (MC1) and (MC2), we call the solutions to these linear equations implicit designs. Computing them also yields the corresponding estimands by (2.1). $\tau$ is minimally causal if and only if an implicit design exists and the true design equals some implicit design.
We conclude this subsection by stating the superpopulation analogue of these results.
Remark 2.1 (Superpopulation).
Suppose we have $(Y_{i}(0),Y_{i}(1),W_{i},X_{i})\overset{\mathrm{i.i.d.}}{\sim}P$ that satisfies unconfoundedness ($Y_{i}(0),Y_{i}(1)\indep W_{i}\mid X_{i}$); unconfoundedness is necessary for the thought experiment in 2.1. Let $\pi^{*}(x)=P(W_{i}=1\mid X_{i}=x)$ denote the true propensity score. Analogous to (2.1), the population regression coefficient equals
$\tau(P)$ $=\Lambda\mathbb{E}_{P}[z(X_{i},W_{i})z(X_{i},W_{i})^{\prime}]^{-1}\mathbb{E}_{P}[z(X_{i},W_{i})Y_{i}(W_{i})]$
$=\mathbb{E}_{P}\left[{\pi_{i}^{*}(X_{i})\Lambda G_{n}^{-1}z(X_{i},1)Y_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\Lambda G_{n}^{-1}z(X_{i},0)Y_{i}(0)}\right]$ (Unconfoundedness)
$\equiv\mathbb{E}_{P}\left[{\pi_{i}^{*}(X_{i})\rho_{i}(1)Y_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)Y_{i}(0)}\right]$
for $G_{n}\equiv\mathbb{E}_{P}[z(X_{i},W_{i})z(X_{i},W_{i})^{\prime}]$.
Let $C_{i}$ be a random variable that satisfies $W_{i}\indep C_{i}\mid X_{i}$. Consider potential outcomes $Y^{\prime}_{i}(w)=Y_{i}(w)+C_{i}$ and let $P^{\prime}$ be the distribution for $(Y_{i}^{\prime}(0),Y_{i}^{\prime}(1),W_{i},X_{i})$. The sampling analogue of the requirement 2.1 is that for all such $P^{\prime}$, the corresponding regression coefficient is invariant: $\tau(P^{\prime})=\tau(P)$. 2.1 thus requires that for all such choices of $C_{i}$,
$\tau(P^{\prime})-\tau(P)=\mathbb{E}[\left\{{\pi_{i}^{*}(X_{i})\rho_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)}\right\}C_{i}]=0.$
Since $\pi_{i}^{*}(X_{i})\rho_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)$ is measurable with respect to $X_{i}$, we may pick $C_{i}=\pi_{i}^{*}(X_{i})\rho_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)$ so that
$\mathbb{E}\left[{\left\{{\pi_{i}^{*}(X_{i})\rho_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)}\right\}^{2}}\right]=0.$
This implies that $\pi_{i}^{*}(X_{i})\rho_{i}(1)+(1-\pi_{i}^{*}(X_{i}))\rho_{i}(0)=0$ for almost every $X_{i}$, which is exactly the sampling analogue of the requirement (2.1). $\blacksquare$
2.2. General setup
Now consider a regression of $\mathbf{Y}_{it}$ on a covariate transform $z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})\in\mathbb{R}^{K}$. The population Gram matrix is
$G_{n}(\bm{\pi}^{*})\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{W}_{i}\sim\pi_{i}^{*}}\left[{z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right].$
Following (i), we treat $(\Lambda,G_{n},\mathbf{z}(\mathbf{x}_{1},\cdot),\ldots,\mathbf{z}(\mathbf{x}_{n},\cdot))$ as known and refer to this tuple as a population regression specification. We make two additional comments in the general setup relevant for panel settings. First, since treating $G_{n}$ as known is motivated by its consistent estimation, we require representing fixed effects through the within-transformation for $\mathbf{z}(\mathbf{x},\cdot)$, rather than through unit-level dummy variables. [^14] That is, individual fixed effects should be incorporated by setting $\sum_{t}z_{t}(\mathbf{x}_{i},\cdot)=0$, rather than by considering unit dummies as covariates. To see this, for unit $i$, let $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{W}_{i})$ denote the covariate transforms that exclude the unit dummy. Assume $z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})$ includes a unit dummy. Then $\sum_{t=1}^{T}\mathbb{E}[\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{W}_{i})]$ is in the Gram matrix (it is the interaction between $\tilde{z}_{t}$ and the unit-$i$ dummy variable). However, this quantity is not consistently estimable as unit $i$ is only observed once. Second, assuming $\mathbf{z}(\mathbf{x},\cdot)$ is known amounts to being able to compute $\mathbf{z}$ for counterfactual treatments. This precludes including mediators (e.g. lagged outcomes) in the right-hand side of the regression, since we do not know counterfactual values of the mediator.
Returning to (ii), analogously write
$\bm{\tau}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\cdot\underbrace{\left[{\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right]}_{\bm{\rho}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}}\mathbf{y}_{i}(\mathbf{w})$ (2.7)
which verifies the representation (1.1). Relative to the simple case (2.1), here we sum over $J+1$ values, and the potential weights are now matrices of dimension $k\times T$, defined as $\bm{\rho}_{i}(\mathbf{w})\equiv\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}.$ [^15] Section D.1 shows that the potential weights for a given contrast do not depend on how the regression is parametrized. For instance, it does not matter if we write $Y_{i}=\alpha+\tau W_{i}+\epsilon_{i}$ instead as $Y_{i}=\mu_{1}\mathbbm{1}(W_{i}=1)+\mu_{0}\mathbbm{1}(W_{i}=0)+\epsilon_{i}$ and consider $\tau=\mu_{1}-\mu_{0}$. Section D.1 also shows that the potential weights are suitably invariant under the Frisch–Waugh–Lovell transform.
For (iii), a natural generalization of 2.1 imposes the invariance of the regression coefficient as one shifts potential outcome paths without changing treatment effects:
Definition 2.2 (Minimally causal).
$\bm{\tau}$ is minimally causal if it is always unchanged when we replace all potential outcomes $\mathbf{y}_{it}(\mathbf{w})$ with $\mathbf{y}_{it}(\mathbf{w})+c_{it}$ for arbitrary $c_{it}\in\mathbb{R}$, fixing $\bm{\pi}^{*},\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$.
For linear estimands, we can again represent the level independence requirement as weighted treatment effects, though more complexly. 2.2 is equivalent to requiring that each individual’s time-$t$ potential outcomes are aggregated by some weights that sum to zero:
$\bm{\tau}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}{\bm{\omega}_{i}(\mathbf{w})}\mathbf{y}_{i}(\mathbf{w})\text{ for some $\bm{\omega}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}$ where }0=\sum_{\mathbf{w}\in\mathcal{W}}\bm{\omega}_{i}(\mathbf{w})$
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq\mathbf{w}_{0}}{\bm{\omega}_{i}(\mathbf{w})}(\mathbf{y}_{i}(\mathbf{w})-\mathbf{y}_{i}(\mathbf{w}_{0}))$ (for any choice of baseline $\mathbf{w}_{0}\in\mathcal{W}$)
Given a choice of a baseline treatment $\mathbf{w}_{0}$, if $\bm{\tau}$ is minimally causal, then it is also a vector of weighted sums of individual treatment effects $\mathbf{y}_{it}(\mathbf{w})-\mathbf{y}_{it}(\mathbf{w}_{0})$.
The analogue to (iv) is then the following equations
$\text{For $i=1,\ldots,n$},\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=0,\quad\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})=1.$
Absorbing the sum-to-one condition, there are $kT$ restrictions for $J$ unknowns. Any solution is an implicit design. Implicit designs are typically unique when they exist, because the number of equations is often at least the number of unknowns ($kT\geq J$). For instance, $J+1$ treatments generate $k=J$ contrasts; panels under staggered adoption admit fewer unique treatment times $(J+1)$ than time horizon $T$. Lemma D.8 proves uniqueness when $T=1$.
For a given implicit design, the corresponding implicit estimand is the following, for $\bm{\omega}_{i}(\pi,\mathbf{w})=\pi_{i}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})$:
$\bm{\tau}(\bm{\pi})$ $\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\omega}_{i}(\bm{\pi},\mathbf{w})\mathbf{y}_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\neq\mathbf{w}_{0}}\bm{\omega}_{i}(\bm{\pi},\mathbf{w})(\mathbf{y}_{i}(\mathbf{w})-\mathbf{y}_{i}(\mathbf{w}_{0})).$ (2.9)
These observations result in the following theorem, which formalizes how implicit designs answer (MC1) and relate to (MC). For a given implicit design $\bm{\pi}$, we call it proper if all $\pi_{i}(\cdot)$ are probability distributions. We say it generates $G_{n}$ if it satisfies (2.1): $G_{n}(\bm{\pi})=G_{n}$.
Theorem 2.1.
$\bm{\tau}$ is minimally causal if and only if
  1. (1)
    Some implicit design $\bm{\pi}$ exists, is proper, and generates $G_{n}$, and
  2. (2)
    The true design $\bm{\pi}^{*}$ is equal to $\bm{\pi}$.
When this happens, the estimand $\bm{\tau}$ is equal to the implicit estimand under $\bm{\pi}$.
Theorem 2.1 separates (MC) into a computable question (MC1) and a scientific question (MC2), representing them as Theorem 2.1(1) and (2). Proper implicit designs that generate $G_{n}$ answer (MC1). If none exist, then $\tau$ cannot be minimally causal. Assessing (MC) further means judging whether the true design is plausibly $\bm{\pi}$ (MC2). That judgment is context-specific. Nevertheless, computing implicit designs makes the task concrete and transparent. Thus, if a researcher interprets $\bm{\tau}$ as a causal effect through quasi-experimental arguments, then they must be prepared to defend why the true assignment process $\bm{\pi}^{*}$ is plausibly one of the implicit designs $\bm{\pi}$.
When $T=J=k=1$, we also formalize Section 2.1 as a corollary.
Corollary 2.2 ($k=T=J=1$).
When $k=T=J=1$, $\tau$ is minimally causal iff
  1. (1)
    $\rho_{i}(1)\rho_{i}(0)\leq 0$ for all $i$. Some implicit design satifying $\pi_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}$ for all $i$ with one of $\rho_{i}(1)$ and $\rho_{i}(0)$ nonzero generates $G_{n}$.
  2. (2)
    For all units $i$ with one of $\rho_{i}(1)$ and $\rho_{i}(0)$ nonzero, $\pi^{*}_{i}=\frac{-\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}$.
When this happens, the implicit estimand is
$\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}^{*}(y_{i}(1)-y_{i}(0))\text{ for }\omega_{i}^{*}\equiv\omega_{i}(\bm{\pi}^{*},1)=\pi_{i}^{*}\rho_{i}(1).$
The unit-level weights $\omega_{i}^{*}$ are negative if and only if $\rho_{i}(1)<0<\rho_{i}(0)$.
In the simple case ($k=T=J=1$), if an implicit design exists, it is uniquely and explicitly defined. The implicit estimand is a weighted average treatment effect. All unit-level weights $\omega_{i}$ are nonnegative provided no unit has $\rho_{i}(1)<0<\rho_{i}(0)$.
2.3. Discussion and practical recommendation
Theorem 2.1 supplies designs and estimands for (MC1). Computing implicit designs clarifies the link between estimation and identification. When researchers invoke treatment-based identification arguments but do not explicitly model the design or define the estimand, our results reveal these hidden choices. If regression is instead primarily viewed as an outcome model, a valid implicit design/estimand makes the regression interpretable to those who challenge the model but are willing to entertain restrictions on treatment assignment.
Our analysis proceeds from the necessary implications of causal interpretation. Answering (MC1) therefore resembles “reverse-engineering” an interpretation for the regression estimator. Echoing Mogstad and Torgovitsky (2024), we stress that an implicit design alone does not guarantee statistical or economic relevance. This relevance ultimately rests on researcher and audience judgment. We therefore propose diagnostics that assess and refine the implicit estimand. Section 5 illustrates the diagnostics in practice.
To implement these diagnostics, practitioners can explicitly compute potential weights and implicit designs—most straightforwardly by replacing $G_{n}$ with $\hat{G}_{n}$. [^16] Certain joint distribution of treatment implies that $\hat{G}_{n}=G_{n}$ almost surely, in which case there is no estimation error in $G_{n}$ to account for and Theorem 2.1 is applicable as-is. See (A.1). We prove estimation consistency in Appendix B. [^17] To give some intuition that may assuage concerns, given a confidence set for $G_{n}$, it is in principle possible to project this confidence set to the set of potential weights and the set of proper implicit designs that generate $G_{n}$. Moreover, if one has a candidate design $\bm{\pi}$ in mind such that $\bm{\pi}$ and $G_{n}(\bm{\pi})$ satisfy (1), it is also straightforward to check whether the implied $G_{n}(\bm{\pi})$ is statistically close enough to $\hat{G}_{n}$.
Probing the implicit design. Suppose (1) reveals a proper implicit design $\bm{\pi}$, we may ask two broad questions:
  1. (1)
    Is $\bm{\pi}$ consistent with economic intuition?
    For instance, does $\bm{\pi}$ covary positively with features that predict selection into treatment? Does it explain too much variance in the observed $W_{i}$ to make us question whether $W_{i}$ is unpredictable?
  2. (2)
    Is $\bm{\pi}$ consistent with the data?
    For instance, is $\bm{\pi}$ well-calibrated? [^18] That is, among those units with $\pi_{i}\approx\pi$, do approximately $100\pi\%$ of units get treated? When we need $\pi^{*}_{i}$ to be linear, Blandhol et al. (2025) suggest using the Ramsey RESET test against the hypothesis that $\mathbb{E}[W_{i}\mid x_{i}]$ is linear. For fitted values $\hat{\pi}_{i}$, the RESET test effectively tests $\beta_{2}=\cdots=\beta_{p}=0$ in the regression $W_{i}=\alpha_{0}+\alpha_{1}^{\prime}x_{i}+\beta_{1}(\hat{\pi}_{i})^{2}+\cdots+\beta_{p}(\hat{\pi}_{i})^{p}+\epsilon_{i}.$ If we replace $\alpha_{0}+\alpha_{1}^{\prime}x_{i}$ with $\hat{\pi}_{i}$, the RESET test is thus similar to a test of the linearity—and thus calibration—of $\mathbb{E}[W_{i}\mid\hat{\pi}_{i}]$. With a binary treatment, are (transformations of) covariates balanced by the propensity scores $\pi_{i}$?
If these checks all pass, then (MC2) continues to be plausibly true—there lacks evidence to the contrary.
Probing the implicit estimand. We may also compute the implicit estimand via (2.9). The resulting weights $\omega$ rarely match canonical treatment contrasts. Similarly, we may check the following broad themes—in addition to Proposition 4.1 for TSLS:
  1. (1)
    Do the weights make sense economically?
    For instance, are any weights negative when $\bm{\tau}$ is not meant to be a difference of treatment effects? How large is the implicit subpopulation whose treatment effect the estimand is measuring (Poirier and Słoczyński, 2024)?
  2. (2)
    Do the weights matter for the bottom line in a given application?
    For instance, one could use the implicit design to construct alternative estimates and check them against the regression estimate: For user-chosen weights $\bm{\omega}_{i}(\mathbf{w})\in\mathbb{R}^{k\times T}$, [^19] In the simple case ($k=T=J=1$), a popular choice may be the trimmed ATE (Crump et al., 2009): $\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}\left({\frac{W_{i}}{\pi_{i}}-\frac{1-W_{i}}{1-\pi_{i}}}\right)Y_{i}\quad\omega_{i}\propto\mathbbm{1}(\epsilon<\pi_{i}<1-\epsilon)\quad\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=1.$ one can form the IPW estimator
    $\hat{\bm{\tau}}(\bm{\pi};\bm{\omega})=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\frac{\mathbbm{1}(\mathbf{W}_{i}=\mathbf{w})}{\hat{\pi}_{i}(\mathbf{w})}\bm{\omega}_{i}(\mathbf{w})\mathbf{y}_{i}(\mathbf{w}).$
    One can also plot estimates of $\mathbb{E}[Y_{i}\mid W_{i}=w,\hat{\pi}_{i}=\pi]$. If these functions are flat in terms of $\pi$, then it is likely that the regression-chosen weights make little difference and alternative user-chosen weights would result in similar estimates.
If both design and estimand pass these checks, then (MC2) is not falsified by the data. This dispels concerns about “negative weights” subjected to assuming (MC2).
Alternatives to regression. If these diagnostics fail, then the quasi-experimental defense for the regression becomes untenable. Practitioners can either use explicitly causal methods to and estimate explicit parameters or rely on outcome-modeling assumptions. For the former, there is a large and growing literature in causal inference that provides explicit methods—rarely are such methods implementable as a simple linear regression (see monographs and reviews by Wager, 2024; Imbens and Rubin, 2015; Ben-Michael et al., 2021; Chernozhukov et al., 2024). Alternatively, we may retreat by defending functional form restrictions and homogeneity on the potential outcomes and treatment effects. [^20] Even with outcome models, identification typically still requires assumptions on the process of treatment assignment. These restrictions are typically of the form that treatment assignment does not predict values of some “structural residual” defined by an outcome model (Andrews et al., 2023).
3. Theoretical applications and examples
This section applies our framework to assess (MC) across a wide swath of regression specifications and discusses them in self-contained vignettes. To emphasize, our results essentially reduce the problem to computing the potential weights and the set of implicit designs. This generates results across starkly distinct settings. Reasoning through $\bm{\pi}^{*}=\bm{\pi}$ theoretically is sometimes bespoke, but can often be aided by numerical methods.
Several specifications have known causal interpretations under specific designs (Angrist, 1998; Goldsmith-Pinkham et al., 2024; Kline, 2011; Imbens and Wooldridge, 2009; Lin, 2013; Athey and Imbens, 2022). Their implicit designs recover those properties and, moreover, supply a converse. Specifically, we show that causal interpretation analyzed in these settings are tenable only under those designs assumed in the literature; they target exactly those estimands found in the literature. Thus the minimal condition (MC) underpins much of the existing work. Imposing (MC) makes their conditions necessary as well.
Applying our framework to less-known specifications reveals new insights: Two classes of specifications admit quasi-experimental interpretations only under stringent conditions. First, cross-sectional regressions with $W\times x$ interactions qualify essentially only when $x$ is saturated (discrete) or $W$ is randomly assigned independently of $x$. Second, TWFE regressions with time-varying covariates or imbalanced panels lack implicit designs whenever treatment timing covaries with covariates or observation patterns.
3.1. A unified analysis of quasi-experimental interpretation in regression
Assume throughout that the population Gram matrix is invertible.
Theorem 3.1.
We compute the implicit designs and estimands of the following regression specifications. In every specification, the implicit design exists and is unique. The implicit design generates the Gram matrix regardless of whether $\bm{\pi}=\bm{\pi}^{*}$ for all specifications except (4).
  1. (1)
    (Angrist, 1998; Blandhol et al., 2025, Section 2.1, $k=T=J=1$) Suppose $x_{i}$ includes a constant. For $Y_{i}=\tau W_{i}+x_{i}^{\prime}\gamma+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      $\pi_{i}=x_{i}^{\prime}\delta$ for $\delta=\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}=x_{i}^{\prime}\delta$
    3. (c)
      $\omega_{i}\equiv\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=\frac{\pi_{i}(1-\pi_{i})}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(1-\pi_{i})}$. When $\pi_{i}^{*}=\pi_{i}$, $\omega_{i}\geq 0$.
  2. (2)
    (Goldsmith-Pinkham et al., 2024, $k=J>1,T=1$) Suppose $\mathcal{W}=\left\{{0,1,\ldots,J}\right\}$ for $J>1$ and $x_{i}$ includes a constant. Let $W_{ij}=\mathbbm{1}(W_{i}=j)$. For $Y_{i}=\sum_{j=1}^{J}\tau_{j}W_{ij}+x_{i}^{\prime}\gamma+\epsilon_{i}$ and $\bm{\tau}=[\tau_{1},\ldots,\tau_{J}]^{\prime}$ the coefficient of interest:
    1. (a)
      $\pi_{i}(j)=x_{i}^{\prime}\delta_{j}$ for $\delta_{j}=\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(j)x_{i}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}(j)=x_{i}^{\prime}\delta_{j}$ for all $j\in[J]$
    3. (c)
      The implicit estimand is shown in (2c). This estimand is generally contaminated (that is, $\omega_{ij}(\bm{\pi},\ell)\neq 0$ for some $j\in[J]$ and $\ell\not\in\left\{{0,j}\right\}$).
  3. (3)
    (Miratrix et al., 2013; Imbens and Wooldridge, 2009; Lin, 2013, $k=T=J=1$) Suppose $x_{i}$ represents saturated discrete covariates taking values in $\left\{{0,\ldots,L}\right\}$, in the sense that $x_{i}=[x_{i1},\ldots,x_{iL}]^{\prime}$ for $x_{i\ell}=\mathbbm{1}(x_{i}=\ell)$. Let $\overline{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$. For $Y_{i}=\alpha_{0}+\gamma_{1}^{\prime}x_{i}+\tau W_{i}+W_{i}(x_{i}-\overline{x})^{\prime}\gamma_{2}+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      For all units $i$ whose covariate value is $\ell$ (i.e., $x_{i\ell}=1$), the implicit design $\pi_{i}=\pi_{\ell}$, where $\pi_{\ell}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(x_{i}=\ell)$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}$ is the same for all units with the same $x_{i}$-value
    3. (c)
      The implicit estimand is the ATE. That is, $\omega_{i}=\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=1$.
  4. (4)
    (Kline, 2011, $k=T=J=1$) Let $\overline{x}_{1}=\frac{\sum_{i=1}^{n}\pi_{i}^{*}x_{i}}{\sum_{i=1}^{n}\pi_{i}^{*}}$. For $Y_{i}=\alpha_{0}+\gamma_{1}^{\prime}x_{i}+\tau W_{i}+W_{i}(x_{i}-\overline{x}_{1})^{\prime}\gamma_{2}+\epsilon_{i}$ and $\tau$ the coefficient of interest:
    1. (a)
      $\pi_{i}=\frac{\delta_{0}+(x_{i}-\overline{x})^{\prime}\delta_{1}}{1+\delta_{0}+(x_{i}-\overline{x})^{\prime}\delta_{1}}$, where $\delta_{0},\delta_{1}$ are equal to the population projection coefficients of $\pi_{i}^{*}/(1-\pi_{i}^{*})$ on $x_{i}-\overline{x}$ and a constant, weighted by $1-\pi_{i}^{*}$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x})$
    3. (c)
      When $\bm{\pi}^{*}=\bm{\pi}$, the implicit estimand is the ATT. That is, $\omega_{i}=\omega_{i}(\bm{\pi},1)=-\omega_{i}(\bm{\pi},0)=\frac{\pi_{i}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}}$.
  5. (5)
    (Athey and Imbens, 2022, $T>1$) Let $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$ be the set of treatment paths such that its nonzero elements are linearly independent vectors whose span excludes the constant vector $1_{T}=[1,\ldots,1]^{\prime}$. [^21] Staggered adoption corresponds to the case $\mathbf{w}_{t}\leq\mathbf{w}_{t+1}$ for all $t$. If $\mathcal{W}$ satisfies staggered adoption and excludes always-treated units, then it contains linearly independent vectors whose span excludes $1_{T}$. For the two-way fixed effects (TWFE) specificaiton $\mathbf{Y}_{it}=\alpha_{i}+\mu_{t}+\tau\mathbf{W}_{it}+\epsilon_{it}$ and the coefficient of interest $\tau$:
    1. (a)
      The implicit design is constant in $i$ and is unique, $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$
    2. (b)
      $\pi_{i}^{*}=\pi_{i}$ if and only if $\pi_{i}^{*}$ is the same for all $i$
    3. (c)
      The implicit estimand is shown in (5c), which matches Theorem 1(ii) in Athey and Imbens (2022) under staggered adoption. [^22] One might wish to impose that the post-treatment weights are nonnegative (i.e., $\bm{\omega}_{t}(\bm{\pi}^{*},\mathbf{w})\geq 0$ if $\mathbf{w}_{t}=1$). Failure of this condition implies that post-treatment units are severely used as comparisons for newly treated units, echoing the “forbidden comparison” issue in the recent difference-in-differences literature (Roth et al., 2023; Borusyak et al., 2024; De Chaisemartin and d’Haultfoeuille, 2020; Goodman-Bacon, 2021). Proposition D.14 shows that when $\mathcal{W}$ only has two elements and includes a never treated unit, all weights post treatment are non-negative, but such forbidden comparisons are possible in all other cases.
Theorem 3.1 computes implicit designs and estimands for specifications considered in the literature. Prior work analyzes each specification separately. Here, simply examining (MC1) shows the implicit design exists, is unique, and matches the form studied; the implicit estimand matches as well. Theorem 3.1 is thus a set of converses to the existing results—the regression estimand is minimally causal only if $\bm{\pi}^{*}=\bm{\pi}$ and the desired causal effect is the implicit estimand. These necessity results are new, to our knowledge, except for Theorem 3.1(1).
Rather than detail every vignette, we highlight two notable findings. First, Theorem 3.1(3) and (4) leave a few puzzles, which the next subsection resolves. Both regressions pick a contrast from the interacted model $Y_{i}=\gamma_{0}+\gamma_{1}^{\prime}x+\tau_{0}W_{i}+\tau_{1}^{\prime}x_{i}W_{i}+\epsilon_{i}$.232323 Theorem 3.1 (3) takes $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ while Theorem 3.1(4) takes $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{1}$. Curiously, (3) requires saturated covariates; (4) does not. Moreover, Theorem 3.1(4) is asymmetric. If we flip the roles of treatment and control, Theorem 3.1(4) would show that the average treatment effect on the untreated (ATU) estimand is minimally causal only if the reciprocal propensity odds $(1-\pi_{i}^{*})/\pi_{i}^{*}$ is linear in $x_{i}$. Thus, worryingly, the same specification yields ATT and ATU interpretations under different designs.
Second, Theorem 3.1(5) shows the TWFE estimand fails to be minimally causal unless treatment timing is fully randomized—the exact design Athey and Imbens (2022) study. Section 3.3 extends this by showing TWFE’s quasi-experimental interpretation is additionally fragile to time-varying covariates and panel imbalance. Section D.6 extends the analysis to one-way FE and event-study designs.
3.2. Forbidden interactions?
Assume $T=J=1$ and split $x_{i}$ into subvectors $x_{1i},x_{2i}$ (possibly overlapping). Consider the specification
$Y_{i}=\gamma_{0}+\tau_{0}W_{i}+\tau_{1}^{\prime}W_{i}x_{1i}+\gamma_{1}^{\prime}x_{2i}+\epsilon_{i}.$
Viewed as a linear outcome model (i.e., $\mathbb{E}[Y_{i}(w)\mid x_{i}]=\gamma_{0}+\gamma_{1}^{\prime}x_{2i}+w(\tau_{0}+\tau_{1}^{\prime}x_{1i})$), this specification models the baseline potential outcome as linear in $x_{2i}$ and the conditional average treatment effect as linear in $x_{1i}$, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid x_{i}]=\tau_{0}+\tau_{1}^{\prime}x_{1i}$. In particular, with a binary $x_{1i}$, $\tau_{1}$ is often interpreted as a difference in treatment effects by $x_{1i}$. Under this outcome model, $\bm{\tau}=(\tau_{0},\tau_{1}^{\prime})$ have causal interpretation.
One might hope these contrasts remain causal under a suitable design—even without the outcome model. Perhaps $x_{1}\mapsto\tau_{0}+\tau_{1}^{\prime}x_{1}$ is some linear approximation of the conditional average treatment effects. Thus, one might conjecture that without the outcome model but with the right design, $\tau_{0}+\tau_{1}^{\prime}x_{1}$ is nevertheless some weighted average of conditional average treatment effects evaluated at values around $x_{1}$.
Surprisingly, Proposition 3.2 shows this hope is misplaced, without full randomization or saturated covariates: Minimally causal interpretation of $\bm{\tau}$ necessitates that both $\pi_{i}^{*}$ and $\pi_{i}^{*}x_{1i}$ be linear in $x_{2i}$. When this fails, some contrast $\tau_{0}+\tau_{1}^{\prime}x_{1}$ does not satisfy level independence. [^24] This result was novel at the time of a working paper draft of this paper (arXiv:2407.21119v2); concurrent and independent work by Zhao et al. (2025) (arXiv:2502.00251) provides a similar result.
Proposition 3.2.
Consider the specification (3.2) and let $\bm{\tau}=(\tau_{0},\tau_{1}^{\prime})^{\prime}$ be the coefficients of interest. Then its corresponding implicit design exists if and only if, for some conformable $\Gamma_{0},\Gamma_{1}$, for all $i$,
$(\delta_{0}+\delta_{1}^{\prime}x_{2i})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$
where $\delta_{0},\delta_{1}$ are population projection coefficients of $\pi_{i}^{*}$ on $x_{2i}$: $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{2i}+\eta_{i}$ and $\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}+\upsilon_{i}$. When this happens, the unique implicit design is $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$. Therefore, if $\bm{\tau}$ is minimally causal, then $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{2i}$ and $\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$.
The necessary condition for interpreting $\bm{\tau}$ as minimally causal is that both the propensity score $\pi_{i}^{*}$ and its interaction with the covariates $\pi_{i}^{*}x_{1i}$ are linear functions of $x_{2i}$. When $x_{1i}$ is included in $x_{2i}$, this condition is unlikely to hold in general, as $\pi_{i}^{*}x_{1i}$ would involve nonlinear transformations of $x_{1i}$ and thus cannot be linear. On the other hand, this condition holds if $\pi_{i}^{*}$ is constant or if $x_{1i}$ represents a saturated categorical variable and $x_{2i}$ contains all other covariates interacted with $x_{1i}$. [^25] That is, $x_{1i}$ contains mutually exclusive binary random variables, and $x_{2i}$ contains $x_{1i}$, some set of other covariates $x_{3i}$, and all interactions $x_{3ik}x_{1i\ell}$.
What is wrong with interpreting $\tau_{0}+\tau_{1}^{\prime}x_{1i}$ as a linear approximation of the conditional average treatment effect? One could think of (3.2) as two regressions, one on the treated $W=1$ and one on the untreated $W=0$. Both regressions are indeed best linear approximations to $\mathbb{E}[Y(1)\mid x,W=1]$ and $\mathbb{E}[Y(0)\mid x,W=0]$, which are equal to the mean potential outcomes $\mathbb{E}[Y(1)\mid x],\mathbb{E}[Y(0)\mid x]$ under unconfoundedness. The contrast $\tau_{0}+\tau_{1}^{\prime}x_{1}$ is then the difference of the fitted values of these two regressions. However, the two regressions are only best linear approximations with respect to different distributions of the covariates ($x_{i}\mid W=1$ vs. $x_{i}\mid W=0$). Thus, their difference is not a best linear approximation to the conditional average treatment effect under any particular distribution of $x$. Shifting $Y(1)$ and $Y(0)$ by the same arbitrary amount therefore causes asymmetric behavior in the two regressions, leading to a failure of level irrelevance.
When $x_{1i}=x_{2i}=x_{i}$, this result supplements Theorem 3.1(3)–(4) by showing different contrasts necessitate incompatible designs. Insisting on all such contrasts being minimally causal imposes a knife-edge condition for the design. As a manifestation of this phenomenon, Theorem 3.1(4) shows that the model-based ATT $\tau_{0}+\tau_{1}^{\prime}\overline{x}_{1}$ maintains its interpretation under the design whose odds is linear, but the ATU requires that the reciprocal propensity odds is instead linear. Theorem 3.1(3) evades this difficulty by requiring that the covariates be saturated.
What about the model-based ATE $\tau_{0}+\tau_{1}^{\prime}\overline{x}$ when covariates are not saturated (cf. Theorem 3.1(3))? Theorem 1 in Chattopadhyay and Zubizarreta (2023) shows that estimating the ATE requires both propensity odds and reciprocal odds to be linear. If we drop the “ATE” requirement and seek any minimally causal contrast, the implicit design exists but is fractional-linear $\pi_{i}=\frac{\theta_{0}+\theta_{1}^{\prime}(x-\overline{x})}{1-\Gamma_{2}^{\prime}(x-\overline{x})}$, a strange model for propensity scores. [^26] The requirement that $\pi_{i}=\pi_{i}^{*}$ then imposes additional restrictions on $(\theta_{0},\theta_{1},\Gamma_{2})$. These constraints are unpleasant in the sense that they depend on the support points $x_{1:n}$. These issues are formalized in D.1 in Section D.5.
Taken together, interacted regressions are less robust in terms of (MC) than the simple regression (2.1), in contrast to the qualitative takeaway in Lin (2013); Negi and Wooldridge (2021). The uninteracted regression introduces variance weighting for the estimand, but maintains validity under a simple design. The interacted regression removes the weighting when the covariates are saturated, but estimates a weighted average treatment effect under a bizarre design in general.
Is there a simple regression that both removes the variance weighting in (2.1) and is quasi-experimental under linear propensity scores? Unfortunately, Proposition D.11 shows that the answer is no, at least not with right-hand-side variables that are linear in $[1,x_{i},W_{i},W_{i}x_{i}]$. [^27] One could estimate the uninteracted regression and weigh by $1/(\pi_{i}(1-\pi_{i}))$ to remove the variance weighting, but this approach requires estimating $\pi_{i}$ separately. As a result, targeting the average treatment effect under the same implicit design as (2.1) necessitates moving beyond regression estimators.
3.3. The fragility of quasi-experimental TWFE
Theorem 3.1(5) already limits TWFE to be minimally causal only under totally randomized treatment paths. This design is restrictive, and we may wonder whether it is possible to modify the regression so that $\pi_{i}^{*}$ are allowed to be functions of observable information. The most straightforward idea—indeed common empirical practice—is to include time-varying covariates in a TWFE regression.
Surprisingly, doing so often destroys quasi-experimental interpretation altogether: We would find no solutions to the equations (1). Because (1) are overdetermined in panel settings, small modifications to the potential weights can in principle make the system of equations unsolvable. With time-varying covariates, this happens precisely when the covariates matter for identification and change the estimand. On the other hand, when adding the covariates do not change the estimand, constant propensity scores remain the unique implicit design, and thus the covariates do not enrich the design either.
Proposition 3.3.
Assume $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$. Consider the regression $\mathbf{Y}_{it}=\alpha_{i}+\gamma_{t}+\tau\mathbf{W}_{it}+\delta^{\prime}\mathbf{x}_{it}$ where $\tau$ is the coefficient of interest. Let $\beta_{w\to x}$ be the population projection coefficient of $W_{it}$ on $x_{it}$ under $\bm{\pi}^{*}$, including individual and time fixed effects. For the set of implicit designs to be non-empty, a necessary condition is
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\})\,\text{ for all $i\in[n]$}.$
When $\beta_{w\to x}=0$, if $\mathcal{W}$ contains linearly independent vectors whose span excludes $1_{T}$, the implicit design is uniquely equal to $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$ for all $i$ as in Theorem 3.1(5).
An implicit design exists only if a demeaned covariate combination lies in the span of $\mathcal{W}$ and $1_{T}$ for every unit. This condition arises because we essentially need that the mean treatment $\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\mathbf{w}$ is exactly described by two-way fixed effects with time-varying covariates, analogous to the intuition for (2.1). This then restricts the space of the covariate vectors, since they need to generate vectors that lie in the linear span of $\mathcal{W}$.
With staggered adoption, $\operatorname{span}\left({\mathcal{W}\cup\left\{{1_{T}}\right\}}\right)$ is the set of vectors that are piecewise constant between adjacent adoption dates. This subspace is highly restrictive if there are relatively few adoption dates. If $\beta_{w\to x}\neq 0$, it is thus knife-edge that $\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}$ happens to be located in that subspace, unless columns of $\mathbf{x}_{i}$ happens to be piecewise constant over $t$ as well. [^28] This is plausible if the time-varying covariates are interactions of fixed covariates with the time fixed effects ($\mathbf{x}_{it}^{\prime}\delta=x_{i}^{\prime}\delta_{t}$). Proposition D.15 shows that for this specification, causal interpretation is possible necessarily under linear generalized propensity scores $\pi_{i}(\mathbf{w})=\delta_{0}(\mathbf{w})+\delta_{1}(\mathbf{w})^{\prime}x_{i}$. On the other hand, if $\beta_{w\to x}$ is zero under $\bm{\pi}^{*}$, including the covariates makes no difference to the coefficient on $W_{it}$ in population. Thus TWFE with time-varying covariates rarely retains a quasi-experimental interpretation. Researchers using such a specification either believe that the covariates do not affect treatment assignment and are irrelevant for identification, or they are embedding outcome modeling assumptions.
Finally, a similar fragility inflicts regressions with imbalanced panels. Such a regression is only design-based when the missingness pattern is uncorrelated with the treatment assignment pattern, in which case the design must again be total randomization of treatment paths. We detail this result in Section D.6.1.
4. Extension: Two-stage least-squares
Similar ideas to Theorem 2.1 extend to two-stage least-squares (TSLS): We can use level irrelevance to recover some design—now a distribution of the excluded instrument $W_{i}$. Interestingly, the implicit estimand also provides necessary conditions on compliance types for TSLS to estimate properly weighted causal effects. For instance, examining the implicit estimand for a binary treatment, binary instrument TSLS regression recovers (strong) monotonicity (Imbens and Angrist, 1994; Słoczyński, 2024) as a necessary condition. Doing so for TSLS with multiple treatments yields a compliance condition in Bhuller and Sigstad (2024). These results are recovered simply by enumerating which compliance types for each unit are consistent with the implicit estimand assigning proper weights to said unit’s potential outcomes.
Consider the following TSLS specification of a scalar outcome on a covariate transform
$Y_{i}=t(D_{i},x_{i})^{\prime}\beta+\epsilon_{i},$
instrumenting $t(D_{i},x_{i})$ with $z(W_{i},x_{i})$. Here, $D_{i}=d^{*}_{i}(W_{i})\in\mathcal{D}$ is the endogenous treatment, $d^{*}_{i}(\cdot)$ is the compliance type for unit $i$, and $t(\cdot,\cdot),z(\cdot,\cdot)$ are again known transforms. As an example, a binary treatment, binary instrument TSLS regression can be represented by $t(D_{i},x_{i})=[1,D_{i}]^{\prime}$ and $z(W_{i},x_{i})=[1,W_{i}]^{\prime}$. Suppose exclusion holds, so that $y_{i}(d_{i}^{*}(w),w)=y_{i}(d_{i}^{*}(w))$.
We extend steps (i)(iv) in Section 2.1. For (i), define the TSLS estimand for a contrast $\bm{\tau}=\Lambda\beta$ as
$\bm{\tau}=\Lambda\left({G_{tz}G_{zz}^{-1}G_{zt}}\right)^{-1}\left({G_{tz}G_{zz}^{-1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}[z(W_{i},x_{i})y_{i}(W_{i})]}\right),$
where $G_{tz}\equiv\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{t(d_{i}^{*}(W_{i}),x_{i})z(W_{i},x_{i})^{\prime}}\right]$, $G_{zt}\equiv G_{tz}^{\prime}$, and $G_{zz}\equiv\allowbreak\frac{1}{n}\sum_{i=1}^{n}\allowbreak\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{z(W_{i},x_{i})z(W_{i},x_{i})^{\prime}}\right]$. Write $H_{n}=\left({G_{tz}G_{zz}^{-1}G_{zt}}\right)^{-1}G_{tz}G_{zz}^{-1}$: $H_{n}$ is the analogue of the inverse Gram matrix $G_{n}^{-1}$. [^29] Indeed, if $t(d_{i}(W_{i}),x_{i})=z(W_{i},x_{i})$ so that the TSLS specification is equivalent to OLS, then $H_{n}$ is exactly the inverse Gram matrix. Like $G_{n}^{-1}$, $H_{n}$ is known in population and consistently estimable in sample. Thus, by the same reasoning, we treat the potential weights as known.
Next, for (ii), write $\bm{\tau}$ in the form of (1.1):
$\bm{\tau}=\frac{1}{n}\sum_{i=1}^{n}\sum_{w\in\mathcal{W}}\pi_{i}^{*}(w)\Lambda H_{n}z(w,x_{i})y_{i}(w)\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{w\in\mathcal{W}}\pi_{i}^{*}(w)\underbrace{\bm{\rho}_{i}(w)}_{k\times 1}y_{i}(d_{i}(w)).$
We define potential weights $\bm{\rho}_{i}(w)\equiv\Lambda H_{n}z(w,x_{i})$.
For (iii)(iv), the requirement that $\bm{\tau}$ is minimally causal (2.1) continues to be reasonable. If $\bm{\tau}$ were to be a comparison of different potential outcomes $y_{i}(d)$, then it should be invariant to shifting all $y_{i}(\cdot)$ by arbitrary $c_{i}$. Maintaining this restriction again yields (1) for $\pi_{i}^{*}(w)$, whose solutions we continue to term implicit designs. They continue to be plausible candidates for the true design $\pi_{i}^{*}(\cdot)$ in the sense of (MC1).
Just-identified settings have enough equations [^30] For an TSLS specification to be non-collinear, an instrument that takes $J$ values can support $k\leq J-1$ endogenous coefficients of interest. Since $\pi_{i}(\cdot)$ is a $J-1$-dimensional unknown vector, we need $k\geq J-1$ restrictions to have a unique implicit design. to pin down an implicit design $\pi_{i}(\cdot)$. If there are more distinct instrument values than coefficients of interest, then we may have too few restrictions on $\pi_{i}^{*}(w)$. However, it may be reasonable to also impose level irrelevance for certain first-stage coefficients, which would add more restrictions to recover a unique implicit design.
The estimand for TSLS depends on units’ compliance types $d_{i}^{*}(\cdot)$. Therefore, interpreting the estimand as a causal effect restricts compliance patterns. This can be operationalized as follows: Given an implicit design $\pi_{i}(\cdot)$, the corresponding implicit estimand $\bm{\tau}$ can be written as a weighted sum of individual potential outcomes:
$\bm{\tau}$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{w}\bm{\omega}_{i}(w;\bm{\pi})y_{i}(d_{i}^{*}(w))\quad\bm{\omega}_{i}(w;\bm{\pi})\equiv\pi_{i}(w)\bm{\rho}_{i}(w)\quad\sum_{w}\bm{\omega}_{i}(w;\bm{\pi})=0.$ (4.1)
$=\frac{1}{n}\sum_{i=1}^{n}\sum_{k\in\mathcal{D}}\left({\sum_{w:d_{i}^{*}(w)=k}\bm{\omega}_{i}(w;\bm{\pi})}\right)y_{i}(k)\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{k\in\mathcal{D}}\bm{\omega}^{*}_{i}(k;\bm{\pi},d_{i}^{*})y_{i}(k).$ (4.2)
(4.1) is an aggregation of $w$-on-$y$ causal effects. (4.2) groups together $w$ values that lead to the same $d_{i}^{*}(w)=k$, thereby translating to $d$-on-$y$ effects. In (4.2), the weight on the $k$th treatment is $\bm{\omega}^{*}_{i}(k;\bm{\pi},d_{i}^{*})$, which is known given $d_{i}^{*}(\cdot)$. If $\bm{\tau}$ has a causal interpretation, we can then enumerate all compliance types $d_{i}^{*}$ for each unit and check which types yield weights $\bm{\omega}^{*}_{i}(k;\bm{\pi},d_{i}^{*})$ that are consistent with the interpretation.
To illustrate, consider a particular class of TSLS specifications: For $x_{i}$ that includes a constant, consider a just-identified specification with $J$ values of an unordered treatment
$t(d,x_{i})$ $=[\mathbbm{1}(d=1),\ldots,\mathbbm{1}(d=J-1),x_{i}^{\prime}]^{\prime}$
$z(w,x_{i})$ $=[\mathbbm{1}(w=1),\ldots,\mathbbm{1}(w=J-1),x_{i}^{\prime}]^{\prime}.$ (4.3)
In this TSLS specification, the coefficients of interest are $\bm{\tau}=(\tau_{1},\ldots,\tau_{J-1})^{\prime}$. $\tau_{k}$ is the coefficient on $\mathbbm{1}(d=k)$, meant to capture the causal effect of $d=k$ relative to $d=0$.
Examining entries in (4.2), we have
$\tau_{k}=\frac{1}{n}\sum_{i=1}^{n}\sum_{k^{\prime}=1}^{J-1}\omega_{i}^{(k,k^{\prime})}(d_{i}^{*})(y_{i}(k^{\prime})-y_{i}(0))\quad\text{ for }\quad\omega_{i}^{(k,k^{\prime})}\equiv(\bm{\omega}^{*}_{i}(k^{\prime};\bm{\pi},d_{i}^{*}))_{k}.$
If $\tau_{k}$ is to be interpreted as a causal effect of $d=k$ relative to $d=0$, then we should at least restrict $\omega_{i}^{(k,k)}\geq 0,\omega_{i}^{(k,k^{\prime})}=0$ for $k\neq k^{\prime}$. This means that $\tau_{k}$ captures a convex aggregation $y_{i}(k)-y_{i}(0)$ that is not contaminated by treatment effects of some other arm $\ell$, $y_{i}(\ell)-y_{i}(0)$. If this is true and if $\bm{\pi}^{*}=\bm{\pi}$, following Bhuller and Sigstad (2024), we say that TSLS assigns proper weights. [^31] When $J=2$, $\bm{\tau}$ is minimally causal and assigns proper weights if and only if it is weakly causal (Blandhol et al., 2025).
Given $\omega_{i}^{(k,k^{\prime})}(\cdot)$, for each unit, we can enumerate all compliance types $d_{i}(\cdot)$ and retain those consistent with proper weights. Analogous to implicit designs, we refer to each element of the following set as an implicit compliance profile:
$\left\{{(d_{1}(\cdot),\ldots,d_{n}(\cdot)):\text{ for all $i$, $k\neq k^{\prime}\in\mathcal{D}$},\omega_{i}^{kk}(d_{i})\geq 0\text{ and }\omega_{i}^{kk^{\prime}}(d_{i})=0}\right\}$
We summarize these results by the following proposition:
Proposition 4.1.
In TSLS, $\bm{\tau}$ is minimally causal if and only if
  1. (1)
    An implicit design $\bm{\pi}$ exists
  2. (2)
    $\bm{\pi}^{*}=\bm{\pi}$.
Additionally, $\bm{\tau}$ from (4.3) also assigns proper weights if and only if the following is true for the implicit estimand under $\bm{\pi}$:
  1. (3)
    An implicit compliance profile $d_{1}(\cdot),\ldots,d_{n}(\cdot)$ in (4) exists
  2. (4)
    Some implicit compliance profile $d_{1}(\cdot),\ldots,d_{n}(\cdot)$ is equal to $d_{i}^{*}(\cdot),\ldots,d_{i}^{*}(\cdot)$.
Like Theorem 2.1, Proposition 4.1 separates requirements for causal interpretation into computable and substantive components. We can directly compute items (1) and (3) since the potential weights, implicit design, and implicit estimand are known in the population. Results from this computation are plausible candidates for items (2) and (4)—if no such candidate is found, then causal interpretation must be rejected.
These computations yield informative answers. To illustrate, simply by computing the implicit design and compliance profiles, we can likewise recover necessary conditions to several recent results in the instrumental variables literature. To introduce these results, let us introduce some compliance notions.
Definition 4.1 (Compliance restrictions).
  • With two treamtents ($J=2$), we say that a profile $d_{1}(\cdot),\ldots,d_{n}(\cdot)$ satisfies strong monotonicity if either $d_{i}(1)\geq d_{i}(0)$ for all $i$ or $d_{i}(1)\leq d_{i}(0)$ for all $i$.
  • With $J>2$, for $k=1,\ldots,J-1$, we say that $d(\cdot)$ is a $k$-always taker if $d(\cdot)=k$; it is a $k$-never taker if $d(\cdot)\neq k$; otherwise we say $d(\cdot)$ is a $k$-complier. We say $d(\cdot)$ is a full complier if it is a $k$-complier for all $k$.
  • We say that a profile $(d_{1}(\cdot),\ldots,d_{n}(\cdot))$ satisfy common compliance if for any $k$ and any two $k$-compliers $d_{i}(\cdot),d_{j}(\cdot)$, we have $d_{i}(w)=k$ if and only if $d_{j}(w)=k$. We say that a compliance profile is instrument-specific if there exists some permutation $f(\cdot)$ of the instrument values $\left\{{0,\ldots,J-1}\right\}$ such that, for all $i$, either (i) $d_{i}(f(w))\in\left\{{0,w}\right\}$ or (ii) $d_{i}(\cdot)$ is constant. [^32] For three instrument values, up to permutation of the instruments, instrument-specific compliance limits $d(\cdot)$ to one of six types $(d(0),d(1),d(2))\in\left\{{(000),(111),(222),(010),(002),(012)}\right\}$—for, respectively, never-taker, always-1-taker, always-2-taker, 1-complier, 2-complier, or full complier (Bhuller and Sigstad, 2024).
Proposition 4.2.
Consider the TSLS specification in (4.3),
  1. (1)
    (Blandhol et al., 2025) The unique implicit design is $\pi_{i}(j)=x_{i}^{\prime}\delta_{j}$ for
    $\delta_{j}=\left({\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\prime}}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(j)x_{i}.$
  2. (2)
    When $\bm{\pi}^{*}=\bm{\pi}$, the implicit compliance profile relative to the implicit design satisfies the following:
    1. (a)
      (Słoczyński, 2024, Corollary 3.4) When $J=2$, all implicit compliance profiles satisfy strong monotonicity.
    2. (b)
      (Bhuller and Sigstad, 2024, Proposition 5/B.1) When $J>2$ and $x_{i}$ includes only a constant, all implicit compliance profiles obey common compliance.
      Furthermore, all implicit compliance profiles which contain a full complier are instrument-specific.
Like how 2.2 supplies a converse for Angrist (1998), Proposition 4.2 recovers several results for TSLS. Proposition 4.2(1) and (2)(a) recover the necessary direction for Corollary 3.4 in Słoczyński (2024) and Theorem 1 in Blandhol et al. (2025): With binary treatment, monotonicity is required for interpreting the TSLS coefficient causally, in the sense that it assigns proper weights. [^33] Theorem 1 in Blandhol et al. (2025) imposes exogeneity and monotonicity and finds that $\bm{\tau}$ is minimally causal and has proper weights if and only if $\pi_{i}^{*}$ is linear. Proposition 4.2(1) and (2)(a) in turn show that, just imposing exogeneity, if $\bm{\tau}$ is minimally causal and has proper weights, then linear propensity scores and monotonicity are satisfied (see Section D.2 for details). Likewise, Corollary 3.4 in Słoczyński (2024) shows that strong monotonicity implies proper weights, but not the converse. Without covariates, this is a converse to Imbens and Angrist (1994).
Proposition 4.2(2)(b) recovers—and corrects—claims in Propositions 5 and B.1 in Bhuller and Sigstad (2024). Proposition 5 in Bhuller and Sigstad (2024) attempts to characterize analogues of monotonicity with multiple treatments. It claims that if TSLS assigns proper weights, then compliance is instrument-specific—that is, up to permutation of the instrument values, we can think of instrument $w$ as an encouragement to take up treatment $w$ and have no effect on other treatment takeup. Unfortunately, just assuming TSLS assigns proper weights is too weak for this conclusion. [^34] See Section D.2 for a counter-example. We are grateful to Henrik Sigstad for discussion. Instead, the essence of their argument implies that compliance profiles satisfy common compliance; their conclusion in turn stands if it is known that some full complier exists. Both implications are captured by Proposition 4.2(2)(b).
5. Empirical application
We illustrate several empirical recommendations (Section 2.3) by reanalyzing Blakeslee et al. (2020). Blakeslee et al. (2020) study the economic impact of water loss in rurual India, viewing borewell failure ($W_{i}$) as treatment. They consider a range of income and employment outcomes and conclude that (i) well failure causes a decline in agricultural income and employment, but reallocation to off-farm offsets the lost income, and (ii) those living in high economic development areas adapt more easily.
The main empirical analyses of Blakeslee et al. (2020) contain estimates of the following regression (their equation (1) reproduced in our notation): For a household $i$, let $v(i)$ denote its village and $t(i)$ denote the year when it drilled its first borewell. Consider the coefficient $\tau$ in
$Y_{i}=\tau W_{i}+\underbrace{\alpha_{0}+x_{i}^{\prime}\gamma+\eta_{1v(i)}+\eta_{2t(i)}}_{\check{x}_{i}^{\prime}\mu}+\epsilon_{i}.$
where $x_{i}$ collects a vector of household characteristics and includes “the age, caste, and literacy of the household head, as well as the total land inherited by the household” (p.208). By Theorem 3.1(1), $\tau$ is minimally causal only if $\pi_{i}^{*}$ is linear in $\check{x}_{i}$. When that happens, the estimand is a weighted average treatment effect with weights proportional to $\pi_{i}^{*}(1-\pi_{i}^{*})$.
In their heterogeneity analysis, Blakeslee et al. (2020) consider the regression that interacts (5) with an indicator $h_{i}$ of whether the village $v(i)$ has high economic development
$Y_{i}=\tau_{0}W_{i}+\tau_{1}W_{i}h_{i}+\underbrace{(\alpha_{0}+x_{i}^{\prime}\gamma_{0}+\eta_{1v(i)}+\eta_{2t(i)}+\alpha_{1}h_{i}+h_{i}x_{i}^{\prime}\gamma_{1}+\kappa_{2t(i)}h_{i})}_{\tilde{x}_{i}^{\prime}\mu}+\epsilon_{i},$
where $\tau_{0}$ is interpreted as a treatment effect for those with $h_{i}=0$ and $\tau_{1}$ is interpreted as a difference of treatment effects among $h_{i}=1$ versus $h_{i}=0$. Our result in Proposition 3.2 show that $\tau_{0}$ and $\tau_{1}$ in (5) are both minimally causal only if $\pi_{i}^{*}$ is linear in $\tilde{x}_{i}$ and $\pi_{i}^{*}h_{i}$ is also linear in $\tilde{x}_{i}$. Because $h_{i}$ is in the linear span of the village fixed effects, it is easy to check that $\pi_{i}^{*}h_{i}$ is linear if $\pi_{i}^{*}$ is linear—and thus the implicit design exists. In this case, we may also compute and verify that the corresponding estimand $\tau_{1}$ takes the form of a difference of two variance-weighted estimands:
$\tau_{1}=\frac{1}{n_{1}}\sum_{i:h_{i}=1}\omega_{1i}(y_{i}(1)-y_{i}(0))-\frac{1}{n_{0}}\sum_{i:h_{i}=0}\omega_{0i}(y_{i}(1)-y_{i}(0))$
such that $\frac{1}{n_{1}}\sum_{i:h_{i}=1}\omega_{1i}=\frac{1}{n_{0}}\sum_{i:h_{i}=0}\omega_{0i}=1$, and $\omega_{1i},\omega_{0i}\propto\pi_{i}^{*}(1-\pi_{i}^{*})$.
Notably, the implicit designs underlying (5) and (5) are different, as (5) involves additional covariates. Since the specifications are nested, one can test $H_{0}:\pi_{i}^{*}$ is linear in $\check{x}_{i}$ against $H_{1}:\pi_{i}^{*}$ is linear in $\tilde{x}_{i}$. A standard Wald test does reject $H_{0}$ in favor of $H_{1}$, providing evidence against the implicit design for (5). In the remainder of the section, we focus on (5), as well as the version of (5) that includes $\tilde{x}_{i}$ as the covariates:
$Y_{i}=\tau W_{i}+\tilde{x}_{i}^{\prime}\mu+\epsilon_{i}.$
Specifications (5) and (5) in turn have the same implicit design.
Figure 1. The distribution of the estimated implicit design for specification (5). 56 out of 786 observations (or 7% of the observations) are outside $[0,1]$.
Figure 2. Calibration of the implicit design. This is a binned scatterplot of $W_{i}$ on $\hat{\pi}_{i}$, with associated pointwise confidence intervals and uniform confidence bands (Cattaneo et al., 2024).
Notes.
“Variance-weighted” refers to weighting with weights proportional to $\pi_{i}(1-\pi_{i})$. “Trimmed” refers to weighting with weights proportional to $\mathbbm{1}(\pi_{i}\in[0.02,0.98])$ (Crump et al., 2009). In either weighting scheme, the weights for $\pi_{i}\not\in[0,1]$ are set to zero. “Patched” treats the implicit design as an estimated propensity score and recalibrates it by binning $\hat{\pi}_{i}$ and computing the empirical treated frequency within each bin. We subclassify on the propensity scores following Chapter 17 in Imbens and Rubin (2015). The binning in the subclassification uses the data-driven procedure in Imbens and Rubin (2015), which recursively partitions the estimated propensity scores until either bins are too small or the mean propensity score is similar among treated and untreated units within a bin. “AIPW” refers to augmented inverse propensity weighting; it uses a logit specification for the propensity score and a linear specification for the outcome means. Both specifications use the same set of covariates as (5). Finally, “regression” estimates (5) and (5) with ordinary least-squares, clustering standard errors by village following Blakeslee et al. (2020) (The estimates in Blakeslee et al. (2020) use weighted least squares with unit-level sampling weights; since their footnote 7 reports that their results are insensitive to these weights, we ignore them and report ordinary least-squares estimates). ∎
Figure 3. Alternative coefficient estimates for $\tau$ in (5) and $\tau_{1}$ in (5)
[…]
These diagnostics on the implicit design undermine a quasi-experimental interpretation of the specification (5), answering (MC1) in the negative. However, at least as an illustration, it may be interesting to understand the regression specification further. Even ignoring the various deficiencies of $\hat{\bm{\pi}}$, we may still be concerned with various choices in the implicit estimand. For instance, the specification (5) includes variance weighting, and we may ask—holding the propensity model fixed—how much the variance weighting alters the estimate. The estimand $\tau_{1}$ in (5) is a difference of two variance-weighted estimands. A priori, we cannot rule out that the difference is driven by the weighting scheme compared to the difference in average treatment effects. The remainder of this section investigates these questions.
We can simply compute a battery of alternative estimates, using the implicit design. These are shown in Figure 3 for both the treatment effect $\tau$ in (5) and the treatment effect difference $\tau_{1}$ in (5). First, the simplest assessment is whether the units with out-of-bounds implicit design feature prominently in the regression estimate. To that end, the variance-weighted estimates $\bm{\times}$ uses the same implicit design targets the same estimand, but removes the out-of-bounds units. For both $\tau_{0}$ and $\tau_{1}$, these $\bm{\times}$ estimates are almost identical to the regression estimates, indicating that the regression estimates put relatively little weight on out-of-bounds units. The estimates $\star$ patches the implicit design by treating it as an estimated propensity score and recalibrating it, as recommended by Zhao and Ding (2022); van der Laan et al. (2024); Imbens and Rubin (2015). This does not meaningfully alter the estimate.
Second, we may assess whether the variance-weighting in the implicit estimand matters by considering weighting schemes that treat units more equally. Since the propensity score estimates are often close or equal to zero and one, overlap violations make estimating the average treatment effect infeasible. Thus, we trim the propensity scores to $[0.02,0.98]$ and construct corresponding estimators for the trimmed average treatment effect (Crump et al., 2009). These estimates—especially the patched estimates $\star$ [^36] The patched implicit design has no $\hat{\pi}_{i}^{\text{patched}}$ within $(0,0.02)\cup(0.98,1)$; thus the trimming only trims out those with overlap failure.—are more different from the regression estimates, though not substantially so compared to sampling noise. On average, the difference between $\star$ and the regression estimates is about 0.5–0.7 times the standard error of the regression estimate.
Finally, moving entirely away from the implicit designs in the regression, we also compute estimates by augmented inverse propensity weighting ($\bm{A}$IPW) by using a simple logit model for the propensity scores and a linear model for the outcome means. These alternative estimates are again similar to the regression estimates, indicating that the outcomes in this application are not so adversarially configured: The implicit design, while clearly rejected, nevertheless produces estimates that are similar to alternative estimates.
(a) Treatment effect $\tau$ in (5) Refer to caption
(b) Heterogeneous treatment effect $\tau_{1}$ in (5) Refer to caption
Figure 4. The correlations of outcomes and implicit designs
Why do alternative weighting schemes not make a difference? Figure 4(a) partitions the implicit designs into 7 bins by quantile, so that each bin within $[0,1]$ contains the same number of units. On each bin, it displays the mean treated and mean untreated outcomes as well as the weight placed on each bin by the implicit estimand. As the bin size become small, computing the difference between the blue and orange curves, weighted by the weights, approximates the regression estimate. As expected, the regression upweights units with implicit design closer to 0.5. In particular, the out-of-bounds units receive a total of about $-3\%$ in weights, explaining their apparent irrelevance to the bottom line. Remarkably, for agricultural employment, the treatment effects hardly vary by the implicit design. Households with low and high propensity to lose water access appear to be affected similarly, with about 10% fewer members engaged in agricultural work, suggesting borewell failure is orthogonal to vulnerability to water loss.
Figure 4(b) considers the analogous decomposition for $\tau_{1}$. [^37] In Figure 4(b), if everyone is treated in a bin, we treat the mean control outcome as zero, and vice versa. Thus the “treatment effects” for $\hat{\pi}_{i}\leq 0$ represent negative mean control outcomes, and the “treatment effects” for $\hat{\pi}_{i}>1$ represent mean treated outcomes. The weighting schemes for $h_{i}=1$ versus $h_{i}=0$ are indeed different in $\tau_{1}$: High employment area ($h_{i}=1$) puts larger weight for households more likely to lose water access—peaking at the bin $(0.605,0.719]$ as opposed to at $(0.435,0.605]$. Thus the apparent treatment effect heterogeneity reflects in part the difference in weighting. But since the differences in conditional average treatment effects are effectively constant and zero, the different weighting again makes little difference to the bottom line estimate.
Our exercise demonstrates a few ways to use implicit designs and estimands followin Section 2.3. One can examine the distribution and calibration properties of the estimated implicit design $\hat{\bm{\pi}}$ (Figures 1 and 2). One can use $\hat{\bm{\pi}}$ to construct implement alternative estimands that maintain the implicit model of treatment assignment (Figure 3). One can also examine how outcomes covary with observed selection (Figure 4). These exercises add transparency and evaluate robustness of the regression estimates, especially if practitioners, like Blakeslee et al. (2020), stake a claim on quasi-experimental causal identification.
6. Conclusion
Linear regressions are ubiquitous. Interpreting their results as causal, thanks to random assignment, is similarly commonplace. This paper studies the necessary implications of so interpreting. We do so by studying the comparisons that regression estimands make under random assignment. For these comparisons to admit a causal interpretation—in the absence of assumptions on the outcomes—the true process of treatment assignment must satisfy certain restrictions. These restrictions are linear in the design and often exactly pin down the distribution of treatment. The resulting set of designs that satisfy these restrictions can be thought of as models of treatment assignment that the regression implicitly estimates. Each design also pinpoints a corresponding estimand that the regression implicitly chooses.
Understanding quasi-experimental interpretation of regressions in this way essentially reduces to mechanical computations that can be scaled and automated. These computations can aid in examining new theoretical properties of particular specifications, itself the subject of a highly influential recent literature. In several theoretical vignettes, these computations unify and strengthen disparate strands of the literature. Additionally, we find that regressions with interactions and with two-way fixed effects have fragile design-based interpretations. This calls for caution and nuance when using them and presenting their results.
Directly computing implicit designs and estimands in practice provides a set of simple diagnostics for practitioners who wish to understand the quasi-experimental properties of a given regression. Doing so makes transparent the statistical and economic choices masked by a regression specification. Having opened up the black box, we can examine each of its components: e.g., evaluating whether the implicit design is plausible, assessing whether the regression targets an economically interesting estimand, and constructing estimates for alternative estimands. Additionally, making these implicit choices transparent may nudge practitioners to wrest the wheel from regression methods and choose methods that are more natively causal.
Appendix A Treating the Gram matrix as known
This section discusses the downstream ramification of treating $G_{n}$ as known in (i) in Section 2.1. This is subtle, since $G_{n}$ itself depend on the true design $\bm{\pi}^{*}$. We can think of (2.1) as a set of nonlinear equations in $\pi_{i}^{*}$:
$\pi_{i}^{*}\rho_{i}(1;\bm{\pi}^{*})+(1-\pi_{i}^{*})\rho_{i}(0;\bm{\pi}^{*})=0$
where $\rho_{i}(\cdot;\bm{\pi}^{*})$ depends on $\bm{\pi}^{*}$ through $G_{n}$. These equations then define a nonlinear surface of designs $\left\{{\bm{\pi}:\pi_{i}\rho_{i}(1;\bm{\pi})+(1-\pi_{i})\rho_{i}(0;\bm{\pi})=0\,\forall\,i}\right\}.$ This surface is defined solely as a function of the regression $(\Lambda,z(x_{i},\cdot))$ and thus does not reflect any information in the data.
The implicit designs can then be thought of as choosing a point near [^38] When a implicit design generates the same Gram matrix as the true design $G_{n}(\bm{\pi})=G_{n}(\bm{\pi}^{*})$, then the implicit design is on the surface. this surface that is “most consistent” with the data by fixing $(G_{n},\rho_{i}(\cdot))$ and solving the corresponding equations. For the regression (2.1), this surface is the span of the column space of the covariates
$\left\{{\bm{\pi}:\pi_{i}=\tilde{x}_{i}^{\prime}\delta,\delta\in\mathbb{R}^{\dim(x_{i})+1}}\right\};$
the implicit design chooses the point on the surface by projecting $\bm{\pi}^{*}$ to the covariates, since all other points are ruled out by large datasets. When $\bm{\pi}^{*}$ belongs to the surface, doing this exactly recovers $\bm{\pi}^{*}$.
Additionally, some joint distributions of treatment is such that $\hat{G}_{n}=G_{n}$ almost surely. When treatment arises from such a joint distribution, there is no estimation noise to ignore. We term those fixed Gram designs and discuss them in the following remark:
Remark A.1 (Fixed Gram designs).
Let $\Pi^{*}$ denote the joint distribution of $\mathbf{W}_{1},\ldots,\mathbf{W}_{n}$. For some $\Pi^{*}$, it is possible that $\hat{G}_{n}=G_{n}(\bm{\pi}^{*})$ with probability one. Under such a design, there is no estimation noise to abstract away. As an example, suppose that we consider $Y_{i}=\alpha+\tau W_{i}+\epsilon_{i}$ with binary $W_{i}$, but $W_{i}$ is assigned in a completely randomized experiment. [^39] That is, there are $n_{1}$ treated units, selected uniformly at random from the set of $\binom{n}{n_{1}}$ subsets. Then, the sample and the population Gram matrices are equal and fixed in every realization of the treatment,
$\hat{G}_{n}=G_{n}=\begin{bmatrix}1&n_{1}/n\\ n_{1}/n&n_{1}/n\end{bmatrix}.$
We term these designs fixed Gram designs for a particular specification. Our results in the population can alternatively be read as exact results for fixed Gram designs. [^40] Regression estimators in the design-based statistical inference literature are frequently analyzed under fixed Gram designs, as they are unbiased for the estimand (see, e.g., Rambachan and Roth, 2020; Athey and Imbens, 2022; Zhao and Ding, 2022; Neyman, 1923/1990; Lin, 2013). $\blacksquare$
Appendix B Estimating potential weights
This section ties up the loose ends in the sampling uncertainty for $G_{n}$. Naturally, we estimate potential weights and implicit designs by plugging in the sample Gram matrix $\hat{G}_{n}$. When we fail to find valid estimated implicit designs, we may wonder whether that is due to sampling noise in $\hat{G}_{n}$ rather than the fault of the regression specification. When we do find a valid estimated implicit design, we may also ask whether $\hat{\bm{\pi}}$ is close to $\bm{\pi}$. To answer these questions, we provide a set of simple estimation and inference guarantees. [^41] Since we view the estimated implicit designs and potential weights as diagnostics and sanity checks, we do not provide particularly strong or optimal guarantees and leave those to future work. In our view, if a practitioner fails to find valid estimated implicit designs, even if that is due to noise, it should nevertheless raise alarms for the interpretation of the regression estimator.
In terms of inference, conceptually, for $\hat{\mathcal{G}}_{n}$ a $(1-\alpha)$-confidence set for $G_{n}$, we may form corresponding confidence sets ^$\hat{\mathcal{R}}_{i}$ for $\bm{\rho}_{i}$ and $\hat{\mathcal{P}}$ for $\bm{\pi}^{*}$ by projection:
^$\hat{\mathcal{R}}_{i}$ =$=\left\{{\Lambda G_{n}^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}:G_{n}\in\hat{\mathcal{G}}_{n}}\right\}$
$\hat{\mathcal{P}}=\bigg{\{}\bm{\pi}:\,\,$ $G\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\in\hat{\mathcal{G}}_{n}$ ($\bm{\pi}$ generates $G_{n}$)
$\forall i,\,\pi_{i}\in\Delta(\mathcal{W})$ (Proper)
$\forall i,\,\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\Lambda G^{-1}\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}=0\bigg{\}}.$ (Level irrelevance)
It is easy to see that, when $\bm{\pi}^{*}$ satisfies (1), $\operatorname{P}[\bm{\pi}^{*}\in\hat{\mathcal{P}}]=\operatorname{P}[G_{n}(\bm{\pi}^{*})\in\hat{\mathcal{G}}_{n}]\geq 1-\alpha$. $\hat{\mathcal{G}}_{n}$ is readily available from, e.g., a simultaneous confidence set for all entries of $G_{n}$. These confidence sets are usually available upon verification of high-probability bounds for $\lVert\hat{G}_{n}-G_{n}\rVert$ or central limit theorems for $\sqrt{n}(\hat{G}_{n,jk}-G_{n,jk})$. Unfortunately, fully computing $\hat{\mathcal{P}}$ is a difficult nonlinear optimization problem. Nonetheless, checking whether a candidate $\bm{\pi}\in\hat{\mathcal{P}}$ is straightforward. Thus, if the researcher has a particular design in mind—perhaps in response to the red flags raised by estimated implicit designs—they can check whether such a design is a plausible justification for the specification in question.
[…]
  1. (1)
    The estimated potential weights $\hat{\bm{\rho}}_{i}(\mathbf{w})$ are consistent for $\bm{\rho}_{i}(\mathbf{w})$.
  2. (2)
    In many cases, the consistency of potential weights implies consistency of estimated implied designs and estimands, though the precise statement depends on further regularity conditions. A key obstruction to a general result is that small perturbations to coefficients in linear equations may result in large perturbations to their solutions.
To discuss asymptotics, we embed our finite population in a sequence of populations. Formally, let $\mathbf{y}_{1}(\cdot),\mathbf{x}_{1},\mathbf{y}_{2}(\cdot),\mathbf{x}_{2},\ldots$ denote a sequence of potential outcomes and covariates. Let $\Pi_{n}^{*}$ denote the joint distribution of treatment assignments for the first $n$ units. We shall consider the behavior as $n\to\infty$, relative to these sequences. We impose a high-level condition on the sequence of populations such that the estimated Gram matrix is consistent for the population Gram matrix. [^42] Assumption B.1 holds whenever the underlying populations are such that a law of large numbers hold. If the treatments are independently assigned, for instance, then Assumption B.1 is true under standard laws of large number (e.g., Theorem 2.2.6 in Durrett (2019)), which we state in Lemma D.4. Section D.3.1 verifies Assumption B.1 for designs that involve sampling without replacement (Hájek, 1964; Rambachan and Roth, 2020).
Assumption B.1.
[…]
Assumption B.2.
The sequence of population Gram matrices $G_{n}$ is such that their minimum eigenvalues are bounded below: For some $\epsilon>0$, for all $n$, $\lambda_{\min}(G_{n})\geq\epsilon>0$.
Assumption B.2 makes sure that the population regression specification is strongly identified, in the sense that the Gram matrix is bounded away from singularity. Under Assumption B.1 and Assumption B.2, we can show that the estimated potential weights are consistent for their population counterparts.
Proposition B.1.
[…]
where $|\cdot|_{\infty}$ takes the entrywise maximum absolute value.
Naturally, we define the set of estimated implicit designs as the set of solutions to the sample counterpart of (2.2): $\hat{\bm{\pi}}$ is an estimated implicit design if for every $i$,
$\sum_{\mathbf{w}\in\mathcal{W}}\hat{\pi}_{i}(\mathbf{w})\hat{\bm{\rho}}_{i}(\mathbf{w})=0,\quad\sum_{\mathbf{w}\in\mathcal{W}}\hat{\pi}_{i}(\mathbf{w})=1.$
Unfortunately, in general, the estimated implicit designs are not always well-behaved. This is because they are the solutions to a system of overdetermined linear equations with noisy coefficients, and small perturbations to the coefficients can result in large changes in the solutions—or even to their existence in the first place. [^44] For instance, it is possible for noise in the coefficients of a underdetermined system to turn the system into an overdetermined one. Thus, it is possible that a population implicit design exists—meaning that (2.2) is under or exactly determined—but no estimated implicit designs do. Failing a general result, we present weaker consistency results that apply in at least in the cross-section ($T=1$). [^45] Proposition D.5 shows that if an estimated implicit design $\hat{\bm{\pi}}$ exists with bounded entries, then it is consistent in the sense that it approximately solves the population level independence equations: $\sum_{\mathbf{w}}\hat{\pi}_{i}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})$ tends to zero.
In particular, if the population implicit design is unique and equal to $\bm{\pi}^{*}$, the estimated implicit designs are consistent for $\bm{\pi}^{*}$. We illustrate with the following result in the cross-sectional case ($T=1$) where the system of equations (2.2) is exactly determined ($k=J$). To state this result, let $R_{i}=[\bm{\rho}_{i}(0),\ldots,\bm{\rho}_{i}(J)]\in\mathbb{R}^{k\times(J+1)}$ be the matrix whose columns are potential weights at each treatment level. Let $R_{i}=\sum_{r=1}^{J}\sigma_{ir}u_{ir}v_{ir}^{\prime}$ be its singular value decomposition, for $\sigma_{1i}\geq\ldots\geq\sigma_{iJ}\geq 0$, $u_{i1},\ldots,u_{iJ}$ orthonormal vectors in $\mathbb{R}^{J}$ and $v_{i1},\ldots,v_{iJ}$ orthonormal vectors in $\mathbb{R}^{J+1}$.
Proposition B.2.
Assume that $k=J$. Suppose that:
  1. (i)
    The estimand $\bm{\tau}$ is a vector of causal contrasts in the sense of 2.2 under $\bm{\pi}^{*}$.
  2. (ii)
    Given $\rho_{i}(\cdot)$, the smallest singular value of $R_{i}$ is uniformly bounded away from zero: For some $\epsilon>0$, $\liminf_{n\to\infty}\min_{i\in[n]}\sigma_{iJ}>\epsilon>0.$
  3. (iii)
    The estimated potential weights are consistent in the sense of (B.1).
Then, for some $C>0$, a unique estimated implicit design $\hat{\bm{\pi}}$ that is bounded by $C$ (that is, $\max_{i}\max_{\mathbf{w}\in\mathcal{W}}|\hat{\pi}_{i}(\mathbf{w})|\leq C$) exists with probability tending to one. Let $\hat{\bm{\pi}}$ be an estimated implicit design if it exists, and otherwise let $\hat{\bm{\pi}}$ be an arbitrary probability vector. Then $\hat{\bm{\pi}}$ is consistent for $\bm{\pi}^{*}$:
[…]
The key condition in Proposition B.2 is (ii), which ensures that the solution to the population level-irrelevance restrictions (2.2) is unique and robust to small perturbations of the potential weights. In the binary treatment, cross-sectional setting, (ii) is equivalent to that the potential weights are bounded away from zero: $\rho_{i}(1)^{2}+\rho_{i}(0)^{2}>\epsilon^{2}$. Lemma D.8 further provides sufficient conditions for (ii) to hold in the $T=1$ case.
Appendix C Proofs
See 2.1
Proof.
We first show that 2.2 is equivalent to the design satisfying (2.2) for a population regression estimand. By (2.7), $\bm{\tau}$ satisfies 2.2 if and only if for all $i,t$
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})c_{it}=0\iff\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=0.$
The sum-to-one condition is automatically satisfied by $\pi_{i}^{*}$. This is then equivalent to (2.2). This shows that $\bm{\tau}$ is minimally causal if and only if $\bm{\pi}^{*}$ is an implicit design. $\bm{\pi}^{*}$ is proper and generates $G_{n}$ by definition. The statement that $\bm{\pi}^{*}$ is an implicit design is equivalent to (1) and (2).
When $\bm{\tau}$ is minimally causal, the estimand $\bm{\tau}$ is by definition the implicit estimand under $\bm{\pi}^{*}$. This proves the second statement. ∎
See 2.2
Proof.
Applying Theorem 2.1, we need to show that 2.2(1) is equivalent to Theorem 2.1(1). Some proper implicit design exists if and only if $\rho_{i}(1)\rho_{i}(0)\leq 0$ for all $i$. When it exists, it satisfies $\pi_{i}=-\rho_{i}(0)/(\rho_{i}(1)-\rho_{i}(0))$. Thus 2.2(1) is equivalent to Theorem 2.1(1). In turn, Theorem 2.1(2) is equivalent to 2.2(2). ∎
See B.1
Proof.
Take $\eta>0$. For the result with fixed $i$, it suffices to show that $\operatorname{P}[\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}>\eta]\to 0$. Note that
$\operatorname{P}\left({\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}>\eta}\right)$ $\leq\operatorname{P}[\lambda_{\min}(\hat{G}_{n})\leq\epsilon/2]$
$+\operatorname{P}\left[{\lVert\Lambda(\hat{G}_{n}^{-1}-G_{n}^{-1})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\rVert_{F}>\eta,\lambda_{\min}(\hat{G}_{n})>\epsilon/2}\right].$
The first term converges to zero by Assumptions B.1 and B.2 and the Hoffman–Wielandt inequality. Note that, by the submultiplicativity of the Frobenius norm, when $\lambda_{\min}(\hat{G}_{n})>\epsilon/2$,
[…]
Thus, for some $C$ dependent on $\epsilon$, $\Lambda$, and $\lVert z_{t}(\mathbf{x}_{i},\mathbf{w})\rVert_{\infty}$,
$\operatorname{P}\left[{\lVert\Lambda(\hat{G}_{n}^{-1}-G_{n}^{-1})\mathbf{z}(\mathbf{x}_{i},\mathbf{w})^{\prime}\rVert_{F}>\eta,\lambda_{\min}(\hat{G}_{n})>\epsilon/2}\right]\leq\operatorname{P}\left[{C\lVert G_{n}-\hat{G}_{n}\rVert_{F}>\eta}\right]\to 0.$
[…]
To prove convergence that is uniform in $i$, note that, if $\lambda_{\min}(\hat{G}_{n})>\epsilon/2$, up to constants,
$\max_{i}|\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})|_{\infty}$ $\lesssim\max_{i}\lVert\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})\rVert_{F}$
$\lesssim\max_{i}\lVert\mathbf{z}(\mathbf{x}_{i},\mathbf{w})\rVert_{F}\cdot\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}$
$\lesssim\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}\cdot\max_{i}\max_{t}\lVert z_{t}(\mathbf{x}_{i},\mathbf{w})\rVert_{\infty}$
$\lesssim\lVert\hat{G}_{n}^{-1}-G_{n}^{-1}\rVert_{F}$
$\lesssim\lVert G_{n}-\hat{G}_{n}\rVert_{F}.$
Therefore,
$\operatorname{P}\left({\max_{i}|\hat{\bm{\rho}}_{i}(\mathbf{w})-\bm{\rho}_{i}(\mathbf{w})|_{\infty}>\eta}\right)$ $\leq\operatorname{P}[\lambda_{\min}(\hat{G}_{n})\leq\epsilon/2]+\operatorname{P}\left[{\lVert G_{n}-\hat{G}_{n}\rVert_{F}\gtrsim\eta}\right]\to 0.$
See B.2
Proof.
Let $\hat{R}_{i}=[\hat{\bm{\rho}}_{i}(0),\ldots,\bm{\rho}_{i}(J)]$ be the sample analogue of $R_{i}$. Both $\hat{R}_{i}$ and $R_{i}$ may depend on $n$. Assume $n$ is sufficiently large such that $\min_{i\in[n]}\sigma_{iJ}>\epsilon/2>0$. Thus, each $R_{i}$ is a $J\times(J+1)$ matrix with rank $J$ such that $R_{i}\pi_{i}^{*}=0$. As a minor abuse of notation, let $\pi_{i},\pi_{i}^{*}$ be represented as vectors in $\mathbb{R}^{J+1}$. Let $u_{i}=\pi_{i}^{*}/\lVert\pi_{i}^{*}\rVert$ be the unit vector in the (one-dimensional) null space of $R_{i}$. Note that $11/(J+1)\leq\lVert\pi_{i}^{*}\rVert_{\infty}\leq\lVert\pi_{i}^{*}\rVert\leq\lVert\pi_{i}^{*}\rVert_{1}=$ 1, and thus $1^{\prime}u_{i}=1/\lVert\pi_{i}^{*}\rVert\in[1,J+1]$ is bounded above and below.
Fix some $C$ to be chosen. Let $E$ be complement to the event that there exists a unique bounded estimated implicit design. Note that $E=\bigcup_{i\in[n]}E_{i}$, where each $E_{i}$ is the complement to the event that $\hat{R}_{i}\hat{\pi}_{i}=0$ has a unique solution $\hat{\pi}_{i}$ with $11^{\prime}\hat{\pi}_{i}=$ 1 where $\lVert\hat{\pi}_{i}\rVert_{\infty}\leq C$. Let $\hat{\sigma}_{i1}\geq\cdots\geq\hat{\sigma}_{iJ}\geq 0$ be the singular values of $\hat{R}_{i}$. Suppose $\hat{\sigma}_{iJ}>0$, then $\hat{R}_{i}$ has full rank and the equation $\hat{R}_{i}\hat{\pi}_{i}=0$ has a unique one-dimensional space of solutions. If that space exists, let $\hat{u}_{i}$ be a unit vector in that space, unique up to sign. If $1^{\prime}\hat{u}_{i}\neq 0$, then we choose the sign of $\hat{u}_{i}$ so that $01^{\prime}\hat{u}_{i}>$ 0. Note that if $|\hat{u}_{i}^{\prime}1|>1/C$, then $\hat{\pi}_{i}=\hat{u}_{i}/\hat{u}_{i}^{\prime}1$ is a bounded estimated implicit design with $\lVert\hat{\pi}_{i}\rVert_{\infty}\leq C$.
Note that
$E_{i}\subset\left\{{\hat{\sigma}_{iJ}\leq\epsilon/4}\right\}\cup\left\{{\hat{\sigma}_{iJ}>\epsilon/4,|\hat{u}_{i}^{\prime}1|<1/C}\right\}.$
We will show that $E_{i}\subset\left\{{\lVert\hat{R}_{i}-R_{i}\rVert_{F}>\kappa}\right\}$ for some $\kappa$. For the event $\left\{{\hat{\sigma}_{iJ}\leq\epsilon/4}\right\}$, the Hoffman–Wielandt inequality implies that
$|\hat{\sigma}_{iJ}-\sigma_{iJ}|^{2}\leq\sum_{k=1}^{J}|\hat{\sigma}_{ik}-\sigma_{ik}|^{2}\leq\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2},$
and thus $\hat{\sigma}_{iJ}\geq\epsilon/2-\lVert R_{i}-\hat{R}_{i}\rVert_{F}$, which further implies that $\left\{{\hat{\sigma}_{iJ}<\epsilon/4}\right\}\subset\left\{{\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\epsilon/4}\right\}.$
Let $\hat{u}_{i}=\hat{c}_{i}u_{i}+\sqrt{1-\hat{c}_{i}^{2}}\hat{u}_{\perp,i}$ where $\hat{c}_{i},\hat{u}_{\perp,i}$ are uniquely chosen so that $\hat{u}_{\perp,i}$ is a unit vector orthogonal to $u_{i}$. Then
$|\hat{u}_{i}^{\prime}1|\geq|\hat{c}_{i}1^{\prime}u_{i}|-\sqrt{1-\hat{c}_{i}^{2}}|1^{\prime}\hat{u}_{\perp,i}|\geq|\hat{c}_{i}|-\sqrt{1-\hat{c}_{i}^{2}}\sqrt{J+1}.$
Now, note that
$\lVert R_{i}\hat{u}_{i}\rVert=\lVert(R_{i}-\hat{R}_{i})\hat{u}_{i}\rVert\leq\lVert R_{i}-\hat{R}_{i}\rVert_{F}.$
On the other hand,
$\lVert R_{i}\hat{u}_{i}\rVert=\sqrt{1-\hat{c}_{i}^{2}}\lVert R_{i}\hat{u}_{\perp,i}\rVert\geq\sqrt{1-\hat{c}_{i}^{2}}\sigma_{iJ}\geq\sqrt{1-\hat{c}_{i}^{2}}\frac{\epsilon}{2},$
and thus
$1-\hat{c}_{i}^{2}\leq\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}\iff|\hat{c}_{i}|\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}}.$
Hence,
$|\hat{u}_{i}^{\prime}1|\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert\hat{R}_{i}-R_{i}\rVert_{F}^{2}}-\frac{2\sqrt{J+1}}{\epsilon}\lVert\hat{R}_{i}-R_{i}\rVert_{F}.$
Suppose $\lVert R_{i}-\hat{R}_{i}\rVert_{F}<\epsilon/(4\sqrt{J+1})$, then
$|\hat{u}_{i}^{\prime}1|\geq\sqrt{1-\frac{1}{J+1}}-\frac{1}{2}\geq\sqrt{1/2}-1/2\geq 0.207.$
Thus, if we pick $C>1/0.207$, then
$E_{i}\subset\left\{{\hat{\sigma}_{iJ}<\epsilon/4}\right\}\cup\left\{{\hat{\sigma}_{iJ}>\epsilon/4,|\hat{u}_{i}^{\prime}1|<1/C}\right\}\subset\left\{{\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\frac{\epsilon}{4\sqrt{J+1}}}\right\}.$
As a result,
$\operatorname{P}\left[{\bigcup_{i}E_{i}}\right]\leq\operatorname{P}\left[{\max_{i\in[n]}\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\frac{\epsilon}{4\sqrt{J+1}}}\right]\to 0$
by assumption. This proves the first part regarding the existence of a unique bounded implicit design.
For the second part, note that when $\hat{\sigma}_{iJ}>0$,
$\hat{\pi}_{i}-\pi_{i}=\frac{\hat{u}_{i}}{1^{\prime}\hat{u}_{i}}-\frac{u_{i}}{1^{\prime}u_{i}}=\frac{1^{\prime}}{1^{\prime}\hat{u}_{i}}(u_{i}-\hat{u}_{i})\frac{\hat{u}_{i}}{1^{\prime}u_{i}}+\frac{1}{1^{\prime}u_{i}}(\hat{u}_{i}-u_{i}).$
Thus
$\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}\leq\frac{1}{0.207}\sqrt{J+1}\lVert u_{i}-\hat{u}_{i}\rVert_{2}+\lVert\hat{u}_{i}-u_{i}\rVert_{2}$
if $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\epsilon/(4\sqrt{J+1})$.
To bound $\lVert\hat{u}_{i}-u_{i}\rVert$, we first show that $\hat{c}_{i}>0$ when $R_{i}$ is close to $\hat{R}_{i}$. When $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, we note that by (C),
$|\hat{c}_{i}|\geq\sqrt{1-\frac{1}{4(J+1)}}\geq\sqrt{7/8}.$
When $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, we also have that, because we pick $\hat{u}_{i}$ to have $1^{\prime}\hat{u}_{i}\geq 0$,
$0.207\leq 1^{\prime}\hat{u}_{i}=\hat{c}_{i}+\sqrt{1-\hat{c}_{i}^{2}}1^{\prime}\hat{u}_{\perp,i}\leq\hat{c}_{i}+\sqrt{1-\hat{c}_{i}^{2}}\sqrt{J+1}\leq\hat{c}_{i}+\frac{2}{\epsilon}\frac{\epsilon}{4\sqrt{J+1}}\sqrt{J+1}=\hat{c}_{i}+\frac{1}{2}.$
Thus $\hat{c}_{i}>0.207-0.5$. Since $|\hat{c}_{i}|>\sqrt{7/8}$, we conclude that $\hat{c}_{i}>\sqrt{7/8}>0$. Thus, when $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$, (C) implies
$c_{i}\geq\sqrt{1-\frac{4}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}}\geq 1-\frac{4}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}.$
Now, when $\lVert R_{i}-\hat{R}_{i}\rVert_{F}\leq\frac{\epsilon}{4\sqrt{J+1}}$,
$\lVert\hat{u}_{i}-u_{i}\rVert_{2}^{2}$ $\leq(1-\hat{c}_{i})^{2}+(1-\hat{c}_{i}^{2})$
$=2-2\hat{c}_{i}\leq\frac{8}{\epsilon^{2}}\lVert R_{i}-\hat{R}_{i}\rVert_{F}^{2}$
Therefore, for some constant $C(J,\epsilon)$,
$\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}\leq C(J,\epsilon)\lVert R_{i}-\hat{R}_{i}\rVert_{F}.$
Hence, for $\eta>0$,
$\left\{{\max_{i\in[n]}\lVert\hat{\pi}_{i}-\pi_{i}\rVert_{2}>\eta}\right\}\subset\left\{{\max_{i\in[n]}\lVert R_{i}-\hat{R}_{i}\rVert_{F}>\max\left({\frac{\epsilon}{4\sqrt{J+1}},\frac{1}{C(J,\epsilon)}}\right)}\right\}$
The probability on the right-hand side converges to zero by assumption. This proves the convergence of $\hat{\pi}_{i}$ to $\pi_{i}$ uniformly in $i$ in $\lVert\cdot\rVert_{2}$. Since $\lVert\cdot\rVert_{\infty}\leq\lVert\cdot\rVert_{2}$ in $\mathbb{R}^{J+1}$, this concludes the proof. ∎
See 3.1
Proof.
We prove the following:
  1. (1)
    By Theorem D.1, the potential weights of this regression are the same as those for the Frisch–Waugh–Lovell-transformed regression of $Y_{i}$ on $W-\delta^{\prime}x_{i}$. For the FWL-transformed regression, its Gram matrix is a positive scalar
    $\tilde{G}_{n}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\delta^{\prime}x_{i})^{2}+(1-\pi_{i}^{*})(\delta^{\prime}x_{i})^{2}>0$
    Thus the potential weights are
    $\rho_{i}(w)=\tilde{G}_{n}^{-1}(w-x_{i}^{\prime}\delta).$
    Since $\tilde{G}_{n}>0$, $\rho_{i}(1)\neq\rho_{i}(0)$ for all $i$.
    1. (a)
      By 2.2, the implicit design is equal to
      $\pi_{i}=-\frac{\rho_{i}(0)}{\rho_{i}(1)-\rho_{i}(0)}=x_{i}^{\prime}\delta.$
      To show that $\bm{\pi}$ generates $G_{n}$, note entries of the Gram matrix of the original regression take the form
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}z_{k}(x_{i},1)z_{\ell}(x_{i},1)+(1-\pi_{i})z_{k}(x_{i},0)z_{\ell}(x_{i},0)$
      where $z_{k}(x_{i},w)z_{\ell}(x_{i},w)$ take one of the following forms: $1,w,wx_{i\ell},x_{ik},x_{i\ell}x_{ik}$. Thus, it suffices to check that
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}$
      $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}x_{i}.$
      Both are true since $\pi_{i}$ is the projection of $\pi_{i}^{*}$ onto $x_{i}$, which contains a constant.
    2. (b)
      Immediately, $\pi_{i}^{*}=\pi_{i}$ implies $\pi_{i}^{*}=x_{i}^{\prime}\delta$. On the other hand, if $\pi_{i}^{*}=x_{i}^{\prime}\delta$, then the corresponding implicit design is equal to $x_{i}^{\prime}\delta$ by (a).
    3. (c)
      We may compute
      $\omega_{i}=\pi_{i}\rho_{i}(1)=\tilde{G}_{n}^{-1}x_{i}^{\prime}\delta(1-x_{i}^{\prime}\delta).$
      and that
      $\tilde{G}_{n}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{W_{i}^{2}-2W_{i}\pi_{i}+\pi_{i}^{2}}\right]$
      $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}}_{=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}\text{ since $x_{i}$ includes a constant}}-2\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}\pi_{i}+\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{2}$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}-2\frac{1}{n}\sum_{i=1}^{n}(\underbrace{\pi_{i}^{*}-\pi_{i}}_{\text{orthogonal to $\pi_{i}$}}+\pi_{i})\pi_{i}+\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{2}$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(1-\pi_{i}).$
      Lastly, when $\pi_{i}=\pi_{i}^{*}\in(0,1)$, $\omega_{i}>0$ since $\tilde{G}_{n}>0$.
  2. (2)
    By Theorem D.1, the potential weights of this regression are the same as those for the FWL-transformed regression
    $Y_{i}=\sum_{j}\tau_{j}(W_{ij}-\delta_{j}^{\prime}x_{i})+\epsilon_{i}$
    The Gram matrix for this regression is a positive definite matrix whose $k\ell$th entry is:
    $\tilde{G}_{n,k\ell}=\begin{cases}-\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(k)(1-\delta_{k}^{\prime}x_{i})\delta_{\ell}^{\prime}x_{i}+\pi_{i}^{*}(\ell)\delta_{k}^{\prime}x_{i}(1-\delta_{\ell}x_{i})+(1-\pi_{k}^{*}-\pi_{\ell}^{*})(\delta_{k}^{\prime}x_{i})(\delta_{\ell}^{\prime}x_{i}),&\text{if }k\neq\ell\\ \frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(k)(1-\delta_{k}^{\prime}x_{i})^{2}+(1-\pi_{i}^{*}(k))(\delta_{k}^{\prime}x_{i})^{2}&\text{if }k=\ell.\end{cases}$
    As a result, the potential weights are of the form
    $\rho_{i}(j)=\tilde{G}_{n}^{-1}\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)$
    where $e_{j}$ is the $j$th standard basis vector and $e_{0}=0_{J}$ is the vector of zeros.
    1. (a)
      Set $\pi_{i}(0)=1-\sum_{j=1}^{J}\pi_{i}(j)$. The equations (2.2) are then equivalent to
      $\tilde{G}_{n}^{-1}\sum_{j=1}^{J}\pi_{i}(j)\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)=0.$
      Since $\tilde{G}_{n}$ is positive definite, we must then have
      $\sum_{j=1}^{J}\pi_{i}(j)\left({e_{j}-\begin{bmatrix}\delta_{1}^{\prime}x_{i}\\ \vdots\\ \delta_{J}^{\prime}x_{i}\end{bmatrix}}\right)=0$
      The $j$th coordinate of the above equations is of the form
      $0=\pi_{i}(j)-\pi_{i}(j)\delta_{j}^{\prime}x_{i}-\delta_{j}^{\prime}x_{i}\sum_{\ell\in\left\{{0,\ldots,J}\right\}\setminus\left\{{j}\right\}}\pi_{i}(\ell)=\pi_{i}(j)-\delta_{j}^{\prime}x_{i}.$
      The implicit design is thus $\pi_{i}(j)=\delta_{j}^{\prime}x_{i}$.
      The proof for $\pi$ generating $G_{n}$ is analogous to the case in (1)(a).
    2. (b)
      Analogous to (1)(b)
    3. (c)
      Inspecting $\omega_{i}(\bm{\pi},\ell)$, we compute that
      $\omega_{i}(\bm{\pi},\ell)=\tilde{G}_{n}^{-1}\pi_{i}(\ell)\left({e_{\ell}-\begin{bmatrix}\pi_{i}(1)\\ \vdots\\ \pi_{i}(J)\end{bmatrix}}\right)$
      Thus the implicit estimand is
      $\begin{bmatrix}\tau_{1}\\ \vdots\\ \tau_{J}\end{bmatrix}=\frac{1}{n}\sum_{i}\sum_{\ell=0}^{J}\omega_{i}(\bm{\pi},\ell)y_{i}(\ell).$
      For the contamination statement, it suffices to show an example. The setup in Example C.1 features a contaminated estimand.
  3. (3)
    The specification is equivalent to the following specification:
    $Y_{i}=\sum_{\ell=0}^{L}\alpha_{\ell}+\tau_{\ell}W_{i}$
    where $\tau=\left({1-\sum_{\ell=1}^{L}\overline{x}_{\ell}}\right)\tau_{0}+\sum_{\ell=1}^{L}\overline{x}_{\ell}\tau_{\ell}\equiv\sum_{\ell=0}^{L}\overline{x}_{\ell}\tau_{\ell}$.
    For this specification (3), let us order the covariates to be
    $[x_{i0},x_{i0}W_{i},x_{i1},x_{i1}W_{i},\ldots,x_{iL},x_{iL}W_{i}].$
    Note that the population Gram matrix is of the form
    $G_{n}=\begin{bmatrix}G_{0}\\ &G_{1}\\ &&\ddots\\ &&&G_{L}\end{bmatrix}$
    where
    $G_{\ell}=\begin{bmatrix}\overline{x}_{\ell}&\overline{x}_{\ell}\pi(\ell)\\ \overline{x}_{\ell}\pi(\ell)&\overline{x}_{\ell}\pi(\ell)\end{bmatrix}$
    where $\pi(\ell)$ is the mean of $\pi_{i}^{*}$ among those with $c_{i}=\ell$. Since $G_{n}$ is block-diagonal, its inverse is similarly block-diagonal with
    $G_{\ell}^{-1}=\frac{1}{\overline{x}_{\ell}\pi(\ell)(1-\pi(\ell))}\begin{bmatrix}\pi(\ell)&-\pi(\ell)\\ -\pi(\ell)&1\end{bmatrix}$
    on the diagonal. Note that $\Lambda=[0,\overline{x}_{0},0,\overline{x}_{1},\ldots,0,\overline{x}_{L}]$. Thus
    $\Lambda G_{n}^{-1}=\left[{\frac{-\pi(0)}{\pi(0)(1-\pi(0))},\frac{1}{\pi(0)(1-\pi(0))},\ldots,\frac{-\pi(L)}{\pi(L)(1-\pi(L))},\frac{1}{\pi(L)(1-\pi(L))}}\right].$
    For someone with $x_{i}=\ell$, the covariate transform is of the form
    $z(x_{i},w)=[0,\ldots,0,1,w,0,\ldots,0]^{\prime}$
    where it is nonzero at the $\ell$th pair. Thus
    $\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)=\frac{w-\pi(\ell)}{\pi(\ell)(1-\pi(\ell))}.$
    1. (a)
      By 2.2, the only implicit design sets $\pi_{i}=\pi(\ell)$ for $c_{i}=\ell$. This design generates $G_{n}$ by our explicit calculation of the Gram matrix.
    2. (b)
      Analogous to (1)(b).
    3. (c)
      Note that the implicit estimand is described by $\omega_{i}=\pi(\ell)\rho_{i}(1)=1.$ Therefore the implicit estimand is the average treatment effect.
  4. (4)
    The proof for this result is rather long and is delegated to self-contained claims. In the notation setup by Proposition D.9, the potential weights are
    $\rho_{i}(w)=\tilde{G}_{n}^{-1}(w-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{21}^{\prime}(w(x_{i}-\overline{x}_{1})-\Gamma_{01}-\Gamma_{11}(x_{i}-\overline{x}_{1})))$
    where
    $\tilde{G}_{n}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{W_{i}\sim\pi_{i}^{*}}\left[{\left\{{W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{21}^{\prime}(W_{i}(x_{i}-\overline{x}_{1})-\Gamma_{01}-\Gamma_{11}(x_{i}-\overline{x}_{1}))}\right\}^{2}}\right].$
    1. (a)
      The implicit design is shown in Proposition D.10. Proposition D.10 shows that the odds for the implicit design is equal to
      $\pi_{i}/(1-\pi_{i})=a_{0}+a_{1}^{\prime}(x-\overline{x}_{0})$
      where $a_{0},a_{1}$ are population projection coefficients of $\pi_{i}^{*}/(1-\pi_{i}^{*})$ on $x-\overline{x}_{0}$, weighted by $1-\pi_{i}^{*}$. Since it does not matter where we center $x_{i}$ for this regression, as it includes an intercept, we can reparametrize to show that $\pi_{i}/(1-\pi_{i})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$ as described.
    2. (b)
      Immediately, $\pi_{i}^{*}=\pi_{i}$ implies that $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$. On the other hand, if $\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})$, then the corresponding implicit design has odds equal to $\pi_{i}/(1-\pi_{i})=\delta_{0}+\delta_{1}^{\prime}(x-\overline{x})=\pi_{i}^{*}/(1-\pi_{i}^{*})$.
    3. (c)
      We calculate, in the notation of Proposition D.9
      $\tilde{G}_{n}\rho_{i}(1)\pi_{i}=\left\{{(1-\Gamma_{21}^{\prime}(x_{i}-\overline{x}_{1}))-\theta_{01}-\theta_{11}^{\prime}(x_{i}-\overline{x})}\right\}\pi_{i}.$
      By the proof of Proposition D.10,
      $(1-\Gamma_{21}^{\prime}(x_{i}-\overline{x}_{1}))-\theta_{01}-\theta_{11}^{\prime}(x_{i}-\overline{x})$
      does not depend on $i$, since $\Gamma_{21}=-\theta_{11}$. As a result, $\rho_{i}(1)\pi_{i}$ is proportional to $\pi_{i}$.
      It suffices to show that $\frac{1}{n}\sum_{i}\rho_{i}(1)\pi_{i}=1$. Showing this is rather tedious by computing $\tilde{G}_{n}$. Alternatively, we note that $\omega_{i}=\rho_{i}(1)\pi_{i}$ do not depend on individual outcomes. When $\pi_{i}^{*}=\pi_{i}$, the regression estimates $\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))$ by 2.2. If the outcomes are such that the model is well-specified: $y_{i}(w)=vw$, then the regression estimates $\tau=v$ exactly for all $v\in\mathbb{R}$. But for this to happen, we must have
      $\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(y_{i}(1)-y_{i}(0))=v\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=v.$
      As a result $\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=1$.
  5. (5)
    Let $\pi(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$. By Theorem D.1, the potential weights of this regression are equal to those of the regression
    $\mathbf{Y}_{it}=\ddot{\mathbf{W}}_{it}+\epsilon_{it}$
    where
    $\ddot{\mathbf{w}}=\mathbf{w}-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{w}-\sum_{\mathbf{w}_{1}\in\mathcal{W}}\pi(\mathbf{w}_{1})\mathbf{w}_{1}+1_{T}\frac{1_{T}^{\prime}}{T}\sum_{\mathbf{w}_{1}\in\mathcal{W}}\pi(\mathbf{w}_{1})\mathbf{w}_{1}.$
    For the FWL-transformed regression, the Gram matrix $\tilde{G}_{n}$ is a positive scalar by Theorem D.1.
    1. (a)
      Any implicit design $\tilde{\pi}_{i}$ then satisfies
      $0=\tilde{G}_{n}^{-1}\cdot\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{w}}=\tilde{G}_{n}^{-1}\left\{{\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\mathbf{w}-1_{T}\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\frac{1_{T}^{\prime}\mathbf{w}}{T}-\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\mathbf{w}+1_{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\frac{1_{T}^{\prime}}{T}\mathbf{w}}\right\}$
      Note that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$ solves this equation, and so $\pi(\mathbf{w})$ is an implicit design. We now show its uniqueness.
      The preceding display implies that the vector
      $\sum_{\mathbf{w}\in\mathcal{W}}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}\in\operatorname{span}(1_{T})$
      However, the only vector in $\operatorname{span}(1_{T})$ and $\operatorname{span}(\mathcal{W})$ is the zero vector by assumption. Thus $\sum_{\mathbf{w}\in\mathcal{W}}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}=\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}\left({\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w})}\right)\mathbf{w}=0$. Since $\mathcal{W}\setminus\left\{{0}\right\}$ is a linearly independent collection of vectors, we conclude that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$ for all $\mathbf{w}\neq 0$. Since both probability vectors sum to one, we conclude that $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$. Therefore, $\pi(\mathbf{w})$ is the unique implicit design.
      To show that $\pi(\cdot)$ generates $G_{n}$, observe that the Gram matrix of the original regression is equal to
      $G_{n}=\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\mathbf{z}(\mathbf{w})^{\prime}\mathbf{z}(\mathbf{w})=\sum_{\mathbf{w}\in\mathcal{W}}\left\{{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})}\right\}\mathbf{z}(\mathbf{w})^{\prime}\mathbf{z}(\mathbf{w})$
      and thus depends on $\bm{\pi}_{i}^{*}$ only through $\pi(\mathbf{w}).$ Therefore $G_{n}(\bm{\pi})=G_{n}$.
    2. (b)
      Immediately, if $\pi_{i}^{*}$ is equal to the implicit design, then it must be constant for all $i$. Conversely, if $\pi_{i}^{*}$ is constant for all $i$, then the implicit design in (a) is equal to it.
    3. (c)
      For $\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$, define
      $\mathbf{w}_{\pi}=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\mathbf{w}\quad g_{t}(\mathbf{w})=\ddot{\mathbf{w}}_{t}=\mathbf{w}_{t}-\mathbf{w}_{\pi,t}-\frac{1}{T}(1_{T}^{\prime}\mathbf{w}-1_{T}^{\prime}\mathbf{w}_{\pi}).$
      Note that
      $\tilde{G}_{n}=\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g_{t}(\mathbf{w})^{2}.$
      Then by definition,
      $\bm{\omega}_{t}(\bm{\pi},\mathbf{w})=\pi(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=\frac{\pi(\mathbf{w})g_{t}(\mathbf{w})}{\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g_{t}(\mathbf{w})^{2}}.$
      Inspecting Theorem 1(ii) in Athey and Imbens (2022), in their expression, for $\mathbf{w}$ having adoption time $a\in\left\{{2,\ldots,T,\infty}\right\}$, the weight put on the potential outcome $y_{t}(\mathbf{w})$ is equal to
      $\gamma_{t,a}=\frac{\pi(\mathbf{w})g(t,a)}{\sum_{t=1}^{T}\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})g(t,a(\mathbf{w}))^{2}}$
      for the expression in their (3.1)
      $g(t,a(\mathbf{w}))$ $=\left({\mathbbm{1}(a(\mathbf{w})\leq t)-\sum_{\mathbf{w}\in\mathcal{W}:a(\mathbf{w})\leq t}\pi(\mathbf{w})}\right)$
      $\quad+\frac{1}{T}\left({a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)-\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})}\right)+\frac{T+1}{T}(\mathbbm{1}(a=\infty)-\pi(0_{T})).$
      Here, we let $a(\mathbf{w})$ denote the adoption time of a treatment path $\mathbf{w}$.
      Now, observe that
      $\mathbbm{1}(a(\mathbf{w})\leq t)$ $=\mathbf{w}_{t}$
      $\sum_{\mathbf{w}\in\mathcal{W}:a(\mathbf{w})\leq t}\pi(\mathbf{w})$ $=\sum_{\mathbf{w}\in\mathcal{W}}\mathbf{w}_{t}\pi(\mathbf{w})=\mathbf{w}_{\pi,t}$
      $a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)$ =$=\begin{cases}0&\mathbf{w}=0_{T}\\ \text{number of untreated periods $+1$}=T+1-1_{T}^{\prime}\mathbf{w}&\text{ otherwise}.\end{cases}$
      $\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})$ $=T+1-1_{T}^{\prime}\mathbf{w}_{\pi}-(T+1)\pi(0_{T})$
      Therefore, $\left({a(\mathbf{w})\mathbbm{1}(a(\mathbf{w})\leq T)-\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0_{T}}\right\}}a(\mathbf{w})\pi(\mathbf{w})}\right)$
      $=1_{T}^{\prime}\mathbf{w}_{\pi}+\begin{cases}(T+1)(\pi(0_{T})-1)&\mathbf{w}=0_{T}\\ -1_{T}^{\prime}\mathbf{w}+(T+1)\pi(0_{T})&\text{ otherwise}\end{cases}$
      Therefore,
      $g(t,a)=\mathbf{w}_{t}-\mathbf{w}_{\pi,t}-\frac{1}{T}\left({1_{T}^{\prime}\mathbf{w}-1_{T}^{\prime}\mathbf{w}_{\pi}}\right)=g_{t}(\mathbf{w})$
      and these expressions match.∎
Example C.1.
It is simple to analyze Theorem 3.1(2) fully numerically, at least given a concrete data-generating process. To do so, we follow the numerical example in Section 2.2 of Goldsmith-Pinkham et al. (2024). Consider the specification
$Y_{i}=\alpha_{0}+\tau_{1}W_{i1}+\tau_{2}W_{i2}+\gamma^{\prime}x_{i}+\epsilon_{i}$
with binary $x_{i}$ and $J=2$. Suppose exactly half the units have $x_{i}=1$. For the units with $x_{i}=0$, $\pi^{*}_{i}(0)=0.5$, $\pi^{*}_{i}(1)=0.05$, and $\pi^{*}_{i}(2)=0.45$. For the units with $x_{i}=1$, $\pi_{i}^{*}(0)=0.1$, $\pi_{i}^{*}(1)=0.45$, and $\pi_{i}^{*}(2)=0.45$. The coefficients of interest are $\bm{\tau}=[\tau_{1},\tau_{2}]^{\prime}$.
Numerically, we can verify that the assignment probabilities $\bm{\pi}^{*}$ are the only solution to (2.2), and thus $\bm{\pi}^{*}$ is the only implicit design for this regression specification. The corresponding implicit estimand for $\tau_{j}$ is then
$\tau_{j}=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=0}^{J}\underbrace{\pi_{i}(k)\rho_{ij}(k)}_{\omega_{ij}(\bm{\pi},k)}y_{i}(k),$
which is the sum of potential outcomes weighted by $\omega_{ij}(\bm{\pi},k)$.
We can calculate the implicit estimand as well, and inspecting the implicit estimand allows us to recover the contamination bias in Goldsmith-Pinkham et al. (2024). For each $x_{i}$ value and each estimand, we tabulate $\omega_{ij}(\cdot)$ for individuals with those $x_{i}$:
$\tau_{1}$ $\tau_{2}$
$\omega_{i1}(0)$ $\omega_{i1}(1)$ $\omega_{i1}(2)$ $\omega_{i2}(0)$ $\omega_{i2}(1)$ $\omega_{i2}(2)$
$x_{i}=0$ $-140/106$ $41/106$ $99/106$ $-160/106$ $9/106$ $151/106$
$x_{i}=1$ $-72/106$ $171/106$ $-99/106$ $-52/106$ $-9/106$ $61/106$
We find that the implicit estimand for $\tau_{j}$ does not solely involve $y_{i}(0)$ and $y_{i}(j)$, contrary to intuition; moreover, $\tau_{j}$ does not measure the same estimand for individuals with $x_{i}=0$ and $x_{i}=1$. This echoes the result by Goldsmith-Pinkham et al. (2024). In fact, under $\bm{\pi}^{*}$, the estimand $\tau_{1}$ is
$\tau_{1}=\frac{1}{2}\left[{\frac{41}{106}\overline{\tau}_{1\mid x=0}+\frac{99}{106}\overline{\tau}_{2\mid x=0}}\right]+\frac{1}{2}\left[{\frac{171}{106}\overline{\tau}_{1\mid x=1}-\frac{99}{106}\overline{\tau}_{2\mid x=1}}\right]$
where $\overline{\tau}_{k\mid x=j}$ is the mean of $y_{i}(k)-y_{i}(0)$ among those with $x_{i}=j$. This decomposition exactly matches the decomposition (7) in Goldsmith-Pinkham et al. (2024). $\blacksquare$
See 3.2
Proof.
By Theorem D.1, we can analyze the regression
$Y_{i}=\tau_{0}(W_{i}-\delta_{0}-\delta_{1}^{\prime}x_{2i})+\tau_{1}^{\prime}\left({W_{i}x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}}\right)+\epsilon_{i}$
instead. Here, $\delta_{0},\delta_{1}$ are population projection coefficients of $\pi_{i}^{*}$ on $1,x_{2i}$ and $\Gamma_{0},\Gamma_{1}$ are population projection coefficients of $\pi_{i}^{*}x_{1i}$ on $1,x_{2i}$. The potential weights for this FWL-transformed regression is
$\rho_{i}(w)=\tilde{G}_{n}^{-1}\begin{bmatrix}w-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ wx_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix},$
where $\tilde{G}_{n}$ is a positive definite matrix by Theorem D.1. Now, the implicit design solves (2.2)
$\tilde{G}_{n}^{-1}\left({\pi_{i}\begin{bmatrix}1-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}+(1-\pi_{i})\begin{bmatrix}-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ -\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}}\right)=0.$
This equation admits a solution if and only if
$\pi_{i}\begin{bmatrix}1-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ x_{1i}-\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}+(1-\pi_{i})\begin{bmatrix}-\delta_{0}-\delta_{1}^{\prime}x_{i2}\\ -\Gamma_{0}-\Gamma_{1}x_{2i}\end{bmatrix}=0,$
which immediately rearranges into
$\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}\quad\pi_{i}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}.$
If the implicit design exists, then $\pi_{i}$ satisfies the above display, meaning that $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$ for all $i$. On the other hand, if $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$, then setting $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$ would solve these equations. Hence an implicit design exists iff $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$.
When this happens, $\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$. Applying Theorem 2.1, if $\bm{\tau}$ is a vector of causal contrasts, then $\pi_{i}^{*}=\pi_{i}=\delta_{0}+\delta_{1}^{\prime}x_{i2}$ and $(\delta_{0}+\delta_{1}^{\prime}x_{i2})x_{1i}=\pi_{i}^{*}x_{1i}=\Gamma_{0}+\Gamma_{1}x_{2i}$, since an implicit design exists. ∎
See 3.3
Proof.
By Theorem D.1, we can focus on the population regression of $\mathbf{y}_{i}$ on $\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x}$. Here,
$\ddot{\mathbf{w}}_{t}=\mathbf{w}_{t}-\frac{1^{\prime}\mathbf{w}}{T}-\frac{1}{n}\sum_{i=1}^{n}\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi_{i}^{*}(\tilde{\mathbf{w}})\tilde{\mathbf{w}}_{t}+\frac{1}{n}\sum_{i=1}^{n}\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi_{i}^{*}(\tilde{\mathbf{w}})\frac{1^{\prime}\tilde{\mathbf{w}}_{t}}{T}$
and
$\ddot{\mathbf{x}}_{it}=\mathbf{x}_{it}-\frac{1^{\prime}\mathbf{x}_{i}}{T}-\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}_{it}+\frac{1}{n}\sum_{i=1}^{n}\frac{1^{\prime}\mathbf{x}_{i}}{T}.$
The potential weight vector is then
$\bm{\rho}_{i}(\mathbf{w})=\frac{\ddot{\mathbf{w}}-\ddot{\mathbf{x}}_{i}\beta_{w\to x}}{\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{W}_{i}\sim\pi_{i}^{*}(\cdot)}\left[{(\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x})^{\prime}(\ddot{\mathbf{W}}_{i}-\ddot{\mathbf{x}}_{i}\beta_{w\to x})}\right]}.$
where the denominator is strictly positive by Theorem D.1.
The condition (2.2) is then
00 0 $=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\rho_{i}(\mathbf{w})\implies\ddot{\mathbf{x}}_{i}\beta_{w\to x}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\ddot{\mathbf{w}}$ (C.5)
The right-hand side is a linear combination of the columns of $\mathcal{W}$ and $1_{T}$. Thus, a necessary condition for the existence of implicit designs is that $\ddot{\mathbf{x}}_{i}\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\}).$ The left-hand side is further more
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}+c1_{T}$
for some scalar $c$. Thus, a necessary condition is furthermore
$\left({\mathbf{x}_{i}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}}\right)\beta_{w\to x}\in\operatorname{span}(\mathcal{W}\cup\left\{{1_{T}}\right\}).\qed$
When we know that $\beta_{w\to x}=0$, then the condition (2.2) is
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\ddot{\mathbf{w}}=0.$
By the same argument as for Theorem 3.1(5)(a), the unique implicit design is $\pi_{i}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$.
See 4.1
Proof.
Formally, since we never precisely defined minimal causality for TSLS, we say that $\bm{\tau}$ is minimally causal if under $\bm{\pi}^{*}$ it is invariant to replacing $y_{i}(\cdot)$ with $y_{i}(\cdot)+c_{i}$. Then exactly analogous to Theorem 2.1, it is minimally causal if and only if an implicit design exists and equals $\bm{\pi}^{*}$.
The estimand $\bm{\tau}$ depends on the profile of compliance types $d_{i}^{*}(\cdot),i=1,\ldots,n$. The estimand is (4). By definition, $\bm{\tau}$ assigns proper weights if and only if $\bm{\pi}^{*}$ belongs to the set of implicit compliance profiles (4). This is in turn equivalent to (3) and (4). ∎
See 4.2
Proof.
  1. (1)
    For (4.3), partition conformably
    $H_{n}^{-1}=G_{zt}=\begin{bmatrix}H_{zt}&H_{zx}\\ H_{xt}&H_{xx}\end{bmatrix}.$
    Let $G_{1|2}=H_{zt}-H_{zx}H_{xx}^{-1}H_{xt}$ be a $(J-1)\times(J-1)$ matrix. It is invertible if $H_{n}$ is invertible. We compute that
    $\rho_{i}(j)=[I_{J-1},0]H_{n}\begin{bmatrix}e_{j}\\ x_{i}\end{bmatrix}=G_{1|2}^{-1}\left({e_{j}-G_{wx}G_{xx}^{-1}x_{i}}\right).$
    The implicit design is therefore $\pi_{i}(\cdot)$ such that
    $0=G_{1|2}^{-1}\left({\begin{bmatrix}\pi_{i}(0)\\ \vdots\\ \pi_{i}(J-1)\end{bmatrix}-H_{wx}H_{xx}^{-1}x_{i}}\right)\iff\pi_{i}(j)=(H_{wx}H_{xx}^{-1}x_{i})_{j}$
    We conclude the proof by observing that $(H_{wx}H_{xx}^{-1}x_{i})_{j}$ is exactly the population projection of $\pi_{i}^{*}(j)$ on $x_{i}$. That is, $(H_{wx}H_{xx}^{-1}x_{i})_{j}=x_{i}^{\prime}\delta_{j}$.
  2. (2)
    1. (a)
      When $J=2$, we can compute
      $G_{1|2}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi^{*}_{i}d_{i}^{*}(1)-\frac{1}{n}\sum_{i=1}^{n}\underbrace{H_{zx}H_{xx}^{-1}x_{i}}_{\pi_{i}}\left\{{\pi_{i}^{*}d_{i}^{*}(1)+(1-\pi_{i}^{*})d_{i}^{*}(0))}\right\}$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i})d_{i}(1)-\frac{1}{n}\sum_{i=1}^{n}\pi_{i}(1-\pi_{i}^{*})d_{i}(0)$
      $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i}^{*})(d_{i}(1)-d_{i}(0)).$
      The estimand is in turn characterized by
      $\omega_{i}(1)=\rho_{i}(1)\pi_{i}^{*}=\frac{1}{G_{1|2}}(1-\pi_{i})\pi_{i}^{*}=\frac{(1-\pi_{i}^{*})\pi_{i}^{*}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i}^{*})(d_{i}(1)-d_{i}(0))},$
      where
      $\tau=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}(1)(y_{i}(d_{i}^{*}(1))-y_{i}(d_{i}^{*}(0))).$
      For this estimand to assign proper weights, $\omega_{i}(1)$ should weakly have the same sign as $d_{i}^{*}(1)-d_{i}^{*}(0)$. Since the sign of $\omega_{i}(1)$ is determined globally by $\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i}^{*})(d_{i}(1)-d_{i}(0))$, this implies that either $d_{i}^{*}(1)\geq d_{i}^{*}(0)$ or the reverse for all $i$.
    2. (b)
      Without covariates, let
      $Q=\left({[0,I_{J-1}]-\begin{bmatrix}\pi(1)\\ \vdots\\ \pi(J-1)\end{bmatrix}1_{J}^{\prime}}\right)\cdot\operatorname{diag}(\pi(0),\ldots,\pi(J)).$
      Because $x_{i}$ is just a constant, the implicit design $\pi(\cdot)$ is the same for every $i$. Observe that with this definition, the $\mathbb{R}^{J-1\times J}$ matrix representing the implicit estimand is equal to
      $[\bm{\omega}(0;\bm{\pi}),\ldots,\bm{\omega}(J-1;\bm{\pi})]=G_{1|2}^{-1}Q.$
      A given compliance type $d(\cdot)$ can be represented with a matrix $C(d)\in\mathbb{R}^{J\times J-1}$, for which
      $C_{kk^{\prime}}(d)=\mathbbm{1}(d(k)=k^{\prime}).$
      For this compliance type, the matrix $\Omega(d)$ whose entries are $\Omega_{kk^{\prime}}(d)=\omega_{i}^{(k,k^{\prime})}(d)$ can be represented as $\Omega(d)=G_{1|2}^{-1}QC(d).$ This matrix does not depend on $i$.
      If $d_{i}=d$ is included in an implicit compliance profile, then $\Omega(d)$ is a positive semidefinite diagonal matrix. When that happens, since
      $QC(d)=G_{1|2}\Omega(d),$
      each column of $QC(d)$ is equal to the corresponding column of $G_{1|2}$, scaled by some nonnegative scalar $c_{d}$. To that end, note that the $k$th column of $Q_{i}C(d)$ is
      $(QC(d))[:,k]=-\left({\sum_{w:d(w)=k}\pi(w)}\right)\begin{bmatrix}\pi(1)\\ \vdots\\ \pi(J-1)\end{bmatrix}+e_{d}\odot\begin{bmatrix}\pi(1)\\ \vdots\\ \pi(J-1)\end{bmatrix}$
      where $e_{d}$ is the binary vector for which $e_{d,m}=1$ if $d(m)=k$, for $m\in\left\{{1,\ldots,J-1}\right\}$.
      We first show common compliance. Take two compliance types $d_{i},d_{j}$ in the implicit compliance profile. Without loss, suppose $i=1,j=2$. Then for all $k=1,\ldots,J-1$, either (i) one of $(QC(d_{\ell}))[:,k]=0$ for some $\ell=1,2$, or (ii)
      $(QC(d_{1}))[:,k]=c(QC(d_{2}))[:,k]$
      for some $c>0$.
      Suppose (i) occurs and, without loss of generality, suppose $(QC(d_{1}))[:,k]=0$. Then
      $e_{d_{1}}=1_{J-1}\left({\sum_{w:d_{1}^{*}(w)=k}\pi(w)}\right)$
      Because $e_{d_{1}}$ is a binary vector, this implies $\sum_{w:d_{1}(w)=k}\pi_{i}(w)\in\left\{{0,1}\right\}$—equivalently, either the treatment is $k$ for all instrument values or it is never $k$ for all instrument values. That is, $d_{1}$ is a $k$-always taker or a $k$-never taker.
      Suppose neither $d_{1},d_{2}$ are $k$-always/never takers, ruling out (i). This in turn implies that
      $(e_{d_{1}^{*}}-ce_{d_{2}^{*}})=\left({-c\left({\sum_{w:d_{2}^{*}(w)=k}\pi(w)}\right)+\left({\sum_{w:d_{1}^{*}(w)=k}\pi(w)}\right)}\right)1_{J-1}$
      for some $c>0$. The right-hand side is proportional to a vector of all ones. For this to happen, either (a) $e_{d_{1}^{*}}=e_{d_{2}^{*}}$ or (b) without loss of generality, $e_{d_{1}^{*}}=1_{J-1}$, $e_{d_{2}^{*}}=0_{J-1}$.
      Suppose (b) occurs, then since neither $d_{1}^{*},d_{2}^{*}$ are always/never-takers, we must have $d_{1}^{*}(0)\neq k$ and $d_{2}^{*}(0)=k$. Then we must have
      $1_{J-1}=(-c\pi(0)+(1-\pi(0)))1_{J-1}$
      Since $-c\pi(0)+(1-\pi(0))<1$, we conclude that (b) is impossible.
      Thus (a) occurs and $e_{d_{1}^{*}}=e_{d_{2}^{*}}$. Then the equation is
      $(1-c)e_{d_{2}^{*}}=\left\{{(1-c)e_{d_{2}^{*}}^{\prime}\begin{bmatrix}\pi(1)\\ \vdots\\ \pi(J-1)\end{bmatrix}+(\mathbbm{1}(d_{1}^{*}(0)=k)-\mathbbm{1}(d_{2}^{*}(0)=k)c)\pi(0)}\right\}1_{J-1}$
      If $c=1$ or if $e_{d_{2}^{*}}=0$, then this forces
      $(\mathbbm{1}(d_{1}(0)=k)-\mathbbm{1}(d_{2}(0)=k))\pi(0)=0\implies\mathbbm{1}(d_{1}^{*}(0)=k)=\mathbbm{1}(d_{2}^{*}(0)=k)$
      This implies that $d_{1}^{*}(w)=k$ if and only if $d_{2}^{*}(w)=k$.
      If $c\neq 1$, then $e_{d_{2}}=1$. We then require
      $(1-c)=(1-c)(1-\pi(0))+(\mathbbm{1}(d_{1}^{*}(0)=k)-\mathbbm{1}(d_{2}^{*}(0)=k)c)\pi(0)$
      which implies
      $(1-c)\pi(0)=(\mathbbm{1}(d_{1}^{*}(0)=k)-\mathbbm{1}(d_{2}^{*}(0)=k)c)\pi(0)$
      The only option is then $\mathbbm{1}(d_{1}^{*}(0)=k)=\mathbbm{1}(d_{2}^{*}(0)=k)=1$. This implies that both $d_{1},d_{2}$ are $k$-always takers, which is a case that is already covered. This proves common compliance.
      Because we know that some compliance type $d_{i}^{*}$ is a full complier. Then $d_{i}^{*}$ is a permutation of $\left\{{0,\ldots,J-1}\right\}$. Set $f=d_{i}^{*}(\cdot)$. Take any $d_{j}^{*}$ and compare with $d_{i}^{*}$.
      Among implicit compliance profiles that contain a full complier $d_{i}(\cdot)$, $f^{-1}(\cdot)=d_{i}(\cdot)$ is a permutation of $\left\{{0,\ldots,J-1}\right\}$. Take any other $d_{j}(\cdot)$. Suppose $d_{j}(\cdot)$ is not constant, then it must be a $k$-complier for some $k$. By common compliance $d_{j}^{-1}(k)=\left\{{f(k)}\right\}$, meaning that $d_{j}(f(k))=k$ and that $d_{j}(w)\neq k$ for all $w\neq f(k)$. Thus if $d_{j}$ is not a $k$-complier, then $d_{j}(f(k))=0$, since any other value $\ell$ is inconsistent with $d_{j}$ being an $\ell$-complier. This shows that $d_{j}(\cdot)$ is either constant or $d_{j}(f(k))\in\left\{{0,k}\right\}$. Thus the compliance profile is instrument-specific. ∎
Appendix D Additional results
D.1. Invariances
Theorem D.1 (Frisch–Waugh–Lovell, in population).
Consider a population regression specification $(\Lambda,G_{n},z_{1},\ldots,z_{n})$. Partition $z_{t}(\mathbf{x}_{i},\mathbf{w})$ into $z_{t1}(\cdot)$ and $z_{t2}(\cdot)$. Suppose $\Lambda=[\Lambda_{1},0]$ loads solely on entries in $z_{t1}$: $\Lambda z_{t}=\Lambda_{1}z_{t1}$. Let
$\Gamma=\left({\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t1}(\mathbf{x}_{i},\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right)\left({\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})z_{t2}(\mathbf{x}_{i},\mathbf{w})^{\prime}}\right)^{-1}$
be the population projection matrix of $z_{t1}$ onto $z_{t2}$. Define $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=z_{t1}(\mathbf{x}_{i},\mathbf{w})-\Gamma z_{t2}(\mathbf{x}_{i},\mathbf{w})$. Consider the regression specification defined by $\Lambda_{1},\tilde{z}_{t}(\cdot),$ and
$\tilde{G}_{n}=\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})^{\prime}.$
Then:
  1. (i)
    $\Gamma,\tilde{G}_{n}$ are functions of the original Gram matrix $G_{n}$. Since $G_{n}$ is assumed to be invertible, $\Gamma$ is well-defined and $\tilde{G}_{n}$ is positive definite.
  2. (ii)
    The potential weights associated with the two population regression specifications are the same.
Proof.
Note that we can partition
$G_{n}=\frac{1}{n}\sum_{i,t}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})z_{t}(\mathbf{x}_{i},\mathbf{w})z_{t}(\mathbf{x}_{i},\mathbf{w})^{\prime}=\begin{bmatrix}G_{11}&G_{12}\\ G_{21}&G_{22}\end{bmatrix}$
conformably where $G_{11}$ is the Gram matrix associated with $z_{t1}$ and $G_{22}$ is the Gram matrix associated with $z_{t2}$. Since $G_{n}$ is assumed to be invertible and thus positive definite, $G_{11},G_{22}$ are invertible and positive definite as well. This implies that $\Gamma=G_{12}G_{22}^{-1}$ and $\tilde{G}_{n}=G_{11}-\Gamma G_{21}$ are well-defined. This also implies that the Schur complement $\tilde{G}_{n}=G_{11}-\Gamma G_{21}$ is invertible (and positive definite). This proves (i).
By the block matrix inversion formula,
$G_{n}^{-1}=\begin{bmatrix}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}&-(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}G_{12}G_{22}^{-1}\\ -G_{22}^{-1}G_{21}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}&G_{22}^{-1}+G_{22}^{-1}G_{21}(G_{11}-G_{12}G_{22}^{-1}G_{21})^{-1}G_{12}G_{22}^{-1}\end{bmatrix}.$
Hence
$\Lambda G_{n}^{-1}=\left[{\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\quad-\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\Gamma}\right]$
Now, the potential weights for the original regression specification are
$\bm{\rho}_{it}(\mathbf{w})$ $=\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}z_{t1}(\mathbf{x}_{i},\mathbf{w})-\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\Gamma z_{t2}(\mathbf{x}_{i},\mathbf{w})$
$=\Lambda_{1}(G_{11}-\Gamma G_{21})^{-1}\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})$
$=\Lambda_{1}\tilde{G}_{n}^{-1}\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w}).$
This proves (ii). ∎
Theorem D.2 (Frisch–Waugh–Lovell, in sample).
Consider a population regression specification $(\Lambda,G_{n},z_{1},\ldots,z_{n})$. Partition $z_{t}(\mathbf{x}_{i},\mathbf{w})$ into $z_{t1}(\cdot)$ and $z_{t2}(\cdot)$. Suppose $\Lambda=[\Lambda_{1},0]$ loads solely in entries in $z_{t1}$: $\Lambda z_{t}=\Lambda_{1}z_{t1}$. Suppose the sample Gram matrix is invertible. Let
$\hat{\Gamma}=\left({\frac{1}{n}\sum_{i,t}z_{t1}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right)\left({\frac{1}{n}\sum_{i,t}z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t2}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}}\right)^{-1}$
be the sample projection matrix of $z_{t1}$ onto $z_{t2}$. Define $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=z_{t1}(\mathbf{x}_{i},\mathbf{w})-\hat{\Gamma}z_{t2}(\mathbf{x}_{i},\mathbf{w})$. Consider the regression specification defined by $\Lambda_{1},\tilde{z}_{t}(\cdot)$. Then the estimated potential weights associated with the two regression specifications are the same.
Proof.
The proof entirely follows from the proof of Theorem D.1 by setting $\pi_{i}^{*}(\mathbf{w})=\mathbbm{1}(\mathbf{w}=\mathbf{W}_{i})$. ∎
Theorem D.3 (Invariance under reparameterization).
Consider a population regression specification $(\Lambda,G_{n},z_{1},\ldots,z_{n})$. Consider an invertible matrix $M$ and another population specification defined by $\tilde{z}_{t}(\mathbf{x}_{i},\mathbf{w})=Mz_{t}(\mathbf{x}_{i},\mathbf{w})$ and $\tilde{\Lambda}=\Lambda M^{\prime}$, so that the two vectors of coefficients represent the same underlying contrasts. Then the potential weights associated with the two regression specifications are the same. The estimated potential weights are also the same.
Proof.
We can easily see that the second regression specification has Gram matrix $\tilde{G}_{n}=MG_{n}M^{\prime}$ (and $\widehat{\tilde{G}_{n}}=M\hat{G}_{n}M^{\prime}$). Thus the potential weights for the second specification are
$\tilde{\bm{\rho}}_{it}(\mathbf{w})=\tilde{\Lambda}(M^{\prime})^{-1}G_{n}^{-1}M^{-1}Mz_{t}(\mathbf{x}_{i},\mathbf{w})=\bm{\rho}_{it}(\mathbf{w}).$
The corresponding equalities similarly hold for the estimated potential weights. ∎
D.2. Relation to Blandhol et al. (2025) and Bhuller and Sigstad (2024)
D.2.1. Relation to Blandhol et al. (2025), Theorem 1
We now discuss the relation to Blandhol et al. (2025) between their Theorem 1 and our Proposition 4.2(1) and (2)(a). In our finite-population setting, the terminology of Blandhol et al. (2025) has analogues:
  1. (1)
    Montonicity / Assumption MON: $d_{i}^{*}(1)\geq d_{i}^{*}(0)$ for all $i$
  2. (2)
    $\tau$ is weakly causal: This is equivalent to that $\tau$ assigns proper weights and can be decomposed into
    1. (a)
      $\tau$ is minimally causal: $\tau=\frac{1}{n}\sum_{i}\omega_{i}^{*}(y_{i}(1)-y_{i}(0))$ for some $\omega_{i}^{*}\in\mathbb{R}$
    2. (b)
      $\tau$ has convex weights $\omega_{i}^{*}\geq 0$.
  3. (3)
    Exogeneity/Assumption EXO: $\pi_{i}^{*}$ not a function of $y_{i}(\cdot),d_{i}(\cdot)$.
Theorem 1 in Blandhol et al. (2025) states that under exogeneity and monotonicity, $\tau$ from (4.3) (with $J=2$) is weakly causal if and only if $\pi_{i}^{*}$ is linear in $x_{i}$. Proposition 4.1 and Proposition 4.2(1) imposes exogeneity and shows that $\tau$ is minimally causal if and only if $\pi_{i}^{*}$ is linear in $x_{i}$. Compared to Blandhol et al. (2025), it allows for $J>2$ but restricts to finitely valued instruments. Proposition 4.2(1) thus clarifies that the equivalence in Theorem 1 of Blandhol et al. (2025) is coming from the equivalence between level independence and linearity of $\pi_{i}^{*}$.
When $\tau$ is minimally causal and $J=2$, Proposition 4.1 and Proposition 4.2(2)(a) shows that $\tau$ is furthermore weakly causal if and only if monotonicity holds. Thus, together, we show the following: Under exogeneity, $\tau$ is weakly causal if and only if $\pi_{i}^{*}$ is linear in $x_{i}$ and monotonicity holds. This moves monotonicity in Blandhol et al. (2025) from an assumption to an implication.
D.2.2. A counterexample to Bhuller and Sigstad (2024)’s Proposition 5
Following Bhuller and Sigstad (2024), we work in a sampling framework. Let $Z$ be an instrument that takes values $\left\{{0,1,2}\right\}$ and $D$ be a treatment that takes values $\left\{{0,1,2}\right\}$. Let $Z_{z}=\mathbbm{1}(Z=z)$ and similarly define $D_{1},D_{2}$. Consider the two-stage least-squares regression
$Y=\alpha+\beta_{1}D_{1}+\beta_{2}D_{2}+\epsilon,$
instrumenting for $[1,D_{1},D_{2}]$ using $[1,Z_{1},Z_{2}]$. Let $s(\cdot)$ denote a response type, mapping from instrument values to treatment values.
Proposition 5 in Bhuller and Sigstad (2024) states that TSLS assigns proper weights in this regression if and only if there exists a one-to-one $f:\left\{{0,1,2}\right\}\to\left\{{0,1,2}\right\}$ such that for all $k\in\left\{{1,2}\right\}$ and response type $s(\cdot)$, either (i) $s(z)\neq k$ for all $z\in\left\{{0,1,2}\right\}$, (ii) $s(z)=k$ for all $z\in\left\{{0,1,2}\right\}$, or (iii) $s(z)=k\iff f(z)=k$.
They then conclude that, up to permuting the instrument values, we must then have response types only being
$(s(0),s(1),s(2))\in\left\{{(000),(111),(222),(010),(002),(012)}\right\}.$
Unfortunately, this claim is not quite correct as stated. Consider instead three types of individuals characterized by $(A,B,C)\equiv(000),(011),(002)$. That is, type $A$ always picks $D=0$. Type $B$ picks $D=1$ when $Z=1$ or $Z=2$. Type $C$ picks $D=2$ only when $Z=2$. These response types do satisfy common compliance—the only 1-complier is type $B$ and the only 2-complier is type $C$. Compliance is not instrument specific: $Z=2$ moves type $C$ to $2$ but type $B$ to $1$.
Indeed, consider the only if clause of Proposition 5 in Bhuller and Sigstad (2024):
  • For $k=1$, types $A$ and $C$ satisfy (i). However, type B satisfying (iii) necessitates $f(1)=f(2)=1$. This contradicts $f(\cdot)$ being one-to-one.
Thus, Proposition 5 in Bhuller and Sigstad (2024) would claim that TSLS does not assign proper weights. However, in this example, one can show that
$\beta_{1}=\mathbb{E}[Y(1)-Y(0)\mid B]\quad\beta_{2}=\mathbb{E}[Y(2)-Y(0)\mid C].$
Thus TSLS does assign proper weights.
D.3. Additional consistency results
Recall the sequence of populations setup in Appendix B. Additionally, let $\pi_{1,n}^{*},\ldots,\pi_{n,n}^{*}$ denote a triangular array of marginal treatment assignment probabilities, and let $\Pi_{n}^{*}$ denote the joint distribution of the treatment assignments $\mathbf{W}_{1},\ldots,\mathbf{W}_{n}$ under the $n$th population with marginals equal to $\pi_{1,n}^{*},\ldots,\pi_{n,n}^{*}$.
Lemma D.4.
Suppose that in each $\Pi_{n}^{*}$, treatments are independently assigned according to $\pi_{i}^{*}=\pi_{i,n}^{*}$. Let $G_{i}=\sum_{t=1}^{T}z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})z_{t}(\mathbf{x}_{i},\mathbf{W}_{i})^{\prime}$. Assume that for all $1\leq j\leq K,1\leq\ell\leq K$, the average second moment of $G_{i,j\ell}$ grows slower than $n$: as $n\to\infty$, $\frac{1}{n}\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\pi_{i,n}^{*}}[G_{i,j\ell}^{2}]}\right)\to 0.$ Then Assumption B.1 holds.
Proof.
Fix a coordinate $(j,k)$. Let $S_{n}=\sum_{i=1}^{n}G_{i,jk}$ and note that $\hat{G}_{n,jk}=\frac{1}{n}S_{n}$. Define $\mu_{n}=\mathbb{E}[S_{n}]$ and $\sigma_{n}^{2}=\operatorname{Var}(S_{n})$. Theorem 2.2.6 in Durrett (2019) states that if $\sigma_{n}^{2}/n^{2}\to 0$ then
[…]
Note that
$\sigma_{n}^{2}\leq\sum_{i=1}^{n}\mathbb{E}[G_{i,jk}^{2}].$
[…]
Proposition D.5.
Suppose that, for some $0<C<\infty$, with probability tending to one, an estimated implicit design $\hat{\bm{\pi}}$ exists and is bounded in the sense that:
$\max_{i}\max_{\mathbf{w}\in\mathcal{W}}|\hat{\pi}_{i}(\mathbf{w})|<C.$
[…]
Proof.
Fix $\eta>0$. Note that
[…]
$+\operatorname{P}\left[{\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\rho}_{i}(\mathbf{w})\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}>\eta,\text{a bounded implicit design exists}}\right]$
Note that if $\hat{\pi}_{i}$ is a bounded implicit design, then
$\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}\bm{\rho}_{i}(\mathbf{w})\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}=\bigg{\lvert}\sum_{\mathbf{w}\in\mathcal{W}}(\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w}))\hat{\pi}_{i}(\mathbf{w})\bigg{\rvert}_{\infty}\leq C(J+1)\max_{\mathbf{w}\in\mathcal{W}}|\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w})|_{\infty}.$
Thus,
[…]
$+\operatorname{P}\left[{C(J+1)\max_{\mathbf{w}\in\mathcal{W}}|\bm{\rho}_{i}(\mathbf{w})-\hat{\bm{\rho}}_{i}(\mathbf{w})|_{\infty}>\eta}\right]\to 0.$
by assumption. The claim of uniform-in-$i$ convergence follows by essentially the same argument, where we take maximum over $i$. ∎
D.3.1. Consistency of $\hat{G}_{n}$ to $G_{n}$ under rejective sampling
This subsection considers a law of large numbers under a version of sampling with replacement (what Hájek, 1964, calls “rejective sampling”), and largely follows Rambachan and Roth (2020). In particular, we assume that the treatment is binary, and there is a sequence of unconditional probabilities $p_{1},p_{2},\ldots$ as well as a sequence of sample sizes $N_{n}$ such that $\Pi^{*}_{n}$ describes the joint distribution of $W_{1},\ldots,W_{n}$ conditional on the event $\sum_{i=1}^{n}W_{i}=N_{n}$, where $W_{i}\sim\operatorname{Bern}(p_{i})$ independently unconditionally. Correspondingly, let $\pi_{i,n}$ be the probability that $W_{i}=1$ under $\Pi^{*}_{n}$.
Lemma D.6.
Write $\pi_{i}(w)=\pi_{i,n}(w)$ and omit the $n$ subscript. Consider $\overline{y}_{w}=\frac{1}{n}\sum_{i}\mathbbm{1}(W_{i}=w)y_{i}$ where $\mathbb{E}[\overline{y}_{w}]=\frac{1}{n}\sum_{i}\pi_{i}(w)y_{i}(w)$. Suppose $C_{n}\equiv\sum_{i=1}^{n}\pi_{i}(w)(1-\pi_{i}(w))\to\infty$ as $n\to\infty$. Then, as $n\to\infty$
$\operatorname{Var}(\overline{y}_{w})=\frac{(1+o(1))}{n}\frac{C_{n}}{n}\sum_{i=1}^{n}\frac{\pi_{i}(w)(1-\pi_{i}(w))}{C_{n}}\left({y_{i}(w)-\frac{\sum_{j=1}^{n}\pi_{j}(w)(1-\pi_{j}(w))y_{j}(w)}{C_{n}}}\right)^{2}$
Proof.
This is a restatement of Theorem 6.1 in Hájek (1964). The notation $y_{i}$ in the theorem corresponds to $\pi_{i}(w)y_{i}(w)$ in our notation. ∎
A sufficient condition for the variance to tend to zero is the following:
Assumption D.1.
For all entries $k$, $z_{k}(x_{i},w)$ is uniformly bounded by $0<M<\infty$ and $C_{n}(w)\equiv\sum_{i=1}^{n}\pi_{i}(w)(1-\pi_{i}(w))\to\infty$ as $n\to\infty$.
The boundedness condition for $z_{k}$ is stronger than needed. In particular, what is needed is that the $\pi_{i}(w)(1-\pi_{i}(w))$-weighted variance of $y_{i}(w)$ is $O(1)$.
Lemma D.7.
[…]
Proof.
We prove this claim by showing all entries converge in probability, since there are finitely many entries. Fix some $j,k$. Note that
$\hat{G}_{n,jk}=\sum_{w\in\left\{{0,1}\right\}}\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(W_{i}=w)z_{j}(x_{i},w)z_{k}(x_{i},w)}_{\overline{z}_{n}(w)}.$
By Lemma D.6 and Assumption D.1 applied to $\overline{z}_{n}$ where $y_{i}(w)$ is taken to be $z_{j}(x_{i},w)z_{k}(x_{i},w)$, we have that $\operatorname{Var}(\overline{z}_{n}(w))\to 0$ as $n\to\infty$, since $C_{n}(w)/n\leq 1$. Since $\mathbb{E}[\overline{z}_{n}(w)]=\frac{1}{n}\sum_{i}\pi_{i}(w)z_{j}(x_{i},w)z_{k}(x_{i},w)$, we have that
[…]
by Chebyshev’s inequality. ∎
D.4. Uniqueness of implicit designs in cross-sections
Consider a cross-sectional setup with $J+1$ treatments where the potential weights are
$\rho_{i}(w)=\tilde{\Lambda}\tilde{G}_{n}^{-1}\tilde{z}(x_{i},w)$
and $\tilde{\Lambda}$ is a $J\times K$ matrix with rank $J$. It is possible to reparametrize the regressors (i.e. choose an invertible $M$ such that $z(\cdot)=M\tilde{z}(\cdot)$) such that
$\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)$
and $\Lambda=[I_{J},0]$. Without essential loss of generality, let us assume the sequence of reparametrized specifications satisfy Assumption B.2.
Partition $z$ into $z_{1}$ and $z_{2}$ where $z_{1}$ is $J$-dimensional. By Theorem D.1, the potential weights are further equivalent to
$\rho_{i}(w)=G_{n,1|2}^{-1}z_{1|2}(x_{i},w)$
where $z_{1|2}(x_{i},w)=z_{1}(x_{i},w)-\Gamma^{\prime}z_{2}(x_{i},w)$ for population projection coefficients $\Gamma$ and
$G_{n,1|2}=\frac{1}{n}\sum_{i,w}\pi_{i}^{*}(w)z_{1|2}(x_{i},w)z_{1|2}(x_{i},w)^{\prime}.$
Assumption D.2.
The residualized covariate transform is non-singular in the sense that the $J\times(J+1)$ matrix
$Z_{i}\equiv[z_{1|2}(x_{i},0),\ldots,z_{1|2}(x_{i},J)]$
whose columns are covariate transforms evaluated at a treatment level has minimum singular value (that is, the $J$th singular value) greater than some $\eta>0$, uniformly for all $i$.
Assumption D.3.
The maximum eigenvalue of $G_{n}$ is bounded above by $M<\infty$ for all $n$.
Lemma D.8.
Under Assumptions D.2, B.2, and D.3, condition (ii) in Proposition B.2 is satisfied for some lower bound $\epsilon>0$ on the minimum (i.e. $J$th) singular value.
Note that since the potential weight matrix $G_{n,1|2}^{-1}Z_{i}$ is $J\times(J+1)$ and has $J$ positive singular values bounded below, if there exists an implicit design, then it must be unique.
Proof.
We first show that Assumption B.2 implies that the minimum eigenvalue of $G_{n,1|2}$ is bounded away from zero and maximum bounded by $M$. Note that $G_{n,1|2}$ is a Schur complement of a submatrix of $G_{n}$, and so $G_{n,1|2}^{-1}$ is a principal submatrix of $G_{n}^{-1}$. The eigenvalues of $G_{n,1|2}^{-1}$ interlace the eigenvalues of $G_{n}^{-1}$ by Cauchy’s interlace theorem (Hwang, 2004). In particular, the spectrum of $G_{n,1|2}^{-1}$ is included in the range of the spectrum of $G_{n}^{-1}$. Hence $\lambda_{\min}(G_{n,1|2})>\epsilon$ by Assumption B.2 and $\lambda_{\max}(G_{n,1|2})<M$ by Assumption D.3.
Next, it suffices to show that $G_{n,1|2}^{-1}Z_{i}$ has minimum singular value bounded below. Write the compact SVD as
$Z_{i}=U\Sigma V^{\prime}$
for diagonal $J\times J$ matrix $\Sigma$ with $UU^{\prime}=U^{\prime}U=V^{\prime}V=I_{J}$. Similarly, write
$G_{n,1|2}^{-1}=QDQ^{\prime}$
for an orthogonal matrix $Q$ and diagonal $D$. Then
$G_{n,1|2}^{-1}Z_{i}Z_{i}^{\prime}G_{n,1|2}^{-1}=QDQ^{\prime}U\Sigma^{2}U^{\prime}QDQ^{\prime}.$
is a real symmetric matrix. The spectrum of this matrix is the same as the spectrum of
$D\underbrace{Q^{\prime}U\Sigma^{2}U^{\prime}Q}_{\text{positive definite matrixwith spectrum $\Sigma^{2}$}}D.$
Since $\lambda_{\min}(AB)\geq\lambda_{\min}(A)\lambda_{\min}(B)$ for two positive definite matrices $A,B$, [^46] To see this, note that the operator norm (largest eigenvalue of a positive semidefinite matrix) is submultiplicative $\lVert AB\rVert_{op}\leq\lVert A\rVert_{op}\lVert B\rVert_{op}$. Apply this inequality to $A^{-1}B^{-1}$. we have that the minimum eigenvalue of $G_{n,1|2}^{-1}Z_{i}Z_{i}^{\prime}G_{n,1|2}^{-1}$ is bounded below by $\frac{\eta^{2}}{\lambda_{\max}(G_{n,1|2})^{2}}$. Thus, the minimum singular value of $G_{n,1|2}^{-1}Z_{i}$ is bounded below by $\eta/M$. ∎
D.5. Additional results for interacted regression
Proposition D.9 provides a unified analyses of the regression (D.9). It implies our results in Theorem 3.1(4) when $\overline{x}_{t}=\overline{x}_{1}$ or $\overline{x}_{0}$, but allows for evaluating on other points in the line connecting $\overline{x}_{1}$ and $\overline{x}_{0}$. The implicit design for the model-based ATE is analyzed when we set $t=\alpha_{0}=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}$.
Inspecting the resulting implicit designs (D.9), we find that they are in general fractional linear, encompassing the linear-odds result of $t=1$ as a special case. Moreover, a necessary implication of $\pi_{i}^{*}=\pi_{i}$ is the condition (D.9), which is very difficult to satisfy unless $t=0$ or $t=1$. In this sense, the model-based ATU and the model-based ATT are the only estimands whose implicit design is reasonable. See D.1 for an example where the implicit model for the ATE depends on the support of $x_{i}$.
Proposition D.9.
Consider the specification indexed by $t\in\mathbb{R}$,
$Y_{i}=\gamma_{0}+W_{i}\tau_{t}+\gamma_{0}+\gamma_{1}^{\prime}x_{i}+W_{i}(x_{i}-\overline{x}_{t})^{\prime}\gamma_{2}+\epsilon_{i}$
where $\overline{x}_{t}=t\overline{x}_{1}+(1-t)\overline{x}_{0}$ for $\overline{x}_{1}=\frac{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}x_{i}}{\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}}$ and $\overline{x}_{0}=\frac{\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})x_{i}}{\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})}$.
Let
$A$ $=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x})(x_{i}-\overline{x})^{\prime}$
$V_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}\quad U_{t}=\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}$
$\alpha_{0}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}$
$\Gamma_{0t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})$
$\alpha_{1}$ $=A^{-1}\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x})=A^{-1}(\Gamma_{0t}+\alpha_{0}(\overline{x}_{t}-\overline{x}))$
$\Gamma_{1t}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x})^{\prime}=\left({V_{t}+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}}\right)A^{-1}$
$\Gamma_{2t}$ $=\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[{W_{i}(x-\overline{x}_{t})\left({W_{i}(x-\overline{x}_{t})-\Gamma_{0t}-\Gamma_{1t}(x_{i}-\overline{x})}\right)^{\prime}}\right]}\right)^{-1}$
×$\quad\quad\times\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[{(W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x}))\left({W_{i}(x_{i}-\overline{x}_{t})-\Gamma_{0t}-\Gamma_{1t}(x-\overline{x})}\right)}\right]$
$=\left({V_{t}-\Gamma_{0t}\Gamma_{0t}^{\prime}-V_{t}\Gamma_{1t}^{\prime}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}}\right)^{-1}\left({\Gamma_{0t}-\alpha_{0}\Gamma_{0t}-V_{t}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}}\right)$
Then, the potential weights satisfy
$\rho_{i}(w)=\tilde{G}_{n}^{-1}\left({w(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))-\underbrace{(\alpha_{0}-\Gamma_{2t}^{\prime}\Gamma_{0t})}_{\theta_{0t}}-\underbrace{(\alpha_{1}-\Gamma_{1t}^{\prime}\Gamma_{2t})^{\prime}}_{\theta_{1t}^{\prime}}(x_{i}-\overline{x})}\right),$
for some scalar $\tilde{G}_{n}>0$. The implicit design $\pi_{i}$ satisfies
$\left({1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t})}\right)\pi_{i}=\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x}).$
Moreover, if $\pi_{i}^{*}=\pi_{i}$, then a necessary condition is that
$tV_{1}(\theta_{1t}+\Gamma_{2t})=(1-t)U_{0}\theta_{1t}.$
Proof.
Observe that by the Frisch–Waugh–Lovell transform, the coefficient $\tau_{t}$ is the same as in the regression
$Y_{i}=\tau_{t}\left({W_{i}-\alpha_{0}-\alpha_{1}^{\prime}(x_{i}-\overline{x})-\Gamma_{2}^{\prime}\left({W_{i}(x_{i}-\overline{x}_{t})-\Gamma_{0}-\Gamma_{1}(x_{i}-\overline{x})}\right)}\right)+\epsilon_{i}.$
The quantities $\alpha_{0},\alpha_{1},\Gamma_{0},\Gamma_{1},\Gamma_{2}$ are defined exactly as these auxiliary regression coefficients. Thus, by Theorem D.1, the potential weights are equivalent to the potential weights of (D.5), which rearrange exactly to the potential weights displayed. The constraint on the implicit design corresponds exactly to
$\rho_{i}(1)\pi_{i}+(1-\pi_{i})\rho_{i}(0)=0.$
For the “moreover” part, observe that we can rewrite $\Gamma_{2t}$ as satisfying
$V_{t}\Gamma_{2t}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}\Gamma_{2t}-\Gamma_{0t}\Gamma_{0t}^{\prime}\Gamma_{2t}-V_{t}\Gamma_{1t}^{\prime}\Gamma_{2t}=\Gamma_{0t}-\alpha_{0}\Gamma_{0t}-V_{t}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}.$
Thus,
00 0 $=\left({\alpha_{0}\Gamma_{0t}-\Gamma_{0t}\Gamma_{0t}^{\prime}\Gamma_{2t}}\right)+\left({V_{t}\alpha_{1}-V_{t}\Gamma_{1t}^{\prime}\Gamma_{2t}}\right)$
$\quad\quad+\left({\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\alpha_{1}-\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\Gamma_{1t}^{\prime}\Gamma_{2t}}\right)+V_{t}\Gamma_{2t}-\Gamma_{0t}$
$=\Gamma_{0t}\theta_{0t}+\underbrace{V_{t}\theta_{1t}+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime}\theta_{1t}}_{\Gamma_{1t}A\theta_{1t}}+V_{t}\Gamma_{2t}-\Gamma_{0t}$
Observe that
$A\theta_{1t}=\Gamma_{0t}+\alpha_{0}(\overline{x}_{t}-\overline{x})-V_{t}\Gamma_{2t}-(\overline{x}_{t}-\overline{x})\Gamma_{0t}^{\prime}\Gamma_{2t}$
and hence
$V_{t}\Gamma_{2t}-\Gamma_{0t}=-A\theta_{1t}+\alpha_{0}(\overline{x}_{t}-\overline{x})-(\overline{x}_{t}-\overline{x})\Gamma_{0t}^{\prime}\Gamma_{2t}=-A\theta_{1t}+\theta_{0t}(\overline{x}_{t}-\overline{x})$
Therefore
00 0 $=\left({\Gamma_{0t}+(\overline{x}_{t}-\overline{x})}\right)\theta_{0t}+(V_{t}-A+\Gamma_{0t}(\overline{x}_{t}-\overline{x})^{\prime})\theta_{1t}$
$=\frac{1}{n}\sum_{i=1}^{n}\{(\pi_{i}^{*}(x_{i}-\overline{x}_{t})-(x_{i}-\overline{x}_{t}))\theta_{0t}$
$\quad\quad+\left({\pi_{i}^{*}(x_{i}-\overline{x}_{t})(x_{i}-\overline{x}_{t})^{\prime}-(x_{i}-\overline{x})(x_{i}-\overline{x})^{\prime}+\pi_{i}^{*}(x_{i}-\overline{x}_{t})(\overline{x}_{t}-\overline{x})^{\prime}}\right)\theta_{1t}\}$
$=-\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x}))$
Thus, if $\pi_{i}^{*}=\pi_{i}$,
$\theta_{0t}+\theta_{1t}^{\prime}(x_{i}-\overline{x})=(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))\pi_{i}^{*}.$
We thus have that
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{t})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))$
$=t\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})(1-\pi_{i}^{*})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))+(1-t)\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{1})\pi_{i}^{*}(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))$
where we may plug in
$\pi_{i}^{*}(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))=\theta_{0t}+\theta_{1t}^{\prime}(x-\overline{x})\quad(1-\pi_{i}^{*})(1-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t}))=1-\theta_{0t}-\Gamma_{2t}^{\prime}(x_{i}-\overline{x}_{t})-\theta_{1t}^{\prime}(x_{i}-\overline{x}).$
Simplifying by noting that $\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})=0=\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})$ yields the condition
$tV_{1}(\theta_{1t}+\Gamma_{2t})=(1-t)U_{0}\theta_{1t}.$
Proposition D.10.
In Proposition D.9, for $t=1$, the implicit design is equal to
$\pi_{i}=\frac{\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x})}{1+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x}_{1})}.$
The associated odds is
$\frac{\pi_{i}}{1-\pi_{i}}=\frac{\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(x-\overline{x})}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}\equiv\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x}_{0}).$
Here,
$\delta_{1}$ $=\left({\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})}\right)^{-1}\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i}^{*})(x_{i}-\overline{x}_{0})\frac{\pi_{i}^{*}}{1-\pi_{i}^{*}}$
$\delta_{0}$ $=\frac{\alpha_{0}}{1-\alpha_{0}}=\frac{1}{n}\sum_{i=1}^{n}\frac{(1-\pi_{i}^{*})}{1-\alpha_{0}}\frac{\pi_{i}^{*}}{1-\pi_{i}^{*}}$
Proof.
Here, we omit the $t$ subscript for objects defined in Proposition D.9. Observe that $\Gamma_{0}=0$. Hence
$\Gamma_{2}=-(I-\Gamma_{1}^{\prime})^{-1}\alpha_{1}\text{ and }\theta_{0}=\alpha_{0}$
and thus
$\theta_{1}=\alpha_{1}-\Gamma_{1}^{\prime}\Gamma_{2}=(I+\Gamma_{1}^{\prime}(I-\Gamma^{\prime})^{-1})\alpha_{1}=(I-\Gamma^{\prime})^{-1}\alpha_{1}.$
It remains to show that $\delta_{0},\delta_{1}$ take the form stated. First, observe that
$\theta_{1}=(I-\Gamma_{1}^{\prime})^{-1}\alpha_{1}=(A-V_{1})^{-1}\alpha_{0}(\overline{x}_{1}-\overline{x})=(A-V_{1})^{-1}\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0}).$
by plugging in $\alpha_{1}=A^{-1}(\alpha_{0}(\overline{x}_{1}-\overline{x}))$. Here,
$A$ $=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x})(x_{i}-\overline{x}_{0})^{\prime}$
$=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}-(\overline{x}-\overline{x}_{0})(\overline{x}-\overline{x}_{0})^{\prime}$
$=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}-\alpha_{0}^{2}(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
$V_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{1})(x_{i}-\overline{x}_{0})$
$=\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{0})(x_{i}-\overline{x}_{0})^{\prime}+(\overline{x}_{0}-\overline{x}_{1})\cdot\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(x_{i}-\overline{x}_{0})^{\prime}$
$=V_{0}+\alpha_{0}(\overline{x}_{0}-\overline{x}_{1})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
Therefore,
$A-V_{1}=U_{0}+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}$
By Sherman–Morrison,
$(A-V_{1})^{-1}=U_{0}^{-1}-\alpha_{0}(1-\alpha_{0})\frac{U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
Then,
$\frac{\theta_{1}}{\alpha_{0}(1-\alpha_{0})}$ $=(A-V_{1})^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$=\left({1-\frac{\alpha_{0}(1-\alpha_{0})\cdot(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}}\right)U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$=\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})$
$\implies\theta_{1}^{\prime}(\overline{x}_{1}-\overline{x}_{0})$ $=\frac{\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
and
$\theta_{1}^{\prime}(\overline{x}-\overline{x}_{1})$ $=-(1-\alpha_{0})\theta_{1}^{\prime}(\overline{x}_{1}-\overline{x}_{0})$
$=-(1-\alpha_{0})\frac{\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}$
Hence,
$1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})=(1-\alpha_{0})\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})},$
and therefore
$\delta_{1}=\frac{\theta_{1}}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}=U_{0}^{-1}\alpha_{0}\left({\overline{x}_{1}-\overline{x}_{0}}\right)=U_{0}^{-1}\frac{1}{n}\sum_{i=1}^{n}(1-\pi_{i})\frac{\pi_{i}}{1-\pi_{i}}(x_{i}-\overline{x}_{0}),$
as claimed.
Now,
$\delta_{0}$ $=\frac{\alpha_{0}+\theta_{1}^{\prime}(\overline{x}_{0}-\overline{x})}{1-\alpha_{0}+((I-\Gamma_{1}^{\prime})^{-1}\alpha_{1})^{\prime}(\overline{x}-\overline{x}_{1})}$
$=\frac{\alpha_{0}-\theta_{1}^{\prime}\alpha_{0}(\overline{x}_{1}-\overline{x}_{0})}{(1-\alpha_{0})\frac{1}{1+\alpha_{0}(1-\alpha_{0})(\overline{x}_{1}-\overline{x}_{0})^{\prime}U_{0}^{-1}(\overline{x}_{1}-\overline{x}_{0})}}$
$=\frac{\alpha_{0}}{1-\alpha_{0}},$
as claimed. ∎
Remark D.1.
Here we consider the ramifications of (D.9) for the model-based ATE ($\tau_{\alpha_{0}}$). For some true design $\bm{\pi}^{*}$ to justify causal interpretation, we must have that this design is fractional linear:
$\pi_{i}^{*}=\frac{\theta_{0}+\theta_{1}^{\prime}(x_{i}-\overline{x})}{1-\Gamma_{2}^{\prime}(x_{i}-\overline{x})}$
for some $\theta_{0},\theta_{1},\Gamma_{2}$. Moreover, the choices $\theta_{0},\theta_{1},\Gamma_{2}$ needs to satisfy (D.9), which depends on the covariates $x_{1},\ldots,x_{n}$. We may thus consider the parameter space as
$\Theta(t;x_{1:n})=\left\{{(\theta_{0},\theta_{1},\Gamma_{2}):\text{ $\tau_{t}$ is a causal contrast under the corresponding $\bm{\pi}^{*}$ under $(\theta_{0},\theta_{1},\Gamma_{2})$ }}\right\}.$
Note that for $t=0$ and $t=1$, this parameter space does not depend on $x_{1:n}$ aside from the obvious support restriction. For instance, if $x_{1:n}$ and $\tilde{x}_{1:n}$ are two sets of covariate values, and
$\pi_{i}^{*}/(1-\pi_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(x_{i}-\overline{x}_{1})\geq 0\quad\tilde{\pi}_{i}^{*}/(1-\tilde{\pi}_{i}^{*})=\delta_{0}+\delta_{1}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}}_{1})\geq 0,$
then the corresponding parameters are in $\Theta(1;x_{1:n})\cap\Theta(1;\tilde{x}_{1:n})$—similarly for $t=0$. This is true by Proposition D.10.
However, for other $t$, and in particular for $t=\alpha_{0}$, it is possible that some $(\theta_{0},\theta_{1},\Gamma_{2})\in\Theta(\alpha_{0};x_{1:n})$, still does not belong to $\Theta(\alpha_{0};\tilde{x}_{1:n})$, even though no support restriction is violated
$\tilde{\pi}_{i}^{*}\equiv\frac{\theta_{0}+\theta_{1}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}})}{1-\Gamma_{2}^{\prime}(\tilde{x}_{i}-\overline{\tilde{x}})}\in[0,1]\text{ for all $i$}.$
It is not very difficult to come up with a numerical example. See LABEL:listing:ex. $\blacksquare$
Listing 1: Example for D.1
1import numpy as np
2
3n = 100
4x = (np.arange(n) / n) ** 3
5
6lin_form = 1 + (x - 1 / 2) * 0.25
7
8# Search for a valid pi* from this value
9pi_init = lin_form / (1 + 2 * lin_form)
10
11# Regression is y ~ 1 + w + x + wx
12z0 = np.c_[np.ones(n), np.zeros(n), x, np.zeros(n)]
13z1 = np.c_[np.ones(n), np.ones(n), x, x]
14
15gram = ((1 - pi_init[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
16 (pi_init[:, None, None]) * z1[:, :, None] * z1[:, None, :]
17).mean(0)
18lmbda = np.array([0, 1, 0, x.mean()]) # for the model-based ATE
19rho0 = lmbda @ np.linalg.inv(gram) @ z0.T
20rho1 = lmbda @ np.linalg.inv(gram) @ z1.T
21pi = -rho0 / (rho1 - rho0)
22
23# Iterate to find a fixed point
24for _ in range(1000):
25 new_gram = ((1 - pi[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
26 (pi[:, None, None]) * z1[:, :, None] * z1[:, None, :]
27 ).mean(0)
28 new_rho0 = lmbda @ np.linalg.inv(new_gram) @ z0.T
29 new_rho1 = lmbda @ np.linalg.inv(new_gram) @ z1.T
30 new_pi = -new_rho0 / (new_rho1 - new_rho0)
31 pi = new_pi
32
33# Decompose to find the fractional linear parameters
34A = ((x - x.mean()) ** 2).mean()
35v = (pi * (x - x.mean()) ** 2).mean()
36gamma1 = v / A
37gamma0 = (pi * (x - x.mean())).mean()
38a0 = pi.mean()
39a1 = gamma0 / A
40gamma2 = 1 / (v - v * gamma1 - gamma0 * gamma0) * (1 - a0 - v * a1 / gamma0) * gamma0
41theta0 = a0 - gamma2 * gamma0
42theta1 = a1 - gamma1 * gamma2
43
44print(theta0, theta1, gamma2) # 0.3254446337090774 0.014146176384094272 0.044263163071495276
45
46# pi $\in$ [0,1]
47assert (pi > 0).all() and (pi < 1).all()
48
49# the fractional linear representation is correct
50assert (
51 np.abs(pi - (theta0 + theta1 * (x - x.mean())) / (1 - gamma2 * (x - x.mean()))) < 1e-11
52).all()
53
54# Now, consider a new set of x’s, will show that the parameters theta0, theta1, gamma2 no
55# longer form a design justifying level independence for the corresponding estimand
56new_x = np.arange(n) / n
57new_z0 = np.c_[np.ones(n), np.zeros(n), new_x, np.zeros(n)]
58new_z1 = np.c_[np.ones(n), np.ones(n), new_x, new_x]
59new_design = (theta0 + theta1 * (new_x - new_x.mean())) / (1 - gamma2 * (new_x - new_x.mean()))
60
[…]
62assert (new_design > 0).all() and (new_design < 1).all()
63
64gram_new_design = ((1 - new_design[:, None, None]) * z0[:, :, None] * z0[:, None, :]).mean(0) + (
65 new_design[:, None, None] * z1[:, :, None] * z1[:, None, :]
66).mean(0)
67new_lmbda = np.array([0, 1, 0, new_x.mean()])
68new_rho0 = new_lmbda @ np.linalg.inv(gram_new_design) @ new_z0.T
69new_rho1 = new_lmbda @ np.linalg.inv(gram_new_design) @ new_z1.T
70
71# Failure of level independence
72print(
73 np.abs(new_design * new_rho1 + (1 - new_design) * new_rho0).max()
74) # Equal to 0.08293240507391064, not zero
Proposition D.11.
Let $n\geq 3$. Consider binary $W_{i}$, covariates $x_{i}\in\mathbb{R}^{d}$, and scalar outcomes. Suppose the true design is linear $\pi_{i}^{*}=\delta_{0}+\delta_{1}^{\prime}x_{i}$ for some $\delta_{0}\in\mathbb{R},\delta_{1}\in\mathbb{R}^{d}$. There is no regression $(\Lambda,z(x,w))$—where $\Lambda$ may [^47] This is to accommodate for estimands like the model-based ATT, where we may consider contrasts that depend on $\overline{x}_{1}=\sum_{i}\pi_{i}^{*}x_{i}/\sum_{i}\pi_{i}^{*}$ depend on $x_{1:n},\pi^{*}_{1:n}$—such that:
  1. (1)
    (Regression is linear in covariates) For all $m$ and all $w$, the $m$th entry of $z(x_{i},w)$ is of the form $a_{m}(w)+b_{m}(w)^{\prime}x_{i}$ for some fixed conformable $a_{m}(\cdot),b_{m}(\cdot)$.
  2. (2)
    ($\Lambda\beta$ is the ATE) The corresponding estimand $\Lambda\beta\in\mathbb{R}$ is equal to the ATE, regardless of the configuration of $d,x_{1:n},\delta_{0},\delta_{1}$ (such that $\pi_{i}^{*}\in[0,1]$ for all $i$)
Proof.
It suffices to show that for certain given choices of $x_{1:n},\delta_{0},\delta_{1}$, there is no regression that satisfies (1) and estimates the ATE.
Suppose $x_{i}\in[-1,1]$ is scalar $(d=1)$ and suppose $\left\{{x_{1},\ldots,x_{n}}\right\}$ includes $\left\{{-1,0,1}\right\}$. Set $\delta_{0}=1/2,\delta_{1}=1/4$ so that $\pi_{i}^{*}\in(0,1)$ for all $i$.
Fix some covariate transform $z(x_{i},w)$. By assumption, its $m$th entry is of the form $a_{mw}+b_{mw}x_{i}$. Suppose this regression estimates the ATE. In particular, it estimates a causal contrast. Thus
$\pi_{i}^{*}\rho_{i}(1)+(1-\pi_{i}^{*})\rho_{i}(0)=0.$
Because the estimand is the ATE, we also have that $\omega_{i}=\rho_{i}(1)\pi_{i}^{*}=1$. Hence, by (D.5), we have that for all $i$
$1+\left({1-\frac{1}{\rho_{i}(1)}}\right)\rho_{i}(0)=0\iff\rho_{i}(1)\rho_{i}(0)=\rho_{i}(1)+\rho_{i}(0)$
Now, the potential weights are of the form
$\rho_{i}(w)=\Lambda G_{n}^{-1}z(x_{i},w)=c_{0w}+c_{1w}x_{i}$
where the coefficients $c_{0w},c_{1w}$ may depend on $x_{1:n},\pi_{1:n}^{*}$, but are the same for all $i$. As a result, for the regression to estimate the ATE, we must have
$(c_{00}+c_{10}x)(c_{01}+c_{11}x)=(c_{00}+c_{10}x)+(c_{01}+c_{11}x)$
for $x=-1,0,1$: That is,
$c_{00}c_{01}=c_{00}+c_{01}$ ($x=0$)
$c_{00}c_{01}+c_{10}c_{11}+c_{10}c_{01}+c_{00}c_{11}=c_{00}+c_{01}+c_{10}+c_{11}$ ($x=1$)
$c_{00}c_{01}+c_{10}c_{11}-c_{10}c_{01}-c_{00}c_{11}=c_{00}+c_{01}-c_{10}-c_{11}$ ($x=-1$)
Now, adding the last two equations yield
$c_{00}c_{01}+c_{10}c_{11}=c_{00}+c_{01}=c_{00}c_{01}\implies c_{10}c_{11}=0\implies\text{ one of $c_{10},c_{11}$ is zero}$
Suppose $c_{10}=0$. Then the bottom two equations simplify to
$c_{00}c_{11}=c_{11}.$
This implies that $c_{00}=1$ or $c_{11}=0$. $c_{00}=1$ contradicts the $x=0$ equation. Thus, we must have $c_{11}=0$. Suppose instead we assumed $c_{11}=0$, by a symmetric argument we would derive $c_{10}=0$. Thus, $c_{10}=c_{11}=0$.
However, in that case, $\rho_{i}(w)=c_{0w}$ does not depend on $x$. For (D.5) to hold (for $x=-1,0,1$) then, we must have $c_{0w}=0$, since $\pi_{i}^{*}$ does depend on $x$. This means $\rho_{1}(w)=0$ regardless of $x$, but this contradicts the fact that we estimate the ATE ($\rho_{1}(1)\pi_{i}^{*}=1$). Thus, this regression does not estimate the ATE. ∎
D.6. Additional results for panel specifications
Proposition D.12 (One-way fixed effects).
Consider the regression $Y_{it}=\alpha_{i}+\tau W_{it}+\epsilon_{it}$. If the span of $\mathcal{W}\subset\left\{{0,1}\right\}^{T}$ excludes $1_{T}$, then the set of implicit designs is empty.
Proof.
Let $\dot{\mathbf{w}}=\mathbf{w}-1_{T}\frac{1_{T}^{\prime}\mathbf{w}}{T}$. Then, by Theorem D.1, the potential weights are equal to
$\rho_{i}(w)=\frac{\dot{\mathbf{w}}}{\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\dot{\mathbf{w}}^{\prime}\dot{\mathbf{w}}}.$
Thus, if $\pi_{i}$ corresponds to some implicit design $\bm{\pi}$,
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}-1_{T}\frac{1_{T}^{\prime}\mathbf{w}}{T}}\right)=0.$
In particular, this implies that $1_{T}$ is in the linear span of $\mathcal{W}$. By assumption, this is not the case. Therefore the set of implicit designs must be empty. ∎
Intuitively, this is because one-way fixed effects fail to account for the case where the treatment path correlates with underlying potential outcomes. If there are no treatment effects whatsoever, but the baseline potential outcome correlates with the treatment path (e.g., later potential outcomes tend to be larger in staggered adoption), then one-way fixed effects would in general estimate a nonzero coefficient, violating level irrelevance.
These results contrast with the results in Arkhangelsky and Imbens (2023), where the one-way fixed effect regression admits a design-based interpretation under within-unit random assignment of $W_{it}$. To reconcile, Arkhangelsky and Imbens (2023) consider a sampling-based setup and the unconfoundedness restriction
$W_{it}\indep(Y_{it}(0),Y_{it}(1))\mid\frac{1}{T}\sum_{s=1}^{T}W_{is},$
which is their (2.2) in our notation. Under staggered adoption, for instance, the associated propensity score $\operatorname{P}\left({W_{it}=1\mid\frac{1}{T}\sum_{s=1}^{T}W_{is}}\right)$ is always degenerate, since $\frac{1}{T}\sum_{s=1}^{T}W_{is}$ perfectly distinguishes which treatment path unit $i$ is assigned. Conversely, if $W_{it}$ is randomly assigned within a unit and that permuting the time index results in valid counterfactual assignments, then $\mathcal{W}$ is large enough to contain $1_{T}$.
Proposition D.13 (Event-study and more complex TWFE).
Consider the regression $Y_{it}=\alpha_{i}+\gamma_{t}+\beta^{\prime}f_{t}(\mathbf{W}_{i})$ with target estimand $\bm{\tau}=\Lambda\beta$ under some design $\bm{\pi}^{*}$. We have:
  1. (i)
    Whether or not $\bm{\tau}$ satisfies level irrelevance under the true design $\bm{\pi}^{*}$, one proper and Gram-consistent implicit design is
    $\pi_{i}(\mathbf{w})\equiv\pi(\mathbf{w})\equiv\frac{1}{n}\sum_{j=1}^{n}\pi_{j}^{*}(\mathbf{w}).$
    Correspondingly, one estimated implicit design sets $\hat{\pi}_{i}(\mathbf{w})$ to be the empirical frequency of treatment path $\mathbf{w}$ for all $i$.
  2. (ii)
    Suppose $\Lambda=I$. If any column $k$ of $f_{t}(\mathbf{W}_{i})$ is such that the multiset
    $\left({\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}:\mathbf{w}\in\mathcal{W}}\right)$
    (a) contains the zero vector at most once, (b) has nonzero vectors that are linearly independent, and (c) does not span the subspace $\left\{{c1_{T}:c\in\mathbb{R}}\right\}$, then $\overline{\pi}$ is also the unique implicit design.
Proof.
  1. (i)
    By Theorem D.1, it suffices to compute potential weights for the two-way residualized specification. That is, let
    $\mathbf{z}_{i}(\mathbf{w})=\begin{bmatrix}f_{1}(\mathbf{w})^{\prime}\\ \vdots\\ f_{T}(\mathbf{w})^{\prime}\end{bmatrix}\equiv\mathbf{z}(\mathbf{w})$
    be the covariate transform. Note that the population residual of projecting $\mathbf{z}_{i}(\mathbf{W}_{i})$ on unit and time fixed effect is
    $\ddot{\mathbf{z}}(\mathbf{w})$ $=\mathbf{z}(\mathbf{w})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w})-\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+1_{T}\frac{1_{T}^{\prime}}{T}\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})$
    $=\mathbf{z}(\mathbf{w})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w})-\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+1_{T}\frac{1_{T}^{\prime}}{T}\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\pi(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})$
    Thus the potential weights are
    $\bm{\rho}_{i}(\mathbf{w})=\Lambda\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\ddot{\mathbf{z}}(\mathbf{w})^{\prime}$
    Note that
    $\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=\Lambda\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\underbrace{\left({\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}}\right)}_{=0}=0.$
    Therefore $\bm{\pi}$ defined by $\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$ is a valid implicit design. It is proper by definition. Now, let $\dot{\mathbf{z}}(\mathbf{w})$ be the within-transformed covariate transform for this regression (which includes the time fixed effects), which does not depend on $i$. Note that the Gram matrix is
    $G_{n}(\bm{\pi}^{*})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\dot{\mathbf{z}}(\mathbf{w})^{\prime}\dot{\mathbf{z}}(\mathbf{w})]=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\dot{\mathbf{z}}(\mathbf{w})^{\prime}\dot{\mathbf{z}}(\mathbf{w})=G_{n}(\bm{\pi}).$
    Thus $\bm{\pi}$ is Gram-consistent.
    The argument for the estimated implicit design is analogous. The sample residuals of $\mathbf{z}(\mathbf{W}_{i})$ regressing on unit and time fixed effects is
    $\mathbf{z}(\mathbf{W}_{i})-1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{W}_{i})-\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\hat{\pi}(\mathbf{w}^{\prime})\mathbf{z}(\mathbf{w}^{\prime})+\sum_{\mathbf{w}^{\prime}\in\mathcal{W}}\hat{\pi}(\mathbf{w}^{\prime})1_{T}\frac{1_{T}^{\prime}}{T}\mathbf{z}(\mathbf{w}^{\prime})$
    where $\hat{\pi}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(\mathbf{W}_{i}=\mathbf{w})$. The estimated potential weights are proportional to these residuals evaluated at $\mathbf{W}_{i}=\mathbf{w}$. Thus $\hat{\bm{\pi}}$ defined by $\hat{\pi}_{i}(\cdot)=\hat{\pi}(\cdot)$ solves (B).
  2. (ii)
    In this case, any implicit design $\tilde{\pi}_{i}$ needs to satisfy
    $0=\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}(\mathbf{w})\bm{\rho}_{i}(\mathbf{w})=\left({\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]}\right)^{-1}\left({\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}}\right)$
    Since $\Lambda=I$ and $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\ddot{\mathbf{z}}(\mathbf{w})^{\prime}\ddot{\mathbf{z}}(\mathbf{w})]$ is positive definite by Theorem D.1, this implies that
    $\sum_{\mathbf{w}\in\mathcal{W}}\tilde{\pi}_{i}(\mathbf{w})\ddot{\mathbf{z}}(\mathbf{w})^{\prime}=0$
    Inspecting this expression, we have
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\mathbf{z}(\mathbf{w})=1_{T}\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\frac{1_{T}^{\prime}\mathbf{z}(\mathbf{w})}{T}$
    Column $k$ of this expression is in turn
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=1_{T}C_{k}$
    for some scalar $C_{k}$. The left hand side is in the span of $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}$, which excludes $1_{T}$ by assumption. Thus,
    $\sum_{\mathbf{w}\in\mathcal{W}}(\tilde{\pi}_{i}(\mathbf{w})-\pi(\mathbf{w}))\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=0.$
    Since at most only one $\mathbf{w}$ leads to $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}=0$, and nonzero vectors $\begin{bmatrix}f_{1k}(\mathbf{w})\\ \vdots\\ f_{Tk}(\mathbf{w})\end{bmatrix}$ are linearly independent, we have that
    $\tilde{\pi}_{i}(\mathbf{w})=\pi(\mathbf{w})$
    for all but one $\mathbf{w}$ by linear independence. Since both designs sum to 1, they also equal for remaining $\mathbf{w}$.
Proposition D.14 (Forbidden comparisons).
Consider the TWFE specification $Y_{it}=\alpha_{i}+\gamma_{t}+\tau W_{it}$ under staggered adoption ($\mathbf{w}_{t}\leq\mathbf{w}_{t+1}$ for all $\mathbf{w}\in\mathcal{W}$ and all $t\in[T]$). Suppose the treatment time is randomly assigned so that $\pi_{i}^{*}(\mathbf{w})=\pi^{*}(\mathbf{w})$. If there are two treatment paths with one being never-treated, $\mathbf{W}=\left\{{0,\mathbf{w}}\right\}$, then $\bm{\rho}_{it}(\mathbf{w})\geq 0$ for all post treatment periods $\mathbf{w}_{t}=1$ and $i\in[n]$. Otherwise, in all other configurations where $|\mathcal{W}|\geq 2$, there exists a choice of $\pi^{*}(\mathbf{w})$ such that the corresponding $\bm{\rho}_{it}(\mathbf{w})<0$ for some treatment path $\mathbf{w}$ with positive assignment probability ($\pi^{*}(\mathbf{w})>0$) and some post-treatment period $t$ ($\mathbf{w}_{t}=1$).
Proof.
The potential weight can be computed in closed form. For some $V>0$,
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left[{\mathbf{w}_{t}-\frac{1^{\prime}}{T}\mathbf{w}-\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi^{*}(\tilde{\mathbf{w}})\tilde{\mathbf{w}}_{t}+\sum_{\tilde{\mathbf{w}}\in\mathcal{W}}\pi^{*}(\tilde{\mathbf{w}})\frac{1^{\prime}}{T}\tilde{\mathbf{w}}}\right].$
When $\mathcal{W}=\left\{{0,\mathbf{w}}\right\}$, then for a post-treatment $t$,
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}(1-\pi^{*}(\mathbf{w}))(1-1^{\prime}\mathbf{w}/T)\geq 0.$
Otherwise, let $\mathbf{w}$ be the treatment path with the earliest adoption date, and consider $t=T$. By assumption, $\mathcal{W}$ contains a path that adopts later than $\mathbf{w}$. Then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({(1-\pi^{*}(\mathbf{w}))(1-1^{\prime}\mathbf{w}/T)-\sum_{\tilde{\mathbf{w}}\neq\mathbf{w}}\pi^{*}(\tilde{\mathbf{w}})\left({1-1^{\prime}\tilde{\mathbf{w}}/T}\right)}\right)$
Pick $\pi^{*}(\mathbf{w})=1/2=\pi^{*}(\tilde{\mathbf{w}})$ where $\tilde{\mathbf{w}}$ adopts later than $\mathbf{w}$. Then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({1/2(1-1^{\prime}\mathbf{w}/T)-1/2\left({1-1^{\prime}\tilde{\mathbf{w}}/T}\right)}\right)=\frac{1}{2}V^{-1}1^{\prime}(\tilde{\mathbf{w}}-\mathbf{w})/T<0$
since $\tilde{\mathbf{w}}$ adopts later than $\mathbf{w}$. ∎
Proposition D.15 (Time-varying covariates by interacting unit-specific covariates with time fixed effect).
Suppose $\mathbf{W}$ satisfies staggered adoption and excludes an always-treated unit. Consider the TWFE specification
$\mathbf{Y}_{it}=\alpha_{i}+\gamma_{t}+\tau\mathbf{W}_{it}+x_{i}^{\prime}\eta_{t}.$
$\tau$ is a causal contrast only if there is some $\delta_{0}(\cdot),\delta_{1}(\cdot)$ such that
$\pi_{i}^{*}(\mathbf{w})=\delta_{0}(\mathbf{w})+\delta_{1}(\mathbf{w})^{\prime}x_{i}$
for all $i\in[n],\mathbf{w}\in\mathcal{W}$.
Proof.
Let $\mathbf{w}(t)$ be the path in $\mathcal{W}$ that is first treated in time $t$. The never-treated unit is denoted by $\mathbf{w}(\infty)$. By excluding the always-treated path, $\mathbf{w}(t)\in\mathcal{W}$ implies that $t\geq 2$.
By the same argument leading to (C.5), we have that if $\tau$ is a causal contrast, then for some $\rho_{i},\lambda_{t},\kappa_{t}$, for all $t\in[T]$ and $i\in[n]$,
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\mathbf{w}_{t}=\rho_{i}+\lambda_{t}+\kappa_{t}^{\prime}x_{i}$
For $s\geq 2$ a treatment time, consider taking the difference of the above equation between $s$ and $s-1$. Because
$\mathbf{w}_{s}-\mathbf{w}_{s-1}=\begin{cases}1,&\text{ if }\mathbf{w}=\mathbf{w}(s)\\ 0&\text{otherwise},\end{cases}$
we have
$\pi_{i}^{*}(\mathbf{w}(s))=\lambda_{s}-\lambda_{s-1}+(\kappa_{s}-\kappa_{s-1})^{\prime}\mathbf{x}_{i}\equiv\delta_{0}(\mathbf{w}(s))+\delta_{1}(\mathbf{w}(s))^{\prime}x_{i}.$
This shows the desired representation for all $i\in[n]$ and all $0\neq\mathbf{w}\in\mathcal{W}$. Lastly,
$\pi_{i}^{*}(0)=1-\sum_{\mathbf{w}\in\mathcal{W}\setminus\left\{{0}\right\}}\pi_{i}(\mathbf{w})$
can also be written as $\delta_{0}(0)+\delta_{1}(0)^{\prime}x_{i}$. ∎
D.6.1. Imbalanced panels
To introduce our result on imbalanced panels, suppose units are observed in some subperiod $\mathcal{T}_{i}\subset\left\{{1,\ldots,T}\right\}$. We consider the TWFE specification
$\mathbf{Y}_{it}=\alpha_{i}+\gamma_{t}+\tau\mathbf{W}_{it}+\epsilon_{it}$
over $i=1,\ldots,n$ and $t\in\mathcal{T}_{i}$. The objects $\mathcal{T}_{1},\ldots,\mathcal{T}_{n}$ are considered fixed and observed.
[…]
Proposition D.16.
Let the population regression specification be characterized by (D.6.1) under $\bm{\pi}^{*}$, where $\tau$ is the coefficient of interest. Suppose $\mathcal{W}$ satisfies staggered adoption and excludes always-treated units. Assume further that there is a commonly observed period $\mathcal{T}\subset\bigcap_{i=1}^{n}\mathcal{T}_{i}$ such that the treatment paths have rich variation in $\mathcal{T}$.
Let $\overline{\pi}(\mathbf{w})\equiv\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})$. Let $Q_{i}(\mathbf{w})=\frac{\sum_{t\in\mathcal{T}_{i}}\mathbf{w}_{t}}{|\mathcal{T}_{i}|}$ be the proportion of treated periods within unit $i$’s observed period for treatment path $\mathbf{w}$. Then an implicit design exists if and only if $\bm{\pi}^{*}$ is uncorrelated with the missingness pattern in the sense that for all $t\in[T]$:
$\underbrace{\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\sum_{i:t\in\mathcal{T}_{i}}(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))}_{\sum_{i:t\in\mathcal{T}_{i}}\mathbb{E}_{\mathbf{w}\sim\overline{\pi}}[(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))]}=\underbrace{\sum_{\mathbf{w}\in\mathcal{W}}\sum_{i:t\in\mathcal{T}_{i}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))}_{\sum_{i:t\in\mathcal{T}_{i}}\mathbb{E}_{\mathbf{w}\sim\pi_{i}^{*}}[(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))]}.$
When this happens, the implicit design is unique and satisfies $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ for all $i$.
Proposition D.16 gives a necessary and sufficient condition for an implicit design to exist, and characterizes the implicit design when it does. In short, an implicit design exists if and only if the missingness patterns are uncorrelated with the treatment timing in a particular sense (D.16). When it exists, the implicit design is the same for all units. Therefore, if $\bm{\pi}^{*}$ justifies $\tau$ as a design-based estimand in the sense of 2.2, then $\bm{\pi}^{*}$ must equal to the implicit design and thus must describe random assignment.
The condition (D.16) describes a situation where for any given time period $t$, among units for whom $t$ is observed, the average expected de-meaned treatment path under $\bm{\pi}^{*}$, $\mathbb{E}_{\mathbf{w}\sim\pi_{i}^{*}}[\mathbf{w}_{t}-Q_{i}(\mathbf{w})]$, is equal to its analogue under the random assignment $\overline{\pi}$. This condition is difficult to satisfy unless the design is constant $\pi_{i}^{*}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ in the first place. As a result, Proposition D.16 shows that TWFE is fragile to imbalanced panels.
Proof of Proposition D.16.
For a given unit $i$ and $t\in\mathcal{T}_{i}$, let $L_{i}=|\mathcal{T}_{i}|$ and let $R_{i}(\mathbf{w})=\sum_{s\in\mathcal{T}_{i}}\mathbf{w}_{s}$ so that $Q_{i}(\mathbf{w})=L_{i}^{-1}R_{i}(\mathbf{w})$. The covariate transform—where we specify the unit fixed effect via within transformation—is
$z_{it}(\mathbf{w})^{\prime}=[\dot{\mathbbm{1}}_{it1},\ldots,\dot{\mathbbm{1}}_{itT},\dot{\mathbf{w}}_{t}]$
where
$\dot{\mathbbm{1}}_{its}=\begin{cases}0&s\not\in\mathcal{T}_{i}\\ 1-\frac{1}{L_{i}}&s=t\\ -\frac{1}{L_{i}}&s\in\mathcal{T}_{i}\setminus\left\{{t}\right\}\end{cases}$
is the demeaned time dummy and
$\dot{\mathbf{w}}_{t}=\mathbf{w}_{t}-Q_{i}(\mathbf{w}).$
By Theorem D.1, it suffices to analyze the potential weights of regressing $Y_{it}$ on $\ddot{W}_{it}$, where
$\ddot{W}_{it}=\dot{W_{it}}-\sum_{s=1}^{T}\delta_{s}\dot{\mathbbm{1}}_{its}=\dot{W_{it}}-\sum_{s\in\mathcal{T}_{i}}\delta_{s}\dot{\mathbbm{1}}_{its}.$
and $\delta_{s}$ are the population projection coefficients of $\dot{W}_{it}$ on $\dot{\mathbbm{1}}_{its}$. Let $V\equiv\frac{1}{n}\sum_{i=1}^{n}\sum_{t\in\mathcal{T}_{i}}\mathbb{E}[\ddot{W}_{it}^{2}]$, which is strictly positive by Theorem D.1. The potential weights are then
$\bm{\rho}_{it}(\mathbf{w})=V^{-1}\left({\dot{\mathbf{w}}_{t}-\sum_{s\in\mathcal{T}_{i}}\delta_{s}\dot{\mathbbm{1}}_{its}}\right)=V^{-1}\left({\dot{\mathbf{w}}_{t}-\delta_{t}+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}}\right).$
An implicit design $\bm{\pi}$ exists if and only if there is some $\pi_{i}(\cdot)$ and $\sum_{\mathbf{w}}\pi_{i}(\mathbf{w})=1$ such that
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\bm{\rho}_{it}(\mathbf{w})=0\text{ for all $i$ and $t\in\mathcal{T}_{i}$}.$
This is further equivalent to that for all $i$, $t\in\mathcal{T}_{i}$,
$\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}_{t}-\frac{R_{i}(\mathbf{w})}{L_{i}}}\right)=\delta_{t}-\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}.$
Suppose first that such a $\bm{\pi}$ exists. Fix $t\in\mathcal{T}$ and consider two units $i,j$, where $\mathcal{T}_{i}\cap\mathcal{T}_{j}\supset\mathcal{T}$ by assumption. Then, by (D.6.1), we have that
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\mathbf{w}_{t}-\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\frac{R_{i}}{L_{i}}+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{j}(\mathbf{w})\mathbf{w}_{t}-\sum_{\mathbf{w}\in\mathcal{W}}\pi_{j}(\mathbf{w})\frac{R_{j}}{L_{j}}+\frac{1}{L_{j}}\sum_{s\in\mathcal{T}_{j}}\delta_{s}.$
Thus, rearranging, we have for all $t\in\mathcal{T}$,
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi_{i}(\mathbf{w})-\pi_{j}(\mathbf{w}))\mathbf{w}_{t}=M(i,j)$
where $M(i,j)$ does not depend on $t$. Viewed as linear combinations for vectors in $\mathcal{W}_{\mathcal{T}}$, the left-hand side collecting over all $t\in\mathcal{T}$ lies in the span of $\mathcal{W}_{\mathcal{T}}$. The right-hand side lies in the span of the constant vector on $|\mathcal{T}|$ dimensions. By assumption, these two spans contain only the zero vector in common. Thus, both sides are equal to zero:
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi_{i}(\mathbf{w})-\pi_{j}(\mathbf{w}))\mathbf{w}_{t}=0\text{ for all $t\in\mathcal{T}$}.$
By linear independence of vectors in $\mathcal{W}_{\mathcal{T}}$, we conclude that $\pi_{i}(\mathbf{w})=\pi_{j}(\mathbf{w})$. Therefore, the implicit design must be constant across units:
$\pi_{i}(\mathbf{w})=\pi(\mathbf{w})$
for some $\pi(\cdot)$.
Next, we show that if $\bm{\pi}$ exists, then it must equal to $\overline{\pi}(\cdot)$. The fact that $\delta_{t}$ are projection coefficients is equivalent to the following orthogonality conditions holding for all $t\in[T]$: For all $s\in\left\{{1,\ldots,T}\right\}$,
00 0 $=\frac{1}{n}\sum_{i=1}^{n}\sum_{t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\ddot{\mathbf{w}}_{it}\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}-\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{it\ell}}\right)\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}}\right)\dot{\mathbbm{1}}_{its}-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{t\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{it\ell}\dot{\mathbbm{1}}_{its}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\sum_{t\in\mathcal{T}_{i}}\left({\mathbf{w}_{t}-R_{i}(\mathbf{w})/L_{i}}\right)\mathbbm{1}(t=s)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{t\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\mathbbm{1}(t=\ell)\dot{\mathbbm{1}}_{its}$ ($\dot{\mathbf{w}}$ and $\dot{\mathbbm{1}}_{i}$ sum to zero over $t\in\mathcal{T}_{i}$)
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-R_{i}(\mathbf{w})/L_{i}}\right)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}\dot{\mathbbm{1}}_{i\ell s}$
$=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-R_{i}(\mathbf{w})/L_{i}}\right)-\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}(\mathbbm{1}(\ell=s)-1/L_{i})$
Rearranging, we have that the orthogonality condition is equivalent to
$\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})\left({\mathbf{w}_{s}-Q_{i}(\mathbf{w})}\right)=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\left({\delta_{s}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}}\right)$
for all $s\in[T]$.
Pick $t\in\mathcal{T}$ and consider the corresponding (D.6.1). Note that $t\in\mathcal{T}_{i}$ for all $i$, and hence
$\delta_{t}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}+\sum_{\mathbf{w}\in\mathcal{W}}\underbrace{\left({\frac{1}{n}\sum_{i=1}^{n}\pi_{i}^{*}(\mathbf{w})}\right)}_{\overline{\pi}(\mathbf{w})}\mathbf{w}_{t}-\frac{1}{n}\sum_{i=1}^{n}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})Q_{i}(\mathbf{w}).$
Given (D.6.1), we can plug in
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}=\sum_{\mathbf{w}\in\mathcal{W}}\pi(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)+\frac{1}{L_{i}}\sum_{s\in\mathcal{T}_{i}}\delta_{s}.$
Rearranging, we obtain that
$\sum_{\mathbf{w}\in\mathcal{W},\mathbf{w}\neq 0}(\pi(\mathbf{w})-\overline{\pi}(\mathbf{w}))\mathbf{w}_{t}=M^{\prime}(i)$
for some $M^{\prime}(i)$ that does not depend $t\in\mathcal{T}$. Again, the left-hand side, over $t\in\mathcal{T}$, is in the span of $\mathcal{W}_{\mathcal{T}}$, whereas the right-hand side is in the span on $1_{\mathcal{T}}$. As a result, both sides are equal to zero by assumption. Since vectors in $\mathcal{W}_{\mathcal{T}}$ are linearly independent, we have that
$\pi(\mathbf{w})=\overline{\pi}(\mathbf{w})\text{ for all $\mathbf{w}\neq 0$}\implies\pi(\mathbf{w})=\overline{\pi}(\mathbf{w})\text{ for all $\mathbf{w}\in\mathcal{W}$}.$
Now, we plug $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ into (D.6.1), and then plug the expression into (D.6.1). We obtain that for all $t\in[T]$,
$\sum_{i:t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w}))=\sum_{i:t\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w})).$
This is (D.16). Thus, we proved that if an implicit design exists, then the above condition holds and it must be uniquely equal to $\overline{\pi}(\cdot)$.
For the reverse direction, let us assume that (D.16) holds. For a given $t$, set
$\delta_{t}=\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\mathbf{w}_{t},$
and thus
$\delta_{t}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}=\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right).$
This means that $\pi_{i}(\mathbf{w})=\overline{\pi}(\mathbf{w})$ satisfies (D.6.1). Thus $\overline{\pi}(\cdot)$ is an implicit design upon checking that (D.6.1) defines the projection coefficients.
Summing over $i$ where $t\in\mathcal{T}_{i}$:
$\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\left\{{\delta_{s}-\frac{1}{L_{i}}\sum_{\ell\in\mathcal{T}_{i}}\delta_{\ell}}\right\}=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\overline{\pi}(\mathbf{w})\left({\mathbf{w}_{t}-Q_{i}(\mathbf{w})}\right)=\frac{1}{n}\sum_{i:s\in\mathcal{T}_{i}}\sum_{\mathbf{w}\in\mathcal{W}}\pi_{i}^{*}(\mathbf{w})(\mathbf{w}_{t}-Q_{i}(\mathbf{w})),$
where the last equality follows from (D.16). This verifies the orthogonality condition (D.6.1) for $t$. Therefore, our choice of (D.6.1) does indeed equal the projection coefficients. This concludes the proof. ∎